A Complete Bibliography of Publications in Probability in the Engineering and Informational Sciences

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

16 October 2023 Version 1.14

Title word cross-reference

(m, n): F [LZ00]. (r, s) [LZ00]. (s, S) [CF06]. (S − 1, S) [SD97]. 1 [vdH87]. 1_{n} [LSSZ16]. 2 [CHL94, vdH87]. ^{X} [CCC05, Eco99, GV10, Dud97, ET87, Fra05, PR08]. ^{X_{1}} [CLC98]. ^{X_{2}} [CLC98]. Y [GV10]. _{K} [Lu09]. _{m} [LSSZ16]. α [SAF22]. C_{p} [Kha99]. C_{pm} [Kha99]. C [MBP07]. L [Bha99]. · [PMB96]. D [GSV16, LN92]. ℓ_{∞} [SSS04]. f [CL06]. G [BA18, Cag17, Fou16, FM17, GS92, Gel93, Mar16, MK17, Mat17, Mat19b, Wan21, DP21, GZDW21, Pek21]. G/M [SZ05]. i [Bar08]. ∞ [CSP12, Eli08, PZ04, PMB96, VY08]. k [AR97, AAC14, Ano21a, BFR99, BPT90, BW14, CH88, CCH99, CHC00, CHJ95, DZH12, HH16, HH00, HY90, Hwa00, KM88, KX10, Kor03, LYF16, LZ00, LHY19, Men94b, POdCW21, RKFI21, SY15, Tam98, WAvHD09, Yao89]. l_{1} [ABI22]. l_{∞} [SSS02, SSS08]. LS [WXWH18]. m [FS23, KM16, VT08]. N [Yec88, Ano21a, BFR99, CH88, CCH99, CHC00, CHJ95, DL10, FS23, HH00, HY90, Jia19, KM88, KX10, Kor03, LYF16, MI10, POdCW21, RKFI21, SY15]. n: F [BPT90, CHL94, DZH12, Yao89]. P [DK13, AEAN23, Ali23, HZ05, HZ06, XH09]. PAR(p) [MD06]. Q [CH15, EK09]. R [Men94b]. σ [Ang97, DR95a, Kha99]. t [ANFPLCSL05, Lin90b, Lin92a, Lin94a, SA88]. × [BvH93]. TP_{2} [DLP90]. U [BR95]. z [Tag00].

-ary [KM16]. -Distribution [ANFPLCSL05]. -limited [BW14, WAvHD09]. -mixtures [SAF22]. -network [DP21, DL10, MK17, Mat17, Mat19b]. -networks [Pek21, GS92, Gel93, Wan21, BA18, Cag17, Fou16, Mar16]. -Orderings [DLP90]. -out-of- [Ano21a, BFR99, BPT90, CH88, CCH99, CHC00, CHL94, CHJ95, DZH12, FS23, HH00, HY90, KM88, KX10, Kor03, LYF16, LZ00, POdCW21, RKFI21, SY15, Yao89, vdH87, Men94b]. -person [MI10]. -Policy [LN92]. -queues [FM17]. -Record [AR97, Tam98, AAC14]. -retrial [GZDW21]. -series [EK09]. -set [Hwa00]. -spacings [AEAN23, Ali23, HZ05, HZ06, XH09]. -spectrum [GSV16]. -Spherical [SSS02, SSS04, SSS08]. -Stage [Yec88]. -Statistics [BR95]. -swap [CL06]. -Tail [Lin92a]. -th [LHY19]. -urn [CH15]. -Within- [LZ00].

/1 [Lu09]. /1/K [GV10]. /1/N [CG96]. /D/1 [CCC05, MC11]. /D/c [CLC98, Fra05]. /G [CG96, GV10]. /G/1 [Dud97]. /g/1/n [PR08]. /G/c [ET87]. /GI [LSSZ16]. /GI/ [PMB96]. /M/1 [Eco99]. /R/1 [CCC05].

0 [Sha98]. 0-201-63481-3 [Sha98].

1 [AGC07, AW96, BPLD21, BB96, BBS97, BdBS20, CCC05, CGM21, CRP93, Dud97, Eco99, EZB06, Elc94, FK96, GZDW21, LN92, LRW09, Lu09, MN05, MK23, MC11, NQ01, ÒLC95, Ros06a, Sha88, SKA09, SK11, SNS19, TS07, Wan02, WW05, XLSZ04, dK89]. 1-Type [PV11]. 1996 [Sha98].

24 [MY11]. 2nd [Joe12].

3 [Sha98].

70th [Pek16].

802.11 [FM16].

abandonment [PM17]. abandonments [RB10, ZCC23]. Aberrant [OGP95]. absolute [BC18]. Absorbing [Han03, MZ17]. Accelerated [ACG95, Lin19]. Access [Rig98, RS90, Ros88b, DP21]. accessing [Cho99]. Account [JS96]. Accurate [HBG21]. Achieve [ALR00]. Achilles [vZS04]. Acknowledgment [Lyn03]. Active [BEP88, FL20a, MMD11, YWL22]. activity [TAM^{+}21]. Actuarial [NPS04, BSF21]. Acyclic [Mad93]. Adam [Man96a]. Adan [Haj15]. Adaptive [RS94b, FS23]. Addison [Sha98]. additional [AY08a]. Additive [LM88, BM11, Gup16, SH19]. adjustable [MAS06]. Adjusted [LK07, IBI08]. Adjustments [Van96b]. Admission [CHL18, HS07, Zie96, Asl22, BB06, BCD^{+}14, CFKP99, LAF99, Pre02]. Advertizing [AL91]. affine [Xu23]. after [BK06, Jia19]. Age [CS92, KG89, Van92b, vZS04, CV22, FW15, LK23, WC00]. Age-Replacement [Van92b]. Ageing [GG96, MMS93, KM19]. Agents [BBS96, BC98]. aggregate [CMH13]. Aggregation [Buc97, NCJ18]. Aggregation/Disaggregation [Buc97]. Aging [BFR99, BN11, CS94, DFS10, MC91, RS98b, SS87, BBCN01, CV22, FPSL03, FM17, HM21, MFN17, RSS09, SH23, TA11]. aid [LFNC23]. airplanes [BEFK04]. al [Pal90]. Alan [Man96a]. Aleliunas [Pal90]. Algebraic [Ano21a, POdCW21]. Algorithm [Bla98, Buc97, DS87, HM94, McN89, Sad96, SR90, WK88, Win97, YZ89, Bor00, DHR00, GM16, HBKK03, LLY22, SVL02]. Algorithmic [Sch96]. Algorithms [Fil98, Reg90, Tem94, JL19]. alliance [Kim22]. Allocation [BEP88, BS91, Fro95, KM88, Li94, LT00, Men94b, WB96, FL20a, HBV10, LZY23, LP23, NCJ18, Wan17a, WWZZ23, WH20]. Allocations [XZP18, MMD11]. Allowance [TS96]. Almost [CHYZ93, YZZ23, CK20]. ALOHA [Spi91, YM92]. ALOHA-system [Spi91]. Alternate [BSS90]. alternating [VA05, XDZ15]. Alternative [Bri14, CCdO01, FG89a, HL92, LC97, Lin90b]. American [FWGD00, MCL22, WWL17, Yan20]. American-style [FWGD00]. Ammunition [Sat97]. among [AKS11, HSB23, KK00, MIK23]. amounts [DKB22, Lu21]. Amplifiers [Smi97b]. Analysis [AL09, AFP88, Ald93, Ang89, ADS88, BLBPV17, BDM15, Bro91, BK03b, CG96, CJ96, Fis91, FH94, FH95, HM87, HKL94, HBvdMB18, HM94, Jia19, LM88, LK08, Man96a, Mat19a, MK19, MP97, NS17, RL93, SB13, Sha88, SY92b, TV89, BW14, Bri14, BHCKS04, CMZ17, CZ03, CCC05, CH87, DP21, DGG22, FM16, FWGD00, GRSV14, HY21, HP03a, HS13, IVA15, KT01, KN00, KM12, LLY22, LXYL21, Miy00, SVL02, SNS19, TP21, VM12, WFM09, XYZC23, ZP16, ZCJ04, ZDS20]. Analytic [BVA13, GW12, BH23, KKL17]. Analytical [AdHBvO14, Wan17b, WÇC22, KN21]. Analytically [CSP12, CGM21]. analyze [Cal10]. Analyzing [YT23, Ros06b]. and/or [PT94]. Andrzej [Bax96]. Annealing [Ald87, CK88, DR97, KT97, RS94b, VBAK89, RBM02]. ant [Gut03]. ante [Ros11]. applicability [AILR08]. Application [CVB05, CFFH98, CK88, GS93, HS92, Miy00, PV11, RMS94, RS94a, RTW09, WWS18, YSM06, AAR11, AY08b, DF08, Fen22, GV11, HH16, HS13, KMS15, LST^{+}22, NK07]. Applications [AFP88, Alz88, BR95, BWY93, DK13, DLP90, GS23, He89, HM20, NZP18, SSS02, SY92a, Smi97b, AHOS23, BSF21, BB00, Bro06, Cag17, FL13, FPOP08, KSvH14, KA04, LS11, LHL08, LK09, PL20, SKA09, TK08, VM04, VHBSW22, Wan02, YHZ06, Joe12]. Applied [ANFPLCSL05, Bax96, BvH93, RW92b, CL06, RB10]. Applying [RL01]. Approach [BvH93, CCdO01, Fan96, GI95, KW89, TV91, VBAK89, AL15, BMSB20, CSL09, Chi23, GK21, GT10, GS12, GT07, GB22, HvdW08, Hei01, HL04, KN21, KKL17, Miy02, PLS23, WH20, Sch96]. approaches [AILR08]. Approximate [DvdLR11, RB10, RB12, Sad96, Van92a, BMSB20, WFM09]. Approximating [AL15, Dod89, Lin92a, OLÖ97, PM17, Ros87a, Ros88a, dK89, vD88, Ang04, Olu06]. Approximation [AdHBvO14, AH96, BB97, BB98, DJ07, Lin90b, Lin90a, Lin92b, Lin94a, Lin95, Pla92, RS94a, SGF20, Smi97a, Van88, Whi88, YZ89, BvHTvO18, BH23, GM16, Idr23, mKK22, MCL22, Röl22, YDJ23]. Approximations [AW95, Lin94b, LK08, MT93, PR94, Ros87b, SA88, Tak96, YS89, CW22, Dim23, FS99, FS20, Pen15, Tij08]. arbitrarily [CK20]. Arbitrary [JS96, NQ01, PdB22]. Archimedean [SH23]. Areas [SA88]. Arising [BC98, HvdHR09, JS89, Li94, BCZZ19, HSB23]. Armed [BK03b, HPR96, BKK17, CK15, EFG17]. Arnold [Joe12]. Arrangement [Li94]. Arrays [WWS18]. Arrival [Bri88, Dav95, MS91, Miy92, MvD97, Ott93a, Ott93b, ÒLC95, TSW97, Fra05, HKK10, Hor01, LWZ16, LW14, NS17, Ros06b]. Arrivals [CP93, CW90a, DJ07, DY90a, ET87, Gou94, PV11, SY92b, Zie96, CY21, EM19, FL22, HLW16, LZ19, PW12a, PW12b, Rig11, SC03]. ary [KM16]. Asian [CW23, PZ19]. Aspects [BC94, Bax96]. Assemblies [DH94]. Asset [AL91, LP23]. assets [SXWW22, YSM06]. assign [AW12]. Assignment [DY90a, Dav95, HK92a, HKL94, Rig87, AAD07, BJ13, BH02, FH13, Haj15, LKH14, Rig11]. Assignments [LM87]. Associated [WWS18, WJ18, BBHK10, Cov10, MK23, Tan06, YHZ06]. Assortativity [BKL22]. assortment [PdB22]. asymmetric [CZ06]. Asymptotic [Ang89, Ang95, BK03b, CBL93, CMH13, Hun91, JZY22, Khe08, LLY22, LM88, MY11, MNH15, MMY20, Rei90, SNS19, Tan06, TV89, VdK87, VMB00, Yan98, ZP16, Zaz07, ZYL10, Ang06, CK20, FH17a, Miy02, WWZ22]. Asymptotically [BKK17, WJ18, XC10, TYZ19]. Asymptotics [ABD23, HL14, KMT18, Man96b, PZ21, Dim23]. asynchronously [AKR^{+}04]. ATM [KW94]. ATO [DJ07]. atoms [RB06]. attachment [FWH21, Zha21a]. attack [BPR21, LnG22]. Author [BZ13a]. Autocorrelated [TSW97]. Automata [Buc97]. Autoregressive [YZ03]. auxiliary [LCYY22]. Availability [Ang01, KWA96, Smi97a, VR89, vdH87, HM15, Mi99, Mi21]. Available [FR92]. Average [CW90a, Ege94, HPR96, MP97, Mi00, Ros88c, Ros99a, Sen89, Sen93a, Sen93b, ARK03, CXY12, Lu10, Sen01, Wak00, XC09]. Average-Reward [HPR96]. averages [SZ99, ZYSZ21]. Aversion [LM88, SBZ16]. aware [BA18]. axiomatic [WH20]. Azuma [Ros95].

B [Joe12, GV10, MO95]. B-Trees [MO95]. Back [Ano87a, Ano87c, Ano87e, Ano87g, Ano88a, Ano88c, Ano88e, Ano88g, Ano89a, Ano89c, Ano89e, Ano89g, Ano90a, Ano90c, Ano90e, Ano90g, Ano91b, Ano91d, Ano91f, Ano91h, Ano92a, Ano92c, Ano92e, Ano92g, Ano93d, Ano94e, Ano95e, Ano96d, Ano97d, Ano09a, Ano09c, Ano09e, Ano09g, Ano10a, Ano10c, Ano10e, Ano10g, Ano11a, Ano11c, Ano11e, Ano11g, Ano12a, Ano12c, Ano12e, Ano12g, Ano13a, Ano13c, Ano13e, Ano13g, Ano14a, Ano14c, Ano14e, Ano14g, Ano15a, Ano17g, Ano18a, Ano18c, Ano18e, Ano18g, Ano19a, Ano19c, Ano19e, Ano19g, Ano20a, Ano20c, Ano20e, Ano20g, Ano21b, Ano21d, Ano21f, Ano22a, Ano22d, Ano22f, Ano15c, Ano15e, Ano15g, Ano16a, Ano16c, Ano16e, Ano16g, Ano17a, Ano17c, Ano17e]. backorders [FL20b]. backup [Rig02, SH00]. Backward [FTC98, Tar17]. bag [HM16]. bag-of-visual [HM16]. Balakrishnan [Kha13, Bur13, Sha13]. Balance [CK88, Bro13, CSL09, HBV10]. Balanced [CHYZ93, Mah13a]. balancing [FS99]. Balking [LK08, GZDW21, GZ09]. Ball [FGR10]. balls [Mah13a]. Band [AK96, Per97, Reg90, BPS14]. Bandit [BK03b, HPR96, Vil18, Wei88, BKK17, CY10, PRZ22]. Bandits [RS98a, CK15, EFG17, Vil16]. bandwidth [DHR00]. Bandwidths [CW95]. Based [BFR99, BB98, BN18, DG94, DM20, KMS20, MS93, NBR97, Rot93, Sad96, SP98, XWTW18, AW12, AAES10, AAK19, CHS87, CHL03, DDK22, DO03, DMM16, FPSL03, FC22, Gut03, Haj15, HKK10, HM16, KA16, KKJ22, MG21, QWGV23, RB12, Roy02, WGI99, YBL22b, ZAE19]. Basu [NK07]. Batch [Bou96, DJ07, ET87, Gel93, Gou94, MvD97, PV11, YZ98, EM19, Fra05, MK17, Mat17, PW12a, PW12b, PC14, Pin07, XMBG02, vdWY03]. batch-service [vdWY03]. batching [KR01]. Bathtub [BJS99, BLS12, NH04, NBR13]. Bathtub-Shaped [BLS12, NH04]. Bay [Dov87]. Bayes [Bla95, BK03b, CS90, YLB96]. Bayesian [GT07, HM20, JC18, SK20, VBAK89]. be [Gra08, ZDS20]. Behavior [AW96, CH87, DR97, MO95, TV89, WJ18, YS89, BBCN01, CMH13, Fra19, OK08, WWZ22, dCHQ09]. Behaviors [SC23, MMY20, SY15]. Behaviour [Hun91]. Belief [Koh88]. Bernoulli [BKL22, HPR96, MMY20, SK11, SAS89, TZS22]. Best [FR92, TS96, LHY19]. Best-Case [FR92]. Best-Choice [TS96]. Better [DG91a, Ege94, JL98]. Between [BZ05, AK96, BFRR01, Cov10, DFS15, DKM16, FRR02, HY21, HZ06, HYL07, HT96, Lin02, Lu10, MPH13, NFMFSA20, SZ99, TYZ19, ZCJL23]. beyond [Nic16]. Bias [LAF99, HH16]. Biased [Alz88, BHP05, Olu99, TE15]. biases [MS06]. biasing [Bro06]. Bicriterion [FK96]. bidegree [BKL22]. Bidimensional [Lef96, FNC20, SGF20]. bike [TP21]. bike-sharing [TP21]. bilateral [MZ17]. Bin [CW90a]. binary [ALM07, BMR05b, BK03a, GSV16, HLM16, Mah14, Nic16, YBL22a]. binary/multi [YBL22a]. binary/multi-state [YBL22a]. Binomial [Abr93, CHH90, Yao88, EK09, KPDR16, SS20, XH11, YDJ23]. Birnbaum [CHC00, ABJ22, CCH99, Gup15, HKA22, Men04]. Birth [DOW93, KR87, Kij94, SSS02, Smi97b, CZ03, JC18, KPDR16, KPH10, KS01, Kyr04, MT16, MZ17, WWZ22, ZCJ04]. birth-and-death [WWZ22]. birth-death [MZ17]. birth/immigration [KPDR16]. birth/immigration-death [KPDR16]. birthday [KM12, ME14, Pek16]. Bit [But89]. Bivariate [GM14, Lin95, Roy90, FL13, GKB13, Gup15, Mon13, NK07, NCJ18]. Black [Dub98, RSZ93, Che09, Che11, Che12]. blackouts [DCL^{+}05]. Blackwell [Kha07, Ros95]. Block [CS92, NK07]. Blockchain [BPR21, Kim22]. Blocking [Bou96, CPS89, LWZ16, Nak01, VMB00]. Blocks [GMW14]. Blood [vSvDdK^{+}20, BLBPV17]. blood-testing [BLBPV17]. Bonds [NZP18]. Bonus [OKA22]. Book [Bax96, Joe12, Man96a, Sch96, Sha98, GD18]. Bootstrap [BS14]. borne [OE10]. borrowing [FL05]. borrowing/storage [FL05]. Bottleneck [HBvdMB18]. bouncing [LKG19]. Bound [Abr93, Din95, Reg90, Ros98, Van92a, Win97, Yao88, NL07, PdB22]. boundaries [MZ17]. Boundary [Han03, SA18, BdlPS15]. Bounded [Rot93, ABD23, BOT09, SD22b]. Bounding [MBP07, Ros06a]. Bounds [BH02, BCS92, BK03a, FR92, GKPR99, Kla00, KL98, LL98, Mei91, MvD97, Miy04, Miy05, RS94a, Sad98, SP98, VS87, YXH19, YS89, vD88, vDTW88, BNT06, BK09, Bro14, CT23, CY06, Eco99, Fro11, FS20, HP03b, KM19, MCL22, MP02, Men02, Nak01, Olu06, PK06, Röl22, THM04, XLSZ04, YZL00, Zha99]. Bowl [DG91a]. Box [RSZ93]. Branch [Win97]. Branching [Bha87, DG94, HV95, Ott93a, Wei88]. Branching-Process-Based [DG94]. Brazilian [MD06]. Breakdown [Fro91, PR88, DKSS06, KN21]. Breakdowns [Gla91, CSZ03, CZ06]. Bregman [SK20]. bridge [LP15]. Bridges [HvdHR09]. Broadcast [RS95, FM16]. Brownian [CLS21, DM03, Fen22, HvdHR09, KPS03, LKG19, PH23, PB06, Per97, Rao15]. Bruijn [CLS21, KA16]. BSDE [Tar17]. bucket [Miy00]. Budget [CH88, XZP18, PS10]. Buffer [Bar96, Gou94, MT93, TV89, Tij92, VdK87, CM00, HS07, Tij08, Zha99]. Buffered [Jaf92]. buffering [FM16]. Buffers [DG91a, KW94, WB96, AAD08, AY08b, CBBG20]. bug [DDK22]. Bugs [PS10]. Building [GMW14]. Bulk [GV10]. Burn [CS90, BJS99]. Burn-In [CS90, BJS99]. Burst [Miy92, SY92b]. bursty [DA03]. Busy [AW95, HL92, AGC07, Bar08, CGM21, PR08, Pek99b]. Busy-Period [AW95].

C [Bax96, Joe12, Sch96, Bar08, CLC98, ET87, Fra05, SS92]. C-MSP [CSP12]. cache [Hil99]. Calculation [VR89]. Calculations [GS98, Kol94]. Calculator [Lin90a, Lin92b]. Call [vDTW88, Bhu09, CFKP99, DK22, HJM09, KNN15, RB10, RB12]. can [Gra08, ZDS20]. candidate [LHY19]. Candidates [DY90a]. Capability [Kha99]. capacitated [KTÇ14, PdB22]. capacities [CFKP99]. Capacity [AF92, AF97, Bre90, DG91b, VT92, WB96, BMSB20, CT02, DO03, FZ15, Kad17, SI04, SLZ04]. Capital [WH20, NCJ18, Xu23]. Capture [LOdBP87, MW94, YM92]. Capture/Recapture [LOdBP87]. care [FKL02]. career [Çağ16]. Carlo [EGL92, FH95, Hei01, Lom94, RW92b]. carousel [BVA13, LAWZ01, wWW04]. carrier [OE10]. carrier-borne [OE10]. Cartesian [DO16]. cascade [ZBA05]. Cascades [Smi97b]. cascading [CH12, DCL^{+}05, DCN05, Lef06]. Case [Fil98, FR92, Ott93a, RSZ93, GKL23, SLZ04, Wak00]. Cases [Gut97, CGM21]. cash [CSL09, SH19]. Casino [Mei14]. casualty [AZR08]. Cat [NZP18]. Catalan [Reu14]. catastrophe [Kyr04, PGZ22]. catastrophes [CZ03, FM17, KPDR16]. catastrophic [Kwo14]. caterpillar [ZD19]. Cauchy [BPG12]. Cayley [SZF18]. cellular [BH02, VMB00]. censored [MG21]. center [DK22, HJM09]. Centers [KS92, Bhu09, KNN15, LKH14, RB10, RB12]. Central [CZL96, RS95, SAS89, Ste87, BDM15]. central-limit [BDM15]. certain [Lin02]. CEV [PZ19]. Chain [Ald87, ACG95, Fil91, GI88, PV11, VS87, CSvD02, Fra99, GT07, JCDJ22, KM07, LY04, Ros13, SC03, SLZ04, SKL12, YSM06]. Chains [AW89, AFP88, BBS88, Bro91, CK88, DKK94, DR97, FHJ94, FTC98, GV94, HJ95, HV95, Ken94, Kij92, MY11, Pla92, PT93b, Ros87a, Ros88a, Sen91, Sen93b, SR90, TV91, Van92a, YS89, ZYL10, vD88, EK09, HHvU07, Hua20, KS12, Kul04, LW19, Lor18, MP02, MBP07, SZF18, TYZ19, Wan21, WWZ15, YZL00, YHZ06, Zaz07, ZYSZ21, dCHQ09]. Change [ZCA18, EK21]. changing [XDZ15]. Channel [BS91, Bre90, Ros88b, vD95, BH02, TAM^{+}21, WNT23]. Channels [RB05a, AAK19, RB05b]. chaos [GCDB20, Bax96]. Chapman [Man96a]. Characteristics [GG96, dK89, BBCN01, DO03, GKB13, SAdK04, WFM09]. Characterization [HN19, Keb94a, MS06, Roy90, SS87]. Characterizations [AK05, BSS90, GN14, MH10]. characterized [AHOS23]. Charts [HT96, Ros99a]. Chebyshev [GMZ10, Mon07]. Chen [Ros23]. Chi [BF91, Lin94b]. Chi-Square [BF91, Lin94b]. Chichester [Sch96]. Choice [TS96, Tur98, Ros13]. Choices [McN88]. Circuit [Whi88]. Circuit-Switched [Whi88]. Circular [MY11, ZYL10, TYZ19]. claim [DKB22, GS23, OKA22]. claims [CL23, DKB22, FL22, QWGV23, SXWW22]. Class [CLC98, EM19, Li94, MBP07, PT94, SB13, Whi89, AKP05, Bha99, BKM10, CF13, CLS21, CHL18, Dim17, GKB13, IVA15, KS12, LX06, Mit18, SD22b, Vil16, WN15, Zha21a, ZCJL23, ZH09]. Classes [LT00, Mad93, MC91, PP89, AK05, Asl22, BS08, BLPS05, CV22, IASB23, Li05, Mat19a, MK19, Mat19b, MK21, RSS09, WC00]. classical [GS23, MK23]. Classification [Roy02, HM16, HBG21]. clearing [AZR08, Jia19, KPS03, WÇC22]. clearings [Sta07]. Clicks [LE94]. climbing [CM08]. Clinical [RTW09, Vil18]. Clocked [Jaf92]. clones [HL01, HL04]. Closed [AY94, BvHvO97, MP02, MBP07, AL15, CGM21, mKK22, PG22, PGZ22, WXL20]. Closed-Form [MBP07, MP02, AL15, CGM21, mKK22, PG22, PGZ22, WXL20]. Closure [GK88]. Cloud [EBTD22, LnG22, Wan17a]. clumpy [wWW04]. Cluster [GS93, CM05, IBI08]. clustering [ABI22, DK09, IBI08]. co [XXN21]. co-jumps [XXN21]. coaster [GV11]. coefficient [BC18]. Coefficients [BS87, FR92]. cognitive [Wan17a]. Coherent [BI88, BEP88, But89, Men94a, ASKZ21, BFRR01, dCBB22, FL20a, GS12, GZY22, HM21, IASB23, LSW21, LZY23, TK08, YWL22]. Cold [VR89]. collapse [KL10]. collecting [BPR08, CR07]. columns [Hwa00]. combination [BR17, NK05]. Combinations [Mei91, KX12]. Combinatorial [Ald87, GS12, Ste87]. combinatorics [HLM16]. Combining [GS98]. Come [BvHvO97]. Comments [Bur13, Kha13, LY13, Sha13, Xu13]. commitment [CK15]. common [CX23, SXWW22, Yam99]. Communication [BS91, KSM88, MW94, Ros88b, TV89, CL06, KT01, Nak01]. Communication/Computer [KSM88]. Communications [Man96a, Man96b]. Commute [AAKN09]. Comparing [BS98, GV94, Men94a, Men04, KX12]. Comparison [BR97, DF08, DK88, Hor01, MT93, PS93, AJK01, Eco03, LSW21, NDZ22, YBL23a]. Comparisons [BCZZ19, BN18, CCH99, CHC00, DDL14, HP00, HH00, KX08, Lin02, ZZZ21, AEAN23, BCMS23, BLRS01, BMR05a, FPOP08, FH17b, HM21, KX07, Koc23, LL03, NS22, Yu21, ZB12, ZH07]. Compensation [BvH93]. Compensators [He89]. competing [AP99]. competition [CR21]. Complete [Gou94, Sie93, WWS18, GZDW21]. completely [Bro09]. Complex [ÇÖ87]. Complexity [BI88]. Component [HM87, YLB96, dCBB22, GS12, KPS03, LYF16, YWL22]. Components [DR95b, Ege94, KX08, KX10, KSM88, Men94b, Wan18a, ABJ22, BKB22, DLB12, DDL14, FL20a, GSV16, GZY22, HM21, HKA22, KX07, LWL19, Men04, NDZ22, NFMFSA20, Rig02, SH23, SY15, SGF20, YWL22, ZDZ20, Zha11]. composed [BMR05b]. Compound [PR04, Ang04, BSF21, CL23, GS23, GLW11, KPS03, NWB13, NYZ04, SZ03, SZ05, XYZC23, XDZ15, YXZ23]. Compression [Sad96, Sad98, KM16]. comprising [ABJ22]. Computation [MBP07, LST^{+}22]. Computational [BI88, CCC05, PL20]. computations [LS11]. compute [HL04, JL19]. Computer [KSM88, LM87, Rig98, Wee90, SH00]. Computing [AB93, AK96, BPvdW91, BK95, CFKP99, Sen01, SR90, Ang01, BR17, GS12, JCDJ22, LnG22, Man96a]. concatenation [SX22]. concave [Rol05]. Concavity [BSF21, SS87, MXH18, XMH21]. Concentrated [MT22]. concentration [LMH12]. concept [FPSL03]. Concepts [CSK89]. Concerning [Pal90]. Concurrency [KS92]. Conditional [Koh88, LW13, XH08, ZYWN06, ASZ17, BBHK10, HJK07, LMZ22, ZY22]. Conditionally [LG89]. Conditioning [Ros88c]. Conditions [GI88, Lyn99, PZ04, Eco03, Lyn03, MFN17, VM04]. Conductivity [Kon02, Kon03]. Congestion [vDTW88, Asl22]. congestion-sensitive [Asl22]. Conjecture [Mül98, Hil99, Zha99]. conjectured [AS00b]. conjectures [AS00a]. Conjugate [MR95]. connection [KP12]. Connectivity [CL91, SAS89, BM02a]. Consecutive [BPT90, CH88, CCH99, CHC00, CHL94, CHJ95, DZH12, DH87, DH94, HY90, SZ87, Yao89]. Consecutive- [CH88, CCH99, CHC00, CHL94, CHJ95, HY90].

Consecutive-2 [DH87, DH94, SZ87]. Conservation [BS93, KW94]. Consistency [WXWH18, XWTW18]. Consistent [GHS91]. Constant [BL87b, CS94, WAI95, Win97, YZ03, CJ96, FNC20, XYZC23]. Constant-Service [WAI95, Win97]. Constrained [Sen91, Sen93b, LC07, Sen01]. Constraint [Wak98, BK09, Bro14, BKK17, LP23, YXH19]. Constraints [CW90a, KS92, LR97, BM02a, FOF07, RBM02, SLZ04]. consumption [SBZ16]. containing [BPG12]. Context [Vil18]. contingent [SXWW22]. Continuous [BBP20, BBS96, BL87a, BL88, CS93, ER95, Fil91, Kij92, PT93b, Ros87a, Ros88a, Tak90, YS89, vD88, Ang01, Bro99, CSvD02, Eli05, FW15, HY21, Sta07]. Continuous-State-Space [Tak90]. Continuous-Time [ER95, Fil91, PT93b, Ros87a, Ros88a, YS89, vD88, Bro99, CSvD02, Eli05, HY21]. continuously [BGDR03]. Continuum [BL89]. contractual [KTÇ14]. contributions [BK06, OST06]. Control [Ang97, BLPS05, BL88, CRP93, CHL18, DJ07, DR95a, Dud97, FZ93, HT96, Key90, PP89, Per97, RB05a, Sen89, YZ98, ZCA18, sADL02, BPLD21, BPS14, BB06, BY99, BP01, BAWD12, Bor00, GHvdL07, GT10, HvdW08, HMW09, HS07, Kyr03, LAF99, LSSZ16, MAS06, ÖK14, Pan08, PC14, RB05b, Ros99a, RB10, Voi09, WS06, XMBG02, XC09, Xu23, vdL11]. Controlled [Sai98, EBTD22]. Controls [Zie96, BEFK04]. Convergence [AH96, Ang04, CHL03, Din95, GV90, KT97, PMB96, WWS18, YZ89, ZL02, BL02, BPG12, Gut03, RBM02]. Converges [GI88]. Converters [HvdHR09]. Convex [FPOP08, MS91, BB00, SI04]. Convexity [Bha87, DP95, Koo04, MS93, SS87, SY92a, AAC14, DLU99, HvdL03, XMH21, YHZ06]. Convexly [Lyn97]. Conveyor [BCG89]. convolution [Bel11, MXH18]. convolutions [AKS11, MHZ10]. Cooled [DKK94, DR97]. cooperative [Bru16, KSvH14]. Coordinated [HKS07, PC14]. coordination [HAA22]. Copula [ANFPLCSL05, NS22, KA16, SH23]. copulas [AAK19]. Correction [Ano91a, KK06, KX08, Mal23, Miy05, RB05a]. Corrections [Ano94a]. Correlated [CCP90, CY21, LKH14, MMY20, Mon13]. Correlation [Wan18b]. Corrigenda [CHC00]. Corrigendum [JLAE18, Ano21a, AO18, PW12b]. Cost [ER95, Sen89, Sen93a, Sen93b, Wak98, WAI95, BHP05, BKK17, CXY12, FL20b, LSW21, PS10, THM04]. Costs [CF06, FY11, LR97, ADGZ09, CSL09, DO03, FW15, Gut14, Pan08, ZCC23]. count [JLL21]. Countable [DF90, Spi15, ZYSZ21]. Counter [HS92]. counterexample [HPS99, RL99]. Counterexamples [AS00a, Zha99]. counterparts [Zog23]. counterparty [CX23, WXL20]. counting [CF13, SS20]. Coupled [Spi91, Tak90, Dim17, HP03a]. coupled-processors [HP03a]. Coupling [Ang95, FTC98, GW96, CS03, WGI99]. Coupon [BPR08, CR07]. Cournot [CVB05]. covariance [LMZ22]. Covariances [Mah23, HYL07]. Cover [Ano87b, Ano87a, Ano87d, Ano87c, Ano87f, Ano87e, Ano87h, Ano87g, Ano88b, Ano88a, Ano88d, Ano88c, Ano88f, Ano88e, Ano88h, Ano88g, Ano89b, Ano89a, Ano89d, Ano89c, Ano89f, Ano89e, Ano89h, Ano89g, Ano90b, Ano90a, Ano90d, Ano90c, Ano90f, Ano90e, Ano90h, Ano90g, Ano91c, Ano91b, Ano91e, Ano91d, Ano91g, Ano91f, Ano91i, Ano91h, Ano92b, Ano92a, Ano92d, Ano92c, Ano92f, Ano92e, Ano92h, Ano92g, Ano93a, Ano93b, Ano93c, Ano93d, Ano94b, Ano94c, Ano94d, Ano94e, Ano95b, Ano95a, Ano95c, Ano95d, Ano95e, Ano96a, Ano96b, Ano96c, Ano96d, Ano97a, Ano97b, Ano97c, Ano97d, Ano98a, Ano98b, Ano98c, Ano98d, Ano09b, Ano09a, Ano09d, Ano09c, Ano09f, Ano09e, Ano09h, Ano09g, Ano10b, Ano10a, Ano10d]. Cover [Ano10c, Ano10f, Ano10e, Ano10h, Ano10g, Ano11b, Ano11a, Ano11d, Ano11c, Ano11f, Ano11e, Ano11h, Ano11g, Ano12b, Ano12a, Ano12d, Ano12c, Ano12f, Ano12e, Ano12h, Ano12g, Ano13b, Ano13a, Ano13d, Ano13c, Ano13f, Ano13e, Ano13h, Ano13g, Ano14b, Ano14a, Ano14d, Ano14c, Ano14f, Ano14e, Ano14h, Ano14g, Ano15b, Ano15a, Ano17h, Ano17g, Ano18b, Ano18a, Ano18d, Ano18c, Ano18f, Ano18e, Ano18h, Ano18g, Ano19b, Ano19a, Ano19d, Ano19c, Ano19f, Ano19e, Ano19h, Ano19g, Ano20b, Ano20a, Ano20d, Ano20c, Ano20f, Ano20e, Ano20h, Ano20g, Ano21c, Ano21b, Ano21e, Ano21d, Ano21g, Ano21f, Ano22b, Ano22a, Ano22c, Ano22e, Ano22d, Ano22g, Ano22f, Ano23a, PR97, Ano15d, Ano15c, Ano15f, Ano15e]. cover [Ano15h, Ano15g, Ano16b, Ano16a, Ano16d, Ano16c, Ano16f, Ano16e, Ano16h, Ano16g, Ano17b, Ano17a, Ano17d, Ano17c, Ano17f, Ano17e]. Cox [Miy05, JQZD23, Miy04]. Coxian [Van88]. Credit [CX23, Bro99, DF08, LS11]. criteria [BJ13, BGG07]. Criterion [HPR96, Las98, CXY12, PC22, Wak00, XC09]. Critical [Sen93a, Yan20]. Criticality [Men94a, HKA22]. Crossing [BPT89, LE94, SA18, BdlPS15, PLS23]. CROSSREG [RL93]. Crystallizer [CH87]. Csiszár’s [Zog23]. cumulant [Pen15]. Cumulative [OLÖ97, SA18, dCBB22, CLN20, JZK20, KP12, Kay16, Whi00, Zog23]. currency [XXN21]. Current [HvdHR09]. curves [Sta11]. Customer [Elc94, FY11, HKL94, RW95, Asl22, FL05, GZ09, PKK13]. Customers [BBS97, HK92a, SB13, ST94, AHW10, AY08a, ADGZ09, CGM21, FM17, KN00, MK17, Mat17, Mat19a, MK19, Mat19b, MK21, ML01, THM04, ZAMB21]. cusum [WN15]. Cut [Men94a, HBKK03]. Cycle [BZ05, Nak01]. Cycles [Joh98]. cyclic [SAdK04]. Cyclical [CBM18]. cylinder [ABD23].

D [CCC05, CLC98, EZB06, Fra05, MC11, TS07, HvdHR09, Lu09]. D/A [HvdHR09]. Dam [BLP93, Kha88]. Dams [BZ05]. DAR [MC11]. Data [Hun91, LOdBP87, Ros87e, Rot93, Sad96, Sad98, ABI22, AY08b, GCDB20, KM16, MG21, NK07, Nad08, SKL12]. Dates [IwW97]. DC [LP23, PC22]. Deadlines [PT93a, Rig87, ST94, Xu91]. Death [DOW93, KR87, Kij94, Smi97b, CZ03, JC18, KPDR16, KPH10, KS01, Kyr04, MT16, MZ17, MOSU10, WWZ22, ZCJ04]. debugging [DDK22]. decay [Miy02]. Decaying [MCB18]. Decision [BBS96, BC98, BK95, CS90, CZB89, HV95, Sen89, Sen91, Sen93b, Wak98, Whi89, YZ98, BGG07, SX22, Wak00, vdL11]. decisions [SvNKC00]. decline [LK23]. decomposability [HM19]. Decomposable [Chy91]. Decomposition [JL19, LK09, HvdW08]. Deconvexification [ER95]. Decrease [FHJ94]. dedicated [AAD07]. Deep [Kol94, TAM^{+}21, Yin21]. Deep-Penetration [Kol94]. default [BJ23, CX23, LQ15]. defects [GK21]. deferral [FKL02]. defined [HS13]. degradation [FC22, UKM12]. degradation-based [FC22]. Degree [DM20, FMR17, KM16, DK09, FH17a, Mah13b, SD22b, ZD19]. Degree-Based [DM20]. Delay [AdHBvO14, Bri88, RS90, Ros88c, SP98, Xu91, YM92, vDTW88, ADGZ09, FW15, GZ09, KKJ22, Wan02]. Delay-Dependent [Bri88, FW15]. delay-risk [GZ09]. Delay-Variance [Ros88c]. Delayed [VT92, ZYSZ21]. Delays [BAWD12, CY10, HLZ23]. Demand [Arr87, BL87b, SD97, BA02, GTDKÖ09, KTÇ14, KAR09, KLT12]. demands [AÖ12, GLW11, PKK13]. denial [LnG22]. Densities [Hes92, SSS02, SSS04, SSS08, JZY22]. Density [Bal92, RL93, HvdL03, TYZ19]. denumerable [YHZ06]. Departure [PW98]. Departures [GSW93, PMB96, RS89]. Dependence [BPT90, CS94, CSK89, DKSS06, FHJ94, HJ95, KK00, KK05, KK06, Koc23, Ott93a, DF08, HP00, HCX06, LHL08]. dependencies [OE10]. Dependent [Bri88, CP95, CH87, DR95b, Dub98, HM87, HV95, Kha88, RSZ93, Sch95, ABJ22, AAK19, APR20, BC13, BB06, BA02, BCD^{+}14, BKKP05, Bro09, CMH13, Che20, DKB22, DCN05, FW15, Fra13, Fra19, FNC20, GFH22, KS02, KA16, LYF16, LRW09, LZY23, MT16, NS17, NFMFSA20, NS22, OKA22, PW12a, PW12b, SH23, SGF20, WAvHD09, XWTW18, Yam00, YWL22]. depletion [ABEK17]. depreciation [CK15]. Depths [DM22]. Derivation [MO95]. Derivative [GHS91, Hei01]. Derivatives [FG91, YWC23]. design [HJM09, LFNC23, Lin19, LCYY22]. Designing [OKA22]. designs [Hwa00, HL01, HL04]. Detect [MR88]. Detecting [Abd16, LE94]. detection [SVL02, TAM^{+}21]. Deteriorating [IwW97, BGDR03, RL05, YK00]. Deterioration [KWA96, BA02, BY99]. Determining [Bel11]. Deterministic [IwW97, PC22, GHvdL07, Rat19, Tij08]. detrimental [Gra08]. Developments [BZ13b, Kha13, Xu13]. Deviation [PW98, CSvD02, GMZ10, HH16, KS01, SD22b, WWH18]. Deviations [DOW93, Man96a, RW92a, HP03a, HP03b, ME14]. device [VM12]. Devices [ÇÖ87, Lin19]. DHR [BL02]. Diagrams [Chy91]. difference [YT23]. differences [BC18]. different [BFRR01, EQ22, Hor01, YBL22a, YBL23a]. Differential [McN88, BHCKS04, CLS21, JZY22, KKJ22, ZCJL23]. Diffusion [Bar96, BL87a, DJ07, Lef01, SS20, Abd17, CL23, PS02, WWL17, Yan20]. Diffusions [AK96, Pla92, STI18]. dikes [SvNKC00]. Dimensional [DP93, Hes92, ABI22, BK03a, CMZ17, DO16, KMS15, STI18]. Dimensioning [BT95]. dimensions [KAJ07]. Direct [Lin94a]. Directed [Mad93, BJ19]. Directional [MS93]. directionally [Rol05, SI04]. Dirichlet [BHP05]. Disaggregation [Buc97]. disappearing [JQZD23]. discarding [DA03]. disciplinary [Çağ16]. Discipline [BvHvO97, Sha88]. Disciplines [ÒLC95]. discontinuous [XXN21]. discount [NYZ04, Tan06]. Discounted [ER95, Sen91, Wak98, CL23, HPS99]. Discounted-Cost [ER95]. discounting [vdWSvN08]. Discrete [AvER95, Ber96, Cai02, CCP90, CG96, CJ96, Din95, Dub98, Eli08, HPS08, Van96b, Bel11, BOR06, Bor00, CLSS13, FS20, GDRV10, HL14, Hor01, Kla00, MC11, PKK13, Wan17b]. Discrete-Time [CCP90, CG96, CJ96, Eli08, BOR06, Bor00, HL14, Hor01, MC11, Wan17b]. discretization [ARK03]. discretized [PZ21]. discrimination [EQ22]. Diseases [Vil18]. Disk [Rig98]. Dispersion [ANFPLCSL05, BHK03, AKS11]. Dispersion-type [BHK03]. Dispersive [LZH16]. Dispersiveness [KK06, KK05]. distance [Cov10, IBI08, LMZ22, YXH19]. Distances [Rot93, KP07]. distinct [Hwa00]. Distinguishable [Fro95, KM88]. Distorted [BN18, YZZ23]. Distributed [GSW93, Kol92, BKB22, Gut14, HM21, PS10]. Distributed-Event [GSW93]. Distribution [AW95, AR97, ANFPLCSL05, Arr87, Bro97, CH87, CHW89, DP93, HW92, Jaf92, LC97, LN92, Lin94a, Lin94b, OLÖ97, RMS94, RS94a, RS90, Roy90, SA88, Smi97a, Tam98, Van88, VS87, ZL02, vdH87, AAC14, Ang04, BHP05, Bel11, BK03a, Bro01, Cha99, CGM21, CM05, DXX16, DK09, EK21, EA06, FE01, GS23, GM14, Gup07, Gup15, JL19, KN21, LZH16, LZ10, LRW09, MG21, Mal22, Mal23, MP02, Mon07, Mon13, NK07, Nad16, Nad17, PL20, Ros06a, RKFI21, XC10, XDZ15, Zha21b, ZPD06, vdMR08]. Distributions [AvER95, Ald87, AFP88, And95, BLS12, BPT89, BS87, BN18, FGR10, GKL87, GM96, KG87, LG89, Lef96, Lin90b, Lyn99, Mah91, MBP07, Mei91, OLÖ97, PT93b, RS98b, AK05, ABJ22, AO18, AO19, AHOS23, BSF21, BS06, BB00, BPG12, BKL22, Bri14, BH23, Bro09, BK09, CCC05, FL13, FH17a, FS20, GK07, GKB13, HJK07, HM19, Kla00, LX06, Lin19, Lyn03, MHZ10, MZ17, MT22, MFN17, NWB13, NCJ18, NBR13, Olu99, Olu06, PS02, PO15, RSS09, Roy02, Sie07, TE15, WC00, ZYWN06, ZH09, ZY22]. divergence [SK20]. divergences [Zog23]. diversification [Mei14]. Diversity [BPT87]. dividend [HLZ23, MZS16, Xu23, YXZ23]. dividend-reinsurance [MZS16]. Division [RS90]. Domain [KSvH14]. dominance [Olu11, YZZ23]. Domination [BSS90]. done [ZDS20]. Dormant [Bor95]. Double [AK96, HL91, NBR97, Per97, BPS14, BDPS11, BPR21, GPM22, GZY22, Han19, WJYL22]. Double-Loop [HL91]. double-spend [BPR21]. doubly [SK23]. Downs [Cal97]. Downs-Thomson [Cal97]. downstream [DMvU07]. DPD [Zog23]. Drawing [Mah13a]. drawings [Che20]. DRHR [AK05]. Drift [DP93]. Driven [KR94, BKM10, CLS21, GKPR99, JZY22, Rao15, UKM12, XDZ15, YWZH21]. drought [NK07]. Dual [MD06, FL13, HLZ23]. Duality [DF90, PP89, Lor18]. Duality-Equivalence [PP89]. Duals [AS96]. Dubins [Mei14]. Due [IwW97, DDK22]. Dueling [PRZ22]. duration [BPR21, PR08]. during [CGM21]. Dynamic [AAD07, ADGZ09, Bhu09, CCdO01, HMW09, HS13, KM88, KSM88, LL90, MD06, PT94, XC09, ZCA18, vdWY03, AP99, BPLD21, BMSB20, Bro00, DKM16, JL19, KMS20, LSSZ16, NDB13, PM17, RB10, RB12, Sen01, YBL23b]. Dynamics [GD18, MW94, DGG22, FLL12, Bax96].

Each [CLC98, BK06, DDK22]. Earliness [For93, PT94, CZ06]. Early [SO94, SVL02]. Echelon [DvdH97, BvHTvO18, ZAMB21]. Echo [BR17]. economic [LnG22]. edge [Mah19]. Edition [Bax96, Joe12]. Editorial [AEL08, Ros87d]. EDoS [LnG22]. Effect [Cal97, Tur98, FWH21]. Effective [CW95, Nak01, AAKN09]. Effects [CH87, EA06]. Efficient [GW96, CMZ17, HBG21, Ros06b, WW05, WNT23]. eigenvalues [MT22]. Eigenvectors [Bro91]. elasticity [PT20]. Electric [SO94, KAA12]. electricity [CVB05]. elliptical [ZY22]. Elsayed [Sha98]. embedded [GT07]. Embedding [VR89]. embeddings [PB06]. emergency [BvHTvO18, Pre02]. emigration [ZCJ04]. Emphasis [BZ13b, Kha13, Xu13, HL14]. empirically [WGI99]. empty [Zaz04]. Encountered [VT08]. end [YS08]. Endogenous [FZ93]. Energy [BA18, DF21, Abd17, Cag17, HBG21, Kad17, Yin17, Zha21b]. Energy-aware [BA18]. energy-efficient [HBG21]. Engineering [MY11, Sha98]. entire [FZ15]. entropies [Kay16, Zog23]. Entropy [AR14, CHS87, CLS21, MG21, MPP93, Hua20, KP12, LZ11, LW19, Qiu19, TYZ19]. Entropy-based [CHS87, MG21]. Environment [AMK18, PP02, SS89, BPS06, SZF18, SD22b, UKM12]. environment-driven [UKM12]. Environments [ÇÖ87, FPOP08, Rat19]. EOQ [BP01, BPS06]. EOQ-type [BP01]. epidemics [LP15, OE10]. Equation [Fan96, FG91, JS89, Sen93a, BHCKS04, KKJ22, YT23]. Equations [Ban92, CT23, CLS21, JZY22]. Equilibrium [EGL92, GZDW21, HH96, HvdW97, HKK10, Jaf92, MI10, VS87, XXN21, BCR05, Bro00, CVB05, Gup07, SD22a]. equilibriums [Che12]. Equipartition [MY11, Yan98, ZYL10]. equipment [JA06, YK00]. equity [PGZ22]. Equivalence [BZ05, Mad93, PP89, Whi89, Bha99, BB00]. Equivalency [YBL22a, YBL23a]. Equivalent [MH10]. era [Mei14]. Erdos [Röl22]. Ergodic [SR90, Hua20]. Ergodicity [Spi91]. Erlang [AHW10, KSvH14, PC14, RS02, Zie96]. Erol [Çağ16, Pek16]. Erratum [MY11, DZB12b]. Error [CZL96, MvD97, Sad98, Van92a, YZL00, vD88, DDK22]. Error-Bound [Van92a]. Errors [DLR87, GM96, WXWH18, XWTW18, SHM09]. Errors-in-variable [WXWH18]. Establishing [PT93b]. Estimate [LOdBP87, vD95]. Estimates [GHS91, Lom94, Tan04]. Estimating [CZL96, DCL^{+}05, FG91, GK21, KW13, OGP95, PR97, Ros89, CM05, GN14, NS04]. Estimation [Ang97, Bro97, EGL92, Jun88, Kha99, Nay89, Rei90, YLB96, Bro01, HBKK03, LCYY22, Lu09, MS06, YWZH21, YXZ23]. Estimator [DR95a, LR91, XWTW18, GR12]. Estimators [DK88, WXWH18, FWGD00, Ros23, WW05, Zaz07]. Euclidean [ADS88]. European [mKK22, PG22, PH23, SXWW22, Wan18b]. Evacuation [BA18].

Evaluated [Vil18]. Evaluating [FS89, FR92, YBL23b]. Evaluation [KSM88, LZ00, AL09, VMB00]. Event [GSW93, ZL02, CMZ17, CM05, KMT18, Lag06, SZ99]. events [CF12, CMH13, Cov10, JQZD23, Mi00]. evidence [LK23]. Evolution [EGL92, Mah19]. evolving [Mah13b]. Exact [GS98, KS02, KMT18, LOdBP87, LK08, PS02, XDZ15, Dim23]. Examples [Bal92, DF90]. excess [KL15, Zar23]. exchange [KKJ22, PG22, WXL20]. Exchangeable [CLSS13]. excursions [Cal10]. Executive [WSX18]. exercise [Yan20]. Exhaustive [BCGW92, BAWD12]. Existence [CCdO01, PT93b, HPS99]. exit [RL03, RL05, SZ03, SZ05]. exotic [RS00]. Expansion [AF92, AL15, Ang06, HPS99, JZY22]. Expansions [NWB13, WN09, HHvU07, LMH12, MNH15]. Expectation [JL98]. Expectations [FG89b, KH91, YZZ23, ZY22]. Expected [Abr93, IwW97, JST01, MO95, BNT06, CSZ03, HL01, MK17, Mat19b, Nak01, PT20, Ros15]. expectiles [MNH15]. experimental [BHCKS04, GCDB20]. Experiments [LE94, Olu99]. Expiring [RS98a]. Explicit [CSP12, HY21, LS11, KSS15]. Exploitation [CK20]. Exploiting [Bro06]. Exploration [Ben90, Ben92, CK20]. Exploration-Exploitation [CK20]. explosiveness [Spi15]. Exponential [BR95, BLS12, Bro97, CPS89, Olu11, Ste87, ZL02, BS06, BKP01, GHvdL07, Han19, KK01, Lin19, MHZ10, MT22, NK07, Olu06, Sie07, Yam99, Yu21, YT23, ZL09, ZB12, vdWSvN08]. Exponentiality [Bha99]. exponentiated [DKC21, DKB22, HN19]. Extended [MK23, GK07, GV11]. Extending [Ros95]. extensions [AAC14, GM14, KP12, KSvH14, VHBSW22]. External [Kij92, CF12]. Extinction [Bha87]. Extinctions [Ros87c]. Extremal [ALM07, TSW97, CMH13]. Extreme [SC23, TE15, ZDS20, BH23, CH11, MIK23, MNH15]. Extremes [Eli05, DKC21]. extropy [EQ22, JZK20, QWW19, YXH19].

face [HP03b, HPS08]. face-homogeneous [HP03b, HPS08]. Facility [PP02, EBTD22, Pre02]. FACS [GN15]. facsimile [KT01]. factor [DF08, mKK22, NYZ04]. Factorization [Ald93]. factors [EA06, Tan06]. Fail [FG89a]. failed [GZY22]. Failure [AENS94, BLS12, BF91, CHJ95, ES91, HY90, Lyn99, Tay87, Yao89, AAD08, AO18, AO19, BSL15, DCL^{+}05, DCN05, FE01, GK07, GV11, Lef06, Lyn03, NH04, ZBA05]. failure-prone [AAD08]. failure-rate [GK07]. Failures [CHL94, HM87, CH12, Kwo14, ÖK14, ZPD06]. fair [Lu09, Lu10]. fairness [AILR08]. fall [TAM^{+}21]. False [HT96]. Family [Ste87]. Fast [BM02b, Tim16, FH17a]. Fatigue [Tay87]. Fault [BPT87, LM87]. Fault-tolerant [LM87]. faults [Hil99]. FB [MN05]. features [MK21]. FEC [XC10]. fed [BM02b]. Feedback [KT97, Yec88]. feedbacks [Jia19]. feeders [ZPD06]. Festschrift [Pek16]. field [DGG22]. Fields [BF90, LY04, SD22b]. Fill [Hil99]. filter [LC07, Miy00]. Filtering [Fan96, LS11]. financial [HL14, PL20]. find [GT07]. Finding [MR95, MK17, Mat17, Mat19b, MK21].

Finite [Bar96, BK03b, CCdO01, DR97, GI88, Gou94, Kad17, Kij92, LOdBP87, Lu13, MS91, Miy92, MT93, PY16, Tij92, VS87, VT92, WB96, vDTW88, AAD08, AGC07, AY08b, BKB22, Bri14, BH23, CM00, CBBG20, Cov10, HHvU07, HS07, JD07, Rat19, SI04, SAF22, Sie07, SNS19, Tan04, Tan06, Tij08, Zha99]. Finite-Buffer [Bar96, Gou94, MT93, Tij92, CM00, Zha99]. Finite-Capacity [WB96, SI04]. Finite-Horizon [BK03b]. finite-source [SNS19]. finite-state [HHvU07]. Finite-Time [DR97, Bri14]. First [AvER95, AR97, BS87, BvHvO97, DP93, Lef96, RL93, SS87, Tam98, VT08, ZL02, vdHHV01, Pin07, SZ03, SZ05, YZZ23]. First-Come [BvHvO97]. first-exit [SZ03, SZ05]. first-order [YZZ23]. First-Passage [Lef96, vdHHV01]. First-Served [BvHvO97]. Fisher [AAK19, Zog23]. Fit [And95, BF91, MG21]. Fitting [AvER95]. Fixed [CH88, Tak96, Yao89, BH23, CSL09, Fro99]. Fleeing [Sat97]. Flexible [ER95, sADL02, AAD07, AAD08, Pan08, Sto05, YWC23]. Floating [Kol92]. Flow [AF92, FS89, LL91, CSZ03, CLS21, GRSV14, KS02, Ros15, Voi09]. Flow-Shop [LL91]. Flows [BBW91, BOT09]. Flowshops [PSC92]. Flowtime [LK96]. fluctuating [GTDKÖ09]. Fluid [BPvdDS99, FZ19, KR94, MR95, Man96b, RW92a, SY92b, ARK03, AJK01, BPS14, BOT09, BPS06, BKP01, BHM09, DA03, EBTD22, GKPR99, KS02, LZ19, Lu21, SZ02]. Fluids [DOW93]. FM [LE94]. following [GS23]. forced [AS18, Pek99a]. Forecast [SHM09]. forecasting [DK22]. forewarning [Kwo14]. fork [LZ10]. fork-join [LZ10]. Form [GS92, KS92, MBP07, RW92b, AL15, BB02, BvHTvO18, CGM21, CX23, Eco99, mKK22, MP02, Mar16, NS21, PG22, PGZ22, WXL20, Yam99, Yam00]. Formal [MO95]. Forms [VV88]. Formula [JS96, mKK22, KKJ22, LST^{+}22, PG22, PGZ22, PH23]. formulas [NCJ18]. Fortune [Dub98]. Fortune-Dependent [Dub98]. forward [CL06]. Four [LL90, FL05]. Four-Station [LL90]. four-threshold [FL05]. Fractal [KAJ07]. Fractals [Bax96]. Fractional [Her95, CLS21, DM03, Rao15, Yan20]. Framework [Fre91a, Wan17b]. Free [AW89, Ken94, BPR10]. frequency [OKA22]. frequency-severity [OKA22]. friendship [BJ19, DRR22]. Fringe [AFP88]. Front [Ano87b, Ano87d, Ano87f, Ano87h, Ano88b, Ano88d, Ano88f, Ano88h, Ano89b, Ano89d, Ano89f, Ano89h, Ano90b, Ano90d, Ano90f, Ano90h, Ano91c, Ano91e, Ano91g, Ano91i, Ano92b, Ano92d, Ano92f, Ano92h, Ano95b, Ano09b, Ano09d, Ano09f, Ano09h, Ano10b, Ano10d, Ano10f, Ano10h, Ano11b, Ano11d, Ano11f, Ano11h, Ano12b, Ano12d, Ano12f, Ano12h, Ano13b, Ano13d, Ano13f, Ano13h, Ano14b, Ano14d, Ano14f, Ano14h, Ano15b, Ano17h, Ano18b, Ano18d, Ano18f, Ano18h, Ano19b, Ano19d, Ano19f, Ano19h, Ano20b, Ano20d, Ano20f, Ano20h, Ano21c, Ano21e, Ano21g, Ano22b, Ano22c, Ano22e, Ano22g, CW90b, Fil98, Ano15d, Ano15f, Ano15h, Ano16b, Ano16d, Ano16f, Ano16h, Ano17b, Ano17d, Ano17f, Cho99, Hil99]. front [YS08]. front-end [YS08]. Frost [OE10]. Full [TS96, LQ15].

Full-Information [TS96]. full/partial [LQ15]. Fully [KW94]. Function [AH96, BSS98, Fis91, Lin92b, MPP93, OLÖ97, Gup16, HvdL03, KSvH14, Mon13, NL07, PK06, RKFI21, Spi15]. Functional [JS89, PW98]. Functionals [HvdHR09, SD22b]. Functions [AENP89, BL89, CS94, Fro95, RS94a, Tem94, AL15, BJS99, BdlPS15, DFS10, FL20b, FE01, GV11, HY21, HYL07, Lu21, Ros06b, XDZ15, YHZ06]. fund [KKL17]. Funding [BK06]. Further [AKP05, BL89, CV22, Qiu19]. futures [Chi23].

G [AGC07, AW96, BB96, BBS97, CG96, CGM21, CRP93, Dud97, ET87, FK96, GV10, LN92, LRW09, LK08, MN05, MK23, NJ95, ÒLC95, PV11, PZ04, Ros06a, SS92, SKA09, SK11, VY08, Wan02, WW05, dK89, Cha99, PR08, Bar08, Eli08, LRW09, SS92]. G/G/1 [LRW09]. G/G/c [SS92]. G/GI/ [Eli08]. G/M/c [Bar08]. Gains [RTW09]. Game [LnG22, McN88, Che11, DGG22, Kay09, Kim22, MI10, Pen10, Ros11, TvdW06]. games [Che09, Che12, KSvH14, SD22a, ZCJL23]. Gamma [BLS12, LG89, Tem94, AKS11, AO18, AO19, HN19, JZY22, KX12, Nad16, PO15, ZZZ21, Zha11, ZHZ15]. gammas [BSL15]. gap [Fra99]. GARCH [GB22, WSX18, WXL20]. GARCH-jump [GB22]. Gated [Bor95, BWY93, BCGW92, PZ04, vdMR08]. Gating [JS89]. Gaussian [AAK19, BF90, BPG12, Bre90, GV11, LC07, dI06]. Gelenbe [Çağ16, Pek16]. General [CS92, Dav95, FG89a, FL98, Fre91a, LR97, Man96b, Van88, ADGZ09, CK15, FW15, GN14, HL04, MK23, ÖAA01, RKFI21, WS06, WN15]. generalization [KL15]. Generalized [CP93, CH11, DLU99, HK90, KW89, LG89, MBP07, ME14, NZP18, OLÖ97, RS90, Van92b, AEAN23, ABJ22, BBHK10, BMR05a, CLN20, CB20, CH15, CLSS13, DFS15, DvdLR11, Dim23, FL13, FRR02, GV11, GKB13, Gut03, HN19, HZ05, HZ06, HJK07, Hua20, Kay16, KL12, KM12, MNH15, Nad16, PO15, SAF22, TA11, XH08, XH09, XZ11, ZY22]. Generalizing [JS96]. generate [GDRV10]. generated [BM11, MOSU10]. Generating [DS87, Kol92, BMR05b, HY21]. generation [BCR05, DO03, LLY22, YSM06]. Generator [But89, MK23, Tar17]. generic [BKM10]. genetic [GN15]. Geo [CLC98, CGM21, CLC98]. Geo/G/1 [CGM21]. geographical [FWH21]. geom [Cha99, CG96]. geom/g/1 [Cha99]. Geometric [BS14, DZZ18, Eco99, LT96, Las98, Spi91, BB02, CF13, CBG15, DZB12a, DZB12b, HY21, LKG19, MHZ10]. Geometric-form [Eco99]. Gerald [BOV95]. Gerber [XYZC23]. GI [BdBS20, Eli08, LSSZ16, NQ01, PMB96, Sha88, XLSZ04, AW96, BBS97, BdBS20, CCC05, CSP12, Eco99, LN92, LSSZ16, NQ01, Sha88, Wan02, WW05, dK89]. GI/C [CSP12]. GI/C-MSP/1/ [CSP12]. GI/G/1 [AW96, BBS97, LN92, Wan02, WW05, dK89]. GI/GI/1 [NQ01, Sha88]. Gibbs [Din95, Ste87]. Gini [SC23, DM20, ZD19]. Gini-Type [SC23]. Gittins [AAR11]. Given [Mei91, Sad98, CW22]. Global [RS94b]. Globally [Bor95, BWY93]. gold [SH00]. gold-mining [SH00]. Goodness [And95, BF91, MG21]. Goodness-of-Fit [And95]. governance [Kim22]. governed [SH23, SC03]. governing [YT23]. gradient [AAES10, EFG17]. gradual [Miy00, SZ02]. gradually [JQZD23]. Graph [BC94, Sie93, Van01, vdHHV01, BKL22, Gut03, Röl22]. graph-based [Gut03]. Graph-Structured [BC94]. Graphs [Ald93, BBW91, BS14, CS93, DH87, Mad93, Pal90, AR14, Ald22, AS00a, DK09, DM20, FHS13, HLM16, Mah13b, Mah14, Mah19, MOSU10, ZBA05]. Greedy [ADS88, BC94, FL98]. Group [AENS94, Abr93, BLSV03, CHS87, CHH90, VT08, Yao88, YH88, AGC07, YDJ23, ZP16, vSvDdK^{+}20]. Group-testing [CHS87, ZP16]. grouping [LWL19]. groups [PY14]. growing [CK20]. Growth [CH87, Bro99, FH17a, KL10, Mah13b]. GSR [vZS04]. guarantee [LP23]. guess [Nic16]. guessing [Pen10].

Hall [Man96a]. hammock [JCDJ22]. Harmonic [JL98]. Harris [FH94]. harvesting [Abd17, GKL23]. Hashing [Ald88]. Having [BLS12, GKL87, Haj15]. Hazard [BSS98, CE96, HH00, KK06, KX08, LR91, Ali23, AAES10, BK09, Bro14, CR01, DZB12a, DZB12b, Gup16, KK05, KX07, NBR13, PT20]. Hazards [JR92, Ros90a, Koc23]. Heavy [BW14, CPR91, FS99, FGR10, IVA15, PB06, VdK87, BM02b, Lu09, Lu13, vdMR08]. heavy-tailed [BM02b]. Heavy-traffic [BW14, IVA15]. hedge [Chi23]. hedging [SXWW22]. heel [vZS04]. height [WWZ22]. Henk [Sch96]. Heston [WJYL22]. Heterogeneous [BZ13b, DZZ18, Kha13, LY13, Wan18a, Xu13, ABJ22, DKB22, DLB12, DDL14, FL20a, HN19, HZW01, LWL19, MHZ10, XH11, Yu21, ZDZ20, ZZZ21, ZL09, Zha11]. Heuristic [ADS88, Jun88, LM87, LAWZ01]. Heuristics [GV90, JLAE18, Tij92, JLAE17]. Hidden [BB97, LP23]. Hierarchical [KS92, FMR17, VMB00]. High [LMZ22, ABI22, CM05]. high-dimensional [ABI22]. High-Order [LMZ22]. high-order-event [CM05]. Higher [BCS92, Pla92]. Higher-Order [Pla92]. Highly [KSM88, NS04]. History [CP95, CHH90, GFH22]. History-Dependent [CP95, GFH22]. Hitting [BOT09, Hes92, MZ17, RS02, STI18]. Hitting-Time [Hes92]. HOL [CLC98]. Holding [FY11]. Holst [Hil99]. homogeneous [HN19, HP03b, HPS08, Hua20, PT20]. hooking [DM22]. Hop [LT00]. Horizon [BK03b, Lu13, SLZ04, SBZ16, Tan06]. horizons [Tan04]. Hrnn4F [TAM^{+}21]. Hunting [Sat97]. hybrid [QWGV23, TAM^{+}21]. Hydrodynamic [GD18]. Hydropower [MD06]. hyper [YT23]. hyper-exponential [YT23]. hyperbolic [LRW99, vdWSvN08]. Hyperexponential [OLÖ97]. Hypergeometric [BLSV03].

ideas [RL01]. Identical [BCD95, JST01]. identically [HM21]. Identification [FGR10, EFG17]. Identifying [BS87, GS93]. Identities [Han03]. Identity [Ros98, CLS21, KA16, Ros23, Wan02]. idle [AW12, AS18, Haj15, Pek99a]. IEEE [FM16]. IFR [BLS12, DFS10, DLP90, KM07]. II [MG21]. Illustration [VR89]. Images [BBS88]. imbeddable [KM07]. IMIT [AK05, AKP05, LX06]. Immigration [Smi97b, CZ03, Kyr03, Kyr04, MOSU10, ZCJ04]. immigration-death [KPDR16]. impact [ZZ13]. impacts [GZ09]. impatience [Cao17]. Impatient [SB13, AY08a, AZR08]. Imperfect [Fre91b, GK21]. implementation [HLZ23]. Implications [FH94]. implied [PZ21]. Impolite [ÒLC95]. Importance [BL89, CCH99, CHC00, Gut97, JR92, Kol94, Ros98, BdBS20, GS12, GR12, HKA22, Men04, NFMFSA20]. Imprecise [dCHQ09]. improve [Pek99a]. Improved [Abr93, Ang97, CT23, DR95a, YK00, Chi21]. Improvement [CE96, McN89, WK88, Yao88, LST^{+}22]. Improvement/Newton [Whi89]. Improving [AS18, BF90, GR12, PR94]. In-Process [WAI95]. inaccuracy [dCBB22, KMS20]. inactivity [DLB12, GZY22, KA04, LL03, LZ11]. incidents [AZR08]. including [YZL00]. Income [Arr87]. Incomplete [BLSV03, BK93, RW95, Tem94, HS07]. Incorporating [DO03, OE10]. Increasing [CHJ95, Lyn99, Tur98, CH11, KP07, Lyn03, Ols01, Rol05]. increments [Fen22, Lu13]. Independence [Mad93, LMZ22, STI18]. Independent [BCD95, HP89, LG89, NZP18, YH88, AKS11, BdBS20, CY21, FL20a, Fen22, KX12]. Independently [RS98a]. Index [LT00, Tam98, AAR11, DM20, HKK10, Vil16, ZD19]. Indexability [CY10, Vil16]. indexed [Hua20, SZF18, WWZ15]. indexes [GS12]. indicators [LDVT22]. Indices [Kha99, FHS13]. Individual [Bro00, YH88, Ros15]. Inequalities [AAK19, KG92, Li94, Mon13, SZ99, GMZ10, Kha07, Olu99, Pap21, Joe12]. Inequality [BR95, Ros95, Mon07]. inert [FM17]. infectives [OE10]. Inference [RSZ93, JC18]. inferences [HM20]. Infinite [AY08a, HPR96, Kha88, PW12a, PW12b, TV91, AW05, BDM15, DPV16, FS99, PY14, SLZ04, SBZ16]. Infinite-Armed [HPR96]. Infinite-server [AY08a, PW12a, PW12b, BDM15]. Infinite-State [TV91]. inflation [PC22]. Influence [Chy91]. Information [Arr87, BLSV03, BBS96, BK93, Fre91b, Has07, HKL94, RW95, TS96, VT92, Yan98, Yao88, AAES10, AAK19, CLN20, CW22, EQ22, HS07, LQ15, LCYY22, SLZ04, WNT23, ZAE19, ZZ13]. Informational [MY11]. Inhomogeneous [CK88, LW19]. injection [Xu23]. Input [ALR00, BKP01, BKKP05, CJ96, DMvU07, DA03, Miy00, SZ02, Whi00]. Input-Queued [ALR00]. inputs [GKPR99]. Insensitivity [KW89, YS08]. insights [AZR08]. Inspected [KWA96]. Inspecting [HP89]. Inspection [DLR87, Ang06, GK21, HZ04, Ros03a, YK00]. Instability [Ros88b]. instant [HM15, Mi21]. Insurance [AK08, NZP18, FS23, FKL02, LP15, QWGV23, Ros03b]. insurers [ZCJL23]. Integral [Ban92, Lin95]. Integrals [BPvdW91, KH91]. Integrated [Hun91]. Integration [CCP90, LST^{+}22]. Intelligent [Ber96]. inter [BH23, XYZC23]. inter-observation [XYZC23]. inter-renewals [BH23]. Interacting [AB93]. Interarrival [CGKS96, Koc96]. Interchangeability [CPS89].

Interchangeable [OS90]. interconnected [JST01]. interepoch [PT20]. Interest [HV95, YZ03, Cai02, FNC20, WJYL22]. Interest-Rate-Dependent [HV95]. Interference [VV88, GN15]. intermittent [BKP01]. intermittently [XDZ15]. Internet [AY08b]. Interpretation [Sha88]. Interpretations [GV90]. Interruptible [SO94]. Interruptions [BWY93, BMK08, LXH04, TZS22]. intersection [DK09]. intersections [BAWD12, HvdW08]. Interval [CL91, Lef01, Smi97a, vdH87, Cov10, HBV10, Mi99, Mi21, MOSU10]. intervals [PT20]. Intrinsic [RSS09]. Introduction [Pek21, Cag17, Kyr03]. Invariant [FTC98, LT96, Las98, CBG15, GMZ10]. Inventory [AF97, BL87b, CF06, FZ93, FH94, PP89, Per97, PP02, ABEK17, AÖ12, BLPS05, BBP20, BPS14, BA02, CXY12, FW15, FL05, FL20b, HKS07, LXH04, NPS04, PC14, PLS23, XC09, ZAMB21, ZK06]. Inverse [Lin92a, Lin92b, Tag00, GV11, MG21]. inversion [dI06]. investment [PC22]. Involving [Bro89, AKP05, Hil99, ÖAA01]. IPA [Miy00]. Irreversible [SO94]. ISBN [Sha98]. isolated [HvdW08]. Israeli [PY14]. issue [Ano87b, Ano87a, Ano87d, Ano87c, Ano87f, Ano87e, Ano87h, Ano87g, Ano88b, Ano88a, Ano88d, Ano88c, Ano88f, Ano88e, Ano88h, Ano88g, Ano89b, Ano89a, Ano89d, Ano89c, Ano89f, Ano89e, Ano89h, Ano89g, Ano90b, Ano90a, Ano90d, Ano90c, Ano90f, Ano90e, Ano90h, Ano90g, Ano91c, Ano91b, Ano91e, Ano91d, Ano91g, Ano91f, Ano91i, Ano91h, Ano92b, Ano92a, Ano92d, Ano92c, Ano92f, Ano92e, Ano92h, Ano92g, Ano93a, Ano93b, Ano93c, Ano93d, Ano94b, Ano94c, Ano94d, Ano94e, Ano95b, Ano95a, Ano95c, Ano95d, Ano95e, Ano96a, Ano96b, Ano96c, Ano96d, Ano97a, Ano97b, Ano97c, Ano97d, Ano98a, Ano98b, Ano98c, Ano98d, Ano09b, Ano09a, Ano09d, Ano09c, Ano09f, Ano09e, Ano09h, Ano09g, Ano10b, Ano10a, Ano10d]. issue [Ano10c, Ano10f, Ano10e, Ano10h, Ano10g, Ano11b, Ano11a, Ano11d, Ano11c, Ano11f, Ano11e, Ano11h, Ano11g, Ano12b, Ano12a, Ano12d, Ano12c, Ano12f, Ano12e, Ano12h, Ano12g, Ano13b, Ano13a, Ano13d, Ano13c, Ano13f, Ano13e, Ano13h, Ano13g, Ano14b, Ano14a, Ano14d, Ano14c, Ano14f, Ano14e, Ano14h, Ano14g, Ano15b, Ano15a, Ano15d, Ano15c, Ano15f, Ano15e, Ano15h, Ano15g, Ano16b, Ano16a, Ano16d, Ano16c, Ano16f, Ano16e, Ano16h, Ano16g, Ano17b, Ano17a, Ano17d, Ano17c, Ano17f, Ano17e, Ano17h, Ano17g, Ano18b, Ano18a, Ano18d, Ano18c, Ano18f, Ano18e, Ano18h, Ano18g, Ano19b, Ano19a, Ano19d, Ano19c, Ano19f, Ano19e, Ano19h, Ano19g, Ano20b, Ano20a, Ano20d, Ano20c, Ano20f, Ano20e, Ano20h]. issue [Ano20g, Ano21c, Ano21b, Ano21e, Ano21d, Ano21g, Ano21f, Ano22b, Ano22a, Ano22c, Ano22e, Ano22d, Ano22g, Ano22f, Ano23a, Cag17, Pek21]. item [AÖ12, LAWZ01]. Items [HP89, YH88, BLPS05, BBP20, BA02]. Iterated [AO18, AO19]. Iteration [dK89, CHL03]. Iterative [GV90, VS87]. Itinerary [Lal90]. Ito [Pla92].

J [BOV95]. Jackson [AW05, Eco03, Lin97, PM17, Zaz04, vD95]. Jobs [BC94, BBCP97, CHW89, FG89a, HW92, IwW97, MCB18, Xu91, ZCA18, AZR08]. jockeying [DKM16, Rav21]. Join [THM04, Dim23, LZ10]. Joint [CL23, YBL22a]. Jump [Bou96, Fan96, BPS14, CY21, CX23, GT10, GB22, Han19, PS02, SXWW22, Tar11, WWL17, Yan20, YWC23]. jump-diffusion [PS02, WWL17, Yan20]. jump-fluid [BPS14]. Jump-Over [Bou96]. jumps [KKJ22, Rao15, XXN21].

K-NN [Rot93]. Kalman [LC07]. Khintchine [JS96]. kin [vSvDdK^{+}20]. knapsack [CR15]. Knockout [CHYZ93, CR19, Fra13]. knowledge [EFG17]. known [LC97]. Koehler [Lin90b]. Kolmogorov [BH21, Röl22]. Koole [Mül98].

l [NJ95]. Labeling [Ang89]. Laboratories [OGP95]. Lag [FHJ94, Kul04]. Lagrangian [HJM09]. lanes [vdHvHE01]. Laplace [Bha99, Keb94a, MCL22, NK05]. Large [AB93, Ben90, BA18, DOW93, HP03a, HP03b, PW98, RW92a, SAS89, SR90, Whi88, BH23, Man96b, ME14, NS04, WWZ15, Man96a]. large-deviations [ME14]. Large-Scale [Whi88, BA18]. Largest [Mei91, HN19, HSB23, ZZZ21, ZB12, ZHZ15, ZDS20]. Lasota [Bax96]. Lateness [LL91]. Lateness-Related [LL91]. latent [HCX06]. lateral [BvHTvO18]. Lattice [LZ00, CH12, DM08, FMR17, GB22]. Law [BS93, SAS89, Bar08, BPG12, BC18, GMZ10, WWZ15, KW13]. law-invariant [GMZ10]. layer [WWZZ23]. layered [PY16]. LCFS [NQ01, Sha88]. LCFS-PR [NQ01]. Lead [DvdH97, FH94, KAR09, PLS23]. Leader [DY90b]. leadtimes [BBP20, PC14]. leaky [Miy00]. leaky-bucket [Miy00]. Learning [JLAE18, Bor00, Cag17, JLAE17, KLT12, Pek21, Tim16, Yin21]. Least [HL92, Tim16]. least-squares [Tim16]. Length [Alz88, SS92, DMvU07, LZ10, Lin00, Olu99]. Length-Biased [Alz88]. Lengths [Mah91, ALM07, KP07, Ros99a]. Leprechaun [Ber96]. Lerch [KPH10]. level [dCBB22, MT16, PLS23, YWL22]. level-crossing [PLS23]. level-dependent [MT16]. lévy [BKM10, DMvU07, CW23, YWZH21]. lévy-driven [BKM10, YWZH21]. Lieberman [BOV95]. Life [BFR99, GKL87, KG89, Vil18, AK05, GV11, Gup16, HSB23, HZW01, LL03, LX06, LZ11, Lin00, Lin19, MFN17, NDB13, Olu11, Roy02, SY15]. life-length [Lin00]. Life-threatening [Vil18]. Lifetime [Kor03, FC22, KP12, Qiu19, WC00]. Lifetimes [CS94, KX10, ASKZ21, ABJ22, BL02, DLB12, IASB23, LYF16]. light [MN05]. light-tailed [MN05]. Likelihood [AEAN23, DZB12a, DZB12b, HZW01, HZ04, LOdBP87, Wan18a, BBHK10, BGR11, HM20]. Limit [CZL96, GD18, HS91, Ols01, RS95, SAS89, Ste87, BY99, BDM15, Che11, FZ19, KPH10, LY04, Lu21, TYZ19, ZYSZ21, dCHQ09]. Limited [AF97, CPR91, BW14, WAvHD09]. Limiting [FG89b, Mah91, WJ18]. Limits [AW95, Bar96, Bla98, Whi00, AJK01, BH02, PB06, SS20, WWZZ23]. Lindley [Roy90]. Line [Che97, CW90a, Fra19]. Linear [Ald88, BPT90, GS98, HS92, KPDR16, LZ00, Mei91, Smi97b, CFKP99, GK07, JA06, KX12, Mah13a, NK05, NL07, ZCJ04]. Lines [AM87, AAD07, AAD08, JD07]. Link [CHL94, DH94]. Lipschitz [Tar17]. liquidation [KTÇ14]. liquidity [CW23]. Lists [CW90b]. Little [KW13]. Load [VdK87, FH17b, FS99, HBV10, PKK13, VHBSW22, XC10]. load-sharing [FH17b]. Loading [Rig92, DCN05]. loading-dependent [DCN05]. local [AL15]. Locally [SXWW22]. location [BKB22, DKC21, DKB22]. location-scale [BKB22, DKC21, DKB22]. Locking [RTW09]. Log [BSF21, Bha87, SS87, Bro01, MXH18, PZ21, XMH21]. Log-Concavity [BSF21, SS87, MXH18, XMH21]. Log-Convexity [Bha87, SS87, XMH21]. log-normal [PZ21]. logconcavity [PSZ00]. logconvexity [AEAN23]. Lomax [Roy90]. Long [CW90a, ARK03, Ang01, FKL02]. long-run [ARK03, Ang01]. Long-Run-Average [CW90a]. long-term-care [FKL02]. Longer [Fla89]. longest [AW12, Haj15]. Longman [Sha98]. look [BDPS11, Rav21]. lookback [Yan20]. Loop [HL91, GHvdL07]. Loss [BT95, BZ05, DP95, GV10, Gou94, HL92, Key90, MS91, Miy92, MT93, RW92b, RS02, Zie96, AHW10, AW12, Bro01, Haj15, HMW09, KSvH14, LWZ16, NS17, Pen15, SBZ16, Zha99]. Losses [MR88, Tan06]. Lost [PLS23]. Lot [Sob95]. low [Ros23]. Lower [Abr93, STI18, XLSZ04, Yao88, AAC14, BCMS23, Che11, Men02, PdB22]. LRD [XLSZ04]. LRD/GI/1 [XLSZ04]. lumping [KS12]. Lyapunov [Las98].

M [Bar08, Dod89, Eco99, Elc94, FY11, HH16, HW92, NJ95, SNS19, BPLD21, GZDW21, AGC07, BB96, CRP93, Dud97, EZB06, ET87, Elc94, FK96, FY11, Fra05, GV10, HH16, LK08, Lu09, MN05, MK23, ÒLC95, PR08, PV11, PZ04, Ros06a, SKA09, SK11, SNS19, TS07, VY08, BPLD21, GZDW21]. M/D [Lu09]. M/D/1 [EZB06, TS07]. M/G/ [PZ04, VY08]. M/G/1 [AGC07, BB96, CRP93, FK96, MN05, MK23, ÒLC95, PV11, Ros06a, SKA09, SK11]. M/G/1-Type [PV11]. M/G/l [NJ95]. M/G/s [LK08]. M/M/1 [Elc94, SNS19, BPLD21, GZDW21]. M/M/1/ [HH16]. M/M/K/N [FY11]. Machine [CHRW94, For93, Fre91b, Fro91, Gla91, LK96, PR88, Cag17, CSZ03, CZ06, CHKL13, Fre91a]. Machines [BBCP97, BCD95, CHW89, Fro95, HW92, Rig92, JST01, Rat19]. Mackey [Bax96]. magneto [RB06]. magneto-optical [RB06]. Maintained [KM88]. Maintenance [BGDR03, CS92, JA06, Wee90, BY99, CHKL13, ÇÖ15, HS13, Jia19, Rig02, SvNKC00]. Majorization [Joe12, FPSL03]. Make [LM88, GTDKÖ09, MAS06]. make-to-stock [GTDKÖ09, MAS06]. Makespan [BBCP97, BCD95, LK96, BHM09, JST01]. Making [Ros06b]. Malus [OKA22]. Management [XZP18, AL09, GKL23, LK09]. Manufacturing [ER95, Sob95, HS13]. Many [BBS96, BC98, GSW93, LM88, BM02b, LZ19]. many-server [LZ19]. Mapping [SW95]. Marginal [Mei91]. markdown [KLT12]. Marked [BBW91]. market [BCR05, CVB05]. markets [KTÇ14]. Markov [MY11, AW89, Abd17, Ald87, AFP88, ACG95, Ang01, AÖ12, BBS88, BR97, BGG07, BB97, BM11, BDM15, BKP01, Bro91, Cal97, ÇÖ15, CCP90, CJ96, CK88, CSvD02, DOW93, DKK94, DP95, EK09, FHJ94, Fil91, FTC98, Fra99, GKPR99, GV94, GI88, GT07, HHvU07, Hes92, HJ95, HP00, HV95, Hua20, JCDJ22, Joh98, KS12, KL10, Ken94, Kij92, KW89, KM07, Kul04, LW19, LY04, LK09, Lor18, LP23, MP02, MBP07, MR95, Man96b, MK21, Miy02, PR97, PV11, Pla92, PT93b, RW92a, Ros87a, Ros88a, Ros13, Sen89, Sen91, Sen93b, SZF18, SC03, SX22, Spi15, SS20, SKL12, SR90, TYZ19, TV91, VS87, Van92a, Wak98, Wak00, WWZ15, Yan98, YZ98, YZL00, YS89, YSM06, YHZ06, Zaz07, ZYL10, ZYSZ21, dCHQ09, vD88, vdL11]. Markov-Modulated [BR97, CJ96, DP95, Abd17, AÖ12, BDM15, KL10, LK09]. Markovian [BK95, EBTD22, FM16, Fra19, Gut97, LST^{+}22, NS21, TZS22]. Marshall [Joe12, FL13, GKB13]. Martingale [GI95, HW90, Ros95, YZ03, Kha07]. martingales [MK23]. Mass [Ros87c, Sha98, AZR08]. Match [DY90a, Dav95]. Matching [ADS88, Sad96, Sch98, BDPS11]. Material [MR88]. Materials [Tay87]. Mathematical [Bax96]. matrices [KS01]. Matrix [Ald93, CSvD02, HY21, HH16, MT22]. matter [Ano87b, Ano87a, Ano87d, Ano87c, Ano87f, Ano87e, Ano87h, Ano87g, Ano88b, Ano88a, Ano88d, Ano88c, Ano88f, Ano88e, Ano88h, Ano88g, Ano89b, Ano89a, Ano89d, Ano89c, Ano89f, Ano89e, Ano89h, Ano89g, Ano90b, Ano90a, Ano90d, Ano90c, Ano90f, Ano90e, Ano90h, Ano90g, Ano91c, Ano91b, Ano91e, Ano91d, Ano91g, Ano91f, Ano91i, Ano91h, Ano92b, Ano92a, Ano92d, Ano92c, Ano92f, Ano92e, Ano92h, Ano92g, Ano93d, Ano94e, Ano95b, Ano95e, Ano96d, Ano97d, Ano09b, Ano09a, Ano09d, Ano09c, Ano09f, Ano09e, Ano09h, Ano09g, Ano10b, Ano10a, Ano10d, Ano10c, Ano10f, Ano10e, Ano10h, Ano10g, Ano11b, Ano11a, Ano11d, Ano11c, Ano11f, Ano11e, Ano11h, Ano11g, Ano12b, Ano12a, Ano12d, Ano12c, Ano12f, Ano12e]. matter [Ano12h, Ano12g, Ano13b, Ano13a, Ano13d, Ano13c, Ano13f, Ano13e, Ano13h, Ano13g, Ano14b, Ano14a, Ano14d, Ano14c, Ano14f, Ano14e, Ano14h, Ano14g, Ano15b, Ano15a, Ano15d, Ano15c, Ano15f, Ano15e, Ano15h, Ano15g, Ano16b, Ano16a, Ano16d, Ano16c, Ano16f, Ano16e, Ano16h, Ano16g, Ano17b, Ano17a, Ano17d, Ano17c, Ano17f, Ano17e, Ano17h, Ano17g, Ano18b, Ano18a, Ano18d, Ano18c, Ano18f, Ano18e, Ano18h, Ano18g, Ano19b, Ano19a, Ano19d, Ano19c, Ano19f, Ano19e, Ano19h, Ano19g, Ano20b, Ano20a, Ano20d, Ano20c, Ano20f, Ano20e, Ano20h, Ano20g, Ano21c, Ano21b, Ano21e, Ano21d, Ano21g, Ano21f, Ano22b, Ano22a, Ano22c, Ano22e, Ano22d, Ano22g, Ano22f]. maturity [PZ19]. Maxima [BZ05, OS90, DZB12a, DZB12b, Tan04]. Maximal [ALR00, BT95, Nad17]. maximization [Tar11]. Maximize [Fro95, RS89]. Maximizing [AAD08, Lef01]. Maximum [BW95, CFFH98, LOdBP87, LM87, SS92, MT16]. may [FG89a]. Maze [GA90]. McMillan [SZF18]. MDP [HvdW08, vdL11]. Mean [Gup16, HSB23, Lyn97, PR97, Ros87a, Ros88a, Ros89, SP98, SA18, BH23, DMM16, DGG22, GV11, HZW01, KA04, KMT18, LCYY22, NDB13, NL07, PC22, Ros99b, ZCJL23]. mean-variance [NL07, PC22, ZCJL23]. Means [Jun88, Ros06b]. measure [dCBB22, CBG15, Men04, Mi99, SK20, YZL00]. Measurement [GM96, EA06]. Measures [FTC98, He89, LL91, Ste87, Van97, CLN20, CBBG20, GT10, GMZ10, HKA22, KMS20, NFMFSA20, Olu11, SH19, WH20, ZAE19, Zog23, SC23]. mechanics [HJM09]. Mechanism [JS89]. Memory [HPR96, BCZZ19]. memoryless [KL12]. merge [HBKK03]. Message [RS90]. metaheuristic [Gut03]. metatheorem [RL99]. Method [Ald87, AK96, BPvdW91, CE96, HW90, YZ03, YWZH21, dK89, ABI22, GDRV10, HY21, JCDJ22, PL20]. Methods [MR88, WN10, CS03, GW12, GRSV14, NS04, QWGV23, Yin21]. Metropolis [RBM02]. Metropolis-type [RBM02]. Michael [Bax96]. Milton [OST06]. Minima [OS90]. Minimal [KG89, KG92, Men94a, Pre02]. Minimax [Bla95]. minimization [JST01]. Minimize [BBCP97, BCD95, WAI95, CSZ03]. Minimizing [BHM09, LK96, Xu91, Ros15, SXWW22]. Minimum [Abr93, Nay89, Cov10, LP23, THM04, ZK06]. mining [SH00]. minus [Fen22]. Mismatch [HvdHR09]. missions [ÇÖ15]. mitigation [EFG17]. mixed [BBCN01, GFH22, HLZ23, IASB23, KN00, LSW21, QWW19, Rao15, Yan20]. Mixing [BBS88, vdL11]. Mixture [BLS12, BPT89, BKB22, BLS03, BSL15, FE01]. Mixtures [GG96, Lyn99, BS08, BS06, DLU99, Li05, Lyn03, NH04, SAF22]. MNL [PdB22]. Mobile [AdHBvO14, LT00, VMB00]. mobiles [Abd16]. mode [Bel11]. Model [AdHBvO14, BC98, BL87a, BL88, BS91, CS92, CHL94, DG94, ER95, EGL92, FG89a, Fla89, FGR10, Fro91, HKL94, KR94, MD06, MW94, NJ95, RTW09, SD97, Spi91, Tak90, Tay87, Van92b, WXWH18, WWS18, XWTW18, YZ03, Abd17, AP99, ABEK17, ABD23, BCR05, BPS14, BOT09, BA02, BPS06, BMK08, CM05, CH15, CY21, CX23, CL23, CH12, DRR22, DCN05, DL10, EK21, FW15, FL20b, FM17, Fra19, FNC20, FL22, GS23, GB22, GV11, Haj15, Han19, HL14, HLZ23, JQZD23, KTÇ14, KS02, KPS03, KL10, KK01, mKK22, Koc23, Lef06, LZ10, LQ15, LC07, Mah13b, MMY20, Nad08, NDB13, OKA22, ÖAA01, PdB22, PS02, PZ19, PZ21, PT11, RB06, Tar11, WXL20, WAvHD09, WJYL22, XYZC23, XXN21, Yan20, YXZ23, YWC23]. Modeling [DvdH97, FE01, GCDB20, GN15, HS88, HS91, DDK22, LnG22, LK23]. Models [AAES10, AF97, BBS96, BR97, BB97, CFM^{+}88, Gut97, HS92, HK96, Keb94b, PP02, RW92a, Sch96, SY92b, Van96a, WSX18, ARK03, AR14, AR00, AL15, ABS12, AS18, BSF21, BCZZ19, BKB22, BBP20, BB02, BP01, BKM10, BDPS11, Çağ16, Cai02, CR21, CF12, CY06, CH11, CLSS13, DXX16, DKC21, DF08, DCL^{+}05, FH17a, GKPR99, GFH22, Gup16, HSB23, HCX06, Idr23, LP15, LWZ16, LXYL21, LZ19, MIK23, NS04, NH04, NS22, NYZ04, NBR13, Ols01, OE10, Pek99a, SK20, SBZ16, Sta11, YWZH21, YSM06, ZB12, ZHZ15]. Moderate [WWH18]. Modes [Dud97, NJ95]. Modular [Her95]. Modulated [BR97, CJ96, DP95, Abd17, AÖ12, BDM15, KL10, LK09, SS20].

Modules [AENS94, ES91, Men94b, YBL22b]. Moment [KG92, Van88, WWS18, dK89, Pap21, Pen15, Spi15, Tag00]. Moment-Iteration [dK89]. Moments [AvER95, PR08, CL23, NWB13, YWZH21]. monitored [BGDR03]. monitoring [ASKZ21, GZY22]. Monotone [AS96, CHYZ93, GKL87, LL98, Bro09, KM19, Ros06b]. Monotonicities [Lin97]. Monotonicity [DLB12, GK88, HJ95, HP89, JD07, Kul04, MS91, Miy04, Miy05, SI04, Van96a, Van97, YHZ06, vDTW88, ASZ17, AO18, AO19, BEFK04, PT11, Rol05]. Monte [EGL92, FH95, Hei01, Lom94, RW92b]. Montmort [Sch98]. mortality [LK23]. Moshe [BCZZ19]. motion [CLS21, DM03, Fen22, PH23, Rao15]. motions [LKG19]. mountain [BPvdDS99]. Move [Cho99, CL06, CW90b, Fil98, Hil99]. Move-forward [CL06]. Move-to-Front [CW90b, Fil98, Cho99, Hil99]. Moving [MJ95, MP97, Ros99a]. Moving-Average [MP97]. MSMPR [CH87]. MSP/1/ [CSP12]. MTBF [Ang01]. Multi [CK15, DvdH97, ET87, EM19, Mit18, PT94, SB13, Ano21a, AÖ12, BKK17, Çağ16, CMZ17, DO16, EFG17, IVA15, NS17, POdCW21, TAM^{+}21, YBL22b, YBL23a, YBL23b, ZAE19]. Multi-armed [CK15, BKK17, EFG17]. multi-channel [TAM^{+}21]. Multi-Class [EM19, PT94, SB13, Mit18, IVA15]. multi-dimensional [CMZ17, DO16]. multi-disciplinary [Çağ16]. Multi-Echelon [DvdH97]. multi-item [AÖ12]. multi-resource [NS17]. Multi-server [ET87]. multi-state [Ano21a, POdCW21, YBL22a, YBL22b, YBL23a, YBL23b, ZAE19]. Multiarmed [RS98a]. multichain [Wak00]. Multichannel [LT00, YM92]. Multiclass [BvHvO97, HK96, RS89, Rig96, SP98, BHM09, Khe08, THM04]. multidimensional [FL22]. multimedia [SKA09]. multinode [Yam99, Yam00]. multipath [XC10]. Multiple [AMK18, Cao17, CW90b, KL98, LT00, Pre94, RS90, ASZ17, BCZZ19, Bro00, CHKL13, CN00, CR07, Che20, DF21, DZB12a, DZB12b, FOF07, Mah19, Mat19a, MK19, Mat19b, MK21, NS21, Ros11, ZB12, ZHZ15]. multiple-outlier [ASZ17, BCZZ19, DZB12a, DZB12b, ZB12, ZHZ15]. multiple-player [Ros11]. multiple-recipe [CHKL13]. multiplicative [Gup16]. Multiscale [BB97]. Multiserver [Che97, KS92, PT96, vDTW88, ADGZ09, JD07, Tij08]. multisets [Mah13a]. multiskill [Bhu09]. multistage [FZ19, KAR09]. Multistate [LL98, HKA22, Men02]. Multistation [AM87]. multitype [AHW10]. Multivariate [ANFPLCSL05, KK06, MC91, ZH07, BBHK10, BGR11, CB20, DFS10, HS13, KK05, LZH16, Mon07, Roy02, WH20]. Multserver [SS92]. mutation [HM19]. mutual [LFNC23, LMZ22]. Myopic [MCB18].

N [Kha13, CG96, CG96, FY11, PR08]. Nakamoto [GPM22]. names [AR14]. Nash [Che12, SD22a]. natural [GKL23]. Navigation [BA18]. NBU [DFS10, Li05]. NBUC [Li05]. NBUE [Fro11]. nearest [Ald22, LAWZ01].

Necessary [Eco03]. Negative [BBS97, CP93, TS07, WJ18, Jia19, Mat19a, Mat19b, SC03, Tim16, XH11, XDZ15, YDJ23]. Negatively [WWS18]. Nested [HBvdMB18, ZP16]. Nets [HS88, HS91]. Network [AY94, BS91, Dod89, EGL92, Hun91, JLAE18, LL90, Lom94, YM92, Abd16, AW05, AGM01, BvHTvO18, BHM09, DP21, DRR22, DL10, HBKK03, HM16, HBG21, JLAE17, JC18, KMS15, LFNC23, MK17, Mat17, Mat19a, MK19, Mat19b, TAM^{+}21, Tim16, VMB00, Wan17a, Yin21]. network-unfriendly [Abd16]. Networks [AF92, AB93, Bar96, BT95, Bou96, Buc97, BvHvO97, CP93, Cha94, CP95, CL91, FS89, GS92, Gel93, HL92, HL91, Key90, Lin97, LT00, PMB96, RS95, RW92b, Sch95, SP98, Tak96, Wan21, Whi88, vD95, AS00b, AY08b, BR17, BH02, BJ19, BA18, Cag17, CM00, CFKP99, DKSS06, DM22, DF21, Eco03, FM16, FH17a, FMR17, FWH21, Fou16, GRSV14, JL19, JCDJ22, Kad17, KT01, KAA12, LW14, Mar16, MK21, Pek21, PM17, Wan17a, Yin17, ZAE19, Zaz04, Zha21b]. Neural [JLAE18, Abd16, BR17, HM16, HBG21, JLAE17, TAM^{+}21, Tim16, Wan17a, Yin21]. neutral [GT10]. Neutrons [Kol94]. Newton [McN89, WK88, Whi89]. next [vSvDdK^{+}20]. NN [Rot93]. No [DG91b, Fre91a, HKL94, BH23]. no-mean [BH23]. Node [CHL94, AW05, CBBG20, Mah13b]. Nodes [Men94a, Abd17, Fou16]. Noise [Bax96, Bre90, AAK19]. Noises [Miy04, Miy05]. noisy [RL03, RL05]. Non [CF06, MCB18, WWS18, XWTW18, BdlPS15, Bru16, FL22, HLW16, IVA15, LC07, MG21, Mat17, MK19, PT20, Spi15, SXWW22, Tar17, Tim16, ZCJL23]. non-cooperative [Bru16]. non-entropy-based [MG21]. non-explosiveness [Spi15]. non-Gaussian [LC07]. non-homogeneous [PT20]. non-negative [Tim16]. non-observable [Bru16]. Non-parametric [WWS18, XWTW18]. non-Poisson [HLW16]. Non-preemptive [MCB18, IVA15]. Non-Quasiconvex [CF06]. non-random [BdlPS15]. non-stationary [FL22, HLW16, Mat17, MK19]. non-tradable [SXWW22]. non-zero-sum [ZCJL23]. Noncentral [LC97]. Nonexponential [MvD97]. Nonhomogeneous [GK88, Koc96, PSZ00, SSS02, Yan98, BLRS01, dCBB22, Hua20, ZYSZ21]. Nonidentical [KX10]. Nonlinear [Fan96, Idr23]. nonmonotone [Kul04]. Nonnegative [CCdO01, Pap21]. Nonnegativity [HYL07]. Nonparametric [YXZ23, RSS09]. Nonpreemptive [Gla91]. nonreporting [PS10]. nonsimple [HPS08]. nonstandard [BGG07]. Nonstationary [Pen15, BY99, MK21]. Nonuniform [Ald88]. Normal [Jun88, Lin90a, Lin92b, Lin95, Bro01, LZH16, Mon07, Mon13, NH04, PZ21, Röl22, Sta11]. Normality [Ang95, CBL93]. normalized [GR12]. Note [Ald93, BBS97, Cha94, Fre91b, Fro91, HvdL03, HH00, Kij92, Mad93, MvD97, Miy04, Miy05, MPP93, NQ01, PR88, Ros88a, Tam98, Tij08, Van96a, vD88, AGC07, BMR05b, CF13, FL13, FWGD00, Kon03, LW19, LZ19, Men02, Rol05, Ros03b, WWZ22, vZS04]. Notification [SO94]. Notion [ANFPLCSL05]. Notions [SS87, DFS10]. Number [Abr93, OGP95, RS89, Ros89, Sen93a, YLB96, AK08, Cha99, CGM21, LLY22, Mi00, Pek99b, PY14, Röl22, RB06]. Numbers [Kol92, SAS89, HL01, WWZ15]. Numerical [BPvdW91, GV90, TV91, dI06]. NWU [BL02].

Object [Ber96]. Objects [Ste87]. observable [BCD^{+}14, Bru16]. observation [Tar11, XYZC23]. observations [LS11]. Observed [NQ01]. obsolescence [PLS23]. obtain [NH04, Ros23]. occupancy [CLSS13, LRW09]. Occupation [Ros87a, Ros88a, Sie93, YZL00]. Occurrence [ZL02, DDK22]. Off [KL98, BKKP05, FZ15, VM04]. offering [FS23]. Offers [DY90a]. Oil [Ben90, Ben92, Sta11]. Oilfields [Ben90]. old [HKS07]. Olkin [Joe12, FL13, GKB13]. On-Line [CW90a]. On-Off [KL98, VM04]. On-road [HM16]. on-times [BKP01]. On/off [BKKP05]. One [BK03b, For93, Fre91a, Fre91b, Lin19, Ros11, vdHvHE01]. One-Armed-Bandit [BK03b]. One-Machine [Fre91b, Fre91a]. one-shot [Lin19]. one-way [vdHvHE01]. Online [BMSB20]. Open [Bar96, LK96, Ali23, GHvdL07, MK21]. open-loop [GHvdL07]. opening [Mei14]. Operating [PP02, Pan08]. Operation [Dud97, MD06, Rig02]. Operational [XZP18, BSF21, DO03, ÖAA01]. Operations [Kol92]. Operator [ER95, KL15]. operators [XMH21]. Optical [Smi97b, RB06]. Optimal [sADL02, Asl22, AL91, BPLD21, BC98, BL88, BB06, Ben90, Bro99, BK95, CCdO01, ÇÖ15, CH88, CXY12, CHS87, CHH90, CS92, DJ07, DG91b, DK22, DH94, Dud97, Elc94, FZ93, FZ15, Fro95, Fro99, GKL23, GS93, GTDKÖ09, HH96, Kay09, Key90, KAR09, KLT12, Kyr03, LKH14, Li94, LFNC23, LHY19, Lin19, LR97, LZY23, MJ95, MAS06, Men94b, MZS16, Nic16, ÖK14, PT94, PT96, Pan08, Pap21, PSC92, Rig96, Rig98, Rig02, RL03, RL05, Sai98, Sat97, SK11, Sie93, SBZ16, ST94, SvNKC00, Sto05, UKM12, VT92, Van97, Van96b, WWZZ23, WB96, Wee90, XMBG02, Xu23, YH88, Yao89, Zha21b, Zie96, vdL11, AAR11, AS00b, BEFK04, BJ23, BCD^{+}14, BKK17, GHvdL07, GN14, HH16, HS13, LWL19, LCYY22, LSSZ16, Ros15, SH00, Sen01, Vil16, WS06, XC10, ZP16, ZK06]. Optimality [BC94, BY99, CHRW94, CF06, FOF07, FR94, FL05, FY11, HK90, LT00, Ngu91, Sen93a, CGK99, LAF99]. Optimally [AMK18, AS00a]. Optimization [BB97, BB98, CF06, DG94, FK96, He89, KNN15, MD06, RS94b, BLBPV17, GN15, JA06, LQ15, PdB22, Pek21, Wan21, ZZ13]. Optimum [GS98, Yin17]. Option [FH95, Hei01, Rao15, DO03, FKL02, GB22, KKJ22, PH23, Yan20]. option-based [DO03]. Options [Wan18b, WSX18, CW23, FL05, FH13, FWGD00, Han19, mKK22, MCL22, PG22, PGZ22, PZ19, RS00, Wan17b, WWL17, WXL20, XXN21]. orbit [Dim17, Dim23]. Order [ANFPLCSL05, BZ13b, BCS92, Bro09, BN18, CLC98, CK88, DG91b, Fro95, GD18, Keb94b, Kha13, LY13, LAWZ01, LMZ22, MMS93, MS91, Ott93b, Pla92, SSS04, SSS08, Tak90, XH11, Xu13, AKP05, AEAN23, ASZ17, AHOS23, BBHK10, BMR05a, BGR11, BNT06, BN11, CM05, CLSS13, FRR02, HN19, HSB23, HZW01, HZ05, HZ06, HJK07, HWZ12, HM20, Koc23, LX06, LMH14, LMH12, MH10, MGM21, PKK13, Pap21, SI04, TA11, WWH18, XH08, XH09, XZ11, YZZ23, Yu21, Zar23, ZZZ21, ZK06, ZL09, ZB12, ZHZ15, ZH07, vdWY03]. Order-book [GD18]. Order-Preserving [Keb94b]. Ordered [YBL23b, Bha99, BB00, HYL07, Lin00, Lor18]. Ordering [BZ13b, CPS89, DZH12, DKC21, DKB22, DZZ18, Fre91a, GV90, HJK07, KK06, Kha13, LY13, MHZ10, SH23, SO94, Wan18a, WAI95, Win97, XH09, Xu13, Ali23, ABJ22, AO18, AO19, BBHK10, BGR11, DKSS06, HN19, HZ04, KA04, KK05, LZH16, MIK23, Roy02, XH08, XZ11, ZK06, ZH09]. Orderings [Alz88, BKB22, DLP90, Keb94a, SW95, YWL22, AK05, BFRR01, BOR06, CV22, DFS15, DLU99, DZB12a, DZB12b, FRR02, HZ06, MH10, MPH13]. Orders [HH00, Kor03, NBR97, SS98, BS06, BHK03, HZW01, wWW04]. organ [BDPS11]. oriented [Zha21a]. Ornstein [KR94, YWZH21]. Orthant [Han03, XWTW18]. Orthant-dependent [XWTW18]. Oscillation [Ang95, CBL93]. other [FS20]. outcome [Lef06]. Outlier [Ang89, ASZ17, BCZZ19, DZB12a, DZB12b, KK01, ZB12, ZHZ15]. Output [YM92, vD95, BKKP05, Cal10, Sto05]. output-queued [Sto05]. outsourcing [BLPS05]. Overflow [Man96b, Ngu91, KNN15]. overhead [ZPD06]. oxide [KN21].

P [Kha13]. Packet [Ott93b, YM92, Cag17, DF21, Kad17, Yin17, Zha21b]. Packing [CFFH98, CW90a, DHR00]. pages [Sch96]. Pair [SAS89]. Pair-Connectivity [SAS89]. paperback [Sch96]. paradigm [OK08]. paradox [Ang06, BJ19, Bro06, HZ04, Ros03a]. Parallel [BC94, BBCP97, BCGW92, CHRW94, CHW89, GV94, GV90, GSW93, HK92a, HK92b, HW92, KX08, LR97, Men94b, Rig92, VR89, Wan18a, WB96, YZ89, ABJ22, BM02a, Bru16, DKM16, DLB12, DDL14, FZ15, FH17b, HLM16, KX07, LL03, LWL19, Mah13b, Mah14, Mah19, NDZ22, XMBG02, YBL22b, ZDZ20, Zha11]. parallel-series [LWL19]. parameter [BHM09, EA06]. Parameterization [Lyn97]. Parameterized [Lyn97]. Parameters [AF97, EGL92, AHOS23, HSB23, PdB22]. Parametric [BB97, BB98, WWS18, XWTW18]. Pareto [BJ23, NZP18, NDZ22]. Pareto-optimal [BJ23]. Parisian [HLZ23]. Part [Her95]. Partial [Bro89, Fre91a, OS90, AAES10, BFRR01, CW22, LQ15, Roy02, Tar11, ZZ13]. partially [Lor18]. particular [FNC20]. partitioning [DO16]. partitions [BHP05]. parts [BvHTvO18, LST^{+}22]. Passage [BS87, DP93, Lef96, RL93, SS87, vdHHV01, Zaz07]. past [ASKZ21, CLN20, Qiu19]. Path [Mah91, Men94a, SY92b, VT08, ALM07, KP07, Pen10]. Paths [CBM18, Dod89, Van01, ZBA05]. patient [AZR08]. Pattern [HT96, Ros99b]. patterns [LK23]. payment [Xu23]. Peaks [Ros87e]. Peer [VT08]. Penalties [For93, PT94, CZ06]. penalty [Xu23]. Penetration [Kol94]. pension [KKL17, LP23, PC22]. Percentage [Lin94b]. percentages [Nad17]. percentile [SY15]. Percentiles [ET87, Lin90b, Lin94a]. Percolation [BS14, vdHHV01]. Peres [LLY22]. Perfect [Fil98, CS03]. Performability [Van96a, vD95]. Performance [Bru16, CBBG20, KT01, KSM88, LL91, LDVT22, Lu09, Man96a, SVL02, Van97, AS18, BH02, CW22, GRSV14, LSW21, LW14, Pek99a, THM04, WÇC22]. performance-per-cost [LSW21]. Period [AW95, AGC07, Bar08, CGM21, Pek99b]. Periodic [Ros87c, DDK22, FLL12, LXH04, vdHvHE01]. periodic-review [LXH04]. Periodically [AF97, vdHvHE01]. periods [Gra08, PR08]. Perishable [Per97, BLPS05, BBP20, GLW11, NPS04]. permanent [KN00]. Permutation [Ang95, CBL93, PSC92, BHP05]. Permutational [CZL96, GRSV14]. person [Che09, Che11, Che12, MI10, TvdW06]. Perspective [BF91, Eli05]. perturbation [CL23, FWGD00]. perturbed [YXZ23, YZL00]. PES [Ano87b, Ano87a, Ano87d, Ano87c, Ano87f, Ano87e, Ano87h, Ano87g, Ano88b, Ano88a, Ano88d, Ano88c, Ano88f, Ano88e, Ano88h, Ano88g, Ano89b, Ano89a, Ano89d, Ano89c, Ano89f, Ano89e, Ano89h, Ano89g, Ano90b, Ano90a, Ano90d, Ano90c, Ano90f, Ano90e, Ano90h, Ano90g, Ano91c, Ano91b, Ano91e, Ano91d, Ano91g, Ano91f, Ano91i, Ano91h, Ano92b, Ano92a, Ano92d, Ano92c, Ano92f, Ano92e, Ano92h, Ano92g, Ano93a, Ano93b, Ano93c, Ano93d, Ano94b, Ano94c, Ano94d, Ano94e, Ano95b, Ano95a, Ano95c, Ano95d, Ano95e, Ano96a, Ano96b, Ano96c, Ano96d, Ano97a, Ano97b, Ano97c, Ano97d, Ano98a, Ano98b, Ano98c, Ano98d, Ano09b, Ano09a, Ano09d, Ano09c, Ano09f, Ano09e, Ano09h, Ano09g, Ano10b, Ano10a, Ano10d]. PES [Ano10c, Ano10f, Ano10e, Ano10h, Ano10g, Ano11b, Ano11a, Ano11d, Ano11c, Ano11f, Ano11e, Ano11h, Ano11g, Ano12b, Ano12a, Ano12d, Ano12c, Ano12f, Ano12e, Ano12h, Ano12g, Ano13b, Ano13a, Ano13d, Ano13c, Ano13f, Ano13e, Ano13h, Ano13g, Ano14b, Ano14a, Ano14d, Ano14c, Ano14f, Ano14e, Ano14h, Ano14g, Ano15b, Ano15a, Ano15d, Ano15c, Ano15f, Ano15e, Ano15h, Ano15g, Ano16b, Ano16a, Ano16d, Ano16c, Ano16f, Ano16e, Ano16h, Ano16g, Ano17b, Ano17a, Ano17d, Ano17c, Ano17f, Ano17e, Ano17h, Ano17g, Ano18b, Ano18a, Ano18d, Ano18c, Ano18f, Ano18e, Ano18h, Ano18g, Ano19b, Ano19a, Ano19d, Ano19c, Ano19f, Ano19e, Ano19h, Ano19g, Ano20b, Ano20a, Ano20d, Ano20c, Ano20f, Ano20e, Ano20h]. PES [Ano20g, Ano21c, Ano21b, Ano21e, Ano21d, Ano21g, Ano21f, Ano22b, Ano22a, Ano22c, Ano22e, Ano22d, Ano22g, Ano22f, Ano23a]. Petri [HS88, HS91]. Phase [Zie96, vdMR08]. Phase-Type [Zie96]. Physically [Bal92]. Picking [wWW04, LAWZ01]. piece [NL07]. Pitfalls [RSZ93]. plan [LP23, PC22]. plane [CBG15, HP03b, Zha21a]. Planning [MD06, BCR05, BMSB20, HKS07]. Plans [CHYZ93]. Play [DY90b]. Play-the-Leader [DY90b]. player [Ros11]. plays [CR19, CR21]. Plots [DK13]. Pocket [Lin90a, Lin92b]. Pocket-Calculator [Lin90a, Lin92b]. Point [BBCP97, BS98, CHW89, DG94, HW90, Kol92, SW95, Tak96, Idr23, LS11, PLS23]. Points [Lin94b]. poison [SD97]. Poisson [BL87b, dCBB22, CL23, Cov10, FG91, FS20, GS23, GLW11, GK88, He89, HLW16, KPS03, Koc96, LRW99, MP97, NWB13, NYZ04, PR94, PSZ00, PT20, RMS94, Ros06b, SD97, SZ03, XYZC23, XDZ15, YXZ23, ZYWN06]. Policies [AK96, BK95, CCdO01, FZ93, FL98, Gla91, HPR96, HK96, MCB18, PP89, Sen93a, SS89, AK08, AGM01, BPLD21, BLPS05, BH02, BY99, BP01, Bhu09, BKK17, CGK99, CK20, FL05, FS23, GTDKÖ09, GN14, KAR09, KNN15, LKH14, Sen01, UKM12, Vil16, ZK06]. Policy [AL91, CRP93, HK90, HK92b, LN92, McN89, Per97, WK88, Whi89, AW12, BGDR03, CXY12, CHL03, EFG17, GHvdL07, HS13, LC07, MK23, Ros15, WWZZ23]. Policy-Improvement [Whi89]. Policy-Improvement/Newton [Whi89]. Pollaczek [JS96]. Pollaczek-Khintchine [JS96]. Polling [AY94, Bla98, Bor95, BWY93, CS93, FL98, Sch95, Van97, VY08, vdMR08, ABS12, AS18, BW14, DPV16, GHvdL07, Ols01, Pek99a, WAvHD09, vdHvHE01, vdWY03]. Pólya [CB20, Che20, Mah23, Nad17]. Polymer [Tay87]. Polynomial [BI88]. polynomials [DM08, GCDB20, LP15]. pool [Hwa00, HL01]. Pooling [Tur98, HL04]. Population [LOdBP87, ASZ17, LCYY22]. Populations [BZ13b, DY90b, Kha13, LY13, Xu13, BBCN01]. Portfolio [LQ15, SBZ16]. portfolios [DKB22, DF08]. Positive [CP93, HJ95, KG87, HL01, Mat19a, MK19, Mat19b, PLS23, SC03, XDZ15]. Positive-Valued [KG87]. possibility [CVB05]. possibly [CH11]. postponement [FH13]. Power [Bla98, HS88, HS91, SO94, CM05, DO03, EA06, WXL20, ZBA05]. Power-Series [Bla98]. powers [Bel11]. pp [Sha98]. pp. [Bax96]. PR [NQ01]. Precedence [LR97, DFS15, MGM21]. predator [Kyr03]. predefined [PKK13]. Prediction [Bla95, ZPD06]. Predictive [GM96, CHKL13]. Preemptive [Fro91, HW92, Mit18, Sha88, CSZ03, Cao17, IVA15, MCB18]. preemptive-repeat [CSZ03]. Preface [Ano23b]. preferential [Zha21a]. Prefix [MO95]. premium [MZS16]. preparation [DPV16]. Presence [DLR87, GM96, HM87]. Preservation [BS08, BC13, BS06, BLS03, IASB23, MXH18, MMS93, SW95, XMH21, KM07]. Preservations [Li05]. Preserving [Keb94b]. Prevention [Kwo14]. preventive [HS13]. price [GT10, Yan20]. Pricing [FH95, GLW11, KKJ22, RMM87, RS00, WWL17, Wan18b, WJYL22, YWC23, BCD^{+}14, CX23, GTDKÖ09, GT10, GB22, HMW09, Hei01, mKK22, KLT12, PC14, PG22, PGZ22, PH23, Rao15, XC09]. principle [WWH18]. Principles [PW98, AILR08, MZS16]. Priorities [CHL18, HvdW97, Mit18, IVA15]. Priority [CLC98, PT93a, SP98, SO94, VM12, Lyn03]. Probabilistic [Ald93, ADS88, Chy91, GV90, IBI08, Keb94a, ABI22, DMM16, DRR22, DCN05]. Probabilities [Ald88, Dub98, GV10, KMT18, LT96, Las98, MS91, Ros87a, Ros98, SA18, Yao89, vD88, BOT09, CM05, CGK99, Cov10, FS20, FL22, GT07, HL04, Kla00, Lu13, Mat17, MK21, Miy02, PT11, Ros22, SGF20, Tan06, TS07, VMB00, vSvDdK^{+}20]. Probability [And95, BCS92, Gou94, HT96, Lin92a, Lin92b, MY11, Miy92, MT93, Sad98, SR90, YZ03, Bri14, BH23, Çağ16, CY06, HY21, KM19, LZ10, NYZ04, PL20, Tan04, Zha99]. Probing [Ald88]. Problem [BC98, Ber96, BWY93, BK03b, FGR10, PT93b, Pre94, RMM87, Rig87, Sai98, Sch98, TS96, Ali23, BJ13, Bro13, CMZ17, CR07, CSL09, CL06, DvdLR11, FOF07, FH13, GN14, KR01, LS11, Lin04, Rig11, Ros03b, Ros13, SH00, Tag00, VA05, Yan20]. Problems [Bro89, Fre91a, Fre91b, GV90, CY10, CR15, HS13, KM12, ME14, PKK13, PRZ22, YXZ23]. Procedure [BLP93, Rei90, KS12]. Procedures [CHS87, CS90, Jun88, BLBPV17, ZP16]. Process [Cal97, CJ96, DG94, Dav95, DP93, Hes92, Kha99, KG89, KG92, KR94, Lef01, Lyn97, Ott93a, Ott93b, PR97, Tay87, WAI95, YZ98, Yec88, YLB96, ZL02, BC13, BM11, BH23, CB20, CX23, Cov10, DDK22, DA03, GT10, HS13, JC18, KPDR16, Kyr03, Kyr04, LS11, MT16, PT20, SS20, WWZ22, XDZ15]. Processes [AFP88, Bal92, BW95, Bha87, BBS97, BS98, BK95, CCP90, DY90a, Fan96, GK88, HW90, Joh98, KR87, Kij94, Koc96, KW89, Lef96, MR95, MP97, McN89, MS93, MS91, Ngu91, PS93, PW98, Sen89, SW95, SSS02, SY92a, Smi97b, SD97, Sti07, TSW97, Wak98, Wei88, Whi89, YLB96, YM92, BS08, BGG07, BLRS01, BOR06, BPvdDS99, BdlPS15, dCBB22, CW23, CN00, CF13, CZ03, CLSS13, HY21, Hor01, HP00, HM19, JZY22, KSS15, KPH10, KS01, LDVT22, MZ17, Miy02, MOSU10, OK08, PSZ00, Rao15, SX22, Spi15, SZ03, SZ05, VHBSW22, Wak00, WWL17, Whi00, ZCJ04]. Processes-Martingale [HW90]. Processing [BC94, GV90, HW92]. Processor [BK93, FM17, HvdW97, Ros90b, Tak90, EZB06, LW13, Lu10, Lu21]. processor-sharing [LW13, Lu21]. Processors [LM87, LR97, Spi91, HP03a]. Product [BCS92, GS92, KS92, Mar16, NS21, RW92b, VV88, Yao89, BB02, BvHTvO18, DO16, Yam99, Yam00]. Product-Form [RW92b, Mar16, NS21, BB02, BvHTvO18]. Product-Type [BCS92]. Production [AM87, AF97, PP02, Van96b, YZ98, AÖ12, BPS14, CXY12, GTDKÖ09, KAR09, PC14, SLZ04, Sta11, XC09, ZAMB21]. Production-Inventory [PP02, BPS14, CXY12]. products [GLW11, HKS07]. profile [FMR17, KM16, ZD19]. profiles [ÖAA01]. Profit [BT95, XC09]. profitability [GPM22]. prognosis [LDVT22]. Programming [MD06, YZ98, BMSB20, GM16, JA06, RB10, RB12, Sen01]. Programs [CCdO01, SH00]. progressively [MG21]. project [RL03, RL05]. projection [KKL17]. prone [AAD08]. Proof [Ang95, LC97, Ros88b]. propagation [DCL^{+}05]. Properties [AAR11, BZ13b, BL89, BFR99, BPT89, BN18, CLN20, CS94, DZZ18, GK88, Kha13, LG89, LY13, LMH12, MMS93, MC91, Ott93b, RMS94, Rei90, Xu13, AAC14, AEAN23, BC13, BLS03, BVA13, BN11, CM00, DLB12, EQ22, GK07, HZ05, HWZ12, IASB23, JD07, KK01, Khe08, Mah13b, Mah14, NBR13, PO15, QWW19, SH23, SAF22, TA11, VHBSW22, Zar23, Zha21a]. Property [CHJ95, DP95, DLP90, MY11, SSS04, SSS08, Yan98, ZYL10, CB20, CLSS13, KL12, KM07, LK09]. Proportional [Che09, CGKS96, KX08, KX07, Koc23, NDB13]. Proportionate [PSC92]. protocol [Lu09, Lu10]. ps [FZ19]. ps-queue [FZ19]. pseudo [RL99]. public [KKL17]. Puiseux [HPS99]. pure [GT10, Tar11]. put [WWL17].

QBD [HY21]. quadrature [dI06]. Quality [Ang97, DR95a, HMW09]. Quantifying [AILR08, ZZ13]. Quantile [KMS20, DMM16, Qiu19]. Quantile-Based [KMS20, DMM16]. quantiles [MNH15, WWH18]. quantity [AÖ12, ZK06]. quarter [CBG15, HP03b]. quarter-plane [CBG15, HP03b]. Quasi [AFP88, DKK94, KW89, Yam99, Yam00, DXX16, MT16]. quasi-birth-death [MT16]. Quasi-product [Yam99, Yam00]. Quasi-Reversibility [KW89]. Quasi-Statically [DKK94]. Quasi-Stationary [AFP88, DXX16]. Quasiconvex [CF06, FL20b]. Queue [BB96, CG96, CRP93, CFM^{+}88, DG91b, Dud97, ET87, Eli08, FK96, FY11, Fla89, GV10, HK90, HK92b, JS89, LN92, Miy92, NQ01, RSZ93, SS92, Sha88, SK11, dK89, AGC07, APR20, BPLD21, BM02b, Bro00, BdBS20, Cha99, CGM21, CBBG20, CW22, DMvU07, Dim23, DA03, EK09, EZB06, EBTD22, FZ15, Fra13, Fra05, FZ19, GZDW21, GZ09, HH16, HS07, IVA15, KN00, LAF99, LZ10, Lu09, MN05, MK23, MC11, PY14, Rav21, Ros06a, SZ02, SKA09, SC03, THM04, TS07, Tij08, WS06, YT23, ZCC23, vdHvHE01]. Queued [ALR00, Sto05]. Queueing [AdHBvO14, BW95, Bou96, BS91, BvHvO97, Cha94, CP95, CSP12, CJ96, CLC98, DP95, Elc94, HH96, HL92, KL98, LL90, Lin97, RS89, Sen89, SP98, ABEK17, AJK01, BR17, CM00, DP21, Eco99, HAA22, Jia19, KMS15, NS21, ÖK14, PY16, SNS19, WFM09, Yam99, Yam00]. Queues [AW95, AdHBvO14, AB93, BCG89, Bar96, BZ05, CPS89, CP93, CGKS96, CPR91, Gou94, GSW93, HvdW97, HK92a, HBvdMB18, Man96a, MS91, MT93, ÒLC95, PT93a, PT94, PV11, PZ04, PMB96, RSZ93, SB13, Tij92, WB96, AY08a, Bar08, BB06, BDM15, BCD^{+}14, BPvdDS99, BDPS11, Bru16, CM00, CCC05, DKM16, Dim17, FM17, HKK10, Hor01, JST01, JD07, Koo04, LW13, LK09, LW14, LSSZ16, Lu21, ML01, Nak01, Pan08, PW12a, PW12b, Pen15, Rol05, SI04, SAdK04, TZS22, VHBSW22, Whi00, XMBG02, XLSZ04, Zha99]. Queuing [Ott93a, sADL02, AL09, AKR^{+}04, AGM01, ADGZ09, AILR08, BM02a, BMK08, DKSS06, Fen22, Has07, LRW09, Lin02, Miy02, RL01, RB10, YHZ06]. Quick [NS04]. Quicker [GV90]. quickest [JL19].

R [CCC05]. Radio [RS95, YM92]. Rainflow [Joh98]. Random [Ald93, BC94, BF90, BS14, BS98, BvHvO97, But89, Cha94, CLC98, ÇÖ87, CFFH98, CW90a, DK09, DS87, FH94, GMW14, Han03, HvdW97, He89, Her95, HM94, Hwa00, JLAE17, JLAE18, JR92, KWA96, KG87, Kol92, Mei91, NZP18, OS90, Pal90, PP02, Reg90, Rig87, Ros87e, Ros88b, Ros90a, SS89, SW95, Sha88, Sie07, Sob95, Ste87, Van01, VY08, WWS18, WJ18, Xu91, Yin21, vdHHV01, Abd16, ALM07, AAKN09, AR14, Ald22, ASKZ21, AKS11, BC13, BM02a, BR17, BPS06, BMR05b, BKL22, BPR10, BdlPS15, BC18, CZ06, CR19, CMH13, CBG15, DP21, FHS13, FH17a, FPOP08, Fro11, GT07, HP03b, HPS08, HZW01, HYL07, HM16, HBG21, HL01, HL04, KA16, KMS15, KX12, Kon02, Kon03, KAJ07, KAA12, LLY22, Lu13, MPH13, NK05, NYZ04, Pap21, PR04, Rat19]. random [Röl22, SVL02, SK23, SZF18, SD22b, Sti07, TAM^{+}21, Tim16, Wan17a, WN10, XH11, ZD19, ZL09, vdHvHE01]. Random-Access [Ros88b]. Random-Bit [But89]. random-time [GT07]. Randomized [DY90b, FR94, ABEK17, FOF07, LP15]. Randomly [RB05a, RB05b, SGF20]. Randomness [Rot93]. Range [Ott93a, ZL09]. ranges [BCZZ19, MH10]. Raphson [McN89, WK88, Whi89]. Rare [CMZ17, KMT18, Lag06, Vil18, ZL02]. Rare-event [KMT18]. Rate [AM87, AH96, BS93, BSS98, BLS12, BF91, CRP93, CJ96, CH87, CHJ95, Din95, Elc94, HH00, HV95, HT96, KK06, Lyn99, RB05a, YZ03, Ali23, AO18, AO19, BSL15, BK09, Bro14, CR01, DZB12a, DZB12b, FE01, GK21, GK07, Gup16, KK05, LW19, Lyn03, MAS06, NH04, NBR13, PM17, PT20, RB05b, Sta11, WJYL22]. Rates [BZ05, HH96, HY90, KX08, AKR^{+}04, Ang04, APR20, BPLD21, BB06, BL02, BKP01, BKKP05, Cai02, FNC20, GV11, KS02, KX07, LWZ16, LW14, NS17, TYZ19]. Ratio [LCYY22, Lu10, PT11, Sad98, Wan18a, AEAN23, BBHK10, BGR11, Chi21, DZB12a, DZB12b, HZW01, HZ04, HM20, ZZ13]. Rayleigh [MG21]. Rayleighs [WN09]. re [Mei14]. re-opening [Mei14]. Reachability [CL91]. reachable [DO16]. Reading [Sha98]. real [MT22, TvdW06]. real-world [TvdW06]. Realizable [Bal92]. Really [Ros87c]. Recall [Sai98]. Recapture [LOdBP87]. Receivers [LE94]. recipe [CHKL13]. reconsidering [Zog23]. Record [AR97, GK88, Tam98, AAC14, BCMS23, BOR06]. records [BCMS23, Cho99]. Rectangle [CFFH98]. Recurrence [CK88, FH94]. Recurrent [Ban92, YBL22b]. Recursions [AS96]. Recursive [BR97, Mah91, Rei90, Tar17, PL20, Zha21a]. Red [Dub98, Che09, Che11, Che12]. Red-and-Black [Dub98, Che09, Che11, Che12]. Redialing [CGK99]. reduced [CX23, HSB23]. Reduction [AW96, Ros88c, Ros90a, RL01]. redundancies [YWL22]. Redundancy [BEP88, FL20a, MMD11, NFMFSA20]. Redundant [LM87]. Reed [OE10]. reentrant [Fra19]. reentrant-line [Fra19]. reflecting [KMS15, OKA22]. refusals [Pek99b]. Regenerating [ZL02]. Regeneration [ACG95, FTC98, CN00]. Regenerative [BLP93, GI95, HM19, NS04]. regime [BDM15, GB22, LXYL21, LP23, ME14, Mat19a, MK19]. regime-switching [GB22, LXYL21, LP23]. Region [Rot93]. Regression [HCX06, WXWH18, WWS18, XWTW18]. regret [CK20, PdB22]. regular [LMH14, LMH12]. Regularity [Ken94, MS93, CB20]. regulation [BJ23, Sti07]. rehabilitation [BMSB20]. reinitialising [Vil16]. reinsurance [BJ23, MZS16, WWZZ23]. Rejection [Gou94]. rejoinder [BZ13a]. rejuvenation [FM17]. Related [LL91, HM19, KMT18, PO15, XDZ15, ZPD06]. Relation [HT96]. Relations [DFS15, KW94, AO18, AO19, HY21]. Relationship [PP89]. Relationships [NFMFSA20, PL20]. Relative [HvdW97, RS98b, HM21, IVA15, MFN17, TYZ19]. Relaxation [BF90, KR87, RBM02]. Release [BLP93, Fre91a, Kha88, LC07, SD22a].

Reliabilities [DH87, HL91]. Reliability [BI88, BL89, BPvdW91, BSS90, BPT87, But89, CE96, ÇÖ87, DG94, DK88, DR95b, FS89, Fis91, GK07, Gup15, HM87, JR92, LM87, LZ00, Lom94, Nay89, RMS94, SSS02, vdH87, Ang01, Ano21a, BS08, DDK22, DM08, Fro11, GSV16, Gup07, GKB13, HBKK03, IASB23, JCDJ22, KA04, LC07, Men02, NBR13, POdCW21, RKFI21, Sha98]. reliability-constrained [LC07]. Reliable [KSM88, AS00a, AS00b, NS04]. Relief [Dov87]. remaining [FC22, Lu21]. remarks [Gut14, Kha07]. Removable [FK96, BPLD21]. Removal [BB96, Gel93, JS96, MK17, Mat17]. removals [GZDW21]. Reneging [LK08, JD07]. Renewal [AH96, Ban92, BL02, CL91, Dav95, EGL92, FLL12, Lyn97, McN89, Ros87b, SD97, vdWSvN08, BS08, BC13, Bri14, BH23, CT23, DDK22, FNC20, LDVT22, Miy02, PK06, PT11, XDZ15]. Renewal-Type [Ban92]. Renewals [Ros89, BH23]. rényi [LZ11, LW19, Röl22]. Reordering [Ros90b]. Repair [KM88, KG89, KG92, PR88, vdH87, DKSS06, Fro99, RKFI21]. repairable [GZDW21, HKA22]. Repairman [BWY93]. Repairmen [Fro95, KM88]. Repeat [Fro91, CSZ03]. Replacement [CS92, SS89, Van92b, CV22, UKM12]. Replacing [AMK18]. replenishment [KTÇ14]. replenishment-liquidation [KTÇ14]. Replication [FG89b]. reporting [PS10]. Representation [BS87, CBG15, DMvU07, Fen22]. representations [NS22]. requirements [AÖ12, DF21]. RES [KSS15]. Reservation [FR94, FY11, Key90, Ngu91, FOF07]. reserve [BK06]. reservoir [BR17]. Residual [BFR99, GKL87, KG89, KX10, MGM21, MPP93, ASKZ21, dCBB22, DLB12, GV11, Gup16, HSB23, HZW01, JZK20, KP12, LL03, LZ11, NDB13, Olu11, SY15]. resilience [HSB23]. resilience-scale [HSB23]. resistance [KAJ07]. resistances [AAKN09]. Resistant [Ang89]. Resolving [Ali23]. Resource [Mit18, Tur98, EM19, GKL23, NS17]. Resources [Li94, NS21]. Respect [HH00, MS91, SI04]. response [CY10]. responsive [DA03]. restless [Vil16]. Restocking [BL87b]. Restricted [HS92, PV11]. Result [HP89, LC97, Pal90, Gut03]. Results [BR97, CSP12, KG87, Koc96, RW92a, VdK87, AKP05, BGR11, BPR10, CR01, CV22, DKC21, DKB22, DvdLR11, HN19, KS02, KX12, LX06, LZ11, Qiu19, SI04, WÇC22, XZ11, ZHZ15, vdMR08]. Resume [Sha88]. Retrial [AY08b, CRP93, HH96, SK11, Tak96, AGC07, Bro00, Dim17, GZDW21, MK23, ML01, SKA09, SNS19]. Retrials [Elc94, RB10]. Return [Elc94]. returns [FL05, RL03, RL05]. revenue [PdB22]. revenues [MK17, Mat19b]. Reversed [BSS98, HH00, CR01]. Reversibility [Che97, KW89, KMS15, SKL12]. Reversible [CHL94, HJ95, AHW10, Fra99]. Review [BZ13b, Bax96, Joe12, Kha13, Man96a, Sch96, Sha98, Xu13, BBP20, FW15, LXH04]. Revisited [Ang97, Ben92, SS92, MHZ10, SD22a, SZ03, XH08, XH09]. Revisiting [BGR11]. Reward [Fro95, HPR96, McN89, ARK03, HS13, Mi00, Wak00]. Rewards [CCdO01, HV95, IwW97, HPS99]. RHR [AK05]. Right [LM88].

Risk [BBS97, GT10, LP15, LM88, VM12, XZP18, AK08, BJ23, Bro99, Cai02, CW23, CX23, CL23, Chi23, DF08, FNC20, FL22, GZ09, HL14, HLZ23, LS11, LMH12, Miy02, PC22, PT11, Ros03b, SXWW22, SH19, WXL20, WH20]. risk-minimizing [SXWW22]. Risk-neutral [GT10]. Risks [LM88, AP99, DKB22, MNH15, NS22]. Risky [McN88]. RNA [GN15]. road [HM16]. robin [Mal22, Mal23]. robot [VM12]. robot-safety [VM12]. robust [CVB05]. Role [AENS94, ES91, Gup07]. Roller [GV11]. Roller-coaster [GV11]. Roots [CSP12]. Ross [AH96, Kha07]. Rotation [LK23]. round [Mal22, Mal23]. round-robin [Mal22, Mal23]. Routing [BEFK04, Bou96, CP95, CHL18, HL92, Sch95, SK11, ST94, Tur98, ADGZ09, Asl22, BM02a, Bhu09, Bro00, Bru16, Cag17, LKH14, Lin04, PKK13, RB12, Sto05, Wan17a]. Rowwise [WWS18]. Ruin [FL22, NYZ04, YZ03, CMZ17, CY06, CL23, Lu13, Tan06, Xu23]. Rule [BLP93, CW90b, Fil98, Kha88, LT00, Cho99, Haj15, Hil99, Pin07, VM12]. Rules [Ang89, DY90b, Ros90b, SP98, CL06, vdL11, vdWY03]. Run [CW90a, ARK03, Ang01, Ros99a]. running [MT16]. Runs [Van96b]. runways [BEFK04].

s [FH94, LK08]. SABR [PZ21]. Safe [McN88]. safety [VM12]. sale [BK06]. sales [PLS23]. Salesman [Lal90]. Sample [CBM18, KMT18, LOdBP87, SY92b, SA18, BCZZ19, LHL08, MH10, WWH18, ZL09]. Sample-mean-related [KMT18]. Sampler [Din95]. Samples [DZZ18, DZB12a, DZB12b, GDRV10, HZ06, TE15, Yu21, ZZZ21]. Sampling [DLR87, DY90b, Fil98, GS93, Gut97, HM94, JR92, Kol94, Ros98, BdBS20, EQ22, GK21, GR12, LCYY22]. Saunders [ABJ22, Gup15]. Savage [Mei14]. scalability [KSvH14]. scalable [Voi09]. Scale [Whi88, ASZ17, BKB22, BA18, Bro13, CS03, DKC21, DKB22, HSB23]. Scaled [Wan18a]. Scaling [VHBSW22]. scan [BK03a]. Scattering [Kol94, HPS08]. schedule [DDK22]. Schedules [KT97, Wee90]. Scheduling [ALR00, AKR^{+}04, AZR08, APR20, BBCP97, BCD95, CHRW94, CHW89, ER95, FG89a, Fre91a, Fre91b, Fro91, Gla91, HW92, IwW97, LL90, LL91, LR97, MCB18, PT93a, PT94, PR88, PSC92, RW95, RS89, Rig96, ST94, VT92, ZCC23, AAR11, CSZ03, CZ06, Cao17, KR01, Pin07]. Scheme [Ald93, BB98, MD06, Tar17]. Schemes [BC94, FG89b, HS92, EQ22, Mah23, YK00]. Schur [AENP89, CS94]. Schur-Constant [CS94]. Sciences [Bax96, MY11]. scientific [OST06]. scores [Mal22, Mal23]. screens [GN15]. Search [AFP88, Ber96, GA90, HS92, MJ95, RS94b, ZBA05, ALM07, BHP05, KM16, Nic16]. search-cost [BHP05]. seasonal [Ros22]. Second [Bax96, Ott93b, Tak90, LMH14, LMH12, Zar23]. Second-Order [Tak90, LMH14, LMH12, Zar23]. secretary [DvdLR11]. see [ABS12]. Seed [CH87]. select [GW12]. Selection [CLC98, LHY19, Olu99, SK20, WNT23]. selective [DA03]. Self [Lal90, HM19]. self-decomposability [HM19]. Self-Similar [Lal90]. Selling [AL91]. Semi [BGG07, KW89, PR97, Sen89, BM11, BKP01, ÇÖ15, GKPR99]. Semi-Markov [BGG07, KW89, PR97, Sen89, BM11, BKP01, ÇÖ15, GKPR99]. Semicrystalline [Tay87]. semidefinite [GM16]. Semimartingales [BS93]. sensing [Gut14, PS10]. sensitive [Asl22]. sensitivities [GZ09]. Sensitivity [Fis91, FH94, FH95, LST^{+}22, Miy00]. sensor [Abd17]. Separation [Kij94]. SEPT [CHRW94]. Sequences [Smi97b, CN00, FLL12]. Sequencing [BK93, For93, PT96, Rig92, Yec88]. Sequential [BN18, DY90a, Dav95, DY90b, LOdBP87, MR88, Rig87, BJ13, BN11, FH13, Rig11, ZH07]. Serial [Sob95]. Series [Bla98, CBM18, DG91b, GSW93, HHvU07, ABJ22, CY06, DLB12, DDL14, EK09, FH17b, Fro99, HLM16, JLL21, LL03, LWL19, Mah13b, Mah14, Mah19, MMD11, YBL22b, ZDZ20]. series-parallel [HLM16, LWL19, Mah13b, Mah14, Mah19]. series/parallel/recurrent [YBL22b]. Served [BvHvO97, CGM21]. Server [AW95, Bor95, BWY93, Bri88, DPV16, FK96, HAA22, MS91, RS89, RSZ93, Sch95, WB96, AW12, AY08a, AGM01, APR20, BPLD21, BM02a, BDM15, Bro00, Cao17, ET87, FM17, Jia19, KT01, KN00, LZ19, ÖK14, PW12a, PW12b, SKA09, SI04, Sto05, Yam99, Yam00, ZCA18]. Servers [AM87, DG91b, AHW10, AW12, sADL02, AAD07, AAD08, FS99, Haj15, Pan08, ZCC23]. Service [BCGW92, BvHvO97, CGKS96, FL98, LT00, MS91, Miy92, NJ95, SO94, WAI95, Win97, AKR^{+}04, Ang01, APR20, BPLD21, BB06, BMK08, CJ96, FZ15, FZ19, HMW09, HLW16, LKH14, Lu21, MN05, MAS06, NS17, PW12a, PW12b, SHM09, VA05, WÇC22, WAvHD09, XMBG02, XLSZ04, vdMR08, vdWY03]. services [Tij08]. Sessions [HBvdMB18]. Set [CW22, SH19, Tem94, Hwa00]. Set-Valued [CW22, SH19]. Sets [Men94a, GW12, GDRV10]. setup [CXY12, FL20b, Yam99, Yam00, vdHvHE01]. setup-cost [FL20b]. setups [Ols01]. Setwise [CSK89]. Several [FY11, Jun88, AHOS23, Cho99, Zha21a]. severity [OKA22]. Shaked [BCZZ19]. Shannon [SZF18]. shape [BSL15]. Shaped [BLS12, AHOS23, NH04, NBR13]. Shapes [DG91a]. Shaping [CBM18]. Shared [AMK18, KW94, Khe08]. Sharing [HvdW97, Mit18, EZB06, EM19, FH17b, FM17, LW13, Lu10, Lu21, SLZ04, TP21, Yam00]. Sharp [BK09, KM19, Mei91]. Sharpe [Chi21, ZZ13]. shelf [FW15]. Shift [CS03]. shipments [BvHTvO18]. Shiu [XYZC23]. Shock [Keb94b, CH11, EK21, GFH22, MIK23]. shocks [CF12]. Shop [LL91]. Shops [LK96]. Short [Ott93a, PZ19, Van96b, QWGV23, YSM06]. Short-Range [Ott93a]. short-term [QWGV23, YSM06]. Shortest [Dod89, HK90, HK92b, VT08, Dim23, Rav21]. Shortfall [SP98]. Shortfall-Based [SP98]. Shot [Miy04, Miy05, Lin19]. showcase [MI10]. showdown [MI10]. Showing [Whi89]. shuffle [vZS04]. Shuttle [VT92]. Shwartz [Man96a]. Sibuya [HM19]. sided [BK09]. Siegmund [Lor18]. signalized [BAWD12, HvdW08]. Signals [Cha94, CP95, Gel93, HT96, Kwo14, MK17, Mat17, MK19]. Signature [ZAE19, TK08, YBL23b]. Signature-based [ZAE19]. Signatures [YBL22b, IASB23, YBL22a, YBL23a]. Similar [Lal90]. Simple [BPvdW91, HT96, Lin95, Reg90, Ros88b, TV91, Van01, vDTW88, vD95, HPS08, Kyr03, Kyr04, Lin19, PH23, XZ11]. Simpler [SA88]. Simulated [Ald87, CBM18, CK88, GV94, KT97, RS94b, VBAK89]. Simulating [AW96, But89, FTC98, KH91, Ros88c]. Simulation [Ald87, AFP88, BB98, BK95, CN00, CE96, FH95, GI95, GW96, GSW93, HS88, JR92, LR91, PR97, Ros89, Ros90a, Sch98, BM02b, CMZ17, Cal10, CHL03, CS03, DM03, Gra08, Hei01, Lag06, Lin00, NS04, RL99, Ros23, WW05, WGI99]. Simulation-Based [BB98, CHL03]. Simulations [ACG95, RL01, Ros06b]. Simultaneous [Jun88, Rei90]. Single [AW95, AL91, Bri88, BK93, CHKL13, Fro91, Gla91, LT00, MS91, PKK13, PR88, RS89, ZCA18, AGM01, APR20, KK01, KN00, SI04, TZS22, YT23]. Single-Hop [LT00]. Single-Machine [Gla91, CHKL13]. single-outlier [KK01]. Single-Server [AW95, Bri88, ZCA18, AGM01, APR20, KN00, SI04]. singletons [KM12]. Singpurwalla [Roy90]. singular [Xu23]. singularity [WFM09]. singularly [YZL00]. SINR [GS98]. SIR [DGG22]. SIS [DXX16, VM04]. Size [BHP05, CH87, LOdBP87, VT08, Bro06, BHCKS04, EK21, HBV10, IBI08, MC11, OKA22, Ros06a]. Size- [CH87]. Size-biased [BHP05]. size-biasing [Bro06]. size-interval [HBV10]. Sizes [Sob95, CY21, GS23, YBL22a, YBL23a]. Skellam [GM14]. skew [PH23, ZY22]. skew-elliptical [ZY22]. skewed [Nad08]. skewness [AKS11]. skill [AW12, Haj15, RB12]. skill-based [AW12, Haj15, RB12]. Skip [AW89, Ken94, BPR10]. Skip-Free [AW89, Ken94, BPR10]. Slotted [YM92]. slow [CK20]. Slowly [DR97]. Small [Ben90, FWH21, PT20, SD22b]. Small-world [FWH21]. smallest [DKB22, Pin07]. Sobel [OST06]. socially [Ros15]. Software [BPT87, DDK22, DK88, Gut97, LM87, Nay89, Wee90, LC07, ÖAA01, SD22a]. Sojourn [AB93, EZB06, MN05, Khe08, LW13, Lu09, Lu10, MC11]. Solitaire [vZS04]. Solution [Ban92, Fan96, Kyr04, Tar17]. Solutions [GV90, KS92, CT23, CGM21, HY21, KSS15]. Solved [ER95]. solvency [BJ23]. Some [AAC14, Ang89, AS00b, BZ13b, BCMS23, BSS90, BPR10, CR01, CHYZ93, DK13, Fro11, FS20, Fro95, Gut14, KP12, Kha13, Kha07, KG92, KG87, Koc96, KX12, LX06, LZ11, LY04, Mah13b, Mah14, MMS93, MFN17, MC91, PS93, PK06, RMS94, RW92a, RSZ93, SS87, SS89, WJ18, XZ11, Xu13, ZHZ15, ZYSZ21, AR14, AHOS23, GKB13, IASB23, Mei91, NBR13, YXZ23, YHZ06]. Sometimes [DG91a]. Sorting [GS93]. Source [Yan98, SNS19]. Sources [CW95, KL98, Man96b, BM02b, VM04]. Space [MW94, Tak90, BB02]. Spaces [DF90, DO16, Lor18]. Spacings [DZZ18, AEAN23, Ali23, BCMS23, FRR02, HZ05, HZ06, KK00, KK01, LHL08, XH09, XZ11]. spare [BvHTvO18]. Sparse [Ald93, AR14]. Spatial [Miy04, Miy05]. Spatiotemporal [SY92a]. special [Cag17, CGM21, Pek21, Yan20]. Spectral [AW89, And95, Bal92, BS87, Bro91, DM03, Fra99]. spectrum [GSV16, Mah19]. Speed [ZL02, KAA12]. spend [BPR21, GPM22]. Spent [Lef01]. Spherical [SSS02, SSS04, SSS08]. spherically [WWZ15]. spiking [HBG21]. splitting [GDRV10]. sporadic [Sta07]. spot [KTÇ14]. Springer [Bax96, Joe12]. Springer-Verlag [Bax96]. Square [BF91, Lin94b]. Squared [BC18, Chi21]. squares [Tim16]. Stability [CW90a, FL98, GS92, PZ04, Sch95, DHR00, DP21, Dim23, GN14, THM04]. stabilize [LWZ16]. Stabilizing [LW14, PM17]. stable [DP21]. Staffing [HLW16, LWZ16, DK22]. Stage [BLP93, HW92, Yec88, sADL02, SLZ04, ZCC23]. staircase [LRW99]. Standby [VR89]. Star [AHOS23]. Star-shaped [AHOS23]. start [DO03]. start-up [DO03]. State [DF90, GHS91, GI88, HKL94, PP02, RSZ93, Sch95, Tak90, TV91, Ano21a, APR20, BR17, BPS06, BCD^{+}14, BKKP05, BdBS20, Cha99, DO16, HHvU07, KS02, Lor18, Mat17, MK21, MS06, NS17, POdCW21, Rat19, Spi15, WAvHD09, Yam00, YBL22a, YBL22b, YBL23a, YBL23b, ZAE19]. State-Dependent [RSZ93, Sch95, APR20, BKKP05, KS02, NS17, WAvHD09, Yam00]. state-independent [BdBS20]. states [YZL00]. Static [CR15, Tay87, LKH14]. Statically [DKK94]. Station [LL90]. Stationarity [Fil91]. Stationary [AFP88, BS93, CCdO01, CW95, DF90, FHJ94, HJ95, Miy04, Miy05, PT93b, CCC05, DXX16, Dim23, FL22, HLW16, HvdL03, HP00, JLL21, KL15, MP02, Mat17, MK19, Miy00, Ros06a]. Stations [WAI95, Win97]. statistical [LK23, LZY23]. statistically [LYF16]. Statistics [BZ13b, BR95, BN18, Kha13, LY13, MMS93, SSS04, SSS08, Xu13, AEAN23, ASZ17, BBHK10, BMR05a, BGR11, BNT06, BK03a, Bro09, BN11, CLSS13, FRR02, HN19, HSB23, HZW01, HZ05, HZ06, HJK07, Koc23, LMH14, MH10, Pap21, TA11, WN15, WWH18, XH08, XH09, XZ11, XH11, Yu21, ZZZ21, ZB12, ZHZ15, ZH07]. Status [ZCA18]. Steady [GHS91, GI88, Cha99]. Steady-State [GHS91, Cha99]. Stein [Ros23]. step [Lin19]. step-stress [Lin19]. Stieltjes [BPvdW91]. Stochastic [ARK03, AF92, AM87, ASZ17, ABJ22, AS96, AÖ12, AF97, BBW91, BJ13, BF90, BCR05, Bax96, BLRS01, BMR05a, BOR06, BB97, BB98, BB00, BBCP97, Buc97, BHCKS04, BN18, CSZ03, CZ06, CF12, CHRW94, CPS89, CF06, CY21, CH87, CZB89, CFM^{+}88, CHW89, DJ07, DvdH97, Dod89, FG89a, FZ93, FL20b, FS89, For93, FRR02, Fre91a, Fro91, Gla91, GV90, HS88, HS91, HZ05, HZ06, HW92, KK01, KX07, KX08, KL98, Kor03, LL91, LK96, Li94, LL03, Lin97, LSSZ16, MD06, MW94, McN88, MS93, NDZ22, NBR97, PT93a, PT94, PS93, PR88, PSC92, Pin07, Rig87, Rig96, Rig11, RB06, SW95, SY92a, SAF22, VV88, Wan18b, YZ89, ZL09, ZB12, ZH09, AR00, AL15, BCMS23, BS06, BOT09, BA02, BP01, Bha99, Bor00, BdlPS15, CSL09, CR15]. stochastic [CLS21, CV22, DFS15, DLU99, Eco03, FH13, FH17b, GM16, GT10, GRSV14, Han19, Idr23, JZY22, KTÇ14, Kay09, KPS03, KAR09, KKL17, KKJ22, KR01, Mah13b, MP02, MIK23, MGM21, PKK13, PG22, Sen01, TP21, TA11, TvdW06, WÇC22, WJYL22, YWZH21, YZZ23, YWC23, Yu21, ZCJL23, ZK06, ZH07, Sch96]. Stochastically [RS89, Voi09, Xu91, Lin00]. Stock [SD97, WSX18, WAI95, BA02, GTDKÖ09, MAS06]. stock-dependent [BA02]. stopped [BC13]. Stopping [BC98, Ben90, Pre94, RS94a, Sai98, GN14, Kay09, Yan20]. Storage [CFM^{+}88, CPR91, BKKP05, EBTD22, FL05, PS02, Sta07]. Storage-Limited [CPR91]. stories [BH21]. Strategic [DKM16, Kim22, ZAMB21]. Strategies [HvdW97, Vil18, ABS12, GZDW21, HKK10, Idr23, SD22a, ZBA05]. Strategy [Dov87, FS23, HLZ23, KLT12, Nic16, PC22, SH00]. stratified [EQ22]. Stream [Ott93a]. streaming [SKA09, XC10]. Streams [Bri88]. Strengthening [Ros95]. stress [Lin19]. Strict [CHL18, PT93a]. String [Sad96]. Strong [DF90, WWZ15, LY04, TYZ19]. Strongly [GHS91, OS90]. Structure [AENP89, BL89, Joh98, KMS15, BB02, CR19, JC18, KM16, YWC23, YBL22b, ZK06]. Structure-reversibility [KMS15]. Structured [BC94]. Structures [Koh88, BFRR01]. Student [Lin92a]. studies [Gup07, Gup15]. Study [ES91, Fil98, GSW93, HM15, Fra13, KMS20]. studying [EK09]. style [FWGD00]. sub [SH19]. sub-additive [SH19]. Subexponential [MOSU10, FL22, Tan04, Tan06, XLSZ04]. Subgroup [DS87]. Subject [BL87a, BL88, Fro91, IwW97, KWA96, PR88, Xu91, Yao89]. subjected [VM12]. Suboptimal [HK96]. subset [CR07]. subsets [Sie07]. subsystems [LZY23]. success [CGK99]. Successful [RS89]. successive [KS12]. sufficient [Eco03, MFN17]. sum [NS22, ZCJL23]. Summation [RW92b]. Sums [Bro89, NZP18, OS90, FPOP08, Fro11, MPH13, SGF20, WN09]. Superimposed [YLB96, LDVT22, Rao15]. superposition [BM02b]. Superpositioned [JLL21]. superpositions [BKL22]. Supplementary [AR00]. Supply [FZ93, AW05, DPV16, LXH04, SLZ04, Wan21]. sure [CK20]. surface [PZ21]. survey [Bur13, Sha13]. Survival [CS94, AAES10, Bha99, CF12, KM19, SK20, YBL23a, Zog23]. susceptibles [OE10]. sustainability [LnG22]. swap [CX23, CL06]. swaps [WJYL22]. Switch [ALR00, BvH93, Jaf92]. Switched [Whi88, vdHvHE01]. Switching [AK96, Joh98, BKKP05, FZ15, GB22, LXYL21, LP23, WNT23]. symmetric [Bro01, Nad08, WWZ15, ZH09]. symmetrizing [WN10]. synchronization [EK09]. Synchronized [Miy92]. System [BL87a, BL88, Bor95, BWY93, BS91, But89, CE96, CCP90, CCH99, CHC00, CSP12, CJ96, CLC98, DR95b, DP95, DM08, Ege94, Elc94, Fis91, HM87, HH96, JR92, KM88, Kor03, LL98, LZ00, MD06, Ott93a, Per97, RS02, VR89, ZCA18, Zie96, vdH87, AHW10, AW12, sADL02, AY08a, AKR^{+}04, AZR08, ADGZ09, BGDR03, BKP01, CXY12, Dim17, Dim23, Eco99, FS99, Fro99, GTDKÖ09, GSV16, GZY22, Gut03, Has07, HLW16, HP03a, HM15, Jia19, LXH04, LRW09, LZY23, MAS06, Men04, Mi99, Mi21, MC11, NS17, OKA22, ÖK14, PLS23, Rav21, Rig02, RKFI21, Spi91, Sta07, SNS19, VM12, XC09, Yam99, Yam00, YS08, YWL22, YBL23b, ZBA05, vdHvHE01]. system-point [PLS23]. Systems [AMK18, AENS94, BI88, BFR99, Bla98, BEP88, BPT90, BN18, CH88, CHL94, CF06, Che97, CHJ95, DZH12, DJ07, DvdH97, DH94, ES91, FZ93, FL98, FH94, GS98, HH00, HY90, KWA96, KX08, KX10, KL98, KSM88, LM87, LK08, Man96b, MW94, Men94a, Men94b, MvD97, PT96, PP89, RS89, Rig96, Sen89, TV89, VV88, VdK87, Van97, VY08, Wan18a, Yao89, vDTW88, AL09, AJK01, ASKZ21, ABJ22, Ang01, Ano21a, AÖ12, AILR08, BLPS05, BFRR01, BMR05b, BW14, BVA13, BKKP05, dCBB22, CM05, CV22, DLB12, DDL14, DPV16, FL20a, FL05, FH17b, GHvdL07, GS12, HMW09, HM21, HAA22, HKA22, IASB23, Khe08, KX07, KM07, LL03, LYF16, LSW21, Lin02, LWL19, LAWZ01, Men02, MMD11, NPS04, NS04, NDZ22, NS21, NFMFSA20, PR08, POdCW21, PY16, PLS23, QWW19, Ros06b]. systems [RB10, SH23, SZ87, SY15, SHM09, Sto05, TP21, TK08, WFM09, WÇC22, YWL22, YBL22a, YBL22b, YBL23a, YBL23b, YHZ06, ZAMB21, ZDZ20, ZK06, Zha11, vdMR08].

t [BH23]. Tail [LT96, Las98, Lin90a, Lin92a, Lin92b, Ros98, ZY22, Dim23, FS20, Kla00, Miy02, PT11, STI18, SGF20, Tan04, SC23]. Tailed [FGR10, BM02b, Lu13, MN05]. tails [EZB06, Tan04]. Tandem [AdHBvO14, BCG89, CPS89, HK92b, HK96, HBvdMB18, PMB96, WAI95, Win97, sADL02, AAD07, AAD08, BdBS20, DMvU07, Koo04, ML01, Nak01, Pan08, SZ02, WS06, Zaz04]. Tardiness [For93, PT94, CZ06]. Target [MJ95]. Targets [Sat97, HMW09]. Tariffing [BT95]. task [HBV10, Wan17a]. Tasks [BCD95, BK93, AY08a]. Taylor [BC18]. TCP [BHCKS04, OK08]. TDM [CCP90]. TDMA [RS90]. Team [Ros22]. Technique [AB93, Reg90, RL93, VR89, AR00]. techniques [RB10, RB12]. Temperature [KT97]. Temporal [CBM18, SX22]. tenable [Mah13a]. term [FKL02, QWGV23, YSM06]. Terminal [Dov87]. terms [CBG15]. ternary [BMR05b, GSV16]. Test [CHH90, DG94, Gut97, Chi21, HWZ12, MG21]. Testing [BLSV03, BR95, CHH90, JL98, RMS94, SKL12, Yao88, YH88, BLBPV17, CHS87, ÖAA01, YDJ23, ZP16]. Tests [Abr93, And95, Rot93, Lin19]. th [Bar08, LHY19]. Their [BBS88, KH91, AJK01, AS96, AY08b, Cag17, IASB23, KSvH14, KSS15, YHZ06, Zog23, dCHQ09]. Theorem [CZL96, RS95, SAS89, Ste87, Van92a, DMM16, Hua20, SZF18, TYZ19]. Theorems [HS91, LY04, Ols01, SD22b, ZYSZ21]. theoretic [LnG22]. Theoretical [PO15]. Theory [AW89, BPvdW91, CZB89, DF90, Joe12, LE94, MS93, Ros87b, SSS02, Bri14, CVB05, KA16, Pek21, vdWSvN08]. thin [KN21]. Thinning [HW90]. Thomson [Cal97]. threatening [Vil18]. Three [HS92, Van88, Che09, mKK22, NL07]. three-factor [mKK22]. Three-Moment [Van88]. three-person [Che09]. three-piece [NL07]. Threshold [AGM01, BJ13, BCD^{+}14, CH11, FL05]. Throughput [AM87, ALR00, HBvdMB18, AAD08, DP21]. Tight [BNT06]. Tijms [Sch96]. Time [AR97, BBS96, Bha87, Bla98, BS87, CCP90, CG96, CBM18, CJ96, CL91, CK88, DR97, DP93, ER95, ET87, Eli08, Fil91, GI88, Hes92, HJ95, HW92, KR87, Kha88, Kij92, Lef01, LN92, ÒLC95, PR97, PT93b, RS95, RS90, Ros87a, Ros88a, RS02, SS92, SB13, Tam98, Van96a, YS89, ZL02, dK89, ABD23, BC13, Bar08, BOR06, Bor00, Bri14, BH23, Bro06, Bro99, Cai02, CSZ03, CY06, CCC05, Che20, CSvD02, EZB06, Eli05, Fra13, Fra19, Fra99, FNC20, GT07, HJM09, HY21, HL14, HKK10, Hor01, HvdL03, HWZ12, JL19, JLL21, KA04, Khe08, KW13, LL03, LZ11, LWZ16, LW19, LRW09, LW14, LZ19, LSSZ16, MZ17, MW94, MC11, Pek99a, PZ21, PLS23, Ros99b, RB10, RKFI21, SZ99, SAdK04, SBZ16, SNS19, VHBSW22, Wan17b, Zaz04, Zaz07, vD88, vdMR08]. Time-Dependent [Kha88, Che20, Fra13, Fra19, FNC20, LRW09]. time-discretized [PZ21]. Time-Inhomogeneous [CK88, LW19]. time-shared [Khe08]. time-to-empty [Zaz04]. time-varying [HJM09, KW13, LWZ16, LW14, LZ19, LSSZ16, RB10, VHBSW22]. Times [AB93, Cha94, CGKS96, Dav95, DvdH97, FTC98, Fre91a, FH94, Koc96, NQ01, RS94a, Ros87a, Ros88a, SS87, VY08, AAKN09, AS18, BOT09, Bha99, BKP01, DLB12, DPV16, Fro99, GZY22, Khe08, KAR09, KW13, LW13, Lu10, MN05, MK23, Nak01, PW12a, PW12b, Ros15, STI18, SZ03, SZ05, XLSZ04, XYZC23, vdHvHE01]. times-corrigendum [PW12b]. Timescale [BB98]. Tolerable [Sad98]. tolerant [LM87]. Total [CE96, Xu91, CZ03, HWZ12]. Tournament [CHYZ93, Mal22, Mal23]. tournaments [CR19]. Tp [LHL08]. tradable [SXWW22]. Traffic [BM11, CPR91, AY08b, BAWD12, BW14, FS99, HvdW08, IVA15, KN00, LKH14, Lu09, PB06, Wan17a, vdHvHE01, vdMR08]. training [LK07]. transfer [EA06]. Transform [Keb94a, HWZ12, Tag00, dI06]. transformation [Bro06]. Transient [AW96, CZ03, CH87, KN00, Kyr04, OK08, WGI99, YS89, ZCJ04, Mat19a, WFM09, Wan02, YZL00]. Transition [Ros87a, vD88, JZY22]. Transitions [PV11, EK09]. Transmission [RB05a, RB05b, FM16]. transplantation [BDPS11]. trap [RB06]. trapezoids [Reu14]. Travelling [Lal90]. treaties [WWZZ23]. treatments [BMSB20]. tree [Hua20, HBKK03, QWGV23, SZF18, SD22b]. tree-based [QWGV23]. treelike [FH17a]. Trees [AFP88, GMW14, Mah91, MO95, SAS89, Ste87, ALM07, AAKN09, KM16, Kon02, Kon03, KAJ07, KAA12, KP07, LY04, SZ87, WWZ15, ZD19, Zha21a]. trees-plane-oriented [Zha21a]. triage [AZR08]. Trials [RTW09, Vil18]. triangles [Röl22]. tries [CM08]. triggered [CF12]. Triggering [Cha94]. Truncated [KR87, SK23]. Truncation [NBR97, CM00]. Trunk [FR94, FY11, Key90, Ngu91, FOF07]. tunable [DK09]. Two [AvER95, AFP88, AK96, Bal92, BLP93, BP01, BBCP97, Bro89, BK93, Che11, CHL18, CHW89, DP93, Dud97, DY90b, Hes92, HW92, LK96, MT93, ML01, NJ95, PP89, PP02, SS98, TSW97, VR89, ZAMB21, AW05, sADL02, BEFK04, BPS06, BB98, BSL15, BVA13, BvHTvO18, BK03a, BK09, Che12, CBBG20, DKB22, DKM16, DK88, Dim17, DZB12a, DZB12b, Fil98, HZ06, Idr23, JST01, KMS15, Men02, MZS16, ÖK14, Pan08, PY16, STI18, SLZ04, TYZ19, TvdW06, ZCC23, ZCJL23, vdHvHE01, vdMR08]. two-carousel [BVA13]. Two-class [CHL18, Dim17]. Two-Dimensional [DP93, Hes92, BK03a, KMS15, STI18]. Two-echelon [ZAMB21, BvHTvO18]. Two-Machine [LK96]. two-node [AW05, CBBG20]. Two-person [Che11, Che12, TvdW06]. two-phase [vdMR08]. Two-Point [BBCP97, CHW89]. two-queue [vdHvHE01]. two-server [ÖK14]. two-sided [BK09]. Two-Stage [BLP93, HW92, sADL02, SLZ04, ZCC23]. Two-State [PP02, BPS06]. Two-Unit [VR89]. Type [Ban92, BCS92, PV11, SC23, Zie96, BHK03, BP01, MG21, RBM02, Zog23]. type-II [MG21]. Types [BK93, FY11, Idr23, MZS16].

Uhlenbeck [KR94, YWZH21]. UMCP [HS13]. unbalance [HBV10]. unbalanced [Idr23]. Unbiased [Nay89]. uncertainties [EA06, OE10]. Uncertainty [LL98, BHM09, Has07, Olu11]. unfriendly [Abd16]. unified [KX12, PKK13]. Uniform [Ald87, BPG12, DS87, SSS04, SSS08, Tan04, Bel11, CLSS13, GDRV10]. Uniformization [Kij92]. Uniformly [Kol92, AS00a, AS00b, SD22b]. Unit [VR89, CT02]. Univariate [BBHK10]. Universal [Sad96]. unknown [WC00]. unreliability [NS04]. Unreliable [PP02, SK11, vD95, Fou16, SKA09]. unresolved [HL01, HL04]. until [CL23]. unvisited [Ald22]. upgrades [DL10]. Upper [CY06, SZ03, Nak01, SZ05]. Upsidedown [BLS12]. urgency [BLPS05]. Uri [AEL08]. Urn [AFP88, FGR10, CH15, Che20, CH12, Idr23, Mah23, Nad17]. urns [Mah13a]. Usage [Gut97, Sat97]. Use [MD06, But89, YBL23a]. Used [MD06, JL98, YBL23b]. Using [BA18, BK95, Cal10, CSP12, FTC98, GSW93, Ros98, Ros23, SS87, WFM09, AS18, GDRV10, HY21, LC07, LCYY22, Pek99a, SK20, YWC23, dI06]. Utility [Tar11].

Vacation [NJ95, BKM10, TZS22, YT23]. vaccination [DGG22]. validation [BHCKS04]. Valuation [Han19, WSX18, DO03, GKL23, NPS04, Wan17b, WXL20, XXN21, YSM06]. Value [Her95, MCB18, Chi23, DMM16, SLZ04, TE15, WNT23, ZDS20]. value-at-risk [Chi23]. Valued [CW22, KG87, SH19, Wak00]. Values [BW95, GK88, AAC14, BOR06, ZDS20]. Valuing [CW23]. Varentropy [SK23]. Variability [BS98, OE10, SS98, SC23, BHK03, ZDZ20]. Variable [Her95, SK20, AK08, AR00, HCX06, SK23, WXWH18]. Variables [BS98, DS87, KG87, Mei91, NZP18, OS90, WWS18, WJ18, AKS11, BBHK10, BMR05b, CMH13, Fro11, HN19, HZW01, HYL07, KA16, KX12, Lin00, MPH13, NK05, Pap21, PR04, WN10, XH11, ZL09]. Variance [AW96, Bro01, Bro14, GKL87, LR91, Nay89, Ros88c, Ros90a, NL07, PC22, Pin07, RL01, Ros23, WJYL22, Zaz07, ZCJL23]. variances [MC11]. Variation [FR92, BC18, LMH14, LMH12]. variational [YXH19]. Variations [Sha88, PT20]. various [HL01, MK21]. Varying [AF97, RB05a, AKR^{+}04, HJM09, KW13, LWZ16, LW14, LZ19, LSSZ16, RB05b, RB10, VHBSW22]. Vector [SR90, Wak00]. Vector-valued [Wak00]. vehicle [HM16, KM16, PKK13]. Verlag [Bax96]. Versatile [Ott93b]. version [YBL23b]. Versions [LM87]. Versus [BvHvO97, JR92, YWL22]. vertex [AR14, Ald22]. vertex-names [AR14]. Via [FG91, BN18, DJ07, GW96, GN15, HJM09, Hei01, Men94a, PR97, PB06, Pen15, PP89, Ros90a, Sha88, WGI99]. Video [Hun91]. Video-Data [Hun91]. view [Idr23]. Visit [VY08, vdWY03]. visit-order [vdWY03]. visual [HM16]. Viterbi [LK07]. VIX [YWC23]. vol [Bax96]. Volatility [LXYL21, AL15, CY21, Han19, PG22, PZ21, WJYL22, YWZH21, YWC23, ZCJL23]. volume [Ano87b, Ano87a, Ano87d, Ano87c, Ano87f, Ano87e, Ano87h, Ano87g, Ano88b, Ano88a, Ano88d, Ano88c, Ano88f, Ano88e, Ano88h, Ano88g, Ano89b, Ano89a, Ano89d, Ano89c, Ano89f, Ano89e, Ano89h, Ano89g, Ano90b, Ano90a, Ano90d, Ano90c, Ano90f, Ano90e, Ano90h, Ano90g, Ano91c, Ano91b, Ano91e, Ano91d, Ano91g, Ano91f, Ano91i, Ano91h, Ano92b, Ano92a, Ano92d, Ano92c, Ano92f, Ano92e, Ano92h, Ano92g, Ano93a, Ano93b, Ano93c, Ano93d, Ano94b, Ano94c, Ano94d, Ano94e, Ano95b, Ano95a, Ano95c, Ano95d, Ano95e, Ano96a, Ano96b, Ano96c, Ano96d, Ano97a, Ano97b, Ano97c, Ano97d, Ano98a, Ano98b, Ano98c, Ano98d, Ano09b, Ano09a, Ano09d, Ano09c, Ano09f, Ano09e, Ano09h, Ano09g, Ano10b, Ano10a, Ano10d]. volume [Ano10c, Ano10f, Ano10e, Ano10h, Ano10g, Ano11b, Ano11a, Ano11d, Ano11c, Ano11f, Ano11e, Ano11h, Ano11g, Ano12b, Ano12a, Ano12d, Ano12c, Ano12f, Ano12e, Ano12h, Ano12g, Ano13b, Ano13a, Ano13d, Ano13c, Ano13f, Ano13e, Ano13h, Ano13g, Ano14b, Ano14a, Ano14d, Ano14c, Ano14f, Ano14e, Ano14h, Ano14g, Ano15b, Ano15a, Ano15d, Ano15c, Ano15f, Ano15e, Ano15h, Ano15g, Ano16b, Ano16a, Ano16d, Ano16c, Ano16f, Ano16e, Ano16h, Ano16g, Ano17b, Ano17a, Ano17d, Ano17c, Ano17f, Ano17e, Ano17h, Ano17g, Ano18b, Ano18a, Ano18d, Ano18c, Ano18f, Ano18e, Ano18h, Ano18g, Ano19b, Ano19a, Ano19d, Ano19c, Ano19f, Ano19e, Ano19h, Ano19g, Ano20b, Ano20a, Ano20d, Ano20c, Ano20f, Ano20e, Ano20h]. volume [Ano20g, Ano21c, Ano21b, Ano21e, Ano21d, Ano21g, Ano21f, Ano22b, Ano22a, Ano22c, Ano22e, Ano22d, Ano22g, Ano22f, Ano23a]. vs [Bla95]. Vulnerable [Wan18b, CW23, Han19, Wan17b, WWL17].

W [Joe12]. W.P.1 [YZ89]. wagering [Pen10]. Wait [ABS12]. Wait-and-see [ABS12]. Waiting [ET87, LN92, ÒLC95, PW98, SS92, SB13, SAdK04, dK89, vdHvHE01, Bar08, Bro06, CCC05, DPV16, HvdL03, JD07, KW13, Ros99b, SNS19, vdMR08]. Waiting-Time [ET87, LN92, dK89, CCC05]. Walk [CFFH98, Han03, Sha88, Ald22, BPR10, KMS15]. Walks [Pal90, CBG15, HP03b, HPS08, KAA12, Lu13]. Warm [Gra08]. Warm-up [Gra08]. warranty [BK06, HKS07, LK09]. Wave [RL93]. Waxman [Van01]. way [vdHvHE01]. Weak [Pla92, Hei01, Kwo14]. weaknesses [EFG17]. wealth [Zar23]. Wear [BL87a, BL88, PS93]. weather [ZPD06]. weather-related [ZPD06]. Web [YS08]. Weibull [AO18, AO19, GG96]. Weight [VT08]. Weighted [KG87, XWTW18, ALM07, BJ19, CSZ03, DM08, KP07, LYF16, Lu21, MPH13, Nad16, NBR13, Olu06, Olu11, PO15, SGF20]. Weiss [Man96a, Haj15]. Wesley [Sha98]. Which [GI88]. White [Bre90]. who [Nic16]. Whose [Bal92]. Widely [XWTW18]. Wiener [DP93, KH91]. Wiley [Sch96]. Win [Dub98, Ros22]. window [BHCKS04]. Winner [CR19, CR21, Mal22, Mal23]. Wireless [GS98, MW94]. Wishart [LC97]. Within [LZ00, AO18, AO19, Bha99, CLC98]. without [Bro91, CB20, HPR96, Kol92, ZDS20]. words [HM16]. Work [BB96, JS96, AW05, TS07]. working [TZS22, YT23]. Workload [BB96, BBS97, BB06, DA03]. workload-dependent [BB06]. workloads [LSSZ16]. world [FWH21, TvdW06]. Worst [FR92]. Worst-Case [FR92]. worth [CT02]. written [SXWW22]. Wrongs [LM88].

xiv [Bax96].

Yechiali [AEL08]. Yields [Sob95]. York [Bax96, Man96a].

Zagreb [FHS13]. Zero [Fre91a, LE94, ZCJL23]. Zero-Crossing-Detecting [LE94]. Zhao [Bur13, Kha13, Sha13]. Zipf [KPH10].

References

Alimohammadi:2014:SCP

[AAC14] Mahdi Alimohammadi, Mohammad Hossein Alamatsaz, and Erhard Cramer. Some convexity properties of the distribution of lower k-record values with extensions. Probability in the Engineering and Informational Sciences, 28(3):389–399, July 2014. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

              tronic). URL https://www.cambridge.org/core/product/
              70F365D9F3E24329F07A02DF6D524CDD.

Andradottir:2007:DAD

[AAD07] Sigrún Andradóttir, Hayriye Ayhan, and Douglas G. Down. Dynamic assignment of dedicated and flexible servers in tandem lines. Probability in the Engineering and Informational Sciences, 21(4):497–538, October 2007. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

              tronic). URL https://www.cambridge.org/core/product/
              BFBFFED5606F659B225EDFA1BC37DD19.

Andradottir:2008:MTT

[AAD08] Sigrún Andradóttir, Hayriye Ayhan, and Douglas G. Down. Maximizing the throughput of tandem lines with flexible failure-prone servers and finite buffers. Probability in the Engineering and Informational Sciences, 22(2):191–211, March 2008. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
71DE1512506807ACC1B44C3843C38500.

Asadi:2010:MBP

[AAES10] Majid Asadi, Somayeh Ashrafi, Nader Ebrahimi, and Ehsan S. Soofi. Models based on partial information about survival and hazard gradient. Probability in the Engineering and Informational Sciences, 24(4):561–584, October 2010. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
1E898E27C5266F241F6A13643886B195.

Asgari:2019:IDG

[AAK19] Fatemeh Asgari, Mohammad Hossein Alamatsaz, and Nayereh Bagheri Khoolenjani. Inequalities for the dependent Gaussian noise channels based on Fisher information and copulas. Probability in the Engineering and Informational Sciences, 33(4):618–657, October 2019. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/inequalities-
for-the-dependent-gaussian-noise-channels-based-on-
fisher-information-and-copulas/E7ED527A27F4AFD190BAD12665F82F9B.

Al-Awadhi:2009:CTE

[AAKN09] Fahimah Al-Awadhi, Mokhtar Konsowa, and Zainab Najeh. Commute times and the effective resistances of random trees. Probability in the Engineering and Informational Sciences, 23(4):649–660, October 2009. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
C5B4431CFEBE0BBE14E008F4F4DF29AB.

Aalto:2011:PGI

[AAR11] Samuli Aalto, Urtzi Ayesta, and Rhonda Righter. Properties of the Gittins index with application to optimal scheduling. Probability in the Engineering and Informational Sciences, 25 (3):269–288, July 2011. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/4232008EBC24C1C432F58BF1E69673A4.

Anantharam:1993:TCS

[AB93] V. Anantharam and M. Benchekroun. A technique for computing sojourn times in large networks of interacting queues. Probability in the Engineering and Informational Sciences, 7(4): 441–464, October 1993. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/C5E43A7B240FBD5BFD2DAB5BC15D8D4E.

Abdelrahman:2016:DNU

[Abd16] Omer H. Abdelrahman. Detecting network-unfriendly mobiles with the random neural network. Probability in the Engineering and Informational Sciences, 30(3):514–531, July 2016. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
9C914FECF82C000B68B0EE588C799EFC.

Abdelrahman:2017:MMD

[Abd17] Omer H. Abdelrahman. A Markov-modulated diffusion model for energy harvesting sensor nodes. Probability in the Engineering and Informational Sciences, 31(4):505–515, October 2017. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
6ED1918F04D9D72606E1D7B9DD56FC94.

Aschenbruck:2023:ATB

[ABD23] Nils Aschenbruck, Stephan Bussmann, and Hanna Döring. Asymptotics of a time bounded cylinder model. Probability in the Engineering and Informational Sciences, 37 (4):1063–1083, October 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/asymptotics-
of-a-time-bounded-cylinder-model/E676F92F8E0B6600495374F22D1259A4.

Albrecher:2017:QMR

[ABEK17] Hansjörg Albrecher, Onno Boxma, Rim Essifi, and Richard Kuijstermans. A queueing model with randomized depletion of inventory. Probability in the Engineering and Informational Sciences, 31(1):43–59, January 2017. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
BBF39D63E4D3EB56D697D9A3324BA78B.

Asamov:2022:PMC

[ABI22] Tsvetan Asamov and Adi Ben-Israel. A probabilistic l_{1} method for clustering high-dimensional data. Probability in the Engineering and Informational Sciences, 36(2):433–448, April 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/probabilistic-1-method-for-clustering-highdimensional-
data/35C272B088F89F19C5CB41C5B07BE436.

Amiri:2022:SOL

[ABJ22] Mehdi Amiri, Narayanaswamy Balakrishnan, and Ahad Jamalizadeh. Stochastic ordering of lifetimes of parallel and series systems comprising heterogeneous dependent components with generalized Birnbaum–Saunders distributions. Probability in the Engineering and Informational Sciences, 36(1):49–65, January 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/stochastic-
ordering-of-lifetimes-of-parallel-and-series-systems-
comprising-heterogeneous-dependent-components-with-
generalized-birnbaumsaunders-distributions/E1AEA8AF87883161BB7CF3C76A44FE26.

Abrahams:1993:ILB

[Abr93] Julia Abrahams. An improved lower bound on the minimum expected number of binomial group tests. Probability in the Engineering and Informational Sciences, 7(1):121–124, January 1993. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
09FB31C7E35CE7C5F4212BD3140CF02E.

Aurzada:2012:WSS

[ABS12] Frank Aurzada, Sergej Beck, and Michael Scheutzow. Waitand-see strategies in polling models. Probability in the Engineering and Informational Sciences, 26(1):17–42, January 2012. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
F14E51C4DC5637097B84C92B39D0D4C2.

Andradottir:1995:ARM

[ACG95] Sigrún Andradóttir, James M. Calvin, and Peter W. Glynn. Accelerated regeneration for Markov chain simulations. Probability in the Engineering and Informational Sciences, 9(4):497– 523, October 1995. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/A831F72C47C0069AA8742566741A0508.

Argon:2009:DRC

[ADGZ09] Nilay Tanik Argon, Li Ding, Kevin D. Glazebrook, and Serhan Ziya. Dynamic routing of customers with general delay costs in a multiserver queuing system. Probability in the Engineering and Informational Sciences, 23(2):175–203, April 2009. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
BCD845DFEBB8351DB3DFBCA26CAE1E13.

AlHanbali:2014:DTQ

[AdHBvO14] Ahmad Al Hanbali, Roland de Haan, Richard J. Boucherie, and Jan-Kees van Ommeren. Delay in a tandem queueing model with mobile queues: an analytical approximation. Probability in the Engineering and Informational Sciences, 28(3): 363–387, July 2014. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/F625A6E001EA007097001A2BFD6EEE29.

Avis:1988:PAG

[ADS88] David Avis, Burgess Davis, and J. Michael Steele. Probabilistic analysis of a greedy heuristic for Euclidean matching. Probability in the Engineering and Informational Sciences, 2(2): 143–156, April 1988. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/BF88214FD88F3A8CDADDD8242AD391F0.

Alimohammadi:2023:LRC

[AEAN23] Mahdi Alimohammadi, Maryam Esna-Ashari, and Jorge Navarro. Likelihood ratio comparisons and logconvexity properties of p-spacings from generalized order statistics. Probability in the Engineering and Informational Sciences, 37(1):86–105, January 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-

engineering-and-informational-sciences/article/likelihood-
ratio-comparisons-and-logconvexity-properties-of-pspacings-
from-generalized-order-statistics/804F59E592403609AD9162E87043535B.

Altman:2008:EUY

[AEL08] Eitan Altman, Iddo Eliazar, and Hanoch Levy. Editorial: Uri Yechiali. Probability in the Engineering and Informational Sciences, 22(4):475–476, October 2008. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
CD8FAD7E1A5A6D222E5A51D72A5E8422.

Abouammah:1989:SSF

[AENP89] A. M. Abouammah, E. El-Neweihi, and F. Proschan. Schur structure functions. Probability in the Engineering and Informational Sciences, 3(4):581–591, October 1989. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
CDC65012DC983E05BB39885D1B867D0C.

Abouammoh:1994:RGM

[AENS94] A. M. Abouammoh, Emad El-Neweihi, and Jayaram Sethuraman. The role of a group of modules in the failure of systems. Probability in the Engineering and Informational Sciences, 8(1):89–101, January 1994. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
1F7D349709212B741D1C810FB3CAA7E9.

Alexopoulos:1992:CES

[AF92] Christos Alexopoulos and George S. Fishman. Capacity expansion in stochastic flow networks. Probability in the Engineering and Informational Sciences, 6(1):99–118, January 1992. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
39C3E69894A17AC3C11299814CFE0A5F.

Aviv:1997:SIM

[AF97] Yossi Aviv and Awi Federgruen. Stochastic inventory models with limited production capacity and periodically varying parameters. Probability in the Engineering and Informational Sciences, 11(1):107–135, January 1997. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
04C7DBDEB33A13AE8BCEEF2637535471.

Aldous:1988:TAU

[AFP88] David Aldous, Barry Flannery, and José Luis Palacios. Two applications of urn processes the fringe analysis of search trees and the simulation of quasi-stationary distributions of Markov chains. Probability in the Engineering and Informational Sciences, 2(3):293–307, July 1988. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
CCAC1BC0DA1AAD0C8C6EA85854FD1E72.

Artalejo:2007:NBP

[AGC07] Jesus R. Artalejo and A. Gómez-Corral. A note on the busy period of the M/G/1 queue with finite retrial group. Probability in the Engineering and Informational Sciences, 21(1):77–82, January 2007. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/7FEA03824DC63D0145627B71DAFE6AEB.

Ansell:2001:TPS

[AGM01] P. S. Ansell, K. D. Glazebrook, and I. Mitrani. Threshold policies for a single-server queuing network. Probability in the Engineering and Informational Sciences, 15(1):15–33, January 2001. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
A31B24F7670E823597DC784743766946.

Angus:1996:RCR

[AH96] John E. Angus and Xiao Hong. On the rate of convergence of the Ross approximation to the renewal function. Probability in the Engineering and Informational Sciences, 10(2): 207–211, April 1996. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/D54150D63DC60502C36F8E872FCE2C1C.

Arab:2023:SSO

[AHOS23] Idir Arab, Milto Hadjikyriakou, Paulo Eduardo Oliveira, and Beatriz Santos. Star-shaped order for distributions characterized by several parameters and some applications. Probability in the Engineering and Informational Sciences, 37(1):49–59, January 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/starshaped-
order-for-distributions-characterized-by-several-parameters-
and-some-applications/375A3C4AE842BEC4D5506106EF9D788B.

Adan:2010:REL

[AHW10] Ivo Adan, Cor Hurkens, and Gideon Weiss. A reversible Erlang loss system with multitype customers and multitype servers. Probability in the Engineering and Informational Sciences, 24(4):535–548, October 2010. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
54DC3F4E3AC5BE746A895B281CCA57A0.

Avi-Itzhak:2008:QFQ

[AILR08] Benjamin Avi-Itzhak, Hanoch Levy, and David Raz. Quantifying fairness in queuing systems: principles, approaches, and applicability. Probability in the Engineering and Informational Sciences, 22(4):495–517, October 2008. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
23322890654F0EF7EF6CF05E6B863860.

Altman:2001:CQS

[AJK01] Eitan Altman, Tania Jiménez, and Ger Koole. On the comparison of queueing systems with their fluid limits. Probability in the Engineering and Informational Sciences, 15(2): 165–178, April 2001. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/5104762A910568CAF9076855F9964DCF.

Avram:1996:MCD

[AK96]  Florin Avram and Fikri Karaesmen.  A method for

computing double band policies for switching between two diffusions. Probability in the Engineering and Informational Sciences, 10(4):569–590, October 1996. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
4AAE48E28D7106263CB7F22FC96991CF.

Ahmad:2005:CRO

[AK05] I. A. Ahmad and M. Kayid. Characterizations of the RHR and MIT orderings and the DRHR and IMIT classes of life distributions. Probability in the Engineering and Informational Sciences, 19(4):447–461, October 2005. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
7AF3A9B81151AAD45FF2EFA34D2A8AB4.

Adan:2008:IRV

[AK08] Ivo Adan and Vidyadhar Kulkarni. Insurance risk with variable number of policies. Probability in the Engineering and Informational Sciences, 22(2):213–219, March 2008. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
D03DF7DE70F22FF38D37FDC20F758055.

Ahmad:2005:FRI

[AKP05] I. A. Ahmad, M. Kayid, and F. Pellerey. Further results involving the MIT order and the IMIT class. Probability in the Engineering and Informational Sciences, 19(3):377–395, July 2005. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
CB17B5318B586FB6B266393FB7202B91.

Andrews:2004:SQS

[AKR^{+}04] Matthew Andrews, Krishnan Kumaran, Kavita Ramanan, Alexander Stolyar, Rajiv Vijayakumar, and Phil Whiting. Scheduling in a queuing system with asynchronously varying service rates. Probability in the Engineering and Informational Sciences, 18(2):191–217, April 2004. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
02930EA6CCD2F0A9BDC909C76D881643.

Amiri:2011:SDA

[AKS11] Leila Amiri, Baha-Eldin Khaledi, and Francisco J. Samaniego. On skewness and dispersion among convolutions of independent gamma random variables. Probability in the Engineering and Informational Sciences, 25(1):55–69, January 2011. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
E760F3FB50E586830A68323D6ABE1EAC.

Assaf:1991:OAP

[AL91] David Assaf and Benny Levikson. Optimal advertizing policy for selling a single asset. Probability in the Engineering and Informational Sciences, 5(1):89–100, January 1991.

CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
64B12F82D95B78180A58CD55F15BA493.

Ahn:2009:AEM

[AL09] Hyun-Soo Ahn and Mark E. Lewis. Analysis, evaluation, and management of queuing systems. Probability in the Engineering and Informational Sciences, 23(2):173–174, April 2009. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
9678B1254E8D248E934DE05CA3B7A647.

An:2015:ALV

[AL15] Yu An and Chenxu Li. Approximating local volatility functions of stochastic volatility models: a closed-form expansion approach. Probability in the Engineering and Informational Sciences, 29(4):547–563, October 2015. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
C18F49B1CF36E40EA3025E2F9F8F1DF9.

Aldous:1987:MCS

[Ald87] David Aldous. On the Markov chain simulation method for uniform combinatorial distributions and simulated annealing. Probability in the Engineering and Informational Sciences, 1(1):33–46, January 1987. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/AA38B1E35EC5787F11BA9F7A6833FF49.

Aldous:1988:HLP

[Ald88] David Aldous. Hashing with linear probing under nonuniform probabilities. Probability in the Engineering and Informational Sciences, 2(1):1–14, January 1988. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
9D2CACA7CC9B73AAEA021B85287BB2C7.

Aldous:1993:NPA

[Ald93] David Aldous. A note on probabilistic analysis of a sparse matrix factorization scheme and random graphs. Probability in the Engineering and Informational Sciences, 7(4):465–469, October 1993. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/521D2428F932C025B2ACDDA5C103BFB0.

Aldous:2022:NUV

[Ald22] David J. Aldous. The nearest unvisited vertex walk on random graphs. Probability in the Engineering and Informational Sciences, 36(3):851–867, July 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/nearest-
unvisited-vertex-walk-on-random-graphs/9FE0E4C052DB835CB1F7E382059484B8.

Alimohammadi:2023:ROP

[Ali23] Mahdi Alimohammadi. Resolving an open problem on the hazard rate ordering of p-spacings. Probability in the Engineering and Informational Sciences, 37(4):1020–1028, October 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/resolving-an-open-problem-on-the-hazard-rate-
ordering-of-pspacings/515A7E87AB0DC47666822D62E35BF48E.

Aguech:2007:EWP

[ALM07] Rafik Aguech, Nabil Lasmar, and Hosam Mahmoud. Extremal weighted path lengths in random binary search trees. Probability in the Engineering and Informational Sciences, 21(1): 133–141, January 2007. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/B06DC72886E27AE1A8E4097C5574D5F2.

Altman:2000:SIQ

[ALR00] Eitan Altman, Zhen Liu, and Rhonda Righter. Scheduling of an input-queued switch to achieve maximal throughput. Probability in the Engineering and Informational Sciences, 14(3): 327–334, July 2000. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/7F35BCFD6F59E8F7AC61E8453E7F30A7.

Alzaid:1988:LBO

[Alz88] Abdulhamid A. Alzaid. Length-biased orderings with applications. Probability in the Engineering and Informational Sciences, 2(3):329–341, July 1988. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/CA894D885C6E9C8721CBA9131C0AB2EF.

Alkaff:1987:TRM

[AM87] Abdullah Alkaff and Eginhard J. Muth. The throughput rate of multistation production lines with stochastic servers. Probability in the Engineering and Informational Sciences, 1(3): 309–326, July 1987. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/A0028A9F80527E2F4108A687ECB38985.

Abdul-Malak:2018:ORM

[AMK18] David T. Abdul-Malak and Jeffrey P. Kharoufeh. Optimally replacing multiple systems in a shared environment. Probability in the Engineering and Informational Sciences, 32(2):179–206, April 2018. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/optimally-
replacing-multiple-systems-in-a-shared-environment/
FB4D943986ACFBF51143A1E7D55C5F39.

Anderson:1995:GFT

[And95] T. W. Anderson. Goodness-of-fit tests for probability distributions and spectral distributions. Probability in the Engineering and Informational Sciences, 9(1):27–37, January 1995. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
809DE66A7C9CE3177AF897CBBB44DE20.

Arias-Nicolas:2005:MDO

[ANFPLCSL05] J. P. Arias-Nicolás, J. M. Fernández-Ponce, P. Luque-Calvo, and A. Suárez-Llorens. Multivariate dispersion order and the notion of copula applied to the multivariate t-distribution. Probability in the Engineering and Informational Sciences, 19 (3):363–375, July 2005. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/99D678E177729E078231304B6247DDFA.

Angus:1989:SAA

[Ang89] John E. Angus. Some asymptotic analysis of resistant rules for outlier labeling. Probability in the Engineering and Informational Sciences, 3(1):157–164, January 1989. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
A37B6ACD1F892AACE758B87B2B98427C.

Angus:1995:CPA

[Ang95] John E. Angus. A coupling proof of the asymptotic normality of the permutation oscillation. Probability in the Engineering and Informational Sciences, 9(4):615–621, October 1995. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
06881F9F781172E3F8615C7F0C73F9F3.

Angus:1997:IEQ

[Ang97] John E. Angus. The improved estimation of σ in quality control, revisited. Probability in the Engineering and Informational Sciences, 11(1):37–42, January 1997. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
356CC17C57BA70F6D78C3485F8BB734C.

Angus:2001:ACS

[Ang01] John E. Angus. Availability of continuous service and computing long-run MTBF and reliability for Markov systems. Probability in the Engineering and Informational Sciences, 15(3): 369–381, July 2001. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/58C3EB1A7AFC1DDE8D9ACDDAC9686231.

Angus:2004:CRA

[Ang04] John E. Angus. Convergence rates in approximating a compound distribution. Probability in the Engineering and Informational Sciences, 18(2):185–190, April 2004. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
10E30F067EA2F435552FCEECF8D5D297.

Angus:2006:AEI

[Ang06] J. E. Angus. An asymptotic expansion for the inspection paradox. Probability in the Engineering and Informational Sciences, 20(1):87–94, January 2006. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
2FA8638D7301C3B914CB07B75EF260CE.

Anonymous:1987:PVIb

[Ano87a] Anonymous. PES volume 1 issue 1 cover and back matter. Probability in the Engineering and Informational Sciences, 1(1):b1–b2, January 1987. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/3B9A45AE606ED3F3ACD250817E1AB7E1.

Anonymous:1987:PVIa

[Ano87b] Anonymous. PES volume 1 issue 1 cover and front matter. Probability in the Engineering and Informational Sciences, 1 (1):f1–f2, January 1987. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/2C35D0649B2D6251E4BFA5540B489D04.

Anonymous:1987:PVId

[Ano87c] Anonymous. PES volume 1 issue 2 cover and back matter. Probability in the Engineering and Informational Sciences, 1 (2):b1–b2, April 1987. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/5CF5C9382DE8094190D9D5A992E1B3ED.

Anonymous:1987:PVIc

[Ano87d] Anonymous. PES volume 1 issue 2 cover and front matter. Probability in the Engineering and Informational Sciences, 1 (2):f1–f2, April 1987. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/F77AB96965669171247AAA72B38B3887.

Anonymous:1987:PVIf

[Ano87e] Anonymous. PES volume 1 issue 3 cover and back matter. Probability in the Engineering and Informational Sciences, 1 (3):b1–b4, July 1987. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/45488D81B577572A8AEAC81632107D9A.

Anonymous:1987:PVIe

[Ano87f] Anonymous. PES volume 1 issue 3 cover and front matter. Probability in the Engineering and Informational Sciences, 1 (3):f1–f2, July 1987. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/4A721A1B4973FAE1B9600D72ED32FC7A.

Anonymous:1987:PVIh

[Ano87g] Anonymous. PES volume 1 issue 4 cover and back matter. Probability in the Engineering and Informational Sciences, 1(4):b1–b7, October 1987. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/14DFBC01D6223679CDA4FE83E1C5B820.

Anonymous:1987:PVIg

[Ano87h] Anonymous. PES volume 1 issue 4 cover and front matter. Probability in the Engineering and Informational Sciences, 1 (4):f1–f2, October 1987. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/CFE06EAC3F8E473BBDF87EC7EBB96443.

Anonymous:1988:PVIb

[Ano88a] Anonymous. PES volume 2 issue 1 cover and back matter. Probability in the Engineering and Informational Sciences, 2(1):b1–b3, January 1988. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/31535D86DF08CDE78B474F75BDABC804.

Anonymous:1988:PVIa

[Ano88b] Anonymous. PES volume 2 issue 1 cover and front matter. Probability in the Engineering and Informational Sciences, 2 (1):f1–f2, January 1988. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/54C9D769E26700167BA3623F5B2A219E.

Anonymous:1988:PVId

[Ano88c] Anonymous. PES volume 2 issue 2 cover and back matter. Probability in the Engineering and Informational Sciences, 2 (2):b1–b2, April 1988. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/FFEEA4D9F36F7E10FC908298E4955791.

Anonymous:1988:PVIc

[Ano88d] Anonymous. PES volume 2 issue 2 cover and front matter. Probability in the Engineering and Informational Sciences, 2 (2):f1–f2, April 1988. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/4E0C7055093A9BFC816D13369CC350F6.

Anonymous:1988:PVIf

[Ano88e] Anonymous. PES volume 2 issue 3 cover and back matter. Probability in the Engineering and Informational Sciences, 2 (3):b1–b2, July 1988. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/A363AEFAA3C5E49DDE1EB924E6659450.

Anonymous:1988:PVIe

[Ano88f] Anonymous. PES volume 2 issue 3 cover and front matter. Probability in the Engineering and Informational Sciences, 2 (3):f1–f2, July 1988. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/F6CA5B65B81B6513EA571F13B06DB382.

Anonymous:1988:PVIh

[Ano88g] Anonymous. PES volume 2 issue 4 cover and back matter. Probability in the Engineering and Informational Sciences, 2(4):b1–b2, October 1988. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/7BD4DB2FBF0872E131A4D8055B871E12.

Anonymous:1988:PVIg

[Ano88h] Anonymous. PES volume 2 issue 4 cover and front matter. Probability in the Engineering and Informational Sciences, 2 (4):f1–f2, October 1988. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/126E0580802D9B512B3BA2CDFFA2318D.

Anonymous:1989:PVIb

[Ano89a] Anonymous. PES volume 3 issue 1 cover and back matter. Probability in the Engineering and Informational Sciences, 3(1):b1–b2, January 1989. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/0BEB621BCF8296134E3FD681B1C8734D.

Anonymous:1989:PVIa

[Ano89b] Anonymous. PES volume 3 issue 1 cover and front matter. Probability in the Engineering and Informational Sciences, 3 (1):f1–f2, January 1989. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/6ED39BCA6D744283EE39939A6FDE8B07.

Anonymous:1989:PVId

[Ano89c] Anonymous. PES volume 3 issue 2 cover and back matter. Probability in the Engineering and Informational Sciences, 3 (2):b1–b2, April 1989. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/D1968D5EC3B695261A2A6D0754725C8B.

Anonymous:1989:PVIc

[Ano89d] Anonymous. PES volume 3 issue 2 cover and front matter. Probability in the Engineering and Informational Sciences, 3 (2):f1–f2, April 1989. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/F6CD2E2A916F8EA771CF2ADA3D4003AB.

Anonymous:1989:PVIf

[Ano89e] Anonymous. PES volume 3 issue 3 cover and back matter. Probability in the Engineering and Informational Sciences, 3 (3):b1–b2, July 1989. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/4DB7CE5F57289FB37B416DEB8AA494E1.

Anonymous:1989:PVIe

[Ano89f] Anonymous. PES volume 3 issue 3 cover and front matter. Probability in the Engineering and Informational Sciences, 3 (3):f1–f2, July 1989. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/DACA5CEE82751457B0E8C713930CE4BE.

Anonymous:1989:PVIh

[Ano89g] Anonymous. PES volume 3 issue 4 cover and back matter. Probability in the Engineering and Informational Sciences, 3(4):b1–b2, October 1989. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/507D03023EE697DDD1A9F73F6BF18BE8.

Anonymous:1989:PVIg

[Ano89h] Anonymous. PES volume 3 issue 4 cover and front matter. Probability in the Engineering and Informational Sciences, 3 (4):f1–f2, October 1989. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/E873DC4C9125AC902699CEDE8358EC08.

Anonymous:1990:PVIb

[Ano90a] Anonymous. PES volume 4 issue 1 cover and back matter. Probability in the Engineering and Informational Sciences, 4(1):b1–b2, January 1990. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/B645BE5ADFF087B7E909C5963376E0F0.

Anonymous:1990:PVIa

[Ano90b] Anonymous. PES volume 4 issue 1 cover and front matter. Probability in the Engineering and Informational Sciences, 4 (1):f1–f2, January 1990. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/05CEABDD267E029535305D426BB381D2.

Anonymous:1990:PVId

[Ano90c] Anonymous. PES volume 4 issue 2 cover and back matter. Probability in the Engineering and Informational Sciences, 4 (2):b1–b2, April 1990. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/A5A985171A9CBEC512701E7CFDE67C48.

Anonymous:1990:PVIc

[Ano90d] Anonymous. PES volume 4 issue 2 cover and front matter. Probability in the Engineering and Informational Sciences, 4 (2):f1–f2, April 1990. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/443002BA1831D180997D625565A06620.

Anonymous:1990:PVIf

[Ano90e] Anonymous. PES volume 4 issue 3 cover and back matter. Probability in the Engineering and Informational Sciences, 4 (3):b1–b2, July 1990. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/374F5682EF283F59EED9F63016863404.

Anonymous:1990:PVIe

[Ano90f] Anonymous. PES volume 4 issue 3 cover and front matter. Probability in the Engineering and Informational Sciences, 4 (3):f1–f2, July 1990. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/EDA24AE1D7017CFAC63C77CC33DD69F3.

Anonymous:1990:PVIh

[Ano90g] Anonymous. PES volume 4 issue 4 cover and back matter. Probability in the Engineering and Informational Sciences, 4(4):b1–b2, October 1990. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/88014445B9B7117B62B473EF585D318F.

Anonymous:1990:PVIg

[Ano90h] Anonymous. PES volume 4 issue 4 cover and front matter. Probability in the Engineering and Informational Sciences, 4 (4):f1–f2, October 1990. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/70184AA39D8DFE0D47FC2D09394E8AAE.

Anonymous:1991:C

[Ano91a] Anonymous. Correction. Probability in the Engineering and Informational Sciences, 5(1):129, January 1991. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
CCA52C5CFF1AF8ECE90A7627BC1174DA.

Anonymous:1991:PVIb

[Ano91b] Anonymous. PES volume 5 issue 1 cover and back matter. Probability in the Engineering and Informational Sciences, 5(1):b1–b2, January 1991. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/4B16494EC891914E87A83D25A24D1BE2.

Anonymous:1991:PVIa

[Ano91c] Anonymous. PES volume 5 issue 1 cover and front matter. Probability in the Engineering and Informational Sciences, 5 (1):f1–f2, January 1991. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/4EBD0DC69EF9C1F296A17A8263693FB9.

Anonymous:1991:PVId

[Ano91d] Anonymous. PES volume 5 issue 2 cover and back matter. Probability in the Engineering and Informational Sciences, 5 (2):b1–b2, April 1991. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/E69355F90228C333957742897985ED13.

Anonymous:1991:PVIc

[Ano91e] Anonymous. PES volume 5 issue 2 cover and front matter. Probability in the Engineering and Informational Sciences, 5 (2):f1–f2, April 1991. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/38807F95EFD2BC3E62041BE06DA864FD.

Anonymous:1991:PVIf

[Ano91f] Anonymous. PES volume 5 issue 3 cover and back matter. Probability in the Engineering and Informational Sciences, 5 (3):b1–b2, July 1991. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/BFF41EE8B1257DFE3ACA3E808D8EC8F2.

Anonymous:1991:PVIe

[Ano91g] Anonymous. PES volume 5 issue 3 cover and front matter. Probability in the Engineering and Informational Sciences, 5 (3):f1–f2, July 1991. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/871B4F8A804128FAF921CE8F2D87CF5D.

Anonymous:1991:PVIh

[Ano91h] Anonymous. PES volume 5 issue 4 cover and back matter. Probability in the Engineering and Informational Sciences, 5(4):b1–b2, October 1991. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/9A01CFBE821CDDF26396F8FB4F882818.

Anonymous:1991:PVIg

[Ano91i] Anonymous. PES volume 5 issue 4 cover and front matter. Probability in the Engineering and Informational Sciences, 5 (4):f1–f2, October 1991. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/57DF6234A5DDD9CEAAB23BDC43754330.

Anonymous:1992:PVIb

[Ano92a] Anonymous. PES volume 6 issue 1 cover and back matter. Probability in the Engineering and Informational Sciences, 6(1):b1–b2, January 1992. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/E3FD3D6ED7BF6248889EF1393DB7D4BF.

Anonymous:1992:PVIa

[Ano92b] Anonymous. PES volume 6 issue 1 cover and front matter. Probability in the Engineering and Informational Sciences, 6 (1):f1–f2, January 1992. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/999E22ECFB16821E30E58BE9E2676AAB.

Anonymous:1992:PVId

[Ano92c] Anonymous. PES volume 6 issue 2 cover and back matter. Probability in the Engineering and Informational Sciences, 6 (2):b1–b2, April 1992. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/A4967870FDC4E3CD8732C48E25894432.

Anonymous:1992:PVIc

[Ano92d] Anonymous. PES volume 6 issue 2 cover and front matter. Probability in the Engineering and Informational Sciences, 6 (2):f1–f2, April 1992. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/D90EF7B8B6CA084F2D87ED08C12DA3A7.

Anonymous:1992:PVIf

[Ano92e] Anonymous. PES volume 6 issue 3 cover and back matter. Probability in the Engineering and Informational Sciences, 6 (3):b1–b2, July 1992. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/18DFC50B5086AA874D73C8E663F87C01.

Anonymous:1992:PVIe

[Ano92f] Anonymous. PES volume 6 issue 3 cover and front matter. Probability in the Engineering and Informational Sciences, 6 (3):f1–f2, July 1992. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/7BA3AE22FFD49FFA30205342A97BCA30.

Anonymous:1992:PVIh

[Ano92g] Anonymous. PES volume 6 issue 4 cover and back matter. Probability in the Engineering and Informational Sciences, 6(4):b1–b2, October 1992. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/26E4C69AA8D5E620B21CB74F01D97C65.

Anonymous:1992:PVIg

[Ano92h] Anonymous. PES volume 6 issue 4 cover and front matter. Probability in the Engineering and Informational Sciences, 6 (4):f1–f2, October 1992. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/4C32F3274C9F6B25D6FEE9947ADCD2FD.

Anonymous:1993:PVIa

[Ano93a] Anonymous. PES volume 7 issue 1 cover. Probability in the Engineering and Informational Sciences, 7(1):b1, January 1993. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
28AA74460292BEEB194DFFCAF5D2A1BC.

Anonymous:1993:PVIb

[Ano93b] Anonymous. PES volume 7 issue 2 cover. Probability in the Engineering and Informational Sciences, 7(2):b1, April 1993. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
6B2E41197704016A6F97ED1326E7FFF7.

Anonymous:1993:PVIc

[Ano93c] Anonymous. PES volume 7 issue 3 cover. Probability in the Engineering and Informational Sciences, 7(3):b1, July 1993. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
54AE4EB7E5E085FFAF79F9D84472A059.

Anonymous:1993:PVId

[Ano93d] Anonymous. PES volume 7 issue 4 cover and back matter. Probability in the Engineering and Informational Sciences, 7(4):b1–b2, October 1993. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/D1CCD9A63B4517CAAFF167176AB16FDD.

Anonymous:1994:C

[Ano94a] Anonymous. Corrections. Probability in the Engineering and Informational Sciences, 8(1):152, January 1994. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
86B4CC3DC0C7224A9CD9A14150779C8A.

Anonymous:1994:PVIa

[Ano94b] Anonymous. PES volume 8 issue 1 cover. Probability in the Engineering and Informational Sciences, 8(1):b1, January 1994. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
927EE67F6FF01014279857038D2D942E.

Anonymous:1994:PVIb

[Ano94c] Anonymous. PES volume 8 issue 2 cover. Probability in the Engineering and Informational Sciences, 8(2):b1, April 1994. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
B91453C046AB432BC14987C41A156A8C.

Anonymous:1994:PVIc

[Ano94d] Anonymous. PES volume 8 issue 3 cover. Probability in the Engineering and Informational Sciences, 8(3):b1, July 1994. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
1B3272AEEBFEF19FE429392EA36A2B00.

Anonymous:1994:PVId

[Ano94e] Anonymous. PES volume 8 issue 4 cover and back matter. Probability in the Engineering and Informational Sciences, 8(4):b1–b2, October 1994. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/0052AC8E82EA3F9F7FB6946C8950173B.

Anonymous:1995:PVIb

[Ano95a] Anonymous. PES volume 9 issue 1 cover. Probability in the Engineering and Informational Sciences, 9(1):b1, January 1995. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
80EAA77557DCCF03D4CA0D68AA18F8F6.

Anonymous:1995:PVIa

[Ano95b] Anonymous. PES volume 9 issue 1 cover and front matter. Probability in the Engineering and Informational Sciences, 9 (1):f1–f2, January 1995. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/D28602F7753A6D072E25369BDA07F569.

Anonymous:1995:PVIc

[Ano95c] Anonymous. PES volume 9 issue 2 cover. Probability in the Engineering and Informational Sciences, 9(2):b1, April 1995. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
8F5488DEA76901B976A64BA3B03CEBC4.

Anonymous:1995:PVId

[Ano95d] Anonymous. PES volume 9 issue 3 cover. Probability in the Engineering and Informational Sciences, 9(3):b1, July 1995. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
2C782B19E65DA461E70F830C0FD89B49.

Anonymous:1995:PVIe

[Ano95e] Anonymous. PES volume 9 issue 4 cover and back matter. Probability in the Engineering and Informational Sciences, 9(4):b1–b2, October 1995. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/E80B7577F4EF7EAEA22CC0932378A174.

Anonymous:1996:PVIa

[Ano96a] Anonymous. PES volume 10 issue 1 cover. Probability in the Engineering and Informational Sciences, 10(1):b1, January 1996. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
C1118282194E616F8EC346BD4D8D9E93.

Anonymous:1996:PVIb

[Ano96b] Anonymous. PES volume 10 issue 2 cover. Probability in the Engineering and Informational Sciences, 10(2):b1, April 1996. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
EB05B61E7B362230D3C05A638EA9EB85.

Anonymous:1996:PVIc

[Ano96c] Anonymous. PES volume 10 issue 3 cover. Probability in the Engineering and Informational Sciences, 10(3):b1, July 1996. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
F88C195A6115029F2696D98FE4B38E41.

Anonymous:1996:PVId

[Ano96d] Anonymous. PES volume 10 issue 4 cover and back matter. Probability in the Engineering and Informational Sciences, 10(4):b1–b2, October 1996. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/768A2711343FF086B8843BCA587DF964.

Anonymous:1997:PVIa

[Ano97a] Anonymous. PES volume 11 issue 1 cover. Probability in the Engineering and Informational Sciences, 11(1):b1, January 1997. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
FDD3DD98AEA45C9C4D87263729138819.

Anonymous:1997:PVIb

[Ano97b] Anonymous. PES volume 11 issue 2 cover. Probability in the Engineering and Informational Sciences, 11(2):b1, April 1997. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
0E45981C1E376A390BC11BF8481D649A.

Anonymous:1997:PVIc

[Ano97c] Anonymous. PES volume 11 issue 3 cover. Probability in the Engineering and Informational Sciences, 11(3):b1, July 1997. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
A24FB85DBA455B40DF27E33224E66E20.

Anonymous:1997:PVId

[Ano97d] Anonymous. PES volume 11 issue 4 cover and back matter. Probability in the Engineering and Informational Sciences, 11(4):b1–b2, October 1997. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/14E2AE19602DEB446311A799EA2DC258.

Anonymous:1998:PVIa

[Ano98a] Anonymous. PES volume 12 issue 1 cover. Probability in the Engineering and Informational Sciences, 12(1):b1, January 1998. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
36EA30992DB5B4B9CF744620827A1BAB.

Anonymous:1998:PVIb

[Ano98b] Anonymous. PES volume 12 issue 2 cover. Probability in the Engineering and Informational Sciences, 12(2):b1, April 1998. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
E0CB855B3CA04568A31CC6B6829644F4.

Anonymous:1998:PVIc

[Ano98c] Anonymous. PES volume 12 issue 3 cover. Probability in the Engineering and Informational Sciences, 12(3):b1, July 1998. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
09739CD1E2583D305A8DD7C600B0FDDF.

Anonymous:1998:PVId

[Ano98d] Anonymous. PES volume 12 issue 4 cover. Probability in the Engineering and Informational Sciences, 12(4):b1, October 1998. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
E76545C9408184D8454C8754225F716A.

Anonymous:2009:PVIb

[Ano09a] Anonymous. PES volume 23 issue 1 cover and back matter. Probability in the Engineering and Informational Sciences, 23(1):b1–b2, January 2009. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/5E25A2B8BDA6C694708E44BC79ABD202.

Anonymous:2009:PVIa

[Ano09b] Anonymous. PES volume 23 issue 1 cover and front matter. Probability in the Engineering and Informational Sciences, 23 (1):f1–f2, January 2009. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/E2D66EBBD5B68EAA339128A37B7C2A1B.

Anonymous:2009:PVId

[Ano09c] Anonymous. PES volume 23 issue 2 cover and back matter. Probability in the Engineering and Informational Sciences, 23 (2):b1–b2, April 2009. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/C8954A1A1A96953F7E042F5AA7C39084.

Anonymous:2009:PVIc

[Ano09d] Anonymous. PES volume 23 issue 2 cover and front matter. Probability in the Engineering and Informational Sciences, 23 (2):f1–f2, April 2009. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/D68B57AC5361C91FCB5D168AB84A879E.

Anonymous:2009:PVIf

[Ano09e] Anonymous. PES volume 23 issue 3 cover and back matter. Probability in the Engineering and Informational Sciences, 23 (3):b1–b2, July 2009. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/4F008E5E87BB4EE608DA32918A9EAA1F.

Anonymous:2009:PVIe

[Ano09f] Anonymous. PES volume 23 issue 3 cover and front matter. Probability in the Engineering and Informational Sciences, 23 (3):f1–f2, July 2009. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/C3B15A52AC388962A94B4DB6B57F118A.

Anonymous:2009:PVIh

[Ano09g] Anonymous. PES volume 23 issue 4 cover and back matter. Probability in the Engineering and Informational Sciences, 23(4):b1–b2, October 2009. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/2A5F0CC3DCCAC0953C70647BE5B57FBC.

Anonymous:2009:PVIg

[Ano09h] Anonymous. PES volume 23 issue 4 cover and front matter. Probability in the Engineering and Informational Sciences, 23 (4):f1–f2, October 2009. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/8879B2A6DB91518DBDC516FF5CFEF139.

Anonymous:2010:PVIb

[Ano10a] Anonymous. PES volume 24 issue 1 cover and back matter. Probability in the Engineering and Informational Sciences, 24(1):b1–b3, January 2010. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/070715AD16230F24C0CDF1FF80D1666A.

Anonymous:2010:PVIa

[Ano10b] Anonymous. PES volume 24 issue 1 cover and front matter. Probability in the Engineering and Informational Sciences, 24 (1):f1–f2, January 2010. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/770BABD0C66B43E6A5CDBFF0559818F1.

Anonymous:2010:PVId

[Ano10c] Anonymous. PES volume 24 issue 2 cover and back matter. Probability in the Engineering and Informational Sciences, 24 (2):b1–b3, April 2010. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/32641AB839C7B9509E672C403CAE8A9A.

Anonymous:2010:PVIc

[Ano10d] Anonymous. PES volume 24 issue 2 cover and front matter. Probability in the Engineering and Informational Sciences, 24 (2):f1–f2, April 2010. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/7D22E757F50A1CF0D0E353D401773ADD.

Anonymous:2010:PVIf

[Ano10e] Anonymous. PES volume 24 issue 3 cover and back matter. Probability in the Engineering and Informational Sciences, 24 (3):b1–b2, July 2010. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/61C1636E8D6E4910A406D9C19E4072DB.

Anonymous:2010:PVIe

[Ano10f] Anonymous. PES volume 24 issue 3 cover and front matter. Probability in the Engineering and Informational Sciences, 24 (3):f1–f2, July 2010. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/7F9C0FCB6699B111DEBABBE77F30113F.

Anonymous:2010:PVIh

[Ano10g] Anonymous. PES volume 24 issue 4 cover and back matter. Probability in the Engineering and Informational Sciences, 24(4):b1–b2, October 2010. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/17D18CC30B48D7404123FA6017710FB0.

Anonymous:2010:PVIg

[Ano10h] Anonymous. PES volume 24 issue 4 cover and front matter. Probability in the Engineering and Informational Sciences, 24 (4):f1–f2, October 2010. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/78C7F13D7C1303F20794B8AA570FED71.

Anonymous:2011:PVIb

[Ano11a] Anonymous. PES volume 25 issue 1 cover and back matter. Probability in the Engineering and Informational Sciences, 25(1):b1–b2, January 2011. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/D61B55AD881496E74F70BE9AC9FB9BCF.

Anonymous:2011:PVIa

[Ano11b] Anonymous. PES volume 25 issue 1 cover and front matter. Probability in the Engineering and Informational Sciences, 25 (1):f1–f2, January 2011. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/FD9EDD84C33DB9AFBBE0284CE152BC81.

Anonymous:2011:PVId

[Ano11c] Anonymous. PES volume 25 issue 2 cover and back matter. Probability in the Engineering and Informational Sciences, 25 (2):b1–b3, April 2011. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/BE34D8616646708407CD80AC196E8272.

Anonymous:2011:PVIc

[Ano11d] Anonymous. PES volume 25 issue 2 cover and front matter. Probability in the Engineering and Informational Sciences, 25 (2):f1–f2, April 2011. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/2673250C1588A2FEBF95846FF84B1BAC.

Anonymous:2011:PVIf

[Ano11e] Anonymous. PES volume 25 issue 3 cover and back matter. Probability in the Engineering and Informational Sciences, 25 (3):b1–b2, July 2011. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/00879903F52138AFD89BDB6A3B3FD84E.

Anonymous:2011:PVIe

[Ano11f] Anonymous. PES volume 25 issue 3 cover and front matter. Probability in the Engineering and Informational Sciences, 25 (3):f1–f2, July 2011. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/0A3FE6EF99E655E3921FB4B474A9C3FE.

Anonymous:2011:PVIh

[Ano11g] Anonymous. PES volume 25 issue 4 cover and back matter. Probability in the Engineering and Informational Sciences, 25(4):b1–b2, October 2011. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/7B8EFC8F7F55EE768123CC1B5C11281D.

Anonymous:2011:PVIg

[Ano11h] Anonymous. PES volume 25 issue 4 cover and front matter. Probability in the Engineering and Informational Sciences, 25 (4):f1–f2, October 2011. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/5E14DE5799EADF5BFF89CBBF2F1078B6.

Anonymous:2012:PVIb

[Ano12a] Anonymous. PES volume 26 issue 1 cover and back matter. Probability in the Engineering and Informational Sciences, 26(1):b1–b3, January 2012. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/333F5CAA199D9FE1C58D37221E7C98CD.

Anonymous:2012:PVIa

[Ano12b] Anonymous. PES volume 26 issue 1 cover and front matter. Probability in the Engineering and Informational Sciences, 26 (1):f1–f2, January 2012. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/B6197C6FAA3BA8357DBBED265B0F9866.

Anonymous:2012:PVId

[Ano12c] Anonymous. PES volume 26 issue 2 cover and back matter. Probability in the Engineering and Informational Sciences, 26 (2):b1–b2, April 2012. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/AC5658C40473314A0B4C82FAA269855D.

Anonymous:2012:PVIc

[Ano12d] Anonymous. PES volume 26 issue 2 cover and front matter. Probability in the Engineering and Informational Sciences, 26 (2):f1–f2, April 2012. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/C6F44B1BA9E675A2BD826F20694DA938.

Anonymous:2012:PVIf

[Ano12e] Anonymous. PES volume 26 issue 3 cover and back matter. Probability in the Engineering and Informational Sciences, 26 (3):b1–b3, July 2012. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/575C7EB1F91C1EC2785072C40798A0A5.

Anonymous:2012:PVIe

[Ano12f] Anonymous. PES volume 26 issue 3 cover and front matter. Probability in the Engineering and Informational Sciences, 26 (3):f1–f2, July 2012. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/8930FB518373AAFC576DEC7141B3CC42.

Anonymous:2012:PVIh

[Ano12g] Anonymous. PES volume 26 issue 4 cover and back matter. Probability in the Engineering and Informational Sciences, 26(4):b1–b3, October 2012. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/A5856FB1225C8D387F6E301110B96FE6.

Anonymous:2012:PVIg

[Ano12h] Anonymous. PES volume 26 issue 4 cover and front matter. Probability in the Engineering and Informational Sciences, 26 (4):f1–f2, October 2012. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/E37F4A14A85F6B6D0F3BE01B3C8098A6.

Anonymous:2013:PVIb

[Ano13a] Anonymous. PES volume 27 issue 1 cover and back matter. Probability in the Engineering and Informational Sciences, 27(1):b1–b2, January 2013. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/84805F1B58365792487B108E7B4B6C3D.

Anonymous:2013:PVIa

[Ano13b] Anonymous. PES volume 27 issue 1 cover and front matter. Probability in the Engineering and Informational Sciences, 27 (1):f1–f2, January 2013. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/6E665A76DDFEE05E9E04E90D54CD01A3.

Anonymous:2013:PVId

[Ano13c] Anonymous. PES volume 27 issue 2 cover and back matter. Probability in the Engineering and Informational Sciences, 27 (2):b1–b2, April 2013. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/13B142367787D29B787415AF72A2C709.

Anonymous:2013:PVIc

[Ano13d] Anonymous. PES volume 27 issue 2 cover and front matter. Probability in the Engineering and Informational Sciences, 27 (2):f1–f2, April 2013. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/78F9C023C28047F85B3996E55C68A623.

Anonymous:2013:PVIf

[Ano13e] Anonymous. PES volume 27 issue 3 cover and back matter. Probability in the Engineering and Informational Sciences, 27 (3):b1–b2, July 2013. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/1241E9417DD1B8AD7AB058236B937D1D.

Anonymous:2013:PVIe

[Ano13f] Anonymous. PES volume 27 issue 3 cover and front matter. Probability in the Engineering and Informational Sciences, 27 (3):f1–f2, July 2013. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/94F538B3F5EA9F1008668B3C9FD080EE.

Anonymous:2013:PVIh

[Ano13g] Anonymous. PES volume 27 issue 4 cover and back matter. Probability in the Engineering and Informational Sciences, 27(4):b1–b2, October 2013. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/0C5723EB9A4282088068DC211BBD192E.

Anonymous:2013:PVIg

[Ano13h] Anonymous. PES volume 27 issue 4 cover and front matter. Probability in the Engineering and Informational Sciences, 27 (4):f1–f2, October 2013. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/87054DA73348A0FC9E3E007B2431FC01.

Anonymous:2014:PVIb

[Ano14a] Anonymous. PES volume 28 issue 1 cover and back matter. Probability in the Engineering and Informational Sciences, 28(1):b1–b2, January 2014. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/C5A0D1E5CBCC5EA7B15DE4C8FCA1777C.

Anonymous:2014:PVIa

[Ano14b] Anonymous. PES volume 28 issue 1 cover and front matter. Probability in the Engineering and Informational Sciences, 28 (1):f1–f2, January 2014. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/8AECA8C501947562DE162E957DB0B76F.

Anonymous:2014:PVId

[Ano14c] Anonymous. PES volume 28 issue 2 cover and back matter. Probability in the Engineering and Informational Sciences, 28 (2):b1–b2, April 2014. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/9862764D39BD63295B6AA385D6CC9D9F.

Anonymous:2014:PVIc

[Ano14d] Anonymous. PES volume 28 issue 2 cover and front matter. Probability in the Engineering and Informational Sciences, 28 (2):f1–f2, April 2014. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/C84B7E85A358C802B3C65738701E6743.

Anonymous:2014:PVIf

[Ano14e] Anonymous. PES volume 28 issue 3 cover and back matter. Probability in the Engineering and Informational Sciences, 28 (3):b1–b2, July 2014. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/DA65FACFE00B2001988085E5BB272827.

Anonymous:2014:PVIe

[Ano14f] Anonymous. PES volume 28 issue 3 cover and front matter. Probability in the Engineering and Informational Sciences, 28 (3):f1–f2, July 2014. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/26FA6CE801E978F65B3253362B78AA81.

Anonymous:2014:PVIh

[Ano14g] Anonymous. PES volume 28 issue 4 cover and back matter. Probability in the Engineering and Informational Sciences, 28(4):b1–b2, October 2014. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/FB95732509B5B466EEE2FD7E80CD2277.

Anonymous:2014:PVIg

[Ano14h] Anonymous. PES volume 28 issue 4 cover and front matter. Probability in the Engineering and Informational Sciences, 28 (4):f1–f2, October 2014. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/71A5611AB379C5466E08BF979F6214F2.

Anonymous:2015:PVIb

[Ano15a] Anonymous. PES volume 29 issue 1 cover and back matter. Probability in the Engineering and Informational Sciences, 29 (1):b1, January 2015. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/A99D1CF9ED10C17C8A6B3152EAED88F5.

Anonymous:2015:PVIa

[Ano15b] Anonymous. PES volume 29 issue 1 cover and front matter. Probability in the Engineering and Informational Sciences, 29 (1):f1, January 2015. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/055FBC59C303CCD943790146C58290BE.

Anonymous:2015:PVId

[Ano15c] Anonymous. PES volume 29 issue 2 cover and back matter. Probability in the Engineering and Informational Sciences, 29 (2):b1–b2, April 2015. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/A1721D1E1666073C7F5C2859C98ED38D.

Anonymous:2015:PVIc

[Ano15d] Anonymous. PES volume 29 issue 2 cover and front matter. Probability in the Engineering and Informational Sciences, 29 (2):f1–f2, April 2015. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/6641D2D4550409A82341A4F53BB0899D.

Anonymous:2015:PVIf

[Ano15e] Anonymous. PES volume 29 issue 3 cover and back matter. Probability in the Engineering and Informational Sciences, 29 (3):b1–b2, July 2015. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/E86F7E60A2D066BABB06FBDC36156A11.

Anonymous:2015:PVIe

[Ano15f] Anonymous. PES volume 29 issue 3 cover and front matter. Probability in the Engineering and Informational Sciences, 29 (3):f1–f2, July 2015. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/0E77C5B3F316F01B08531E6853ECBFE0.

Anonymous:2015:PVIh

[Ano15g] Anonymous. PES volume 29 issue 4 cover and back matter. Probability in the Engineering and Informational Sciences, 29(4):b1–b2, October 2015. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/85B610E6D87FF341F2DA68B36B45D543.

Anonymous:2015:PVIg

[Ano15h] Anonymous. PES volume 29 issue 4 cover and front matter. Probability in the Engineering and Informational Sciences, 29 (4):f1–f2, October 2015. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/3F948D756ADD1EA30E93AF19A59C0B59.

Anonymous:2016:PVIb

[Ano16a] Anonymous. PES volume 30 issue 1 cover and back matter. Probability in the Engineering and Informational Sciences, 30(1):b1–b2, January 2016. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/7A3913D03DF3879C3C1700DE611A6FEA.

Anonymous:2016:PVIa

[Ano16b] Anonymous. PES volume 30 issue 1 cover and front matter. Probability in the Engineering and Informational Sciences, 30 (1):f1–f2, January 2016. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/69FA6ADE8C467F2AE6FFA85B7635F3F8.

Anonymous:2016:PVId

[Ano16c] Anonymous. PES volume 30 issue 2 cover and back matter. Probability in the Engineering and Informational Sciences, 30 (2):b1–b2, April 2016. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/B7C421C75FE12EB02463B5187711B9BA.

Anonymous:2016:PVIc

[Ano16d] Anonymous. PES volume 30 issue 2 cover and front matter. Probability in the Engineering and Informational Sciences, 30 (2):f1–f2, April 2016. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/309B49A3329558F29D93215E6E01B0CE.

Anonymous:2016:PVIf

[Ano16e] Anonymous. PES volume 30 issue 3 cover and back matter. Probability in the Engineering and Informational Sciences, 30 (3):b1–b2, July 2016. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/DFC60B5C9A7E1A7FEBA731686602CC79.

Anonymous:2016:PVIe

[Ano16f] Anonymous. PES volume 30 issue 3 cover and front matter. Probability in the Engineering and Informational Sciences, 30 (3):f1–f2, July 2016. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/8FDFC79A0BD25F2DD873CCD4A7B01466.

Anonymous:2016:PVIh

[Ano16g] Anonymous. PES volume 30 issue 4 cover and back matter. Probability in the Engineering and Informational Sciences, 30(4):b1–b2, October 2016. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/7731B38450A48C8088352D18DBCFA9A1.

Anonymous:2016:PVIg

[Ano16h] Anonymous. PES volume 30 issue 4 cover and front matter. Probability in the Engineering and Informational Sciences, 30 (4):f1–f2, October 2016. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/FA40CB86B7F71AAAF2417AED111FFED0.

Anonymous:2017:PVIb

[Ano17a] Anonymous. PES volume 31 issue 1 cover and back matter. Probability in the Engineering and Informational Sciences, 31(1):b1–b2, January 2017. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/318C53548C8002C54FACC7745DAFEC42.

Anonymous:2017:PVIa

[Ano17b] Anonymous. PES volume 31 issue 1 cover and front matter. Probability in the Engineering and Informational Sciences, 31 (1):f1–f2, January 2017. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/B2076719A3E73735D32059D99780E0E0.

Anonymous:2017:PVId

[Ano17c] Anonymous. PES volume 31 issue 2 cover and back matter. Probability in the Engineering and Informational Sciences, 31 (2):b1–b2, April 2017. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/4FA6BD7FA0B4BC3A8F27C39A4E7A3D40.

Anonymous:2017:PVIc

[Ano17d] Anonymous. PES volume 31 issue 2 cover and front matter. Probability in the Engineering and Informational Sciences, 31 (2):f1–f2, April 2017. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/F00D7BA95FA6FAEC097FFAF7B0A03F18.

Anonymous:2017:PVIf

[Ano17e] Anonymous. PES volume 31 issue 3 cover and back matter. Probability in the Engineering and Informational Sciences, 31 (3):b1–b2, July 2017. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/BB4B273618440AE771BC6F5E4DE7DD4C.

Anonymous:2017:PVIe

[Ano17f] Anonymous. PES volume 31 issue 3 cover and front matter. Probability in the Engineering and Informational Sciences, 31 (3):f1–f2, July 2017. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/37011E54EB8D5D94B0AD79DC44867E43.

Anonymous:2017:PVIh

[Ano17g] Anonymous. PES volume 31 issue 4 cover and back matter. Probability in the Engineering and Informational Sciences, 31(4):b1–b2, October 2017. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/1431A90F9553147846A90C5710562E9F.

Anonymous:2017:PVIg

[Ano17h] Anonymous. PES volume 31 issue 4 cover and front matter. Probability in the Engineering and Informational Sciences, 31 (4):f1–f2, October 2017. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/5E401E869996D2302E8A819D4FD19DBE.

Anonymous:2018:PVIb

[Ano18a] Anonymous. PES volume 32 issue 1 cover and back matter. Probability in the Engineering and Informational Sciences, 32(1):b1–b2, January 2018. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/pes-
volume-32-issue-1-cover-and-back-matter/BF11E3E4CCA73DE6852D54F7ECB75886.

Anonymous:2018:PVIa

[Ano18b] Anonymous. PES volume 32 issue 1 cover and front matter. Probability in the Engineering and Informational Sciences, 32(1):f1–f2, January 2018. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/pes-
volume-32-issue-1-cover-and-front-matter/43214063E898AC15F5A9F0601E232EE5.

Anonymous:2018:PVId

[Ano18c] Anonymous. PES volume 32 issue 2 cover and back matter. Probability in the Engineering and Informational Sciences, 32(2):b1–b2, April 2018. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/pes-
volume-32-issue-2-cover-and-back-matter/5600E47EF646D678BCD4468BC1098895.

Anonymous:2018:PVIc

[Ano18d] Anonymous. PES volume 32 issue 2 cover and front matter. Probability in the Engineering and Informational Sciences, 32(2):f1–f2, April 2018. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/pes-
volume-32-issue-2-cover-and-front-matter/7103731E53A765F455853C293ACBFC3E.

Anonymous:2018:PVIf

[Ano18e] Anonymous. PES volume 32 issue 3 cover and back matter. Probability in the Engineering and Informational Sciences, 32(3):b1–b2, July 2018. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/pes-
volume-32-issue-3-cover-and-back-matter/5B6AFCB0A8115BB0004E80D307556E8F.

Anonymous:2018:PVIe

[Ano18f] Anonymous. PES volume 32 issue 3 cover and front matter. Probability in the Engineering and Informational Sciences, 32(3):f1–f2, July 2018. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/pes-
volume-32-issue-3-cover-and-front-matter/F7E436F1863F097E4201CBEA196DCBFA.

Anonymous:2018:PVIh

[Ano18g] Anonymous. PES volume 32 issue 4 cover and back matter. Probability in the Engineering and Informational Sciences, 32(4):b1–b2, October 2018. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/pes-
volume-32-issue-4-cover-and-back-matter/F2DBEF81AEA45A3167690E5C8C6926B4.

Anonymous:2018:PVIg

[Ano18h] Anonymous. PES volume 32 issue 4 cover and front matter. Probability in the Engineering and Informational Sciences, 32(4):f1–f2, October 2018. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-

engineering-and-informational-sciences/article/pes-
volume-32-issue-4-cover-and-front-matter/9E736E4014767A456DAC8313CD902A59.

Anonymous:2019:PVIb

[Ano19a] Anonymous. PES volume 33 issue 1 cover and back matter. Probability in the Engineering and Informational Sciences, 33(1):b1–b2, January 2019. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/pes-
volume-33-issue-1-cover-and-back-matter/D0E15A9EFFB1ADA8D003193A2F5F5695.

Anonymous:2019:PVIa

[Ano19b] Anonymous. PES volume 33 issue 1 cover and front matter. Probability in the Engineering and Informational Sciences, 33(1):f1–f2, January 2019. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/pes-
volume-33-issue-1-cover-and-front-matter/4CDD4A534644AF419E5934DC496C4568.

Anonymous:2019:PVId

[Ano19c] Anonymous. PES volume 33 issue 2 cover and back matter. Probability in the Engineering and Informational Sciences, 33(2):b1–b2, April 2019. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/pes-
volume-33-issue-2-cover-and-back-matter/7A49E89E08FE3C3A9E0CF4AADDF90019.

Anonymous:2019:PVIc

[Ano19d] Anonymous. PES volume 33 issue 2 cover and front matter. Probability in the Engineering and Informational Sciences, 33(2):f1–f2, April 2019. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/pes-
volume-33-issue-2-cover-and-front-matter/20C3304EF263394B7F11E3C24739DE41.

Anonymous:2019:PVIf

[Ano19e] Anonymous. PES volume 33 issue 3 cover and back matter. Probability in the Engineering and Informational Sciences, 33(3):b1–b2, July 2019. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/pes-
volume-33-issue-3-cover-and-back-matter/A59FFF6B2C3CE02E2F67B1DC1A38A338.

Anonymous:2019:PVIe

[Ano19f] Anonymous. PES volume 33 issue 3 cover and front matter. Probability in the Engineering and Informational Sciences, 33(3):f1–f2, July 2019. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/pes-
volume-33-issue-3-cover-and-front-matter/BC127566F51B08E1953763112AA94EE1.

Anonymous:2019:PVIh

[Ano19g] Anonymous. PES volume 33 issue 4 cover and back matter. Probability in the Engineering and Informational Sciences, 33(4):b1–b2, October 2019. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/pes-
volume-33-issue-4-cover-and-back-matter/73B6787639B2CABBD22FBF218DCAF2DE.

Anonymous:2019:PVIg

[Ano19h] Anonymous. PES volume 33 issue 4 cover and front matter. Probability in the Engineering and Informational Sciences, 33(4):f1–f2, October 2019. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/pes-
volume-33-issue-4-cover-and-front-matter/DD6004CEF65B940B36FF43B1791A8180.

Anonymous:2020:PVIb

[Ano20a] Anonymous. PES volume 34 issue 1 cover and back matter. Probability in the Engineering and Informational Sciences, 34(1):b1–b2, January 2020. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/pes-
volume-34-issue-1-cover-and-back-matter/4F5D4314ED701CA2116AC1F875E08EE4.

Anonymous:2020:PVIa

[Ano20b] Anonymous. PES volume 34 issue 1 cover and front matter. Probability in the Engineering and Informational Sciences, 34(1):f1–f2, January 2020. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/pes-
volume-34-issue-1-cover-and-front-matter/8E6AC87DFCF3AE8DD44B222A8F71DD11.

Anonymous:2020:PVId

[Ano20c] Anonymous. PES volume 34 issue 2 cover and back matter. Probability in the Engineering and Informational Sciences, 34(2):b1–b2, April 2020. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/pes-
volume-34-issue-2-cover-and-back-matter/E64134B1E6AD700D17B69B442C4980EF.

Anonymous:2020:PVIc

[Ano20d] Anonymous. PES volume 34 issue 2 cover and front matter. Probability in the Engineering and Informational Sciences, 34(2):f1–f2, April 2020. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/pes-
volume-34-issue-2-cover-and-front-matter/4ED8097D3AEC24FF9DD06777C971A6A4.

Anonymous:2020:PVIf

[Ano20e] Anonymous. PES volume 34 issue 3 cover and back matter. Probability in the Engineering and Informational Sciences, 34(3):b1–b2, July 2020. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/pes-
volume-34-issue-3-cover-and-back-matter/CD38E5D208FF8CB82CE14B4640B06DC3.

Anonymous:2020:PVIe

[Ano20f] Anonymous. PES volume 34 issue 3 cover and front matter. Probability in the Engineering and Informational Sciences, 34(3):f1–f2, July 2020. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/pes-
volume-34-issue-3-cover-and-front-matter/79BCDDDBB426A7D3BDDEC1AE1DFD27C5.

Anonymous:2020:PVIh

[Ano20g] Anonymous. PES volume 34 issue 4 cover and back matter. Probability in the Engineering and Informational Sciences, 34(4):b1–b2, October 2020. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/pes-
volume-34-issue-4-cover-and-back-matter/B23B94CAD1B07A086E317F23439558DA.

Anonymous:2020:PVIg

[Ano20h] Anonymous. PES volume 34 issue 4 cover and front matter. Probability in the Engineering and Informational Sciences, 34(4):f1–f2, October 2020. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/pes-
volume-34-issue-4-cover-and-front-matter/FD1B5110968BE48A00EA6A01952E14D0.

Anonymous:2021:ARM

[Ano21a] Anonymous. Algebraic reliability of multi-state k-out-of-n systems — corrigendum. Probability in the Engineering and Informational Sciences, 35(4):1005, October 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/algebraic-
reliability-of-multistate-koutofn-systemscorrigendum/
59CEFF878AE1FCA5E9B731AE4F611E42. See [POdCW21].

Anonymous:2021:PVIb

[Ano21b] Anonymous. PES volume 35 issue 1 cover and back matter. Probability in the Engineering and Informational Sciences, 35(1):b1–b2, January 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/pes-
volume-35-issue-1-cover-and-back-matter/E5630D293797E6B11DB975EFD4145503.

Anonymous:2021:PVIa

[Ano21c] Anonymous. PES volume 35 issue 1 cover and front matter. Probability in the Engineering and Informational Sciences, 35(1):f1–f2, January 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-

engineering-and-informational-sciences/article/pes-
volume-35-issue-1-cover-and-front-matter/16E570B09B928262EFF8B256565D0750.

Anonymous:2021:PVId

[Ano21d] Anonymous. PES volume 35 issue 2 cover and back matter. Probability in the Engineering and Informational Sciences, 35(2):b1–b2, April 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/pes-
volume-35-issue-2-cover-and-back-matter/0484F71342CFFA3DB9B3C727581CD1EA.

Anonymous:2021:PVIc

[Ano21e] Anonymous. PES volume 35 issue 2 cover and front matter. Probability in the Engineering and Informational Sciences, 35(2):f1–f2, April 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/pes-
volume-35-issue-2-cover-and-front-matter/1C0DA405D536AF57E62A31F45839EBEC.

Anonymous:2021:PVIf

[Ano21f] Anonymous. PES volume 35 issue 4 cover and back matter. Probability in the Engineering and Informational Sciences, 35(4):b1–b2, October 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/pes-
volume-35-issue-4-cover-and-back-matter/921E0D9CE2CA07C5E3C6C8991B37DA8C.

Anonymous:2021:PVIe

[Ano21g] Anonymous. PES volume 35 issue 4 cover and front matter. Probability in the Engineering and Informational Sciences, 35(4):f1–f2, October 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/pes-
volume-35-issue-4-cover-and-front-matter/8B0DA766BBDE650B74D7CA70307CC4C5.

Anonymous:2022:PVIb

[Ano22a] Anonymous. PES volume 36 issue 1 cover and back matter. Probability in the Engineering and Informational Sciences, 36(1):b1–b2, January 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/pes-
volume-36-issue-1-cover-and-back-matter/31F8D0238146ED9B8F959C23BEB4CB8C.

Anonymous:2022:PVIa

[Ano22b] Anonymous. PES volume 36 issue 1 cover and front matter. Probability in the Engineering and Informational Sciences, 36(1):f1–f2, January 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/pes-
volume-36-issue-1-cover-and-front-matter/E23381E506E74888F25482DEDB9BB91A.

Anonymous:2022:PVIc

[Ano22c] Anonymous. PES volume 36 issue 2 cover and front matter. Probability in the Engineering and Informational Sciences, 36(2):f1–f2, April 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/pes-
volume-36-issue-2-cover-and-front-matter/0B889BCC7296752B0EE298B8CF08A56A.

Anonymous:2022:PVIe

[Ano22d] Anonymous. PES volume 36 issue 3 cover and back matter. Probability in the Engineering and Informational Sciences, 36(3):b1–b2, July 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/pes-
volume-36-issue-3-cover-and-back-matter/2ACB68FF45552AD55298EA1DC75D09AD.

Anonymous:2022:PVId

[Ano22e] Anonymous. PES volume 36 issue 3 cover and front matter. Probability in the Engineering and Informational Sciences, 36(3):f1–f2, July 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/pes-
volume-36-issue-3-cover-and-front-matter/BAFFBFC59EF21FA1AB80F5223CE5874D.

Anonymous:2022:PVIg

[Ano22f] Anonymous. PES volume 36 issue 4 cover and back matter. Probability in the Engineering and Informational Sciences, 36(4):b1–b2, October 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/pes-
volume-36-issue-4-cover-and-back-matter/A9A22B2670C0F13BB3A0235B2D08010D.

Anonymous:2022:PVIf

[Ano22g] Anonymous. PES volume 36 issue 4 cover and front matter. Probability in the Engineering and Informational Sciences, 36(4):f1–f2, October 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/pes-
volume-36-issue-4-cover-and-front-matter/C6F36A3C51025939E3505F7C37FDA238.

Anonymous:2023:PVI

[Ano23a] Anonymous. PES volume 37 issue 1 cover. Probability in the Engineering and Informational Sciences, 37(1):f1, January 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/pes-volume-37-issue-1-cover/E8D2E71E87DD66ACFFAA3C0A4E83195C.

Anonymous:2023:P

[Ano23b] Anonymous. Preface. Probability in the Engineering and Informational Sciences, 37:322–323, April 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/preface/
690B63D7849D539366DA2EDC10EAFBF0.

Atali:2012:SMI

[AÖ12] Aykut Atali and Özalp Özer. Stochastic multi-item inventory systems with Markov-modulated demands and production quantity requirements. Probability in the Engineering and Informational Sciences, 26(2):263–293, April 2012. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
A337825C18745FBB8858D3E51F2C27AD.

Arab:2018:IFR

[AO18] Idir Arab and Paulo Eduardo Oliveira. Iterated failure rate monotonicity and ordering relations within gamma and Weibull distributions — corrigendum. Probability in the Engineering and Informational Sciences, 32(4):640–641, October 2018. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/iterated-failure-rate-monotonicity-and-ordering-
relations-within-gamma-and-weibull-distributions-corrigendum/
5BC086888C6C37509B183A779F0A383E. See [AO19].

Arab:2019:IFR

[AO19] Idir Arab and Paulo Eduardo Oliveira. Iterated failure rate monotonicity and ordering relations within gamma and Weibull distributions. Probability in the Engineering and Informational Sciences, 33(1):64–80, January 2019. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/iterated-
failure-rate-monotonicity-and-ordering-relations-within-
gamma-and-weibull-distributions/B2FF4F2927A1B7C2117645609B88563C.

See corrigendum [AO18].

Agustin:1999:DCR

[AP99] Marcus A. Agustin and Edsel A. Peña. A dynamic competing risks model. Probability in the Engineering and Informational Sciences, 13(3):333–358, July 1999. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
12E23BCE543BEF58B8DF7BD898C5E140.

Ayesta:2020:SSS

[APR20] Urtzi Ayesta, Balakrishna Prabhu, and Rhonda Righter. Scheduling in a single-server queue with state-dependent service rates. Probability in the Engineering and Informational Sciences, 34(4):507–521, October 2020. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/scheduling-
in-a-singleserver-queue-with-statedependent-service-
rates/DAE6F8C2CB91E45FCF28D0FD24D117F6.

Adler:1997:DTF

[AR97] Ilan Adler and Sheldon M. Ross. Distribution of the time of the first k-record. Probability in the Engineering and Informational Sciences, 11(3):273–278, July 1997.

CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
9C7D58B1F7E19D5AEFAFDAE49C705BBD.

Alfa:2000:SVT

[AR00] Attahiru Sule Alfa and T. S. S. Srinivasa Rao. Supplementary variable technique in stochastic models. Probability in the Engineering and Informational Sciences, 14(2):203–218, April 2000. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
8D477603E117B6051DCF814359C342D8.

Aldous:2014:ESM

[AR14] David J. Aldous and Nathan Ross. Entropy of some models of sparse random graphs with vertex-names. Probability in the Engineering and Informational Sciences, 28(2):145–168, April 2014. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
613DF900D75BA3DF43A3E6B0D102A124.

Adan:2003:SDL

[ARK03] I. J. B. F. Adan, J. A. C. Resing, and V. G. Kulkarni. Stochastic discretization for the long-run average reward in fluid models. Probability in the Engineering and Informational Sciences, 17(2):251–265, April 2003. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
998EB27A5A94EF594D875B603D051A36.

Arrow:1987:DID

[Arr87] Kenneth J. Arrow. The demand for information and the distribution of income. Probability in the Engineering and Informational Sciences, 1(1):3–13, January 1987. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
C1DC68B0103ABEB0213D3CFFD04121B7.

Asmussen:1996:MSR

[AS96] Søren Asmussen and Karl Sigman. Monotone stochastic recursions and their duals. Probability in the Engineering and Informational Sciences, 10(1):1–20, January 1996. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
7642CDAFDAC130D3D31AE92C5AFA034A.

Ath:2000:CCU

[AS00a] Yontha Ath and Milton Sobel. Counterexamples to conjectures for uniformly optimally reliable graphs. Probability in the Engineering and Informational Sciences, 14(2):173–177, April 2000. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
FB6D8FFC8EEAFE8A7AA423E51BB8BB48.

Ath:2000:SCU

[AS00b] Yontha Ath and Milton Sobel. Some conjectured uniformly optimal reliable networks. Probability in the Engineering and Informational Sciences, 14(3):375–383, July 2000. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
25E30E8D2954612119F509FF2EE0CFC2.

Aurzada:2018:IPP

[AS18] Frank Aurzada and Sebastian Schwinn. Improving the performance of polling models using forced idle times. Probability in the Engineering and Informational Sciences, 32(4):580–602, October 2018. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/improving-
the-performance-of-polling-models-using-forced-idle-
times/58B243323B57B4EABC79EA25BB3F1AF6.

Amini-Seresht:2021:RPL

[ASKZ21] Ebrahim Amini-Seresht, Maryam Kelkinnama, and Yiying Zhang. On the residual and past lifetimes of coherent systems under random monitoring. Probability in the Engineering and Informational Sciences, 35(3):465–480, July 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/on-the-residual-and-past-lifetimes-of-coherent-
systems-under-random-monitoring/53FFA20629858F894AE2D86923580CB2.

Aslan:2022:OAR

[Asl22] Ayse Aslan. Optimal admission and routing with congestionsensitive customer classes. Probability in the Engineering and Informational Sciences, 36(3):774–798, July 2022.

CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/optimal-admission-and-routing-with-congestionsensitive-
customer-classes/B0E025784505E19AE79208D5DEDBA0FA.

Amini-Seresht:2017:SMC

[ASZ17] Ebrahim Amini-Seresht and Yiying Zhang. Stochastic monotonicity of conditional order statistics in multipleoutlier scale population. Probability in the Engineering and Informational Sciences, 31(3):366–380, July 2017. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
6041AF618C34568974580FDFD40D1A07.

Adan:1995:FDD

[AvER95] Ivo Adan, Michel van Eenige, and Jacques Resing. Fitting discrete distributions on the first two moments. Probability in the Engineering and Informational Sciences, 9(4):623–632, October 1995. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/D3A59E8257A551724DB6DF2B1E505BA4.

Abate:1989:STS

[AW89] Joseph Abate and Ward Whitt. Spectral theory for skipfree Markov chains. Probability in the Engineering and Informational Sciences, 3(1):77–88, January 1989. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
1F773D5C8AB5F739895B747991F11A75.

Abate:1995:LAB

[AW95] Joseph Abate and Ward Whitt. Limits and approximations for the busy-period distribution in single-server queues. Probability in the Engineering and Informational Sciences, 9(4):581–602, October 1995. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/5A2DD3622BF70819E8DAE2C16CF08557.

Asmussen:1996:VRS

[AW96] Søren Asmussen and Chia-Li Wang. Variance reduction for simulating transient GI/G/1 behavior. Probability in the Engineering and Informational Sciences, 10(2):197–205, April 1996. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
E4DEC14CBCCF819049FD1F09E1A8475F.

Adan:2005:TNJ

[AW05] Ivo Adan and Gideon Weiss. A two-node Jackson network with infinite supply of work. Probability in the Engineering and Informational Sciences, 19(2):191–212, April 2005. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
C5972AAEA343EABA33F1AE37C55AE1C0.

Adan:2012:LSS

[AW12] Ivo Adan and Gideon Weiss. A loss system with skill-based servers under assign to longest idle server policy. Probability in the Engineering and Informational Sciences, 26(3):307– 321, July 2012. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/A111B87EB059E1DE940006F74A6C068A.

Altman:1994:PCN

[AY94] Eltan Altman and Uri Yechiali. Polling in a closed network. Probability in the Engineering and Informational Sciences, 8 (3):327–343, July 1994. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/D5C9C47E3E0FD307BA1A3E40C4BD26B3.

Altman:2008:ISQ

[AY08a] Eitan Altman and Uri Yechiali. Infinite-server queues with system’s additional tasks and impatient customers. Probability in the Engineering and Informational Sciences, 22(4):477–493, October 2008. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/DDFF74549DA02372128A9EDDFD24DFE3.

Avrachenkov:2008:RNF

[AY08b] Kostia Avrachenkov and Uri Yechiali. Retrial networks with finite buffers and their application to Internet data traffic. Probability in the Engineering and Informational Sciences, 22(4): 519–536, October 2008. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/D8AA783C93ADBCB08C4F81C609AE5675.

Argon:2008:SIJ

[AZR08] Nilay Tanik Argon, Serhan Ziya, and Rhonda Righter. Scheduling impatient jobs in a clearing system with insights on patient triage in mass casualty incidents. Probability in the Engineering and Informational Sciences, 22(3):301–332, July 2008. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
E0922CAC56E1479BA0F74C5DEE3B0F1D.

Benkherouf:2002:SIM

[BA02] Lakdere Benkherouf and Lakhdar Aggoun. On a stochastic inventory model with deterioration and stock-dependent demand items. Probability in the Engineering and Informational Sciences, 16(2):151–165, April 2002. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
D64D6980DACDAADD33FF09C8A7641225.

Bi:2018:EAN

[BA18] Huibo Bi and Omer H. Abdelrahman. Energy-aware navigation in large-scale evacuation using G-networks. Probability in the Engineering and Informational Sciences, 32(3):340–352, July 2018. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/energyaware-navigation-in-largescale-evacuation-
using-gnetworks/14A3C6BCD86377350EAA3021C9F82C77.

Balakrishnan:1992:TEP

[Bal92] A. V. Balakrishnan. Two examples of physically realizable processes whose spectral density is not “physically realizable”. Probability in the Engineering and Informational Sciences, 6 (2):257–259, April 1992. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/CCB2AE75B0833727BC9BB6014AC8E411.

Banjevic:1992:RSR

[Ban92] Dragan Banjević. On the recurrent solution of renewaltype integral equations. Probability in the Engineering and Informational Sciences, 6(2):261–270, April 1992. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
8FF2D03EA6DD0F10255BA6857CAF4A9F.

Bardhan:1996:DLO

[Bar96] Indrajit Bardhan. Diffusion limits for open networks of finite-buffer queues. Probability in the Engineering and Informational Sciences, 10(3):341–361, July 1996. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
235A0CC9B2C5F27F15282B5E4FD5ECF2.

Baron:2008:LWT

[Bar08] Opher Baron. On the law of the i th waiting time in a busy period of G/M/c queues. Probability in the Engineering and Informational Sciences, 22(1):75–80, January 2008. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
D94123BF5379C7852FCFCAE83DBD3FB2.

Boon:2012:DSI

[BAWD12] M. A. A. Boon, I. J. B. F. Adan, E. M. M. Winands, and D. G. Down. Delays at signalized intersections with exhaustive traffic control. Probability in the Engineering and Informational Sciences, 26(3):337–373, July 2012. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
F9D188A9084195A1B569991DF038AE39.

Baxter:1996:BRC

[Bax96] Laurence A. Baxter. Book review: Chaos, Fractals, and Noise: Stochastic Aspects of Dynamics, Second Edition, 1994, by Lasota Andrzej and Mackey Michael C., Applied Mathematical Sciences, vol. 97 (New York: Springer-Verlag), xiv + 472 pp. Probability in the Engineering and Informational Sciences, 10 (2):311–313, April 1996. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/8E1FEEBAA6B4915BB311D298223C2FED.

Boucherie:1996:WMG

[BB96] Richard J. Boucherie and Onno J. Boxma. The workload in the M/G/1 queue with work removal. Probability in the Engineering and Informational Sciences, 10(2):261–277, April 1996. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
45A3E82575035D8143B7CF652A23DF44.

Bhatnagar:1997:MSA

[BB97] Shalabh Bhatnagar and Vivek S. Borkar. Multiscale stochastic approximation for parametric optimization of hidden Markov models. Probability in the Engineering and Informational Sciences, 11(4):509–522, October 1997. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
815CFCD36E68C226CAB6EE9ECE1FA315.

Bhatnagar:1998:TTS

[BB98] Shalabh Bhatnagar and Vivek S. Borkar. A two timescale stochastic approximation scheme for simulation-based parametric optimization. Probability in the Engineering and Informational Sciences, 12(4):519–531, October 1998. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
3336D0E04552ACC56FB5C00A9F467961.

Bhattacharjee:2000:SEC

[BB00] M. C. Bhattacharjee and R. N. Bhattacharya. Stochastic equivalence of convex ordered distributions and applications. Probability in the Engineering and Informational Sciences, 14(1): 33–48, January 2000. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/C55483F85769A04B49DE822E40EF40E1.

Bayer:2002:SSG

[BB02] Nimrod Bayer and Richard J. Boucherie. On the structure of the space of geometric product-form models. Probability in the Engineering and Informational Sciences, 16(2):241–270, April 2002. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
CCB70B5E26175A7D970CD70A7C469D30.

Bekker:2006:OAC

[BB06] René Bekker and Sem C. Borst. Optimal admission control in queues with workload-dependent service rates. Probability in the Engineering and Informational Sciences, 20(4):543–570, October 2006. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/215C41699880E5D73BD385FE3ED0DD80.

Badia:2001:BAC

[BBCN01] F. G. Badía, M. D. Berrade, C. A. Campos, and M. A. Navascués. On the behavior of aging characteristics in mixed populations. Probability in the Engineering and Informational Sciences, 15(1):83–94, January 2001. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
11818845D5F2E86626BF22BC0562F2D3.

Borst:1997:STP

[BBCP97] Sem Borst, John Bruno, E. G. Coffman, and Steven Phillips. Scheduling two-point stochastic jobs to minimize the makespan on two parallel machines. Probability in the Engineering and Informational Sciences, 11(1):95–105, January 1997. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
924C05817EE4539474DFAFCB9A0AC608.

Balakrishnan:2010:UML

[BBHK10] Narayanaswamy Balakrishnan, Félix Belzunce, Nasrin Hami, and Baha-Eldin Khaledi. Univariate and multivariate likelihood ratio ordering of generalized order statistics and associated conditional variables. Probability in the Engineering and Informational Sciences, 24(3):441–455, July 2010. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
476DFBB24220C9F44E6FD6A13F01656B.

Baron:2020:CRI

[BBP20] Opher Baron, Oded Berman, and David Perry. Continuous review inventory models for perishable items with leadtimes. Probability in the Engineering and Informational Sciences, 34(3):317–342, July 2020. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/continuous-
review-inventory-models-for-perishable-items-with-leadtimes/
F7A9335D61329ACCA9A3FC6A2C31FE11.

Barnsley:1988:MMC

[BBS88] Michael F. Barnsley, Marc A. Berger, and H. Meté Soner. Mixing Markov chains and their images. Probability in the Engineering and Informational Sciences, 2(4):387–414, October 1988. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
7EDD4F649C3AF9B8241F41BA04CF79FD.

Bassan:1996:ICT

[BBS96] Bruno Bassan, Monica Brezzi, and Marco Scarsini. Information in continuous time decision models with many agents. Probability in the Engineering and Informational Sciences, 10(4): 543–555, October 1996. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/DCD1F9822B822958169864A9C7EA2670.

Boucherie:1997:NNC

[BBS97] Richard J. Boucherie, Onno J. Boxma, and Karl Sigman. A note on negative customers, GI/G/1 workload, and risk processes. Probability in the Engineering and Informational Sciences, 11(3):305–311, July 1997. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
13A9ADF3B3054CFD7229063967237A19.

Baccelli:1991:FSM

[BBW91] F. Baccelli, N. Bambos, and J. Walrand. On flows in stochastic marked graphs. Probability in the Engineering and Informational Sciences, 5(2):145–157, April 1991. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
8B7F0F8794333267F3DBFEEB86985C0B.

Bambos:1994:OAG

[BC94] Nicholas Bambos and Shou C. Chen. Optimality aspects of greedy schemes in parallel processing of random graph-structured jobs. Probability in the Engineering and Informational Sciences, 8(2):229–243, April 1994. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
F3C1AB13645A2BA88BEEB226CA6433F9.

Bassan:1998:OSP

[BC98] Bruno Bassan and Claudia Ceci. An optimal stopping problem arising from a decision model with many agents. Probability in the Engineering and Informational Sciences, 12(3): 393–408, July 1998. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/8E344FB40A64EF382DC0AFA4451DD173.

Badia:2013:PPR

[BC13] F. G. Badía and Ji Hwan Cha. Preservation properties of a renewal process stopped at a random dependent time. Probability in the Engineering and Informational Sciences, 27(2): 163–175, April 2013. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/E3761FCDA17381D3E2EBDE361C296144.

Brown:2018:SCV

[BC18] Mark Brown and Joel E. Cohen. Squared coefficient of variation of Taylor’s law for random absolute differences. Probability in the Engineering and Informational Sciences, 32(4):483–494, October 2018. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/squared-
coefficient-of-variation-of-taylors-law-for-random-
absolute-differences/103A2A6CE47BAA0D9E87ADB68AC15F1A.

Bruno:1995:SIT

[BCD95] John Bruno, Edward G. Coffman, and Peter Downey. Scheduling independent tasks to minimize the makespan on identical machines. Probability in the Engineering and Informational Sciences, 9(3):447–456, July 1995. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
EBED69AF731D4756C195566852C39C17.

Borgs:2014:OAT

[BCD^{+}14] Christian Borgs, Jennifer T. Chayes, Sherwin Doroudi, Mor Harchol-Balter, and Kuang Xu. The optimal admission threshold in observable queues with state dependent pricing. Probability in the Engineering and Informational Sciences, 28(1): 101–119, January 2014. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/1F641237B90462375C990D81B53A5E10.

Baccelli:1989:TCQ

[BCG89] F. Baccelli, E. G. Coffman, and E. N. Gilbert. Tandem conveyor queues. Probability in the Engineering and Informational Sciences, 3(4):517–536, October 1989. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
376B29219B1759414059E74E8C23CA3E.

Browne:1992:GEP

[BCGW92] Sid Browne, E. G. Coffman, Jr., E. N. Gilbert, and Paul E. Wright. Gated, exhaustive, parallel service. Probability in the Engineering and Informational Sciences, 6(2):217–239, April 1992. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
E9842E72C272280B49BE6CA4C400948E.

Balakrishnan:2023:SSC

[BCMS23] Narayanaswamy Balakrishnan, Antonia Castaño-Martínez, and Miguel A. Sordo. Some stochastic comparisons of lower records and lower record spacings. Probability in the Engineering and Informational Sciences, 37(1):192–205, January 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/some-stochastic-comparisons-of-lower-records-
and-lower-record-spacings/65A89A655D32657498D94AB78DBDF78A.

Barquin:2005:SME

[BCR05] J. Barquín, E. Centeno, and J. Reneses. Stochastic market equilibrium model for generation planning. Probability in the Engineering and Informational Sciences, 19(4):533–546, October 2005. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/48862CED10E3897D492A1350A329D838.

Block:1992:PTP

[BCS92] Hentry W. Block, Tim Costigan, and Allan R. Sampson. Product-type probability bounds of higher order. Probability in the Engineering and Informational Sciences, 6(3):349– 370, July 1992. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/15ECFC57E4DD7460C7B59CD1BD4047B8.

Balakrishnan:2019:CSR

[BCZZ19] Narayanaswamy Balakrishnan, Jianbin Chen, Yiying Zhang, and Peng Zhao. Comparisons of sample ranges arising from multiple-outlier models: in memory of Moshe Shaked. Probability in the Engineering and Informational Sciences, 33(1):28–49, January 2019. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/comparisons-
of-sample-ranges-arising-from-multipleoutlier-models-
in-memory-of-moshe-shaked/B8E83F105AA5E762973FAE174C7FB6EB.

Buijsrogge:2020:SII

[BdBS20] Anne Buijsrogge, Pieter-Tjerk de Boer, and Werner R. W. Scheinhardt. On state-independent importance sampling for the GI—GI—1 tandem queue 1. Probability in the Engineering and Informational Sciences, 34(1):131–156, January 2020. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/on-stateindependent-importance-sampling-for-
the-gigi1-tandem-queue1/2A8A04A870BE1EC7DFE567CF180D3CB1.

Brown:2015:BCN

[BdlPS15] Mark Brown, Victor de la Peña, and Tony Sit. From boundary crossing of non-random functions to boundary crossing of stochastic processes. Probability in the Engineering and Informational Sciences, 29(3):345–359, July 2015. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
FE0F0BD42DE51FC14853C23DD6797A1C.

Blom:2015:AMM

[BDM15] Joke Blom, Koen De Turck, and Michel Mandjes. Analysis of Markov-modulated infinite-server queues in the central-limit regime. Probability in the Engineering and Informational Sciences, 29(3):433–459, July 2015. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
E48B00709AB74015E40316BB468F5C70.

Boxma:2011:NLO

[BDPS11] Onno J. Boxma, Israel David, David Perry, and Wolfgang Stadje. A new look at organ transplantation models and double matching queues. Probability in the Engineering and Informational Sciences, 25(2):135–155, April 2011. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
872CA4377C9B70F3638D30B38EDAB0C2.

Bauerle:2004:RAT

[BEFK04] N. Bäuerle, O. Engelhardt-Funke, and M. Kolonko. Routing of airplanes to two runways: monotonicity of optimal controls. Probability in the Engineering and Informational Sciences, 18(4):533–560, October 2004. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
D545C623590B2F590E066CFD0AEA3487.

Belbachir:2011:DMC

[Bel11] Hacène Belbachir. Determining the mode for convolution powers of discrete uniform distribution. Probability in the Engineering and Informational Sciences, 25(4):469–475, October 2011. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
80608BC00D756A04A3CF2A8232D19511.

Benkherouf:1990:OSO

[Ben90] Lakdere Benkherouf. Optimal stopping in oil exploration with small and large oilfields. Probability in the Engineering and Informational Sciences, 4(3):399–411, July 1990. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
46EBB24FF717D78567BB27CB5E963C84.

Benkherouf:1992:OER

[Ben92] Lakdére Benkherouf. Oil exploration: Revisited. Probability in the Engineering and Informational Sciences, 6(2):241–243, April 1992. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/7875094E2C14C359B702439E5A2C89F9.

Boland:1988:ARA

[BEP88] Philip J. Boland, Emad El Neweihi, and Frank Proschan. Active redundancy allocation in coherent systems. Probability in the Engineering and Informational Sciences, 2(3):343–353, July 1988. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/31F231F7204AE93B4990EBD915FEF422.

Berry:1996:DSI

[Ber96] Scott Berry. Discrete search for an intelligent object: The Leprechaun’s problem. Probability in the Engineering and Informational Sciences, 10(3):363–375, July 1996. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
3785408D9F7D45D5A1A3CFAB2216AE27.

Barone:1990:ISR

[BF90] Piero Barone and Arnolodo Frigessi. Improving stochastic relaxation for Gaussian random fields. Probability in the Engineering and Informational Sciences, 4(3):369–389, July 1990. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
E726DFF8E43C4C08B3D4CCC139B25ECB.

Brown:1991:CSG

[BF91] Mark Brown and Marcia H. Flicker. Chi-square goodness of fit: A failure rate perspective. Probability in the Engineering and Informational Sciences, 5(3):273–284, July 1991. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
0BA56376732103179F008F838C3F1EB4.

Belzunce:1999:APB

[BFR99] Félix Belzunce, Manuel Franco, and José M. Ruiz. On aging properties based on the residual life of k-out-of-n systems. Probability in the Engineering and Informational Sciences, 13 (2):193–199, April 1999. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/EA6BEA27D5836BEAE20F2905ACF62F46.

Belzunce:2001:POB

[BFRR01] Félix Belzunce, Manuel Franco, José-Maria Ruiz, and M. Carmen Ruiz. On partial orderings between coherent systems with different structures. Probability in the Engineering and Informational Sciences, 15(2):273–293, April 2001. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
9BD8944C3C8E117980D8BB77C14FDEB8.

Berenguer:2003:MPC

[BGDR03] C. Bérenguer, A. Grall, L. Dieulle, and M. Roussignol. Maintenance policy for a continuously monitored deteriorating system. Probability in the Engineering and Informational Sciences, 17(2):235–250, April 2003. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
FF9A55832F6E808EE1090BAB3776311E.

Baykal-Gursoy:2007:SMD

[BGG07] M. Baykal-Gürsoy and K. Gürsoy. Semi-Markov decision processes: nonstandard criteria. Probability in the Engineering and Informational Sciences, 21(4):635–657, October 2007. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
A6F36E6388AC5E9687E08821503D9237.

Belzunce:2011:RML

[BGR11] Félix Belzunce, Selma Gurler, and José M. Ruiz. Revisiting multivariate likelihood ratio ordering results for order statistics. Probability in the Engineering and Informational Sciences, 25(3):355–368, July 2011. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
E271E4DAEAC97DAAAE76086AF3C4AD5A.

Bauerle:2002:BPL

[BH02] Nicole Bäuerle and Anja Houdek. Bounds and performance limits of channel assignment policies in cellular networks. Probability in the Engineering and Informational Sciences, 16(1): 85–100, January 2002. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/90C97F1587FA19BD891D58E3DCC370F9.

Belyaev:2021:KS

[BH21] Yu. K. Belyaev and Asaf H. Hajiyev. Kolmogorov stories. Probability in the Engineering and Informational Sciences, 35(3):355–368, July 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/kolmogorov-
stories/5481B504635D7F14034AC1215D15D377.

Brill:2023:AAF

[BH23] Percy H. Brill and Mei Ling Huang. On approximation of the analytic fixed finite time large t probability distributions in an extreme renewal process with no-mean interrenewals. Probability in the Engineering and Informational Sciences, 37(3):695–710, July 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/on-approximation-
of-the-analytic-fixed-finite-time-large-t-probability-
distributions-in-an-extreme-renewal-process-with-nomean-
interrenewals/F39B0B19B1D3EDD13CFE2752AB64E96B.

Bhattacharjee:1987:TEB

[Bha87] M. C. Bhattacharjee. The time to extinction of branching processes and log-convexity: I. Probability in the Engineering and Informational Sciences, 1(3):265–278, July 1987. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
70B1257C0B9847211583472402B34707.

Bhattacharjee:1999:EWC

[Bha99] M. C. Bhattacharjee. Exponentiality within class L and stochastic equivalence of Laplace ordered survival times. Probability in the Engineering and Informational Sciences, 13(2): 201–207, April 1999. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/045834B40D5461931ADF329002CC12FF.

Budhiraja:2004:SDE

[BHCKS04] A. Budhiraja, F. Hernández-Campos, V. G. Kulkarni, and F. D. Smith. Stochastic differential equation for TCP window size: analysis and experimental validation. Probability in the Engineering and Informational Sciences, 18(1):111–140, January 2004. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/2F0F662FF95E25E18810D06AF6CE4D16.

Belzunce:2003:DTV

[BHK03] Félix Belzunce, Taizhong Hu, and Baha-Eldin Khaledi. Dispersion-type variability orders. Probability in the Engineering and Informational Sciences, 17(3):305–334, July 2003. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
4713B2792BA9A3B2F5094D754A4F3D8A.

Buke:2009:MMM

[BHM09] Burak Büke, John J. Hasenbein, and David P. Morton. Minimizing makespan in a multiclass fluid network with parameter uncertainty. Probability in the Engineering and Informational Sciences, 23(3):457–480, July 2009. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
787E15DB8410BCE0E073CCBD9CFB94A4.

Barrera:2005:SBP

[BHP05] Javiera Barrera, Thierry Huillet, and Christian Paroissin. Size-biased permutation of Dirichlet partitions and searchcost distribution. Probability in the Engineering and Informational Sciences, 19(1):83–97, January 2005. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
E7093464F2376FAD025F1B7E8E2A41F5.

Bhulai:2009:DRP

[Bhu09] Sandjai Bhulai. Dynamic routing policies for multiskill call centers. Probability in the Engineering and Informational Sciences, 23(1):101–119, January 2009. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
51D1BB1DB11063C2E214789843D1D2A8.

Barlow:1988:CCC

[BI88] R. E. Barlow and S. Iyer. Computational complexity of coherent systems and the reliability polynomial. Probability in the Engineering and Informational Sciences, 2(4):461–469, October 1988. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/539CBA5F2F376D249269157E90B6D59A.

Baharian:2013:SSA

[BJ13] Golshid Baharian and Sheldon H. Jacobson. Stochastic sequential assignment problem with threshold criteria. Probability in the Engineering and Informational Sciences, 27(3): 277–296, July 2013. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/B74A8B43996B127779E7385CE0CD71C1.

Berenhaut:2019:FPW

[BJ19] Kenneth S. Berenhaut and Hongyi Jiang. The friendship paradox for weighted and directed networks. Probability in the Engineering and Informational Sciences, 33(1):136–145, January 2019. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/friendship-paradox-for-weighted-and-directed-
networks/5A489320AE3DDE214D136DC17E922E52.

Boonen:2023:POR

[BJ23] Tim J. Boonen and Wenjun Jiang. Pareto-optimal reinsurance with default risk and solvency regulation. Probability in the Engineering and Informational Sciences, 37:518–545, April 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/paretooptimal-reinsurance-with-default-risk-
and-solvency-regulation/EDF2F0B2BC6C811C592993630A44DC55.

Block:1999:BFB

[BJS99] Henry W. Block, Yi-Kuan Jong, and Thomas H. Savits. Bathtub functions and burn-in. Probability in the Engineering and Informational Sciences, 13(4):497–507, October 1999. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
9E452EB2C035ED7DCF51AF5AAFFC00ED.

Burnetas:1993:STT

[BK93] Apostolos N. Burnetas and Michael N. Katehakis. On sequencing two types of tasks on a single processor under incomplete information. Probability in the Engineering and Informational Sciences, 7(1):85–119, January 1993. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
A4C92A4056634F2E90502DAA46EEE687.

Burnetas:1995:COP

[BK95] Apostolos N. Burnetas and Michael N. Katehakis. Computing optimal policies for Markovian decision processes using simulation. Probability in the Engineering and Informational Sciences, 9(4):525–537, October 1995. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
4520575062290B0DF16463BABE0C3498.

Boutsikas:2003:BDT

[BK03a] Michael V. Boutsikas and Markos V. Koutras. Bounds for the distribution of two-dimensional binary scan statistics. Probability in the Engineering and Informational Sciences, 17(4): 509–525, October 2003. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/05426543C113DBC21C67215BBF1B788A.

Burnetas:2003:ABA

[BK03b] Apostolos N. Burnetas and Michael N. Katehakis. Asymptotic Bayes analysis for the finite-horizon one-armed-bandit problem. Probability in the Engineering and Informational Sciences, 17(1):53–82, January 2003. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/4148BEF7EE4DCCAFE92DE0F564C03AA8.

Buczkowski:2006:FWR

[BK06] Peter S. Buczkowski and Vidyadhar G. Kulkarni. Funding a warranty reserve with contributions after each sale. Probability in the Engineering and Informational Sciences, 20(3): 497–515, July 2006. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/73461CE5D28B1F477F16520890FD5EA2.

Brown:2009:STS

[BK09] Mark Brown and J. H. B. Kemperman. Sharp two-sided bounds for distributions under a hazard rate constraint. Probability in the Engineering and Informational Sciences, 23(1): 31–36, January 2009. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/852E7358B7928679BB992CD1D1AEADEB.

Barmalzan:2022:OFM

[BKB22] Ghobad Barmalzan, Sajad Kosari, and Narayanaswamy Balakrishnan. Orderings of finite mixture models with locationscale distributed components. Probability in the Engineering and Informational Sciences, 36(2):461–481, April 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/

article/orderings-of-finite-mixture-models-with-locationscale-
distributed-components/5477A3C88246BD0E2A545B3234BDF2C3.

Burnetas:2017:AOM

[BKK17] Apostolos Burnetas, Odysseas Kanavetas, and Michael N. Katehakis. Asymptotically optimal multi-armed bandit policies under a cost constraint. Probability in the Engineering and Informational Sciences, 31(3):284–310, July 2017. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
9605D61DC502C6705CF6C8CCE06EF08C.

Boxma:2005:SSS

[BKKP05] Onno Boxma, Haya Kaspi, Offer Kella, and David Perry. On/off storage systems with state-dependent input, output, and switching rates. Probability in the Engineering and Informational Sciences, 19(1):1–14, January 2005. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
A4488E39395F23D5C0391B087FFDB30F.

Bloznelis:2022:ABD

[BKL22] Mindaugas Bloznelis, Joona Karjalainen, and Lasse Leskelä. Assortativity and bidegree distributions on Bernoulli random graph superpositions. Probability in the Engineering and Informational Sciences, 36(4):1188–1213, October 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/assortativity-and-bidegree-distributions-on-
bernoulli-random-graph-superpositions/2C9BFA1C0DFEDA8572661EADFD0DDE36.

Boxma:2010:GCL

[BKM10] Onno Boxma, Offer Kella, and Michel Mandjes. On a generic class of lévy-driven vacation models. Probability in the Engineering and Informational Sciences, 24(1):1–12, January 2010. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
72204341E855C217FF26C18D87029A92.

Boxma:2001:IFS

[BKP01] Onno Boxma, Offer Kella, and David Perry. An intermittent fluid system with exponential on-times and semi-Markov input rates. Probability in the Engineering and Informational Sciences, 15(2):189–198, April 2001. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
A7F117FEA66CCCA4C81D054B2682BAFC.

Baxter:1987:DMS

[BL87a] Laurence A. Baxter and Eui Yong Lee. A diffusion model for a system subject to continuous wear. Probability in the Engineering and Informational Sciences, 1(4):405–416, October 1987. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
B2D7C5C84F5A83EFFFE0B4378C45A074.

Baxter:1987:ICD

[BL87b] Laurence A. Baxter and Eui Yong Lee. An inventory with constant demand and Poisson restocking. Probability in the Engineering and Informational Sciences, 1(2):203–210, April 1987. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
F4D5D64D4773DD224DFE34B1BD728986.

Baxter:1988:OCM

[BL88] Laurence A. Baxter and Eui Yong Lee. Optimal control of a model for a system subject to continuous wear. Probability in the Engineering and Informational Sciences, 2(3):321– 328, July 1988. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/28890FD43AC8FA8F3862A3023E539773.

Baxter:1989:FPR

[BL89] Laurence A. Baxter and Seung Min Lee. Further properties of reliability importance for continuum structure functions. Probability in the Engineering and Informational Sciences, 3(2): 237–246, April 1989. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/33220988C09F52893EC54B92FE5269C1.

Berenhaut:2002:RCR

[BL02] Kenneth S. Berenhaut and Robert Lund. Renewal convergence rates for DHR and NWU lifetimes. Probability in the Engineering and Informational Sciences, 16(1):67–84, January 2002. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
F2CA1D939434DB7845C0F8A73BE848F3.

Blackwell:1995:MVB

[Bla95] D. Blackwell. Minimax vs. Bayes prediction. Probability in the Engineering and Informational Sciences, 9(1):53–58, January 1995. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
CED0C85AA55766F495849934D2792877.

Blanc:1998:PSA

[Bla98] J. P. C. Blanc. The power-series algorithm for polling systems with time limits. Probability in the Engineering and Informational Sciences, 12(2):221–237, April 1998. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
F305A021829B314E1CF5D8F14E782C94.

Bar-Lev:2017:AOB

[BLBPV17] Shaul K. Bar-Lev, Onno Boxma, David Perry, and Lazaros P. Vastazos. Analysis and optimization of blood-testing procedures. Probability in the Engineering and Informational Sciences, 31(3):330–344, July 2017. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/989632B903CFFA90B48F04F5E887170C.

Bar-Lev:1993:TSR

[BLP93] Shaul K. Bar-Lev and David Perry. Two-stage release rule procedure in a regenerative dam. Probability in the Engineering and Informational Sciences, 7(4):571–588, October 1993. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
DFF43F8C5D6FB8702B6F77C30583DEE8.

Bar-Lev:2005:CPI

[BLPS05] Shaul K. Bar-Lev, David Perry, and Wolfgang Stadje. Control policies for inventory systems with perishable items: outsourcing and urgency classes. Probability in the Engineering and Informational Sciences, 19(3):309–326, July 2005. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
8F7AB3F9900CFD8999588C0FD96679CC.

Belzunce:2001:SCN

[BLRS01] Félix Belzunce, Rosa E. Lillo, José-Maria Ruiz, and Moshe Shaked. Stochastic comparisons of nonhomogeneous processes.

Probability in the Engineering and Informational Sciences, 15 (2):199–224, April 2001. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/71A37AFAF50C02B49760B4DF947D748A.

Block:2003:PPU

[BLS03] Henry W. Block, Yulin Li, and Thomas H. Savits. Preservation of properties under mixture. Probability in the Engineering and Informational Sciences, 17(2):205–212, April 2003. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
B50BA4A7EB4901AA9EF75FBC8E906BCB.

Block:2012:MEI

[BLS12] Henry W. Block, Naftali A. Langberg, and Thomas H. Savits. A mixture of exponential and IFR gamma distributions having an upsidedown bathtub-shaped failure rate. Probability in the Engineering and Informational Sciences, 26(4):573–580, October 2012. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/F9D8A208C069236812B621E561E719C1.

Bar-Lev:2003:HGT

[BLSV03] Shaul K. Bar-Lev, Wolfgang Stadje, and Frank A. Van der Duyn Schouten. Hypergeometric group testing with incomplete information. Probability in the Engineering and Informational Sciences, 17(3):335–350, July 2003. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
AF95830B68D25DB24BD479EE5CA4F9C5.

Bambos:2002:PQR

[BM02a] Nicholas Bambos and George Michailidis. On parallel queuing with random server connectivity and routing constraints. Probability in the Engineering and Informational Sciences, 16(2): 185–203, April 2002. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/B99B35DF49689F9CA4C8CB734AE0D120.

Boots:2002:FSQ

[BM02b] Nam Kyoo Boots and Michel Mandjes. Fast simulation of a queue fed by a superposition of many (heavy-tailed) sources. Probability in the Engineering and Informational Sciences, 16 (2):205–232, April 2002. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/662FD45FB20F27F8308CB3179CF9F564.

Blom:2011:TGS

[BM11] Joke Blom and Michel Mandjes. Traffic generated by a semi-Markov additive process. Probability in the Engineering and Informational Sciences, 25(1):21–27, January 2011. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
DC7CE4204ED03B4A3F40CCC095125622.

Boxma:2008:QMS

[BMK08] Onno Boxma, Michel Mandjes, and Offer Kella. On a queuing model with service interruptions. Probability in the Engineering and Informational Sciences, 22(4):537–555, October 2008. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
8AEA783E88AC4B135CD8F5CE63523CE1.

Belzunce:2005:SCG

[BMR05a] Félix Belzunce, José-Angel Mercader, and José-María Ruiz. Stochastic comparisons of generalized order statistics. Probability in the Engineering and Informational Sciences, 19(1): 99–120, January 2005. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/976A324BE18F44411B8D2AB9CD281F59.

Block:2005:NGS

[BMR05b] Henry Block, Mainak Mazumdar, and Sheldon M. Ross. A note on generating systems composed of binary and ternary random variables. Probability in the Engineering and Informational Sciences, 19(2):213–217, April 2005. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
9099079A4DF62ECC460B05142E52568B.

Bikker:2020:OCP

[BMSB20] Ingeborg A. Bikker, Martijn R. K. Mes, Antoine Sauré, and Richard J. Boucherie. Online capacity planning for rehabilitation treatments: an approximate dynamic programming approach. Probability in the Engineering and Informational Sciences, 34(3):381–405, July 2020. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/online-
capacity-planning-for-rehabilitation-treatments-an-
approximate-dynamic-programming-approach/4E57D3194D81BE459530F21D154F40DC.

Burkschat:2011:APS

[BN11] M. Burkschat and J. Navarro. Aging properties of sequential order statistics. Probability in the Engineering and Informational Sciences, 25(4):449–467, October 2011. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
5BF06294FE9E5FBD3F08441B69B1BA4E.

Burkschat:2018:SCS

[BN18] M. Burkschat and J. Navarro. Stochastic comparisons of systems based on sequential order statistics via properties of distorted distributions. Probability in the Engineering and Informational Sciences, 32(2):246–274, April 2018. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/stochastic-
comparisons-of-systems-based-on-sequential-order-statistics-
via-properties-of-distorted-distributions/2C0AF316CE8A0ACC242D0DC1C7DADDE7.

Bertsimas:2006:TBE

[BNT06] Dimitris Bertsimas, Karthik Natarajan, and Chung-Piaw Teo. Tight bounds on expected order statistics. Probability in the Engineering and Informational Sciences, 20(4):667–686, October 2006. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/FA30F965984A3AE7B21D4524BAB61476.

Borst:1995:GGP

[Bor95] S. C. Borst. A globally gated polling system with a dormant server. Probability in the Engineering and Informational Sciences, 9(2):239–254, April 1995. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
3573195A51FD1EDF0FC5C686BCEF5165.

Borkar:2000:LAD

[Bor00] V. S. Borkar. A learning algorithm for discrete-time stochastic control. Probability in the Engineering and Informational Sciences, 14(2):243–258, April 2000. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
086AE3506E62639669577D537C85AD75.

Belzunce:2006:SOD

[BOR06] Félix Belzunce, Eva-María Ortega, and José M. Ruiz. Stochastic orderings of discrete-time processes and discrete record values. Probability in the Engineering and Informational Sciences, 20(3):447–464, July 2006. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
F39946660A33B6DEC8A95BD67C3804AF.

Bean:2009:HPH

[BOT09] Nigel G. Bean, Małgorzata M. O’Reilly, and Peter G. Taylor. Hitting probabilities and hitting times for stochastic fluid flows: the bounded model. Probability in the Engineering and Informational Sciences, 23(1):121–147, January 2009. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
514BD609F8444A7847E72ECAABB23892.

Boucherie:1996:BRQ

[Bou96] Richard J. Boucherie. Batch routing queueing networks with jump-over blocking. Probability in the Engineering and Informational Sciences, 10(2):287–297, April 1996. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
83212ECD1AFB0CDD2032BF996543F95B.

Bowker:1995:GJL

[BOV95] Albert H. Bowker, Ingram Olkin, and Arthur F. Veinott. Gerald J. Lieberman. Probability in the Engineering and Informational Sciences, 9(1):3–26, January 1995. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
0CA17EE46402AEB1BD21C5069D148E64.

Berman:2001:TCP

[BP01] Oded Berman and David Perry. Two control policies for stochastic EOQ-type models. Probability in the Engineering and Informational Sciences, 15(4):445–463, October 2001.

CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
4BC56E55170D5EB705373E723EBFD546.

Blanchet:2012:UCL

[BPG12] Jose H. Blanchet and Carlos G. Pacheco-González. Uniform convergence to a law containing Gaussian and Cauchy distributions. Probability in the Engineering and Informational Sciences, 26(3):437–448, July 2012. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
B41D20509557DE102B8FC6F8DB56EE2A.

Badian-Pessot:2021:OCP

[BPLD21] Pamela Badian-Pessot, Mark E. Lewis, and Douglas G. Down. Optimal control policies for an m/m/1 queue with a removable server and dynamic service rates. Probability in the Engineering and Informational Sciences, 35(2):189–209, April 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/optimal-control-policies-for-an-mm1-queue-with-
a-removable-server-and-dynamic-service-rates/AEA9643A2F073E5C0AE9782E7BBD1113.

Brown:2008:CC

[BPR08] Mark Brown, Erol A. Peköz, and Sheldon M. Ross. Coupon collecting. Probability in the Engineering and Informational Sciences, 22(2):221–229, March 2008. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/F066C4EF67EDA38E39E78A213C2A6EC9.

Brown:2010:SRS

[BPR10] Mark Brown, Erol A. Peköz, and Sheldon M. Ross. Some results for skip-free random walk. Probability in the Engineering and Informational Sciences, 24(4):491–507, October 2010. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
C2BE08A54128BCFF4925B8A61962C9FB.

Brown:2021:BDS

[BPR21] Mark Brown, Erol Peköz, and Sheldon Ross. Blockchain double-spend attack duration. Probability in the Engineering and Informational Sciences, 35(4):858–866, October 2021.

CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/blockchain-doublespend-attack-duration/EB8E30325D6BB203E1F15CEF6327565C.

Berman:2006:FEM

[BPS06] Oded Berman, David Perry, and Wolfgang Stadje. A fluid EOQ model with a two-state random environment. Probability in the Engineering and Informational Sciences, 20(2):329–349, April 2006. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
665B44070E698703CF20343E6AB27826.

Barron:2014:JFP

[BPS14] Yonit Barron, David Perry, and Wolfgang Stadje. A jumpfluid production-inventory model with a double band control. Probability in the Engineering and Informational Sciences, 28 (3):313–333, July 2014. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/EE753BBC11D233743518B9DA160DFF60.

Boland:1987:FDS

[BPT87] Philip J. Boland, Frank Proschan, and Y. L. Tong. Fault diversity in software reliability. Probability in the Engineering and Informational Sciences, 1(2):175–187, April 1987. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
F3FCE9F07C6BE0D065753AAF64501C8E.

Boland:1989:CPM

[BPT89] Philip J. Boland, Frank Proschan, and Y. L. Tong. Crossing properties of mixture distributions. Probability in the Engineering and Informational Sciences, 3(3):355–366, July 1989. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
48097D25BD6CE1587F6E6F4F1EEC75E4.

Boland:1990:LDC

[BPT90] Philip J. Boland, Frank Proschan, and Y. L. Tong. Linear dependence in consecutive k-out-of-n: F systems. Probability in the Engineering and Informational Sciences, 4(3):391– 397, July 1990. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/80103556B7A420F9D39977C87FEBFCC7.

Boxma:1999:FQM

[BPvdDS99] O. J. Boxma, D. Perry, and F. A. van der Duyn Schouten. Fluid queues and mountain processes. Probability in the Engineering and Informational Sciences, 13(4):407–427, October 1999. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
D2146134F0EA322FF3D411D8918646F7.

Boehme:1991:SNM

[BPvdW91] T. K. Boehme, W. Preuss, and V. van der Wall. On a simple numerical method for computing Stieltjes integrals in reliability theory. Probability in the Engineering and Informational Sciences, 5(1):113–128, January 1991. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
357463FAC57C386A9DBB58DA5FFDA6ED.

Bickel:1995:EIS

[BR95] Peter J. Bickel and Ya’acov Ritov. An exponential inequality for U-statistics with applications to testing. Probability in the Engineering and Informational Sciences, 9(1):39–52, January 1995. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
0CA0D6EE245DDA78515B96476CF28149.

Bauerle:1997:CRM

[BR97] Nicole Bäuerle and Ulrich Rieder. Comparison results for Markov-modulated recursive models. Probability in the Engineering and Informational Sciences, 11(2):203–217, April 1997. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
7F059D91EDE236821CF1196506A9A072.

Basterrech:2017:ESQ

[BR17] Sebastián Basterrech and Gerardo Rubino. Echo state queueing networks: a combination of reservoir computing and random neural networks. Probability in the Engineering and Informational Sciences, 31(4):457–476, October 2017. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
B61D5741A5AD03BC3B6433BA139940D4.

Bretienbach:1990:CWG

[Bre90] Jerome R. Bretienbach. The capacity of the white Gaussian noise channel. Probability in the Engineering and Informational Sciences, 4(3):345–353, July 1990. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
A6F31B81CD759EB79843F8F2B49F825B.

Brill:1988:SSD

[Bri88] P. H. Brill. Single-server with delay-dependent arrival streams. Probability in the Engineering and Informational Sciences, 2 (2):231–247, April 1988. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/2606826FA4F25E816D0C6C4D278123FE.

Brill:2014:AAF

[Bri14] Percy H. Brill. Alternative analysis of finite-time probability distributions of renewal theory. Probability in the Engineering and Informational Sciences, 28(2):183–201, April 2014. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
E31F52C0D3E0DB1D3EFE70319AAD7A52.

Brown:1989:TPI

[Bro89] Mar Brown. On two problems involving partial sums. Probability in the Engineering and Informational Sciences, 3(4):511– 516, October 1989. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/D5572920EB88E5EAF5876FCD129DDBEE.

Brown:1991:SAE

[Bro91] Mark Brown. Spectral analysis, without eigenvectors, for Markov chains. Probability in the Engineering and Informational Sciences, 5(2):131–144, April 1991. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
9C6CA7CDA7C80CAF913BC1B5C205D2E6.

Brown:1997:EED

[Bro97] Mark Brown. Estimation of an exponential distribution. Probability in the Engineering and Informational Sciences, 11(3): 341–359, July 1997. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/C5D192F89EEAD1E14174020DDF167A49.

Browne:1999:OGC

[Bro99] Sid Browne. Optimal growth in continuous-time with credit risk. Probability in the Engineering and Informational Sciences, 13(2):129–145, April 1999. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/06F0D1565261243409350114B0D0576B.

Brooms:2000:IED

[Bro00] Anthony C. Brooms. Individual equilibrium dynamic routing in a multiple server retrial queue. Probability in the Engineering and Informational Sciences, 14(1):9–26, January 2000. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
0B67E704436BB641274BEA7A2A718275.

Brown:2001:VEN

[Bro01] Mark Brown. Variance estimation for the normal distribution under log symmetric loss. Probability in the Engineering and Informational Sciences, 15(3):351–367, July 2001. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
D6B6FFCD8F8C3DEE24F010215BA880AD.

Brown:2006:EWT

[Bro06] Mark Brown. Exploiting the waiting time paradox: applications of the size-biasing transformation. Probability in the Engineering and Informational Sciences, 20(2):195–230, April 2006. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
DB5A0F707F45ABB3721921024336A5EE.

Brown:2009:OSD

[Bro09] Mark Brown. Order statistics from dependent, completely monotone distributions. Probability in the Engineering and Informational Sciences, 23(3):449–455, July 2009. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
2E0C785204200F6DDC28BA211A67E26C.

Brown:2013:BSP

[Bro13] Mark Brown. A balance scale problem. Probability in the Engineering and Informational Sciences, 27(2):141–146, April 2013. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
A5454BCFACF52AE0E43643C681FC819E.

Brown:2014:VBU

[Bro14] Mark Brown. Variance bounds under a hazard rate constraint. Probability in the Engineering and Informational Sciences, 28 (2):203–208, April 2014. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/9530E44F8CFB909D681A0BE697078AC7.

Brun:2016:PNC

[Bru16] Olivier Brun. Performance of non-cooperative routing over parallel non-observable queues. Probability in the Engineering and Informational Sciences, 30(3):455–469, July 2016. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
E24C8281FD1D2DC82F2151217CE64CB2.

Brown:1987:ICS

[BS87] Mark Brown and Yi-Shi Shao. Identifying coefficients in the spectral representation for first passage time distributions. Probability in the Engineering and Informational Sciences, 1(1):69–74, January 1987. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/63F8B676090B8921A13EF799A0B8CECF.

Brandt:1991:QNM

[BS91] A. Brandt and H. Sulanke. A queueing network model for a communication system with channel allocation. Probability in the Engineering and Informational Sciences, 5(2):171– 183, April 1991. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/286B2868671F82A33122EFD5A52E5457.

Bardhan:1993:RCL

[BS93] Indrajit Bardhan and Karl Sigman. Rate conservation law for stationary semimartingales. Probability in the Engineering and Informational Sciences, 7(1):1–17, January 1993.

CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
145D399140926AD595C468FA467BEBC3.

Brown:1998:CVR

[BS98] Mark Brown and J. George Shanthikumar. Comparing the variability of random variables and point processes. Probability in the Engineering and Informational Sciences, 12(4):425–444, October 1998. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/8CF8984FD8FA1979AFCA0C08332AFC15.

Bartoszewicz:2006:PSO

[BS06] Jaroslaw Bartoszewicz and Magdalena Skolimowska. Preservation of stochastic orders under mixtures of exponential distributions. Probability in the Engineering and Informational Sciences, 20(4):655–666, October 2006. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
D7CC03F411713C4D01B859BC7111DA70.

Badia:2008:PRC

[BS08] F. G. Badía and C. Sangüesa. Preservation of reliability classes under mixtures of renewal processes. Probability in the Engineering and Informational Sciences, 22(1):1–17, January 2008. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
C672712B5DEA31975B967A0B01C7B4DE.

Bradonjic:2014:BPR

[BS14] Milan Bradonjić and Iraj Saniee. Bootstrap percolation on random geometric graphs. Probability in the Engineering and Informational Sciences, 28(2):169–181, April 2014. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
9D466E6276655175381ED250CB102079.

Badia:2021:LCC

[BSF21] F. G. Badía, C. Sangüesa, and A. Federgruen. Log-concavity of compound distributions with applications in operational and actuarial models. Probability in the Engineering and Informational Sciences, 35(2):210–235, April 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-

engineering-and-informational-sciences/article/logconcavity-
of-compound-distributions-with-applications-in-operational-
and-actuarial-models/E91407C89E0C3B9CABBC241206D9BD39.

Block:2015:FRS

[BSL15] Henry W. Block, Thomas H. Savits, and Naftali A. Langberg. The failure rate shape for a mixture of two gammas. Probability in the Engineering and Informational Sciences, 29(2): 253–264, April 2015. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/DF0A9542D5A971C5FB62E71BE218BC46.

Boesch:1990:SAC

[BSS90] F. T. Boesch, A. Satyanarayana, and C. L. Suffel. Some alternate characterizations of reliability domination. Probability in the Engineering and Informational Sciences, 4(2):257–276, April 1990. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/8B9B0C78A79730022B507758C5568CB3.

Block:1998:RHR

[BSS98] Henry W. Block, Thomas H. Savits, and Harshinder Singh. The reversed hazard rate function. Probability in the Engineering and Informational Sciences, 12(1):69–90, January 1998. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
4255CB93F95BFF03B49D101C05858D00.

Bean:1995:MPD

[BT95] N. G. Bean and P. G. Taylor. Maximal profit dimensioning and tariffing of loss networks. Probability in the Engineering and Informational Sciences, 9(3):323–340, July 1995. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
9B63EEEFE51486624FA93B12113EB7AF.

Buchholz:1997:ADA

[Buc97] Peter Buchholz. An aggregation/disaggregation algorithm for stochastic automata networks. Probability in the Engineering and Informational Sciences, 11(2):229–253, April 1997. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
047BD366A320BA0D67936150062739DE.

Burkschat:2013:CSB

[Bur13] Marco Burkschat. Comments on the survey by Balakrishnan and Zhao. Probability in the Engineering and Informational Sciences, 27(4):465–466, October 2013. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
E4F6C2F00B1BEA074C5E797ADCDA1B4B. See [BZ13b].

Butler:1989:RBG

[But89] David A. Butler. A random-bit generator for use in simulating the reliability of a coherent system. Probability in the Engineering and Informational Sciences, 3(1):141–147, January 1989. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
2C2964E730008624C67C1262123E0D36.

Bossier:2013:APT

[BVA13] Ruben Bossier, Maria Vlasiou, and Ivo Adan. Analytic properties of two-carousel systems. Probability in the Engineering and Informational Sciences, 27(1):57–84, January 2013. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
76CE116ABC308E5B555BF1C64D29B4E3.

Boxma:1993:CAA

[BvH93] O. J. Boxma and G. J. van Houtum. The compensation approach applied to a 2 × 2 switch. Probability in the Engineering and Informational Sciences, 7(4):471–493, October 1993. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
EE1108EC8F7CAB27D175445C2B93B2D0.

Boucherie:2018:TES

[BvHTvO18] Richard J. Boucherie, Geert-Jan van Houtum, Judith Timmer, and Jan-Kees van Ommeren. A two-echelon spare parts network with lateral and emergency shipments: a product-form approximation. Probability in the Engineering and Informational Sciences, 32(4):536–555, October 2018. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/twoechelon-
spare-parts-network-with-lateral-and-emergency-shipments-
a-productform-approximation/1A91D5F872625A7CB73C83E0AAA126CD.

Buitenhek:1997:FCF

[BvHvO97] Ronald Buitenhek, Geert-Jan van Houtum, and Jan-Kees van Ommeren. On first-come first-served versus random service discipline in multiclass closed queueing networks. Probability in the Engineering and Informational Sciences, 11(3):313– 326, July 1997. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/EA8C3E6D5280E72CFBC9929D318FF5AB.

Berger:1995:MVQ

[BW95] Arthur W. Berger and Ward Whitt. Maximum values in queueing processes. Probability in the Engineering and Informational Sciences, 9(3):375–409, July 1995. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
080C0C0354ABA07CB74866585FEF1BD3.

Boon:2014:HTA

[BW14] M. A. A. Boon and E. M. M. Winands. Heavy-traffic analysis of k-limited polling systems. Probability in the Engineering and Informational Sciences, 28(4):451–471, October 2014. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
24735233197A2A4DB1FC647A7DFD800E.

Boxma:1993:GGP

[BWY93] O. J. Boxma, J. A. Weststrate, and U. Yechiali. A globally gated polling system with server interruptions, and applications to the repairman problem. Probability in the Engineering and Informational Sciences, 7(2):187–208, April 1993. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
571F6A9F84769E9634BC26CAA42F3733.

Benyamini:1999:OCL

[BY99] Zvi Benyamini and Uri Yechiali. Optimality of control limit maintenance policies under nonstationary deterioration. Probability in the Engineering and Informational Sciences, 13(1): 55–70, January 1999. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/32AF274AFCDE025118E4C1A1F51456BC.

Bekker:2005:EBL

[BZ05] René Bekker and Bert Zwart. On an equivalence between loss rates and cycle maxima in queues and dams. Probability in the Engineering and Informational Sciences, 19(2):241–255, April 2005. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
51C2EC0F861A43E5ED029F0E7249B727.

Balakrishnan:2013:AR

[BZ13a] N. Balakrishnan and Peng Zhao. Authors’ rejoinder. Probability in the Engineering and Informational Sciences, 27(4): 467–469, October 2013. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/56F6C70965FB0CD13D38730D2B8E4A57.

Balakrishnan:2013:OPO

[BZ13b] N. Balakrishnan and Peng Zhao. Ordering properties of order statistics from heterogeneous populations: a review with an emphasis on some recent developments. Probability in the Engineering and Informational Sciences, 27(4):403–443, October 2013. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
5A1D2D4335FFB061F2C3426685BE6173. See comments [Bur13,

Kha13, LY13, Sha13, Xu13].

Caglayan:2016:EGC

[Çağ16] M. Ufuk Çağlayan. Erol Gelenbe: a career in multidisciplinary probability models. Probability in the Engineering and Informational Sciences, 30(3):308–325, July 2016. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
1C800D22C861623326F6E6EB5FB433F0.

Caglayan:2017:GNT

[Cag17] Mehmet Ufuk Caglayan. G-networks and their applications to machine learning, energy packet networks and routing: introduction to the special issue. Probability in the Engineering and Informational Sciences, 31(4):381–395, October 2017. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
A326B513C832079444829FE52BD8EF9D.

Cai:2002:DTR

[Cai02] Jun Cai. Discrete time risk models under rates of interest. Probability in the Engineering and Informational Sciences, 16 (3):309–324, July 2002. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/673F1098077A1A19A8D44356280B70CE.

Calvert:1997:DTE

[Cal97] Bruce Calvert. The downs-Thomson effect in a Markov process. Probability in the Engineering and Informational Sciences, 11 (3):327–340, July 1997. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/72E25F8D4739F1297A62082DD47258C6.

Calvin:2010:UEA

[Cal10] James M. Calvin. Using excursions to analyze simulation output. Probability in the Engineering and Informational Sciences, 24(2):201–218, April 2010. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
F2303F9171AFFB7CB5640495757FF4A3.

Cao:2017:MSP

[Cao17] Yang Cao. Multiple server preemptive scheduling with impatience. Probability in the Engineering and Informational Sciences, 31(2):226–238, April 2017. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/15D1D09FEF2D6887417C40C0A309FBD0.

Cha:2020:MGP

[CB20] Ji Hwan Cha and F. G. Badía. On a multivariate generalized Pólya process without regularity property. Probability in the Engineering and Informational Sciences, 34(4):484–506, October 2020. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
journals/probability-in-the-engineering-and-informational-
sciences/article/on-a-multivariate-generalized-polya-
process-without-regularity-property/5D4BC12538EC4323102D8C6CCBC55A9C.

Chen:2020:PMT

[CBBG20] Yanting Chen, Xinwei Bai, Richard J. Boucherie, and Jasper Goseling. Performance measures for the two-node queue with finite buffers. Probability in the Engineering and Informational Sciences, 34(4):522–549, October 2020. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/performance-
measures-for-the-twonode-queue-with-finite-buffers/
C106C6DB9DBC97A8F8DBB744F6C91FA5.

Chen:2015:IMR

[CBG15] Yanting Chen, Richard J. Boucherie, and Jasper Goseling. The invariant measure of random walks in the quarter-plane: representation in geometric terms. Probability in the Engineering and Informational Sciences, 29(2):233–251, April 2015. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
5FCE7848ACDB3AC6882F6B7ACA05D9C3.

Chao:1993:ANO

[CBL93] Chern-Ching Chao, Zhidong Bai, and Wen-Qi Liang. Asymptotic normality for oscillation of permutation. Probability in the Engineering and Informational Sciences, 7(2):227–235, April 1993. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
76BC5DE2F6B0E2682CD68DC64B87A926.

Chen:2018:TSS

[CBM18] Weiwei Chen, Alok Baveja, and Benjamin Melamed. Temporal shaping of simulated time series with cyclical sample paths. Probability in the Engineering and Informational Sciences, 32(1):126–143, January 2018. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/temporal-
shaping-of-simulated-time-series-with-cyclical-sample-
paths/3012A7300B31EF8620E04DDF5D2016AC.

Chaudhry:2005:CAS

[CCC05] Mohan L. Chaudhry, Dae W. Choi, and Kyung C. Chae. Computational analysis of stationary waiting-time distributions of GI^{X} /R/1 and GI^{X} /D/1 queues. Probability in the Engineering and Informational Sciences, 19(1):121–140, January 2005. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
33E02C754B75CD3DDA0BB2FC906BA05B.

Cavazos-Cadena:2001:EOS

[CCdO01] Rolando Cavazos-Cadena and Raúl Montes de Oca. Existence of optimal stationary policies in finite dynamic programs with nonnegative rewards: an alternative approach. Probability in the Engineering and Informational Sciences, 15(4):557–564, October 2001. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/915A0F63E1222238FB56FCF1959316B3.

Chang:1999:NCB

[CCH99] Gerard J. Chang, Lirong Cui, and Frank K. Hwang. New comparisons in Birnbaum importance for the consecutivek-out-of-n system. Probability in the Engineering and Informational Sciences, 13(2):187–192, April 1999. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
6B929643546ADB33B1B85BDC9C5EB555. See corrigenda [CHC00].

Chang:1990:IDT

[CCP90] Cheng-Shang Chang, Xiuli Chao, and Michael Pinedo. Integration of discrete-time correlated Markov processes in a TDM system. Probability in the Engineering and Informational Sciences, 4(1):29–56, January 1990. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
AA30871C9BFEB377955BEE3FE967495B.

Cancela:1996:ITH

[CE96] Héctor Cancela and Mohamed El Khadiri. An improvement to the total hazard method for system reliability simulation. Probability in the Engineering and Informational Sciences, 10 (2):187–196, April 1996. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/604B788F1F07AF5A41EC0C69804FB6E2.

Chen:2006:OOS

[CF06] Frank Y. Chen and Y. Feng. Optimization and optimality of (s, S) stochastic inventory systems with non-quasiconvex costs. Probability in the Engineering and Informational Sciences, 20 (2):287–306, April 2006. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/E0B783B41C4AB8DCAE9A7FAD4F36222A.

Cha:2012:SSM

[CF12] Ji Hwan Cha and Maxim Finkelstein. Stochastic survival models with events triggered by external shocks. Probability in the Engineering and Informational Sciences, 26(2):183–195, April 2012. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
BEEA4D46F01469A770419FB5C333EA7E.

Cha:2013:NCG

[CF13] Ji Hwan Cha and Maxim Finkelstein. A note on the class of geometric counting processes. Probability in the Engineering and Informational Sciences, 27(2):177–185, April 2013. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
FA114BDD697CB72FF8BE6604E0A02F55.

Coffman:1998:MRW

[CFFH98] E. G. Coffman, Philippe Flajolet, Leopold Flatto, and Micha Hofri. The maximum of a random walk and its application to rectangle packing. Probability in the Engineering and Informational Sciences, 12(3):373–386, July 1998. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
7A279164867055FFC9B825E2AEF5AB5A.

Coffman:1999:CCA

[CFKP99] E. G. Coffman, Anja Feldmann, Nabil Kahale, and Bjorn Poonen. Computing call admission capacities in linear networks. Probability in the Engineering and Informational Sciences, 13(4):387–406, October 1999. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
62F4BCC5DCA091D767F7E965726410B0.

Coffman:1988:SMQ

[CFM^{+}88] E. G. Coffman, L. Flatto, I. Mitrani, L. A. Shepp, and C. Knessl. Stochastic models of queue storage. Probability in the Engineering and Informational Sciences, 2(1):75–93, January 1988. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/077CA19F824D5AD0CED177DE9F666FA8.

Chaudhry:1996:ADT

[CG96] M. L. Chaudhry and U. C. Gupta. On the analysis of the discrete-time Geom(n)/G(n)/1/N queue. Probability in the Engineering and Informational Sciences, 10(3):415–428, July 1996. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
2AB1F0428024667A686E0E16644804B3.

Coffman:1999:RPO

[CGK99] E. G. Coffman, E. N. Gilbert, and Y. A. Kogan. Redialing policies: optimality and success probabilities. Probability in the Engineering and Informational Sciences, 13(1):37–53, January 1999. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
A22F8FF003664A195952530837A446DF.

Cidon:1996:QIT

[CGKS96] Israel Cidon, Roch Guréin, Asad Khamisy, and Moshe Sidi. On queues with interarrival times proportional to service times. Probability in the Engineering and Informational Sciences, 10(1):87–107, January 1996. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
67B1B07173FF3839E136D11F2A32E179.

Chaudhry:2021:ACF

[CGM21] M. L. Chaudhry, Veena Goswami, and Abdalla Mansur. Analytically closed-form solutions for the distribution of a number of customers served during a busy period for special cases of the Geo/G/1 queue. Probability in the Engineering and Informational Sciences, 35(3):650–679, July 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/analytically-
closedform-solutions-for-the-distribution-of-a-number-
of-customers-served-during-a-busy-period-for-special-
cases-of-the-geog1-queue/F2B19116865BCED66C62ADB2BD7D7C1E.

Chou:1987:SAT

[CH87] S. T. Chou and J. P. Hsu. Stochastic analysis of the transient behavior of an MSMPR crystallizer: Effects of the seed size distribution and size- dependent growth rate. Probability in the Engineering and Informational Sciences, 1(4):383–404, October 1987. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/EF2350FEC85578F1599E4AEE578414F2.

Chang:1988:OCS

[CH88] G. J. Chang and F. K. Hwang. Optimal consecutive-k-out-of-n systems under a fixed budget. Probability in the Engineering and Informational Sciences, 2(1):63–73, January 1988. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
DABF7141D1E9DF1E11D77B5762F13010.

Cirillo:2011:GES

[CH11] Pasquale Cirillo and Jürg Hüsler. Generalized extreme shock models with a possibly increasing threshold. Probability in the Engineering and Informational Sciences, 25(3):419–434, July 2011. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
6BCFEAE2EE2CDE7EDAF219768D9FA12C.

Cirillo:2012:UMC

[CH12] Pasquale Cirillo and Jürg Hüsler. An urn model for cascading failures on a lattice. Probability in the Engineering and Informational Sciences, 26(4):509–534, October 2012. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
C19A321A9B630A8E4C033A9A7BE826EE.

Chen:2015:GUM

[CH15] May-Ru Chen and Shoou-Ren Hsiau. On a generalized Q-urn model. Probability in the Engineering and Informational Sciences, 29(1):99–115, January 2015. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
E26F6C8350164D143428FF24FF140BEB.

Chao:1994:NQN

[Cha94] Xiuli Chao. A note on queueing networks with signals and random triggering times. Probability in the Engineering and Informational Sciences, 8(2):213–219, April 1994. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
BF70ABA6D9D12F4FA1E3D055A0E99539.

Chaudhry:1999:SSD

[Cha99] M. L. Chaudhry. On the steady-state distribution of number in the queue geom/g/1. Probability in the Engineering and Informational Sciences, 13(1):121–125, January 1999. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
D1FC6A5069544CF768D400061FE159B4.

Chang:2000:CNC

[CHC00] Gerard J. Chang, Frank K. Hwang, and Lirong Cui. Corrigenda on “New Comparisons in Birnbaum Importance for the Consecutive-k-out-of-n System”. Probability in the Engineering and Informational Sciences, 14(3):405, July 2000. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
26B45606BEF7EF8B0A7A0B0046B78A48. See [CCH99].

Cheng:1997:LRM

[Che97] Dinah W. Cheng. Line reversibility of multiserver systems. Probability in the Engineering and Informational Sciences, 11 (2):177–188, April 1997. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/D72DCDEC1C3AC923955E489698C3B274.

Chen:2009:PTP

[Che09] May-Ru Chen. Proportional three-person red-and-black games. Probability in the Engineering and Informational Sciences, 23(1):37–50, January 2009. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/33631F8F0280E43A6B28B138AE99BF34.

Chen:2011:TPR

[Che11]  May-Ru Chen.  Two-person red-and-black game with

lower limit. Probability in the Engineering and Informational Sciences, 25(1):119–133, January 2011. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
B97C1A87E30686F9776330EF67D1C7C1.

Chen:2012:NET

[Che12] May-Ru Chen. Nash equilibriums in two-person redand-black games. Probability in the Engineering and Informational Sciences, 26(3):323–336, July 2012. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
FCE923B92C7FD68BC979071EC0C2A593.

Chen:2020:TDP

[Che20] May-Ru Chen. A time-dependent Pólya urn with multiple drawings. Probability in the Engineering and Informational Sciences, 34(4):469–483, October 2020. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/timedependent-
polya-urn-with-multiple-drawings/55B201841942006B4A711E5CBFBBCF6D.

Chen:1990:OBG

[CHH90] B. Chen, X. D. Hu, and F. K. Hwang. Optimal binomial group testing with a test history. Probability in the Engineering and Informational Sciences, 4(4):523–530, October 1990. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
F6A0E0355A93BE1AC5145BD5E7A7BB7F.

Chiu:2021:ITS

[Chi21] Wan-Yi Chiu. An improved test of the squared Sharpe ratio. Probability in the Engineering and Informational Sciences, 35(3):404–431, July 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/an-improved-
test-of-the-squared-sharpe-ratio/5778993FF88178B28275907D18E7AA81.

Chiu:2023:VRA

[Chi23] Wan-Yi Chiu. A value-at-risk approach to futures hedge. Probability in the Engineering and Informational Sciences, 37(3): 818–832, July 2023. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/journals/probability-in-the-engineering-and-informational-
sciences/article/valueatrisk-approach-to-futures-hedge/
92A600825D8437D017366321879B0426.

Cui:1995:IFR

[CHJ95] Lirong Cui, Alan G. Hawkes, and Assad Jalali. The increasing failure rate property of consecutive-k-out-of-n systems. Probability in the Engineering and Informational Sciences, 9(2): 217–225, April 1995. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/4DE2B298BCD5DAF2637320C1CAD9755B.

Cai:2013:SMM

[CHKL13] Yiwei Cai, John J. Hasenbein, Erhan Kutanoğlu, and Melody Liao. Single-machine multiple-recipe predictive maintenance. Probability in the Engineering and Informational Sciences, 27 (2):209–235, April 2013. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/D2183CEDE7A0D449A5C1EA3F0BE3B157.

Chen:1994:RMC

[CHL94] Robert W. Chen, F. K. Hwang, and Wen-Ching Winnie Li. A reversible model for consecutive-2-out-of-n: F systems with node and link failures. Probability in the Engineering and Informational Sciences, 8(2):189–200, April 1994. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
F1CCD0237EDCBD76F3F3CA8CDFF49EAC.

Cooper:2003:CSB

[CHL03] William L. Cooper, Shane G. Henderson, and Mark E. Lewis. Convergence of simulation-based policy iteration. Probability in the Engineering and Informational Sciences, 17(2):213–234, April 2003. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/2B9838A317088FAA2938C10A9A6A0F66.

Chong:2018:TCR

[CHL18] Kenneth C. Chong, Shane G. Henderson, and Mark E. Lewis. Two-class routing with admission control and strict priorities. Probability in the Engineering and Informational Sciences, 32(2):163–178, April 2018. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/twoclass-
routing-with-admission-control-and-strict-priorities/
AF296F8E664452DA2B8996E6C53B7A90.

Chong:1999:MFR

[Cho99] King Sing Chong. Move-to-front rule for accessing several records. Probability in the Engineering and Informational Sciences, 13(1):75–93, January 1999. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
B4BDB3AB9D1FAFD48744E69C0D9499E9.

Chang:1994:SOS

[CHRW94] Cheng-Shang Chang, Arie Hordijk, Rhonda Righter, and Gideon Weiss. The stochastic optimality of SEPT in parallel machine scheduling. Probability in the Engineering and Informational Sciences, 8(2):179–188, April 1994. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
9598CE551059D4CC0A44F4B2DD896CF6.

Chen:1987:EBO

[CHS87] Pinyuen Chen, Lifang Hsu, and Milton Sobel. Entropy-based optimal group-testing procedures. Probability in the Engineering and Informational Sciences, 1(4):497–509, October 1987. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
25BB4AAAAC38105C26800188B78DDCC8.

Coffman:1989:SSJ

[CHW89] E. G. Coffman, M. Hofri, and G. Weiss. Scheduling stochastic jobs with a two-point distribution on two parallel machines. Probability in the Engineering and Informational Sciences, 3(1):89–116, January 1989. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
FBCA6A440CD16ECEF62B4FE71405F463.

Chyu:1991:DPI

[Chy91] C. C. Chyu. Decomposable probabilistic influence diagrams. Probability in the Engineering and Informational Sciences, 5 (2):229–243, April 1991. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/C6B95EA26C25DC7B633A04D6803301FC.

Chen:1993:SMA

[CHYZ93] R. W. Chen, F. K. Hwang, Y. C. Yao, and A. Zame. Some monotone almost balanced knockout tournament plans. Probability in the Engineering and Informational Sciences, 7(3): 361–368, July 1993. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/AE478BE0342F6441E8A2AC1B0DE92EB1.

Choi:1996:ADT

[CJ96] Woo-Yong Choi and Chi-Hyuck Jun. Analysis of a discretetime queueing system with a Markov-modulated input process and constant service rate. Probability in the Engineering and Informational Sciences, 10(3):429–441, July 1996. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
4DFCC0E836354F00321A309641112B7D.

Connors:1988:BRO

[CK88] D. P. Connors and P. R. Kumar. Balance of recurrence order in time-inhomogeneous Markov chains with application to simulated annealing. Probability in the Engineering and Informational Sciences, 2(2):157–184, April 1988. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
DD40434B4CE6AF89FE9FA01A3086958B.

Cowan:2015:MAB

[CK15] Wesley Cowan and Michael N. Katehakis. Multi-armed bandits under general depreciation and commitment. Probability in the Engineering and Informational Sciences, 29(1):51–76, January 2015. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/CE9A226238311D80DD60E6A241C7F836.

Cowan:2020:EEP

[CK20] Wesley Cowan and Michael N. Katehakis. Explorationexploitation policies with almost sure, arbitrarily slow growing asymptotic regret. Probability in the Engineering and Informational Sciences, 34(3):406–428, July 2020. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/explorationexploitation-
policies-with-almost-sure-arbitrarily-slow-growing-
asymptotic-regret/E099DD0167AAB99CFF010EC2984FFA54.

Colbourn:1991:RNC

[CL91] Charles J. Colbourn and Michael V. Lomonosov. Renewal networks: Connectivity and reachability on a time interval. Probability in the Engineering and Informational Sciences, 5(3): 361–368, July 1991. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/56B8672CECEEDABC6CC61B4593BBD16B.

Chong:2006:MFR

[CL06] King Sing Chong and Kin Lam. Move-forward rules and fswap rules applied to a communication problem. Probability in the Engineering and Informational Sciences, 20(2):271–286, April 2006. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/DEA1A551620404B10CC1BC64525DE4D0.

Cheung:2023:JMD

[CL23] Eric C. K. Cheung and Haibo Liu. Joint moments of discounted claims and discounted perturbation until ruin in the compound Poisson risk model with diffusion. Probability in the Engineering and Informational Sciences, 37:387–417, April 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/joint-
moments-of-discounted-claims-and-discounted-perturbation-
until-ruin-in-the-compound-poisson-risk-model-with-
diffusion/9D3EA0F741D876DF7BCA81B09138FCD9.

Choi:1998:GGD

[CLC98] Bong Dae Choi, Yutae Lee, and Doo Il Choi. Geo^{X_{1}}, Geo^{X_{2}} /D/c HOL priority queueing system with random order selection within each priority class. Probability in the Engineering and Informational Sciences, 12(1):125–139, January 1998. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
EBBB271C3CD57DC55A4C6EC56C7ECB15.

Cali:2020:PGC

[CLN20] Camilla Calì, Maria Longobardi, and Jorge Navarro. Properties for generalized cumulative past measures of information. Probability in the Engineering and Informational Sciences, 34(1):92–111, January 2020. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/properties-
for-generalized-cumulative-past-measures-of-information/
34187C1467EFDEFC0AC9F7DA8D1C6D1F.

Choi:2021:EFB

[CLS21] Michael C. H. Choi, Chihoon Lee, and Jian Song. Entropy flow and de Bruijn’s identity for a class of stochastic differential equations driven by fractional Brownian motion. Probability in the Engineering and Informational Sciences, 35(3):369–380, July 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/entropy-
flow-and-de-bruijns-identity-for-a-class-of-stochastic-
differential-equations-driven-by-fractional-brownian-
motion/7DA3FD3922416F26655B818B62F38136.

Collet:2013:EOM

[CLSS13] Francesca Collet, Fabrizio Leisen, Fabio Spizzichino, and Florentina Suter. Exchangeable occupancy models and discrete processes with the generalized uniform order statistics property. Probability in the Engineering and Informational Sciences, 27(4):533–552, October 2013. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
293C685319AB3B0E0640A11961671593.

Chao:2000:TPF

[CM00] Xiuli Chao and Masakiyo Miyazawa. On truncation properties of finite-buffer queues and queueing networks. Probability in the Engineering and Informational Sciences, 14(4):409–423, October 2000. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/58D735A646278AD780BAB6967004472A.

Chen:2005:CDM

[CM05] Qiming Chen and James D. McCalley. A cluster distribution as a model for estimating high-order-event probabilities in power systems. Probability in the Engineering and Informational Sciences, 19(4):489–505, October 2005. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
02D28AE99CD4615D46F938659FE32D8F.

Christophi:2008:CT

[CM08] Costas Christophi and Hosam Mahmoud. On climbing tries. Probability in the Engineering and Informational Sciences, 22(1):133–149, January 2008. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/0882E9D1151A8B0ED19858E814E4533D.

Chen:2013:ABE

[CMH13] Die Chen, Tiantian Mao, and Taizhong Hu. Asymptotic behavior of extremal events for aggregate dependent random variables. Probability in the Engineering and Informational Sciences, 27(4):507–531, October 2013. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
C85B4B27625F29FF318FB89A80CDAB99.

Cahen:2017:REA

[CMZ17] Ewan Jacov Cahen, Michel Mandjes, and Bert Zwart. Rare event analysis and efficient simulation for a multidimensional ruin problem. Probability in the Engineering and Informational Sciences, 31(3):265–283, July 2017. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
26CE76A723C76DAC143DB8B4D5A78083.

Calvin:2000:SPM

[CN00] James M. Calvin and Marvin K. Nakayama. Simulation of processes with multiple regeneration sequences. Probability in the Engineering and Informational Sciences, 14(2):179–201, April 2000. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
3E84AD7797E25FCE9A0521ADD6072442.

Cinlar:1987:RCD

[ÇÖ87] E. Çinlar and S. Özekici. Reliability of complex devices in random environments. Probability in the Engineering and Informational Sciences, 1(1):97–115, January 1987. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
2BA48BAFE714D06BAB96DA20ADDD8F4E.

Cekyay:2015:OMS

[ÇÖ15] Bora Çekyay and Süleyman Özekici. Optimal maintenance of semi-Markov missions. Probability in the Engineering and Informational Sciences, 29(1):77–98, January 2015.

CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
CBA55DAE56E276349C865D8D23180B15.

Covo:2010:PAM

[Cov10] Shai Covo. On probabilities associated with the minimum distance between events of a Poisson process in a finite interval. Probability in the Engineering and Informational Sciences, 24 (3):423–439, July 2010. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/EDD3422062C4D620E6079731E39627B3.

Chao:1993:GNQ

[CP93] Xiuli Chao and Michael Pinedo. On generalized networks of queues with positive and negative arrivals. Probability in the Engineering and Informational Sciences, 7(3):301–334, July 1993. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
011750C3115DA2A1E54BD47A4A76825E.

Chao:1995:QNS

[CP95] Xiuli Chao and Michael Pinedo. On queueing networks with signals and history-dependent routing. Probability in the Engineering and Informational Sciences, 9(3):341–354, July 1995. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
96239347235C98F6A58587E77E0F10D8.

Coffman:1991:SLQ

[CPR91] E. G. Coffman, A. A. Puhalskii, and M. I. Reiman. Storagelimited queues in heavy traffic. Probability in the Engineering and Informational Sciences, 5(4):499–522, October 1991. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
22632EF08F292B8829E5648DD07B39F0.

Chao:1989:ISO

[CPS89] Xiuli Chao, Michael Pinedo, and Karl Sigman. On the interchangeability and stochastic ordering of exponential queues in tandem with blocking. Probability in the Engineering and Informational Sciences, 3(2):223–236, April 1989. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
88EB4CBDF228852E74DD42CA1F3EA4B8.

Chandra:2001:SRR

[CR01] Nimai Kumar Chandra and Dilip Roy. Some results on reversed hazard rate. Probability in the Engineering and Informational Sciences, 15(1):95–102, January 2001. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
052ACCB211EA2C3D1829B8DE1C99EC87.

Chang:2007:MSC

[CR07] Kuang-Chao Chang and Sheldon M. Ross. The multiple subset coupon collecting problem. Probability in the Engineering and Informational Sciences, 21(3):435–440, July 2007. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
B7AA45852DF17FCBEA0AE032D7E6E264.

Chen:2015:SSK

[CR15] Kai Chen and Sheldon M. Ross. Static stochastic knapsack problems. Probability in the Engineering and Informational Sciences, 29(4):527–546, October 2015. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
FC1B974EE7BBF25690BBC616BF09EEF5.

Cao:2019:WPS

[CR19] Yang Cao and Sheldon M. Ross. Winner plays structure in random knockout tournaments. Probability in the Engineering and Informational Sciences, 33(4):500–510, October 2019. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/winner-plays-structure-in-random-knockout-tournaments/
3D9280EDE46BDB0978B32550DC2FA583.

Cao:2021:WPC

[CR21] Yang Cao and Sheldon M. Ross. Winner plays competition models. Probability in the Engineering and Informational Sciences, 35(2):236–241, April 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/winner-
plays-competition-models/5AD6F79C90A9A803F7C7935C0A17D5B7.

Choi:1993:MGR

[CRP93] Bong Dae Choi, Kyung Hyune Rhee, and Kwang Kyu Park. The M/G/1 retrial queue with retrial rate control policy. Probability in the Engineering and Informational Sciences, 7(1): 29–46, January 1993. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/3247FC33D38806427A3D8EB3A598EEAF.

Clarotti:1990:BBD

[CS90] C. A. Clarotti and F. Spizzichino. Bayes burn-in decision procedures. Probability in the Engineering and Informational Sciences, 4(4):437–445, October 1990. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
5452B5C72398496E0A911321868E0861.

Chen:1992:OAB

[CS92] C. S. Chen and Thomas H. Savits. Optimal age and block replacement for a general maintenance model. Probability in the Engineering and Informational Sciences, 6(1):81–98, January 1992. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
A0922A56A831EB8737C5CAA787454A02.

Coffman:1993:CPG

[CS93] E. G. Coffman and Aleksandr Stolyar. Continuous polling on graphs. Probability in the Engineering and Informational Sciences, 7(2):209–226, April 1993. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/3FE3D4C5A6B2D1E7549BEA2729941EE4.

Caramellino:1994:DAP

[CS94] Lucia Caramellino and Fabio Spizzichino. Dependence and aging properties of lifetimes with Schur-constant survival functions. Probability in the Engineering and Informational Sciences, 8(1):103–111, January 1994. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
A472C5C7C2ABAEF337BC338F5166E579.

Corcoran:2003:SSC

[CS03] J. N. Corcoran and U. Schneider. Shift and scale coupling methods for perfect simulation. Probability in the Engineering and Informational Sciences, 17(3):277–303, July 2003. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
2E2BBA6D18BE59318ADDA2A40E2BC0DE.

Chhetry:1989:CSD

[CSK89] Devendra Chhetry, Allan R. Sampson, and George Kimeldorf. Concepts of setwise dependence. Probability in the Engineering and Informational Sciences, 3(3):367–380, July 1989. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
7F359A7A2428DBBF58465D1E5CF878CA.

Chen:2009:NAS

[CSL09] Xin Chen and David Simchi-Levi. A new approach for the stochastic cash balance problem with fixed costs. Probability in the Engineering and Informational Sciences, 23(4):545–562, October 2009. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/F2F0DF41615B39217207627AC497678F.

Chaudhry:2012:AER

[CSP12] M. L. Chaudhry, S. K. Samanta, and A. Pacheco. Analytically explicit results for the GI/C-MSP/1/∞ queueing system using roots. Probability in the Engineering and Informational Sciences, 26(2):221–244, April 2012. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
1ED5FB5B13AE662BFF23693122ED9604.

Coolen-Schrijner:2002:DMC

[CSvD02] Pauline Coolen-Schrijner and Erik A. van Doorn. The deviation matrix of a continuous-time Markov chain. Probability in the Engineering and Informational Sciences, 16(3):351– 366, July 2002. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/4B4CBA64C5FEEAE6A887BE05C8807153.

Cai:2003:SSP

[CSZ03] Xiaoqiang Cai, Xiaoqian Sun, and Xian Zhou. Stochastic scheduling with preemptive-repeat machine breakdowns to minimize the expected weighted flow time. Probability in the Engineering and Informational Sciences, 17(4):467–485, October 2003. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/F4CAED7D1612BCA8A6EC8F1A97095240.

Chiera:2002:WUC

[CT02] B. A. Chiera and P. G. Taylor. What is a unit of capacity worth? Probability in the Engineering and Informational Sciences, 16(4):513–522, October 2002. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
48FD60282864AA65B0AAECB5D33C8656.

Chadjiconstantinidis:2023:IBS

[CT23] Stathis Chadjiconstantinidis and George Tzavelas. Improved bounds for the solutions of renewal equations. Probability in the Engineering and Informational Sciences, 37(3):740–777, July 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/improved-bounds-for-the-solutions-of-renewal-
equations/6A0EC816E95333CD5B6043798C7DCD2C.

Corujo:2022:FRS

[CV22] Josué Corujo and José E. Valdés. Further results on stochastic orderings and aging classes in systems with age replacement. Probability in the Engineering and Informational Sciences, 36(3):660–689, July 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/further-
results-on-stochastic-orderings-and-aging-classes-in-
systems-with-age-replacement/B962335EFDB2F1D31E3DC5C42BAD948D.

Campos:2005:APT

[CVB05] F. A. Campos, J. Villar, and J. Barquín. Application of possibility theory to robust Cournot equilibrium in the electricity market. Probability in the Engineering and Informational Sciences, 19(4):519–531, October 2005. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
62187344DEE8DA2344BDFCB40F4CB5BF.

Courcoubetis:1990:SLP

[CW90a] Coastas Courcoubetis and Richard Weber. Stability of on-line bin packing with random arrivals and long-run-average constraints. Probability in the Engineering and Informational Sciences, 4(4):447–460, October 1990. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
144B51CD23274E98F81DFF0CE233DC46.

Courcoubetis:1990:MFR

[CW90b] Costas Courcoubetis and Richard R. Weber. The move-to-front rule for multiple lists. Probability in the Engineering and Informational Sciences, 4(1):19–27, January 1990. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
1709B866B64EC22E4FA93AF73D90036F.

Courcoubetis:1995:EBS

[CW95] Costas Courcoubetis and Richard Weber. Effective bandwidths for stationary sources. Probability in the Engineering and Informational Sciences, 9(2):285–296, April 1995. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
BA3762BBF5724EFD8C917EBEBE0CFDB6.

Chen:2022:SVP

[CW22] Yan Chen and Ward Whitt. Set-valued performance approximations for the queue given partial information. Probability in the Engineering and Informational Sciences, 36(2):378–400, April 2022. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
journals/probability-in-the-engineering-and-informational-
sciences/article/setvalued-performance-approximations-
for-the-gigik-queue-given-partial-information/3D5EF2850B5C7041CB4548354A51775B.

Cai:2023:VVA

[CW23] Chengyou Cai and Xingchun Wang. Valuing vulnerable Asian options with liquidity risk under Lévy processes. Probability in the Engineering and Informational Sciences, 37(3):653–673, July 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/valuing-
vulnerable-asian-options-with-liquidity-risk-under-
levy-processes/906FEA7F7FB1EFD4C7476B14ED35A9B5.

Chen:2023:CDS

[CX23] Yu Chen and Yu Xing. Credit default swap pricing with counterparty risk in a reduced form model with a common jump process. Probability in the Engineering and Informational Sciences, 37(1):275–293, January 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/credit-
default-swap-pricing-with-counterparty-risk-in-a-reduced-
form-model-with-a-common-jump-process/ADF79CB813DE6DE27ED637A23A3A6586.

Chao:2012:OPP

[CXY12] Xiuli Chao, Yifan Xu, and Baimei Yang. Optimal policy for a production-inventory system with setup cost and average cost criterion. Probability in the Engineering and Informational Sciences, 26(4):457–481, October 2012. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
76DB4BCB13288B211C3749296A14177F.

Chan:2006:UBR

[CY06] Gary K. C. Chan and Hailiang Yang. Upper bounds for ruin probability under time series models. Probability in the Engineering and Informational Sciences, 20(3):529–542, July 2006. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
D42576CFFF6EFBC205A9524D0C539A46.

Caro:2010:IBP

[CY10] Felipe Caro and Onesun Steve Yoo. Indexability of bandit problems with response delays. Probability in the Engineering and Informational Sciences, 24(3):349–374, July 2010. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
DC002FB2397C619F78641EFB2FCB48CF.

Chen:2021:SVM

[CY21] Pengzhan Chen and Wuyi Ye. Stochastic volatility model with correlated jump sizes and independent arrivals. Probability in the Engineering and Informational Sciences, 35(3):513–531, July 2021. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
journals/probability-in-the-engineering-and-informational-

sciences/article/stochastic-volatility-model-with-correlated-
jump-sizes-and-independent-arrivals/FDEB2C0A934A76D37241F0493ED4BFE1.

Chao:2003:TAI

[CZ03] Xiuli Chao and Yuxi Zheng. Transient analysis of immigration birth–death processes with total catastrophes. Probability in the Engineering and Informational Sciences, 17(1):83–106, January 2003. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/216959303AAD75F20B132075447D5A40.

Cai:2006:SSA

[CZ06] Xiaoqiang Cai and Xian Zhou. Stochastic scheduling with asymmetric earliness and tardiness penalties under random machine breakdowns. Probability in the Engineering and Informational Sciences, 20(4):635–654, October 2006. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
F6E129414F9D91DA060DF57F2FCD936B.

Cockett:1989:SDT

[CZB89] J. Robin B. Cockett, Jalel Zrida, and J. Douglas Birdwell. Stochastic decision theory. Probability in the Engineering and Informational Sciences, 3(1):13–54, January 1989. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
F6426DEBF4E27B088A39050D76CFE97F.

Chao:1996:EEP

[CZL96] Chern-Ching Chao, Lincheng Zhao, and Wen-Qi Liang. Estimating the error of a permutational central limit theorem. Probability in the Engineering and Informational Sciences, 10(4):533–541, October 1996. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/B07E5D27DF5A36E420C1E8EDE6203082.

Dube:2003:WPF

[DA03] Parijat Dube and Eitan Altman. On the workload process in a fluid queue with a responsive bursty input and selective discarding. Probability in the Engineering and Informational Sciences, 17(4):527–543, October 2003. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
C3F13BA97E72BAF0392CD50ED5164B5A.

David:1995:SAM

[Dav95] Israel David. A sequential assignment match process with general renewal arrival times. Probability in the Engineering and Informational Sciences, 9(3):475–492, July 1995. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
42403B5920ED7EE9D3C155DC9CED8A24.

Bueno:2022:CRI

[dCBB22] Vanderlei da Costa Bueno and Narayanaswamy Balakrishnan. A cumulative residual inaccuracy measure for coherent systems at component level and under nonhomogeneous Poisson processes. Probability in the Engineering and Informational Sciences, 36(2):294–319, April 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/cumulative-
residual-inaccuracy-measure-for-coherent-systems-at-
component-level-and-under-nonhomogeneous-poisson-processes/
4B2B0F5E0D75BD2F95290357C6F5C137.

deCooman:2009:IMC

[dCHQ09] Gert de Cooman, Filip Hermans, and Erik Quaeghebeur. Imprecise Markov chains and their limit behavior. Probability in the Engineering and Informational Sciences, 23(4):597–635, October 2009. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/DF646257C294C16C2733F7C563F73730.

Dobson:2005:EFP

[DCL^{+}05] Ian Dobson, Benjamin A. Carreras, Vickie E. Lynch, Bertrand Nkei, and David E. Newman. Estimating failure propagation in models of cascading blackouts. Probability in the Engineering and Informational Sciences, 19(4):475–488, October 2005. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
597E0243FA60E5B91E7495FF25466838.

Dobson:2005:LDM

[DCN05] Ian Dobson, Benjamin A. Carreras, and David E. Newman. A loading-dependent model of probabilistic cascading failure. Probability in the Engineering and Informational Sciences, 19(1):15–32, January 2005. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
70CF31827E64E4F3785A3EA234EE5613.

Das:2022:SRB

[DDK22] Sudipta Das, Anup Dewanji, and Subrata Kundu. Software reliability based on renewal process modeling for error occurrence due to each bug with periodic debugging schedule. Probability in the Engineering and Informational Sciences, 36(1):87–104, January 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/software-
reliability-based-on-renewal-process-modeling-for-error-
occurrence-due-to-each-bug-with-periodic-debugging-
schedule/37DD74EAE0126519C82EAE8633F1F621.

Ding:2014:CSP

[DDL14] Weiyong Ding, Gaofeng Da, and Xiaohu Li. Comparisons of series and parallel systems with heterogeneous components. Probability in the Engineering and Informational Sciences, 28(1):39–53, January 2014. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/B537E37C9A58E79FC2D3C1CE1E6A2F11.

Diaconis:1990:ETS

[DF90] Persi Diaconis and James Allen Fill. Examples for the theory of strong stationary duality with countable state spaces. Probability in the Engineering and Informational Sciences, 4(2): 157–180, April 1990. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/B7F63B81E44F493F6D250B0A617EE887.

Denuit:2008:CDF

[DF08] Michel Denuit and Esther Frostig. Comparison of dependence in factor models with application to credit risk portfolios. Probability in the Engineering and Informational Sciences, 22(1): 151–160, January 2008. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/162D79C181FE4200754A41823817A9E1.

Doncel:2021:EPN

[DF21] Josu Doncel and Jean-Michel Fourneau. Energy packet networks with multiple energy packet requirements. Probability in the Engineering and Informational Sciences, 35(1):92–110, January 2021. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
journals/probability-in-the-engineering-and-informational-
sciences/article/energy-packet-networks-with-multiple-
energy-packet-requirements/2B2DABDB90882262AEE9ABF86CD968AE.

Durante:2010:AFM

[DFS10] Fabrizio Durante, Rachele Foschi, and Fabio Spizzichino. Aging functions and multivariate notions of NBU and IFR. Probability in the Engineering and Informational Sciences, 24(2): 263–278, April 2010. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/A85D308F267EC8252A58C25269633B16.

DeSantis:2015:RBS

[DFS15] Emilio De Santis, Fabio Fantozzi, and Fabio Spizzichino. Relations between stochastic orderings and generalized stochastic precedence. Probability in the Engineering and Informational Sciences, 29(3):329–343, July 2015. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
91931F811A4FA157C159AC71D41B376D.

Ding:1991:BSB

[DG91a] Jie Ding and Betsy S. Greenberg. Bowl shapes are better with buffers — sometimes. Probability in the Engineering and Informational Sciences, 5(2):159–169, April 1991. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
586C00A94ABF6D62F70962193DE361FB.

Ding:1991:OOS

[DG91b] Jie Ding and Betsy S. Greenberg. Optimal order for servers in series with no queue capacity. Probability in the Engineering and Informational Sciences, 5(4):449–461, October 1991. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
84823802B8AE8DE41A0687AFE303E98B.

Danninger:1994:TPO

[DG94] Gabriele Danninger and Walter J. Gutjahr. Test point optimization in a branching-process-based reliability model. Probability in the Engineering and Informational Sciences, 8(4):

591–609, October 1994. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/932EA6C00D6E746121625D97DC20793B.

Doncel:2022:MFG

[DGG22] Josu Doncel, Nicolas Gast, and Bruno Gaujal. A mean field game analysis of SIR dynamics with vaccination. Probability in the Engineering and Informational Sciences, 36(2):482–499, April 2022. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
journals/probability-in-the-engineering-and-informational-
sciences/article/mean-field-game-analysis-of-sir-dynamics-
with-vaccination/08625C59A0892A49F7E0C63F1385437D.

Du:1987:RCG

[DH87] D. Z. Du and F. K. Hwang. Reliabilities of consecutive-2 graphs. Probability in the Engineering and Informational Sciences, 1(3):293–298, July 1987. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/5326038BFD74DAB73BC3F30E3F288FED.

Du:1994:OAC

[DH94] D. Z. Du and F. K. Hwang. Optimal assemblies of consecutive-2 link systems. Probability in the Engineering and Informational Sciences, 8(4):511–520, October 1994. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
D6656D8B04DAEA0E6EEDD50E12DC1874.

Dantzer:2000:SBP

[DHR00] Jean-François Dantzer, Mostafa Haddani, and Philippe Robert. On the stability of a bandwidth packing algorithm. Probability in the Engineering and Informational Sciences, 14(1):57–79, January 2000. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/DA0C2D7B69D22A0495CE8B1E878E3E59.

denIseger:2006:NTI

[dI06] Peter den Iseger. Numerical transform inversion using Gaussian quadrature. Probability in the Engineering and Informational Sciences, 20(1):1–44, January 2006. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
ADB4E45A3ADF44C5F7808F14C7A45B5D.

Dimitriou:2017:TCR

[Dim17] Ioannis Dimitriou. A two-class retrial system with coupled orbit queues. Probability in the Engineering and Informational Sciences, 31(2):139–179, April 2017. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
A57363FEFDE65E80DB878EF3D747D755.

Dimitriou:2023:GJS

[Dim23] Ioannis Dimitriou. The generalized join the shortest orbit queue system: stability, exact tail asymptotics and stationary approximations. Probability in the Engineering and Informational Sciences, 37(1):154–191, January 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/generalized-
join-the-shortest-orbit-queue-system-stability-exact-
tail-asymptotics-and-stationary-approximations/088F0095A5B89F42F56D0A37F624F9FC.

Dinwoodie:1995:BRC

[Din95] I. H. Dinwoodie. A bound on the rate of convergence for the discrete Gibbs sampler. Probability in the Engineering and Informational Sciences, 9(2):211–215, April 1995. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
AEF3D3C4A6DF288036362E50D0ECA2E5.

Dai:2007:SOC

[DJ07] Wanyang Dai and Qian Jiang. Stochastic optimal control of ATO systems with batch arrivals via diffusion approximation. Probability in the Engineering and Informational Sciences, 21 (3):477–495, July 2007. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/BF6219D4690811DE53C9A730434117A6.

Derman:1988:CTS

[DK88] Cyrus Derman and Showlong P. Koh. On the comparison of two software reliability estimators. Probability in the Engineering and Informational Sciences, 2(1):15–21, January 1988. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
53640738F86E8BDD75FCB407DFD90286.

deKok:1989:MIM

[dK89] A. G. de Kok. A moment-iteration method for approximating the waiting-time characteristics of the GI/G/1 queue. Probability in the Engineering and Informational Sciences, 3(2): 273–287, April 1989. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/99BA144501C3D361F6330362884867E7.

Deijfen:2009:RIG

[DK09] Maria Deijfen and Willemien Kets. Random intersection graphs with tunable degree distribution and clustering. Probability in the Engineering and Informational Sciences, 23(4): 661–674, October 2009. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/32954CFB72F5B470A207042D05492991.

Dewan:2013:SNA

[DK13] Isha Dewan and Subhash Kochar. Some new applications of P –P plots. Probability in the Engineering and Informational Sciences, 27(3):353–366, July 2013. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
281522853DDA4865F34DED16D7903E36.

Ding:2022:OCC

[DK22] Sihan Ding and Ger Koole. Optimal call center forecasting and staffing. Probability in the Engineering and Informational Sciences, 36(2):254–263, April 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/optimal-
call-center-forecasting-and-staffing/52796A0DC4D999FDAF019E03DAD100BD.

Das:2022:ORS

[DKB22] Sangita Das, Suchandan Kayal, and N. Balakrishnan. Ordering results for smallest claim amounts from two portfolios of risks with dependent heterogeneous exponentiated locationscale claims. Probability in the Engineering and Informational Sciences, 36(4):1116–1137, October 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/ordering-

results-for-smallest-claim-amounts-from-two-portfolios-
of-risks-with-dependent-heterogeneous-exponentiated-
locationscale-claims/80AF4B5F18BD07CC96607B59667FA3BB.

Das:2021:ORE

[DKC21] Sangita Das, Suchandan Kayal, and Debajyoti Choudhuri. Ordering results on extremes of exponentiated location-scale models. Probability in the Engineering and Informational Sciences, 35(2):331–354, April 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/ordering-
results-on-extremes-of-exponentiated-locationscale-
models/ECD11748DFE198752A1886269EF770AA.

Desai:1994:QSC

[DKK94] Madhav Desai, Sunil Kumar, and P. R. Kumar. Quasistatically cooled Markov chains. Probability in the Engineering and Informational Sciences, 8(1):1–19, January 1994. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
E72F88C4456574AB7DEF41EDFCEAE070.

Dehghanian:2016:SDJ

[DKM16] Amin Dehghanian, Jeffrey P. Kharoufeh, and Mohammad Modarres. Strategic dynamic jockeying between two parallel queues. Probability in the Engineering and Informational Sciences, 30(1):41–60, January 2016. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
64009FF6A5709174976919787EBAC616.

Daduna:2006:DOQ

[DKSS06] Hans Daduna, Rafał Kulik, Cornelia Sauer, and Ryszard Szekli. Dependence ordering for queuing networks with breakdown and repair. Probability in the Engineering and Informational Sciences, 20(4):575–594, October 2006. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
78F550494C31EDD5D4084AA3E95116A0.

Down:2010:NMU

[DL10] Douglas G. Down and Mark E. Lewis. The n-network model with upgrades. Probability in the Engineering and Informational Sciences, 24(2):171–200, April 2010. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
90D6B8149EBDFE38F4711B3BCFB7CB7B.

Ding:2012:MPR

[DLB12] Weiyong Ding, Xiaohu Li, and Narayanaswamy Balakrishnan. Monotonicity properties of residual lifetimes of parallel systems and inactivity times of series systems with heterogeneous components. Probability in the Engineering and Informational Sciences, 26(1):61–75, January 2012. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
901BEB094EB8D81AD2431AE2FA4AB3A8.

Durham:1990:OIP

[DLP90] S. Durham, J. Lynch, and W. J. Padgett. TP_{2}-orderings and the IFR property with applications. Probability in the Engineering and Informational Sciences, 4(1):73–88, January 1990. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
31ABBA7895F11F947DC6A68617D0EF9A.

Derman:1987:SIP

[DLR87] Cyrus Derman, Gerald J. Lieberman, and Sheldon M. Ross. On sampling inspection in the presence of inspection errors. Probability in the Engineering and Informational Sciences, 1 (2):237–249, April 1987. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/D757504297752AF7AAEDF0B7F2643628.

Denuit:1999:GSC

[DLU99] Michel Denuit, Claude Lefèvre, and Sergey Utev. Generalized stochastic convexity and stochastic orderings of mixtures. Probability in the Engineering and Informational Sciences, 13 (3):275–291, July 1999. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/8186325823C47CB48C76E5F1117B8E72.

Dieker:2003:SSF

[DM03] A. B. Dieker and M. Mandjes. On spectral simulation of fractional Brownian motion. Probability in the Engineering and Informational Sciences, 17(3):417–434, July 2003.

CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
38D5A221D899318B206369DD96D4FE34.

Dukhovny:2008:SRW

[DM08] Alexander Dukhovny and Jean-Luc Marichal. System reliability and weighted lattice polynomials. Probability in the Engineering and Informational Sciences, 22(3):373–388, July 2008. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
D515BB8F34F9A078B926A57FDFE855FC.

Domicolo:2020:DBG

[DM20] Carly Domicolo and Hosam Mahmoud. Degree-based Gini index for graphs. Probability in the Engineering and Informational Sciences, 34(2):157–171, April 2020. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/degreebased-
gini-index-for-graphs/A6F47E7A8624FF61247FB8D9DBC29AFD.

Desmarais:2022:DHN

[DM22] Colin Desmarais and Hosam Mahmoud. Depths in hooking networks. Probability in the Engineering and Informational Sciences, 36(4):941–949, October 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/depths-
in-hooking-networks/5FC28C8EF358F22F9A930CD6A1944B1D.

DiCrescenzo:2016:QBP

[DMM16] Antonio Di Crescenzo, Barbara Martinucci, and Julio Mulero. A quantile-based probabilistic mean value theorem. Probability in the Engineering and Informational Sciences, 30(2): 261–280, April 2016. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/B70EA646FFF9DD096CBBC7AD2AB53551.

Debicki:2007:TQL

[DMvU07] Krzysztof Debicki, Michel Mandjes, and Miranda van Uitert. A tandem queue with lévy input: a new representation of the downstream queue length. Probability in the Engineering and Informational Sciences, 21(1):83–107, January 2007.

CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
1A687EE678AE781D4B48536C7E68F6F1.

Deng:2003:IOC

[DO03] Shi-Jie Deng and Shmuel S. Oren. Incorporating operational characteristics and start-up costs in option-based valuation of power generation capacity. Probability in the Engineering and Informational Sciences, 17(2):155–181, April 2003. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
A4B8115C0F2BA1BEA7CE2EA0DC8B2C0D.

Dayar:2016:CPP

[DO16] Tuğrul Dayar and M. Can Orhan. Cartesian product partitioning of multi-dimensional reachable state spaces. Probability in the Engineering and Informational Sciences, 30(3): 413–430, July 2016. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/54ED8D6599252681630BBDF134634436.

Dodin:1989:ASN

[Dod89] Bajis Dodin. Approximating the stochastic network by its m shortest paths. Probability in the Engineering and Informational Sciences, 3(3):435–451, July 1989. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
8E6F22127F847101778658BBD6F902FB.

Doverspik:1987:TBR

[Dov87] R. D. Doverspik. A terminal Bay relief strategy. Probability in the Engineering and Informational Sciences, 1(4):481–496, October 1987. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/98F652CD496352657BD6519A66362FF5.

DeVeciana:1993:LDB

[DOW93] G. De Veciana, C. Olivier, and J. Walrand. Large deviations of birth death Markov fluids. Probability in the Engineering and Informational Sciences, 7(2):237–255, April 1993. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
A337F073B587A98139F1D5C84463C3BD.

Domine:1993:FPT

[DP93] Marco Dominé and Volkmar Pieper. First passage time distribution of a two-dimensional Wiener process with drift. Probability in the Engineering and Informational Sciences, 7(4):545– 555, October 1993. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/D39A150AD3AB9D579572863C9135E791.

Du:1995:CPM

[DP95] Charles Du and Michael Pinedo. A convexity property of a Markov-modulated queueing loss system. Probability in the Engineering and Informational Sciences, 9(2):193–199, April 1995. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
A73B13456F298864C9E494BF193C66F3.

Dimitriou:2021:RAG

[DP21] Ioannis Dimitriou and Nikolaos Pappas. A random access g-network: stability, stable throughput, and queueing analysis. Probability in the Engineering and Informational Sciences, 35(1):111–137, January 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/random-
access-gnetwork-stability-stable-throughput-and-queueing-
analysis/AF44727D1DF966002BD93B89F6230B15.

Dorsman:2016:SWT

[DPV16] Jan-Pieter L. Dorsman, Nir Perel, and Maria Vlasiou. Server waiting times in infinite supply polling systems with preparation times. Probability in the Engineering and Informational Sciences, 30(2):153–184, April 2016. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
012BEF2B1019E1CAD1F9E2B80DE5F9D9.

Derman:1995:IEQ

[DR95a] Cyrus Derman and Sheldon Ross. An improved estimator of σ in quality control. Probability in the Engineering and Informational Sciences, 9(3):411–415, July 1995. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
0ACF68EFB0541F1CF1E799B247995377.

Derman:1995:RSD

[DR95b] Cyrus Derman and Sheldon Ross. Reliability for a system with dependent components. Probability in the Engineering and Informational Sciences, 9(1):59–63, January 1995. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
8A38EE4A3FBDF21AE0331E03027CE82E.

Desai:1997:FTB

[DR97] Madhav P. Desai and Vasant B. Rao. Finite-time behavior of slowly cooled annealing chains. Probability in the Engineering and Informational Sciences, 11(2):137–176, April 1997. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
F1D59EDFC1BDA3B8C21CCDCB2D5B18E1.

Dizon-Ross:2022:PFN

[DRR22] Rebecca Dizon-Ross and Sheldon M. Ross. A probabilistic friendship network model. Probability in the Engineering and Informational Sciences, 36(1):1–16, January 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/probabilistic-friendship-network-model/DB32B9939580C01450058F3054340B44.

Diaconis:1987:SAG

[DS87] Persi Diaconis and Mehrdad Shahshahani. The subgroup algorithm for generating uniform random variables. Probability in the Engineering and Informational Sciences, 1(1):15–32, January 1987. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/5F15097B60994F362F6926689FA8A465.

Dubins:1998:DRB

[Dub98] Lester E. Dubins. Discrete red-and-black with fortunedependent win probabilities. Probability in the Engineering and Informational Sciences, 12(4):417–424, October 1998. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
6B57AB14E7FDAE889B97A75C983BD95C.

Dudin:1997:OCM

[Dud97] Alexander Dudin. Optimal control for an M^{x} /G/1 queue with two operation modes. Probability in the Engineering and Informational Sciences, 11(2):255–265, April 1997. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
3554944C33AFBC08D9702E38E1CDC892.

Diks:1997:MSL

[DvdH97] E. B. Diks and M. C. van der Heijden. Modeling stochastic lead times in multi-echelon systems. Probability in the Engineering and Informational Sciences, 11(4):469–485, October 1997. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
45F25BAEDD96CB9C5E32AC289308908C.

Dietz:2011:ARG

[DvdLR11] Chris Dietz, Dinard van der Laan, and Ad Ridder. Approximate results for a generalized secretary problem. Probability in the Engineering and Informational Sciences, 25(2):157–169, April 2011. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/1C955CDFCB89D0E140A9B5C59EE56F4B.

Da:2016:QSD

[DXX16] Gaofeng Da, Maochao Xu, and Shouhuai Xu. On the quasistationary distribution of SIS models. Probability in the Engineering and Informational Sciences, 30(4):622–639, October 2016. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
9F724E72DBBD36471F7D86535F54196E.

David:1990:SAM

[DY90a] Israel David and Uri Yechiali. Sequential assignment match processes with arrivals of candidates and offers. Probability in the Engineering and Informational Sciences, 4(4):413–430, October 1990. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/2FDF35E878F25776157043CE8ED7393A.

Durham:1990:RPL

[DY90b] S. D. Durham and K. F. Yu+. Randomized play-the-leader rules for sequential sampling from two populations. Probability in the Engineering and Informational Sciences, 4(3):355– 367, July 1990. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/6E27D8072E69A56E056EF6B09585C30B.

Du:2012:LRHa

[DZB12a] Baojun Du, Peng Zhao, and N. Balakrishnan. Likelihood ratio and hazard rate orderings of the maxima in two multiple-outlier geometric samples. Probability in the Engineering and Informational Sciences, 26(3):375–391, July 2012. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/EDBA497CEAA8A0FABE62767D3B9EA00A. See erratum

[DZB12b].

Du:2012:LRHb

[DZB12b] Baojun Du, Peng Zhao, and N. Balakrishnan. Likelihood ratio and hazard rate orderings of the maxima in two multipleoutlier geometric samples — erratum. Probability in the Engineering and Informational Sciences, 26(4):613, October 2012. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
07E32DA41822A2626BA06817EB0D949C. See [DZB12a].

Da:2012:OCS

[DZH12] Gaofeng Da, Ben Zheng, and Taizhong Hu. Ordering consecutive k-out-of-n: F systems. Probability in the Engineering and Informational Sciences, 26(2):147–158, April 2012. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
5B634DD3B6E7AFF8FCA3D89B1066742D.

Ding:2018:OPS

[DZZ18] Weiyong Ding, Yiying Zhang, and Peng Zhao. Ordering properties of spacings from heterogeneous geometric samples. Probability in the Engineering and Informational Sciences, 32(2):306–322, April 2018. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/ordering-
properties-of-spacings-from-heterogeneous-geometric-
samples/23D1CC7C3336587513DE232DD2EE353E.

Evrenosoglu:2006:EMP

[EA06] Cansin Y. Evrenosoğlu and Ali Abur. Effects of measurement and parameter uncertainties on the power transfer distribution factors. Probability in the Engineering and Informational Sciences, 20(1):175–182, January 2006. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
52973741C0F1C45B591818A2A6BA06D1.

El-Baz:2022:CSF

[EBTD22] A. H. El-Baz, A. M. K. Tarabia, and A. M. Darwiesh. Cloud storage facility as a fluid queue controlled by Markovian queue. Probability in the Engineering and Informational Sciences, 36(2):237–253, April 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/cloud-
storage-facility-as-a-fluid-queue-controlled-by-markovian-
queue/29952EBF48A7C91B5FECE8EE533A7F1E.

Economou:1999:GFB

[Eco99] Antonis Economou. Geometric-form bounds for the GI^{X} /M/1 queueing system. Probability in the Engineering and Informational Sciences, 13(4):509–520, October 1999. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
346409E89A75D02246B24073E3D856FC.

Economou:2003:NSC

[Eco03] Antonis Economou. Necessary and sufficient conditions for the stochastic comparison of Jackson networks. Probability in the Engineering and Informational Sciences, 17(1):143–151, January 2003. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/1552A4E82BFCB5BB091C4A0CA12BEF43.

Edwards:2017:IMW

[EFG17] James Edwards, Paul Fearnhead, and Kevin Glazebrook. On the identification and mitigation of weaknesses in the knowledge gradient policy for multi-armed bandits. Probability in the Engineering and Informational Sciences, 31(2):239–263, April 2017. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
F040D75B80CBEAD864EBEB90860BB787.

Egeland:1994:WSB

[Ege94] Thore Egeland. When is a system better than the average of its components? Probability in the Engineering and Informational Sciences, 8(3):431–434, July 1994. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
FFC3F0F8464388C643B1D3F857E48C2B.

Elperin:1992:EMM

[EGL92] T. Elperin, I. Gertsbakh, and M. Lomonosov. An evolution model for Monte Carlo estimation of equilibrium network renewal parameters. Probability in the Engineering and Informational Sciences, 6(4):457–469, October 1992. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
D5D0C3B23410BEEC38C55C965CBCFC8E.

Economou:2009:SMC

[EK09] Antonis Economou and Stella Kapodistria. Q-series in Markov chains with binomial transitions: studying a queue with synchronization. Probability in the Engineering and Informational Sciences, 23(1):75–99, January 2009. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
A3F1008A20B8193730ADF85973AF5630.

Eryilmaz:2021:NSM

[EK21] Serkan Eryilmaz and Cihangir Kan. A new shock model with a change in shock size distribution. Probability in the Engineering and Informational Sciences, 35(3):381–395, July 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/new-shock-model-with-a-change-in-shock-size-
distribution/17E0DEC0109C7DCD5980278FAD55092B.

Elcan:1994:OCR

[Elc94] Amie Elcan. Optimal customer return rate for an M/M/1 queueing system with retrials. Probability in the Engineering and Informational Sciences, 8(4):521–539, October 1994.

CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
20A23E2E9D33DBE8024F5FC42D2DCC80.

Eliazar:2005:ECT

[Eli05] Iddo Eliazar. Extremes: a continuous-time perspective. Probability in the Engineering and Informational Sciences, 19(3): 289–308, July 2005. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/950B59FF8735E3B86691655828B62DE5.

Eliazar:2008:DTG

[Eli08] Iddo Eliazar. On the discrete-time G/GI/∞ queue. Probability in the Engineering and Informational Sciences, 22(4):557–585, October 2008. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/A5E87D271153F772A82FE247BF27F786.

Ezhilchelvan:2019:MCR

[EM19] Paul Ezhilchelvan and Isi Mitrani. Multi-class resource sharing with batch arrivals. Probability in the Engineering and Informational Sciences, 33(3):348–366, July 2019. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/multiclass-
resource-sharing-with-batch-arrivals/E7ADE4BA55BBAB80C93C297AAC6FC729.

Eftekharian:2022:EPD

[EQ22] Abbas Eftekharian and Guoxin Qiu. On extropy properties and discrimination information of different stratified sampling schemes. Probability in the Engineering and Informational Sciences, 36(3):644–659, July 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/on-extropy-
properties-and-discrimination-information-of-different-
stratified-sampling-schemes/F5AEAAE49F4D608B7AFD836BBFC3620E.

Eaves:1995:DCC

[ER95] B. Curtis Eaves and Uriel G. Rothblum. A discounted-cost continuous-time flexible manufacturing and operator scheduling model solved by deconvexification over time. Probability in the Engineering and Informational Sciences, 9(1):65–98, January 1995. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/B7DC4CBC64524B558F28EB475F8CFFBA.

ElNeweihi:1991:SRM

[ES91] Emad El Neweihi and Jayaram Sethuraman. A study of the role of modules in the failure of systems. Probability in the Engineering and Informational Sciences, 5(2):215–227, April 1991. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
B2366D75CE12926D5C5A34C692BC0C2B.

Eikeboom:1987:WTP

[ET87] A. M. Eikeboom and H. C. Tijms. Waiting-time percentiles in the multi-server M^{x} /G/c queue with batch arrivals. Probability in the Engineering and Informational Sciences, 1(1): 75–96, January 1987. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/709DD1E7B340EE35CCB93C52C8954C37.

Egorova:2006:STT

[EZB06] Regina Egorova, Bert Zwart, and Onno Boxma. Sojourn time tails in the M/D/1 processor sharing queue. Probability in the Engineering and Informational Sciences, 20(3):429–446, July 2006. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
03EC7F9104AEE1DC3EE569920167FC8E.

Fan:1996:NAS

[Fan96] Kaisheng Fan. On a new approach to the solution of the nonlinear filtering equation of jump processes. Probability in the Engineering and Informational Sciences, 10(1):153–163, January 1996. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
429EA32CF904BE705388E95AE923488E.

Finkelstein:2022:DBR

[FC22] Maxim Finkelstein and Ji Hwan Cha. On degradation-based remaining lifetime. Probability in the Engineering and Informational Sciences, 36(3):812–823, July 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/on-degradationbased-
remaining-lifetime/1EF00FCAC79A047D81CAE5A3430B76D3.

Finkelstein:2001:MFR

[FE01] Max S. Finkelstein and Veronica Esaulova. Modeling a failure rate for a mixture of distribution functions. Probability in the Engineering and Informational Sciences, 15(3):383–400, July 2001. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
29EC5FED71EA1421579947BD16E81E10.

Fendick:2022:BMM

[Fen22] Kerry Fendick. Brownian motion minus the independent increments: representation and queuing application. Probability in the Engineering and Informational Sciences, 36(1):144–168, January 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/brownian-
motion-minus-the-independent-increments-representation-
and-queuing-application/2DD2991D85C2722A53C2A1AC7445F755.

Fay:1989:GMS

[FG89a] N. A. Fay and K. D. Glazebrook. A general model for the scheduling of alternative stochastic jobs that may fail. Probability in the Engineering and Informational Sciences, 3(2): 199–221, April 1989. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/1AE57E73528420291381DF48D4EB5660.

Fox:1989:RSL

[FG89b] Bennett L. Fox and Peter W. Glynn. Replication schemes for limiting expectations. Probability in the Engineering and Informational Sciences, 3(3):299–318, July 1989. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
4909F8C5E408F616154478DA94348F46.

Fox:1991:EDP

[FG91] Bennett L. Fox and Paul Glasserman. Estimating derivatives via Poisson’s equation. Probability in the Engineering and Informational Sciences, 5(4):415–428, October 1991. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
3BEA71BBA170916FF57AFE4E6090F7DF.

Fricker:2010:IPU

[FGR10] Christine Fricker, Fabrice Guillemin, and Philippe Robert. An identification problem in an urn and ball model with heavy tailed distributions. Probability in the Engineering and Informational Sciences, 24(1):77–97, January 2010. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
C069F83D8752FB481938B6D1D1EACF1B.

Fu:1994:ISR

[FH94] Michael C. Fu and Jian-Qiang Hu. ( s, s ) inventory systems with random lead times: Harris recurrence and its implications in sensitivity analysis. Probability in the Engineering and Informational Sciences, 8(3):355–376, July 1994. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
D6E92BEEEEE31BC326560BE61D8B1345.

Fu:1995:SAM

[FH95] Michael C. Fu and Jian-Qlang Hu. Sensitivity analysis for Monte Carlo simulation of option pricing. Probability in the Engineering and Informational Sciences, 9(3):417–446, July 1995. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
B27F770BCBBCC528B3E97578F040057B.

Feng:2013:SSA

[FH13] Tianke Feng and Joseph C. Hartman. The sequential stochastic assignment problem with postponement options. Probability in the Engineering and Informational Sciences, 27(1):25–51, January 2013. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/8A18D73EF152CE0F4BCA4681ED3406B5.

Feng:2017:ADD

[FH17a] Qunqiang Feng and Zhishui Hu. The asymptotic degree distributions of random fast growth models for treelike networks. Probability in the Engineering and Informational Sciences, 31 (2):180–195, April 2017. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/23E43F27BC56F2A49FEF2CEC10EA4711.

Finkelstein:2017:SCL

[FH17b] Maxim Finkelstein and Nil Kamal Hazra. On stochastic comparisons for load-sharing series and parallel systems. Probability in the Engineering and Informational Sciences, 31(3): 311–329, July 2017. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/797465F7485C38398BFD050FB8219F3E.

Fang:1994:DDL

[FHJ94] Zhaoben Fang, Taizhong Hu, and Harry Joe. On the decrease in dependence with lag for stationary Markov chains. Probability in the Engineering and Informational Sciences, 8(3): 385–401, July 1994. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/C4B1FC9D35567768BE7544C3D4247084.

Feng:2013:ZIR

[FHS13] Qunqiang Feng, Zhishui Hu, and Chun Su. The Zagreb indices of random graphs. Probability in the Engineering and Informational Sciences, 27(2):247–260, April 2013. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
4AC19C2F16F57817574C2D96662B126E.

Fill:1991:TSC

[Fil91] James Allen Fill. Time to stationarity for a continuoustime Markov chain. Probability in the Engineering and Informational Sciences, 5(1):61–76, January 1991. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
BDCCC1463C6223EDA28BB166D4538862.

Fill:1998:MFR

[Fil98] James Allen Fill. The move-to-front rule: A case study for two perfect sampling algorithms. Probability in the Engineering and Informational Sciences, 12(3):283–302, July 1998. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
494E9284D5687CA086CED1C280C74D69.

Fishman:1991:SAS

[Fis91] George S. Fishman. Sensitivity analysis for the system reliability function. Probability in the Engineering and Informational Sciences, 5(2):185–213, April 1991. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
6F5EB1660E0F22A02D83614C08CBCC17.

Feinberg:1996:BOM

[FK96] Eugene A. Feinberg and Dong J. Kim. Bicriterion optimization of an M/G/1 queue with a removable server. Probability in the Engineering and Informational Sciences, 10(1):57–73, January 1996. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
D08AE3FB13C6214A706F4DCFA4FA0251.

Frostig:2002:DOL

[FKL02] Esther Frostig, Doron Kliger, and Benny Levikson. The deferral option in long-term-care insurance. Probability in the Engineering and Informational Sciences, 16(1):101–119, January 2002. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
1FE0378BF5468B276E73DCAB86891201.

Foss:1998:SGP

[FL98] Serguei Foss and Günter Last. On the stability of greedy polling systems with general service policies. Probability in the Engineering and Informational Sciences, 12(1):49–68, January 1998. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
8CDC52D41B17937DD698AA13D3D75C87.

Feinberg:2005:OFT

[FL05] Eugene A. Feinberg and Mark E. Lewis. Optimality of fourthreshold policies in inventory systems with customer returns and borrowing/storage options. Probability in the Engineering and Informational Sciences, 19(1):45–71, January 2005. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
B92D5FCA0E55EDBD1A5EA3E1E36210E0.

Fang:2013:NBD

[FL13] Rui Fang and Xiaohu Li. A note on bivariate dual generalized Marshall–Olkin distributions with applications. Probability in the Engineering and Informational Sciences, 27(3): 367–374, July 2013. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/DC62339EFBBE3E7D76C3C09D7D945B9C.

Fang:2020:ARA

[FL20a] Rui Fang and Xiaohu Li. Active redundancy allocation for coherent systems with independent and heterogeneous components. Probability in the Engineering and Informational Sciences, 34(1):72–91, January 2020. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/active-
redundancy-allocation-for-coherent-systems-with-independent-
and-heterogeneous-components/B1BC2833C37909D4A17D0F0D4DC48ADF.

Feinberg:2020:SSC

[FL20b] Eugene A. Feinberg and Yan Liang. Stochastic setup-cost inventory model with backorders and quasiconvex cost functions. Probability in the Engineering and Informational Sciences, 34(3):429–468, July 2020. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/stochastic-
setupcost-inventory-model-with-backorders-and-quasiconvex-
cost-functions/CB66A8CC22F400F16D2AD1BA7D441236.

Fu:2022:RPM

[FL22] Ke-Ang Fu and Yang Liu. Ruin probabilities for a multidimensional risk model with non-stationary arrivals and subexponential claims. Probability in the Engineering and Informational Sciences, 36(3):799–811, July 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/ruin-
probabilities-for-a-multidimensional-risk-model-with-
nonstationary-arrivals-and-subexponential-claims/3A304D45A8848B7581FC88D0662F613E.

Flatto:1989:LQM

[Fla89] Leopold Flatto. The longer queue model. Probability in the Engineering and Informational Sciences, 3(4):537–559, October 1989. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
12A1220517AB7CF1B1BBC7FDF4948962.

Fralix:2012:RSP

[FLL12] Brian Fralix, James Livsey, and Robert Lund. Renewal sequences with periodic dynamics. Probability in the Engineering and Informational Sciences, 26(1):1–15, January 2012. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
A838364CA9F298B1154D53F105EF0B85.

Fayolle:2016:MAI

[FM16] Guy Fayolle and Paul Muhlethaler. A Markovian analysis of IEEE 802.11 broadcast transmission networks with buffering. Probability in the Engineering and Informational Sciences, 30 (3):326–344, July 2016. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/CC811C0FFAB30FF1E68AB7FAF742B560.

Fourneau:2017:PSG

[FM17] J.-M. Fourneau and Y. Ait El Majhoub. Processor sharing G-queues with inert customers and catastrophes: a model for server aging and rejuvenation. Probability in the Engineering and Informational Sciences, 31(4):420–435, October 2017. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
326208E4627E3BA61D182174C8E6C450.

Feng:2017:DPH

[FMR17] Yarong Feng, Hosam Mahmoud, and Ludger Rüschendorf. Degree profile of hierarchical lattice networks. Probability in the Engineering and Informational Sciences, 31(1):60–82, January 2017. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
45CE692305BD448B63D242CC932B3281.

Fu:2020:PBT

[FNC20] Ke-Ang Fu, Chang Ni, and Hao Chen. A particular bidimensional time-dependent renewal risk model with constant interest rates. Probability in the Engineering and Informational Sciences, 34(2):172–182, April 2020. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/particular-
bidimensional-timedependent-renewal-risk-model-with-
constant-interest-rates/912BFD968413819DEA882BF53421D6BD.

Fan-Orzechowski:2007:ORT

[FOF07] Xiaofei Fan-Orzechowski and Eugene A. Feinberg. Optimality of randomized trunk reservation for a problem with multiple constraints. Probability in the Engineering and Informational Sciences, 21(2):189–200, April 2007. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
9792ED45F34498BA6179806AAE671F95.

Forst:1993:SSO

[For93] Frank G. Forst. Stochastic sequencing on one machine with earliness and tardiness penalties. Probability in the Engineering and Informational Sciences, 7(2):291–300, April 1993. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
FEDDFACA182661BC8C384A3194FA82FB.

Fourneau:2016:NUN

[Fou16] Jean-Michel Fourneau. G-networks of unreliable nodes. Probability in the Engineering and Informational Sciences, 30(3): 361–378, July 2016. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/CF1C9B60B9E94E811666888FFDEFAC4A.

Fernandez-Ponce:2008:CCR

[FPOP08] José María Fernández-Ponce, Eva María Ortega, and Franco Pellerey. Convex comparisons for random sums in random environments and applications. Probability in the Engineering and Informational Sciences, 22(3):389–413, July 2008. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
8BD90CEFC3775319236AAAF390B40DCC.

Fernandez-Ponce:2003:ACB

[FPSL03] J. M. Fernández-Ponce and A. Suárez-Llorens. An aging concept based on majorization. Probability in the Engineering and Informational Sciences, 17(1):107–117, January 2003. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
EB938C5A3E6FCAEA2A926194D84EE549.

Fishman:1992:EBC

[FR92] George S. Fishman and David S. Rubin. Evaluating best-case and worst-case coefficients of variation when bounds are available. Probability in the Engineering and Informational Sciences, 6(3):309–322, July 1992. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
0B07EC8054D917AB34114A0DBE493A54.

Feinberg:1994:ORT

[FR94] Eugene A. Feinberg and Martin I. Reiman. Optimality of randomized trunk reservation. Probability in the Engineering and Informational Sciences, 8(4):463–489, October 1994. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
FCC0F473C500D7BD7D58F1854685C4B5.

Francois:1999:SGT

[Fra99] Olivier François. On the spectral gap of a time reversible Markov chain. Probability in the Engineering and Informational Sciences, 13(1):95–101, January 1999. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
F7AC7D38AE1D4C8AAD9373E68B8C80E5.

Franx:2005:MDC

[Fra05] Geert Jan Franx. The M^{x} /D/c batch arrival queue. Probability in the Engineering and Informational Sciences, 19(3): 345–349, July 2005. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/23D86010758A60A23EC0412E3A893131.

Fralix:2013:TDS

[Fra13] Brian Fralix. A time-dependent study of the knockout queue. Probability in the Engineering and Informational Sciences, 27 (3):309–317, July 2013. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/E7684EA0FEC93130C3D49875630DE5E6.

Fralix:2019:TDB

[Fra19] Brian Fralix. On the time-dependent behavior of a Markovian reentrant-line model. Probability in the Engineering and Informational Sciences, 33(3):367–386, July 2019. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/on-the-
timedependent-behavior-of-a-markovian-reentrantline-
model/D21A9A089CE9AFCA33071731019A55B1.

Frenk:1991:GFS

[Fre91a] J. B. G. Frenk. A general framework for stochastic onemachine scheduling problems with zero release times and no partial ordering. Probability in the Engineering and Informational Sciences, 5(3):297–315, July 1991. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
286AEFDA810B3D2C48E605F63557A2DE.

Frenk:1991:NOM

[Fre91b] J. B. G. Frenk. A note on one-machine scheduling problems with imperfect information. Probability in the Engineering and Informational Sciences, 5(3):317–331, July 1991. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
F61CBB617964271D77EE2B7B5032F4C1.

Frostig:1991:NSS

[Fro91] Esther Frostig. A note on stochastic scheduling on a single machine subject to breakdown — the preemptive repeat model. Probability in the Engineering and Informational Sciences, 5 (3):349–354, July 1991. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/702FC1E76C2FD1F1DD4CAA1E33D5999C.

Frostig:1995:OAM

[Fro95] Esther Frostig. Optimal allocation of machines to distinguishable repairmen in order to maximize some reward functions. Probability in the Engineering and Informational Sciences, 9(4):633–645, October 1995. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
BB788373736450F8ECAC60914E084B15.

Frostig:1999:ORS

[Fro99] Esther Frostig. Optimal repair of a series system with fixed repair times. Probability in the Engineering and Informational Sciences, 13(4):477–487, October 1999. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
1D8830142716ADD6554FB5BDBFB24BAF.

From:2011:SNR

[Fro11] Steven G. From. Some new reliability bounds for sums of NBUE random variables. Probability in the Engineering and Informational Sciences, 25(1):83–102, January 2011. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
D10244D00BA904992FCA612B13B070BE.

Franco:2002:SOB

[FRR02] Manuel Franco, José M. Ruiz, and M. Carmen Ruiz. Stochastic orderings between spacings of generalized order statistics. Probability in the Engineering and Informational Sciences, 16(4):471–484, October 2002. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
5FA4F3EB615206C4A90B3BD4EBFF8DC9.

Fishman:1989:ERS

[FS89] George S. Fishman and Tien-Yi Danny Shaw. Evaluating reliability of stochastic flow networks. Probability in the Engineering and Informational Sciences, 3(4):493–509, October 1989. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
EB72EE52414CD0ADFFD2FA6FF4F5A1EF.

Fleming:1999:HTA

[FS99] Philip J. Fleming and Burton Simon. Heavy traffic approximations for a system of infinite servers with load balancing. Probability in the Engineering and Informational Sciences, 13 (3):251–273, July 1999. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/82388EC616A625A10FF2AE3B104C5C45.

From:2020:SNB

[FS20] Steven G. From and Andrew W. Swift. Some new bounds and approximations on tail probabilities of the Poisson and other discrete distributions. Probability in the Engineering and Informational Sciences, 34(1):53–71, January 2020. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/some-
new-bounds-and-approximations-on-tail-probabilities-
of-the-poisson-and-other-discrete-distributions/5D150AD3C15A0F8219340C7437513C3D.

Fishman:2023:ASO

[FS23] George S. Fishman and Shaler Stidham. An adaptive strategy for offering m-out-of-n insurance policies. Probability in the Engineering and Informational Sciences, 37(1):106–134, January 2023. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
journals/probability-in-the-engineering-and-informational-
sciences/article/an-adaptive-strategy-for-offering-
moutofn-insurance-policies/891596AF4CEB6FAE7087EC6882A22A4F.

Foss:1998:SIM

[FTC98] S. G. Foss, R. L. Tweedie, and J. N. Corcoran. Simulating the invariant measures of Markov chains using backward coupling at regeneration times. Probability in the Engineering and Informational Sciences, 12(3):303–320, July 1998. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
C632E375DEA69AA22EAA320B8FBBC42A.

Federgruen:2015:CRM

[FW15] Awi Federgruen and Min Wang. A continuous review model with general shelf age and delay-dependent inventory costs. Probability in the Engineering and Informational Sciences, 29(4):507–525, October 2015. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
1DEC01A650619BA0C76BF8840B54A0F6.

Fu:2000:NPA

[FWGD00] Michael C. Fu, Rongwen Wu, Gül Gürkan, and A. Yonca Demir. A note on perturbation analysis estimators for American-style options. Probability in the Engineering and Informational Sciences, 14(3):385–392, July 2000. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
312A1E892996938F9907573774FE8D71.

Feng:2021:SWE

[FWH21] Qunqiang Feng, Yongkang Wang, and Zhishui Hu. Small-world effect in geographical attachment networks. Probability in the Engineering and Informational Sciences, 35(2):276–296, April 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/

probability-in-the-engineering-and-informational-sciences/
article/smallworld-effect-in-geographical-attachment-
networks/B811DCC696AEE8CC482BB53BF2BFE95C.

Feinberg:2011:OTR

[FY11] Eugene A. Feinberg and Fenghsu Yang. Optimality of trunk reservation for an M/M/K/N queue with several customer types and holding costs. Probability in the Engineering and Informational Sciences, 25(4):537–560, October 2011. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
B3677BD08A5AD259579264DD464C0A17.

Federgruen:1993:OCP

[FZ93] A. Federgruen and Y.-S. Zheng. Optimal control policies for stochastic inventory systems with endogenous supply. Probability in the Engineering and Informational Sciences, 7(2): 257–272, April 1993. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/82EC39BCB510A0BAC8F1A59E86599394.

Feinberg:2015:OSE

[FZ15] Eugene A. Feinberg and Xiaoxuan Zhang. Optimal switching on and off the entire service capacity of a parallel queue. Probability in the Engineering and Informational Sciences, 29(4): 483–506, October 2015. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/5E2BC802137845B0052A41AC9B635BAD.

Frolkova:2019:FLP

[FZ19] Maria Frolkova and Bert Zwart. Fluid limit of a ps-queue with multistage service. Probability in the Engineering and Informational Sciences, 33(1):1–27, January 2019. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/fluid-
limit-of-a-psqueue-with-multistage-service/7507BB3D4E0A8FD12BE7B6E54FAAB44E.

Gal:1990:SM

[GA90] S. Gal and E. J. Anderson. Search in a maze. Probability in the Engineering and Informational Sciences, 4(3):311–318, July 1990. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
9BC82D10F41F29F6CA503171520BEB69.

Guo:2022:LAO

[GB22] Zhiyu Guo and Yizhou Bai. A lattice approach for option pricing under a regime-switching GARCH-jump model. Probability in the Engineering and Informational Sciences, 36(4):1138–1170, October 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/lattice-
approach-for-option-pricing-under-a-regimeswitching-
garchjump-model/A3B8BF75BDC93E1C73326F8A73C0F526.

Gayrard:2020:MED

[GCDB20] Emeline Gayrard, Cédric Chauvière, Hacène Djellout, and Pierre Bonnet. Modeling experimental data with polynomials chaos. Probability in the Engineering and Informational Sciences, 34(1):14–26, January 2020. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/modeling-
experimental-data-with-polynomials-chaos/066B350464EDBC62D760FBF9F27A71C2.

Gao:2018:HLO

[GD18] Xuefeng Gao and S. J. Deng. Hydrodynamic limit of orderbook dynamics. Probability in the Engineering and Informational Sciences, 32(1):96–125, January 2018. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/hydrodynamic-
limit-of-orderbook-dynamics/92B6A9FEC3792B2DF8DADDF26BD9E748.

Glynn:2010:HGU

[GDRV10] Peter W. Glynn, Andrey Dolgin, Reuven Y. Rubinstein, and Radislav Vaisman. How to generate uniform samples on discrete sets using the splitting method. Probability in the Engineering and Informational Sciences, 24(3):405–422, July 2010. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
0B3C256EA8973B413777A5121374AE05.

Gelenbe:1993:NSB

[Gel93] Erol Gelenbe. G-networks with signals and batch removal. Probability in the Engineering and Informational Sciences, 7 (3):335–342, July 1993. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/5E6E9172810894893FB71F5265B45379.

Goyal:2022:HDM

[GFH22] Dheeraj Goyal, Maxim Finkelstein, and Nil Kamal Hazra. On history-dependent mixed shock models. Probability in the Engineering and Informational Sciences, 36 (4):1080–1097, October 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/on-historydependent-
mixed-shock-models/54D3CAAC5CABC6C1A89FFB8AF2403E3B.

Gupta:1996:ACW

[GG96] Pushpa L. Gupta and Ramesh C. Gupta. Ageing characteristics of the Weibull mixtures. Probability in the Engineering and Informational Sciences, 10(4):591–600, October 1996. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
657109A8486BE96B507AF216C6F0A5D7.

Glasserman:1991:SCS

[GHS91] Paul Glasserman, Jian-Qiang Hu, and Stephen G. Strickland. Strongly consistent steady-state derivative estimates. Probability in the Engineering and Informational Sciences, 5(4):391– 413, October 1991. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/8DF2FD9F14309DC5BB12C2FD012BC33A.

Gaujal:2007:OOL

[GHvdL07] Bruno Gaujal, Arie Hordijk, and Dinard van der Laan. On the optimal open-loop control policy for deterministic and exponential polling systems. Probability in the Engineering and Informational Sciences, 21(2):157–187, April 2007. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
CEB827B5D7D01247D302FC11E6A53F00.

Glynn:1988:CUW

[GI88] Peter W. Glynn and Donald L. Iglehart. Conditions under which a Markov chain converges to its steady state in finite time. Probability in the Engineering and Informational Sciences, 2(3):377–382, July 1988. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
0E7FEB0A34737B9F29EB2AE61B56C77B.

Glynn:1995:MAR

[GI95] Peter W. Glynn and Donald L. Iglehart. A martingale approach to regenerative simulation. Probability in the Engineering and Informational Sciences, 9(1):123–131, January 1995. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
CCBBF15AA438B5EA5DFAC4EB9E02AE61.

Gupta:1988:CMP

[GK88] Ramesh C. Gupta and S. N. U. A. Kirmani. Closure and monotonicity properties of nonhomogeneous Poisson processes and record values. Probability in the Engineering and Informational Sciences, 2(4):475–484, October 1988. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
B880B8225D74F415C68690F762721ECC.

Ghitany:2007:RPE

[GK07] M. E. Ghitany and Samuel Kotz. Reliability properties of extended linear failure-rate distributions. Probability in the Engineering and Informational Sciences, 21(3):441–450, July 2007. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
757ADB7E0A569F4A4EAC8283AFB2408A.

Gadrich:2021:ERD

[GK21] Tamar Gadrich and Guy Katriel. Estimating the rate of defects under imperfect sampling inspection — a new approach. Probability in the Engineering and Informational Sciences, 35(2):242–257, April 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/estimating-
the-rate-of-defects-under-imperfect-sampling-inspectiona-
new-approach/394921056712DEBBDBBEA9F532013BD7.

Gupta:2013:CGM

[GKB13] Ramesh C. Gupta, S. N. U. A. Kirmani, and N. Balakrishnan. On a class of generalized Marshall–Olkin bivariate distributions and some reliability characteristics. Probability in the Engineering and Informational Sciences, 27(2):261–275, April 2013. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
39AF9FEF69C6740E4AFD795E38B73F2D.

Gupta:1987:LDH

[GKL87] Ramesh C. Gupta, S. N. U. A. Kirmani, and Robert L. Launer. On life distributions having monotone residual variance. Probability in the Engineering and Informational Sciences, 1(3): 299–307, July 1987. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/02CBD8E171C389D7CBFD9FB8B7272817.

Gaigi:2023:OMV

[GKL23] M’hamed Gaïgi, Idris Kharroubi, and Thomas Lim. Optimal management and valuation of a natural resource: the case of optimal harvesting. Probability in the Engineering and Informational Sciences, 37(3):674–694, July 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/optimal-
management-and-valuation-of-a-natural-resource-the-
case-of-optimal-harvesting/93D77349914BA9AA06CAB05CC8256F6A.

Gautam:1999:BFM

[GKPR99] N. Gautam, V. G. Kulkarni, Z. Palmowski, and T. Rolski. Bounds for fluid models driven by semi-Markov inputs. Probability in the Engineering and Informational Sciences, 13(4): 429–475, October 1999. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/C90DE5F58C2562CF02D5B8DFEF8DDC8A.

Glazebrook:1991:NPS

[Gla91] Kevin D. Glazebrook. On nonpreemptive policies for stochastic single-machine scheduling with breakdowns. Probability in the Engineering and Informational Sciences, 5(1):77–87, January 1991. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
2EFFAB7F6A89500417876E06E47A2DDE.

Guo:2011:PPP

[GLW11] Pengfei Guo, Zhaotong Lian, and Yulan Wang. Pricing perishable products with compound Poisson demands. Probability in the Engineering and Informational Sciences, 25(3):289– 306, July 2011. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/96F9D78EEB064E25CD14CE3F5A3A90E9.

Guttman:1996:PDP

[GM96] Irwin Guttman and Ulrich Menzefricke. Predictive distributions in the presence of measurement errors. Probability in the Engineering and Informational Sciences, 10(4):601–611, October 1996. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/3D972A6D68F09DF25A297E339E30ACE3.

Genest:2014:BES

[GM14] Christian Genest and Mhamed Mesfioui. Bivariate extensions of Skellam’s distribution. Probability in the Engineering and Informational Sciences, 28(3):401–417, July 2014. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
C7E37AD82480FDF1D0CCFAEEB84E0279.

Gaujal:2016:SAA

[GM16] Bruno Gaujal and Panayotis Mertikopoulos. A stochastic approximation algorithm for stochastic semidefinite programming. Probability in the Engineering and Informational Sciences, 30(3):431–454, July 2016. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/C4888BCA21C1C9CC6A4B8DA2BD405F20.

Gopaladesikan:2014:BRT

[GMW14] M. Gopaladesikan, H. Mahmoud, and M. D. Ward. Building random trees from blocks. Probability in the Engineering and Informational Sciences, 28(1):67–81, January 2014. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
0F9F80B1AFECDE903315AC0980755642.

Grechuk:2010:CIL

[GMZ10] Bogdan Grechuk, Anton Molyboha, and Michael Zabarankin. Chebyshev inequalities with law-invariant deviation measures. Probability in the Engineering and Informational Sciences, 24(1):145–170, January 2010. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/B83382FD72DD451DF9D680D4F3FE5527.

Gordienko:2014:COP

[GN14] Evgueni Gordienko and Andrey Novikov. Characterizations of optimal policies in a general stopping problem and stability estimating. Probability in the Engineering and Informational Sciences, 28(3):335–352, July 2014. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
54492AE418796A1B0C8DF2B67886F7CC.

Goldberg:2015:MOG

[GN15] Yair Goldberg and Yuval Nov. Modeling and optimization of genetic screens via rna interference and FACS. Probability in the Engineering and Informational Sciences, 29(1):131–145, January 2015. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/65A07D5CE1C4B08BE580C3115233E43F.

Gouweleeuw:1994:LPF

[Gou94] Frank N. Gouweleeuw. The loss probability in finite-buffer queues with batch arrivals and complete rejection. Probability in the Engineering and Informational Sciences, 8(2):221– 227, April 1994. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/7CA08DA7C7D37229E7A2A33C30C9FC19.

Grunspan:2022:PND

[GPM22] Cyril Grunspan and Ricardo Pérez-Marco. On profitability of Nakamoto double spend. Probability in the Engineering and Informational Sciences, 36(3):732–746, July 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/on-profitability-
of-nakamoto-double-spend/4DF05998AA2F76EAB5D3AF2D181AE584.

Ghamami:2012:INI

[GR12] Samim Ghamami and Sheldon M. Ross. Improving the normalized importance sampling estimator. Probability in the Engineering and Informational Sciences, 26(4):569–572, October 2012. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
3EBF35FFBD9FD3C6B4AA5D6862189695.

Grassmann:2008:WPS

[Gra08] Winfried K. Grassmann. Warm-up periods in simulation can be detrimental. Probability in the Engineering and Informational Sciences, 22(3):415–429, July 2008. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
395C5FB4DFCF5ACE1124EBEEA4BF8C70.

Gertsbakh:2014:PMP

[GRSV14] Ilya Gertsbakh, Reuven Rubinstein, Yoseph Shpungin, and Radislav Vaisman. Permutational methods for performance analysis of stochastic flow networks. Probability in the Engineering and Informational Sciences, 28(1):21–38, January 2014. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
28DEE6C0B070CA3BF91B0F79676B8819.

Gelenbe:1992:SPF

[GS92] Erol Gelenbe and Rolf Schassberger. Stability of product form G-networks. Probability in the Engineering and Informational Sciences, 6(3):271–276, July 1992. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
BD3D0426C3BDEB416CBD13C9AAE6A8A5.

Gal:1993:OSI

[GS93] Shmuel Gal and Dafna Sheinwald. Optimal sampling for identifying a cluster: An application in sorting. Probability in the Engineering and Informational Sciences, 7(3):387–407, July 1993. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
8614E0C02EEEB17B658C7067AAD77010.

Gao:1998:ESC

[GS98] Hongsheng Gao and Peter J. Smith. Exact SINR calculations for optimum linear combining in wireless systems. Probability in the Engineering and Informational Sciences, 12(2):261– 281, April 1998. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/F3B065C6882056BC8D12C4420B213E24.

Gertsbakh:2012:CAC

[GS12] Ilya B. Gertsbakh and Yoseph Shpungin. Combinatorial approach to computing component importance indexes in coherent systems. Probability in the Engineering and Informational Sciences, 26(1):117–128, January 2012. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
2E0EAA71BD2436CBEB663B64A6AA4BE2.

Gao:2023:ACC

[GS23] Dechen Gao and Kristina P. Sendova. Applications of the classical compound Poisson model with claim sizes following a compound distribution. Probability in the Engineering and Informational Sciences, 37:357–386, April 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/applications-
of-the-classical-compound-poisson-model-with-claim-
sizes-following-a-compound-distribution/34A091B77C7DA92F83B725A366D31640.

Gertsbakh:2016:SRB

[GSV16] Ilya B. Gertsbakh, Yoseph Shpungin, and Radislav Vaisman. D-spectrum and reliability of a binary system with ternary components. Probability in the Engineering and Informational Sciences, 30(1):25–39, January 2016. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
24E7691CA6D870EEE74EE2682724306F.

Greenberg:1993:UDE

[GSW93] Albert G. Greenberg, Otmar Schlunk, and Ward Whitt. Using distributed-event parallel simulation to study departures from many queues in series. Probability in the Engineering and Informational Sciences, 7(2):159–186, April 1993. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
FDBE80C0BE4144DF181EF2DB0627D771.

Grassmann:2007:BAF

[GT07] Winfried K. Grassmann and Javad Tavakoli. A Bayesian approach to find random-time probabilities from embedded Markov chain probabilities. Probability in the Engineering and Informational Sciences, 21(4):551–556, October 2007. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
59B3BE47F9C30ADEDFB8E51DC8A5FC3E.

Gerardi:2010:RNM

[GT10] Anna Gerardi and Paola Tardelli. Risk-neutral measures and pricing for a pure jump price process: a stochastic control approach. Probability in the Engineering and Informational Sciences, 24(1):47–76, January 2010. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
81DE1570126347E23143672B964B0971.

Gayon:2009:OPP

[GTDKÖ09] Jean-Philippe Gayon, Isilay Talay-Degirmenci, Fikri Karaesmen, and E. Lerzan Örmeci. Optimal pricing and production policies of a make-to-stock system with fluctuating demand. Probability in the Engineering and Informational Sciences, 23 (2):205–230, April 2009. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/AA38F473EF9988B2F05E8D7C236AFBF6.

Gupta:2007:RED

[Gup07] Ramesh C. Gupta. Role of equilibrium distribution in reliability studies. Probability in the Engineering and Informational Sciences, 21(2):315–334, April 2007. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
954C35E2944E1E8EB60CF8A616F2304F.

Gupta:2015:RSB

[Gup15] Ramesh C. Gupta. Reliability studies of bivariate Birnbaum– Saunders distribution. Probability in the Engineering and Informational Sciences, 29(2):265–276, April 2015. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
6DBD1823FA77CBEBA17F0F4A42101E17.

Gupta:2016:MRL

[Gup16] Ramesh C. Gupta. Mean residual life function for additive and multiplicative hazard rate models. Probability in the Engineering and Informational Sciences, 30(2):281–297, April 2016. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
A7CBE23E0BB0F4F7E543521F3515631A.

Gutjahr:1997:IST

[Gut97] Walter J. Gutjahr. Importance sampling of test cases in Markovian software usage models. Probability in the Engineering and Informational Sciences, 11(1):19–36, January 1997. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
579517AAAD7E87CD78A840FA6C519392.

Gutjahr:2003:GCR

[Gut03] Walter J. Gutjahr. A generalized convergence result for the graph-based ant system metaheuristic. Probability in the Engineering and Informational Sciences, 17(4):545–569, October 2003. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
681D16E6FBCEBEB23AF380A9495A882A.

Gut:2014:SRD

[Gut14] Allan Gut. Some remarks on distributed sensing with costs. Probability in the Engineering and Informational Sciences, 28 (2):271–277, April 2014. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/672E28672759F62B5EC16BB67118740E.

Greenberg:1990:QCI

[GV90] Albert G. Greenberg and Robert J. Vanderbei. Quicker convergence for iterative numerical solutions to stochastic problems: Probabilistic interpretations, ordering heuristics, and parallel processing. Probability in the Engineering and Informational Sciences, 4(4):493–521, October 1990. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
CD1737AB33C042CDE00A66ED38039921.

Glasserman:1994:CMC

[GV94] Paul Glasserman and Pirooz Vakili. Comparing Markov chains simulated in parallel. Probability in the Engineering and Informational Sciences, 8(3):309–326, July 1994. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
DAB12883E180055703938EF2540C8CC5.

Germs:2010:LPM

[GV10] Remco Germs and Nicky Van Foreest. Loss probabilities for the M^{X} /G^{Y} /1/K + B bulk queue. Probability in the Engineering and Informational Sciences, 24(4):457–471, October 2010. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
931333CE3D25FB371371E0DDCF11579B.

Gupta:2011:RCF

[GV11] Ramesh C. Gupta and Weston Viles. Roller-coaster failure rates and mean residual life functions with application to the extended generalized inverse Gaussian model. Probability in the Engineering and Informational Sciences, 25(1):103–118, January 2011. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/2B37E9F124A2D3699B69D3ED8A32CAC3.

Glynn:1996:ESC

[GW96] Peter W. Glynn and Eugene W. Wong. Efficient simulation via coupling. Probability in the Engineering and Informational Sciences, 10(2):165–186, April 1996. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
8159085F3EFC4C81BDE4EC918B0F57FF.

Gaither:2012:AMS

[GW12] J. Gaither and M. D. Ward. Analytic methods for select sets. Probability in the Engineering and Informational Sciences, 26(4):561–568, October 2012. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
966E3CC67E2F8B2AE5413B40C23657C8.

Guo:2009:ICD

[GZ09] Pengfei Guo and Paul Zipkin. The impacts of customers’ delayrisk sensitivities on a queue with balking. Probability in the Engineering and Informational Sciences, 23(3):409–432, July 2009. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
250E7C69A18B1A569DCF50D53B72EA5C.

Gao:2021:EBS

[GZDW21] Shan Gao, Deran Zhang, Hua Dong, and Xianchao Wang. Equilibrium balking strategies in the repairable m/m/1 gretrial queue with complete removals. Probability in the Engineering and Informational Sciences, 35(1):138–157, January 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/equilibrium-balking-strategies-in-the-repairable-
mm1-gretrial-queue-with-complete-removals/E7FBE5B391E017566FE7179249066CE5.

Guo:2022:ITF

[GZY22] Zhouxia Guo, Jiandong Zhang, and Rongfang Yan. On inactivity times of failed components of coherent system under double monitoring. Probability in the Engineering and Informational Sciences, 36(3):923–940, July 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/on-inactivity-
times-of-failed-components-of-coherent-system-under-
double-monitoring/7832C95F5153B3F71E2449D4CC71BD70.

Hu:2022:SCQ

[HAA22] Junqi Hu, Sigrún Andradóttir, and Hayriye Ayhan. Server coordination in queueing systems: when and how? Probability in the Engineering and Informational Sciences, 36(3):868– 922, July 2022. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
journals/probability-in-the-engineering-and-informational-
sciences/article/server-coordination-in-queueing-systems-
when-and-how/468E51413BCCF6A07840938B3FB81D3C.

Haji:2015:AWL

[Haj15] Babak Haji. On the Adan–Weiss loss model having skill-based servers and longest idle assignment rule. Probability in the Engineering and Informational Sciences, 29(2):181–189, April 2015. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
27D29D8B0A9E1ECB6DFD29AE82125508.

Haneveld:2003:IRW

[Han03] Leendert A. Klein Haneveld. On identities for a random walk in an orthant with absorbing boundary. Probability in the Engineering and Informational Sciences, 17(3):411–415, July 2003. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
3356B4AD5846A3678B43272F017B5EB8.

Han:2019:VVO

[Han19] Xingyu Han. Valuation of vulnerable options under the double exponential jump model with stochastic volatility. Probability in the Engineering and Informational Sciences, 33(1):81–104, January 2019. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/valuation-
of-vulnerable-options-under-the-double-exponential-
jump-model-with-stochastic-volatility/260EB606F1520AEA5909046BA36000BB.

Hassin:2007:IUQ

[Has07] Refael Hassin. Information and uncertainty in a queuing system. Probability in the Engineering and Informational Sciences, 21(3):361–380, July 2007. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/0FF4417B847EC741B51F43EA513B2BBE.

Hussain:2021:AEE

[HBG21] Khaled F. Hussain, Mohamed Yousef Bassyouni, and Erol Gelenbe. Accurate, energy-efficient classification with spiking random neural network. Probability in the Engineering and Informational Sciences, 35(1):51–61, January 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/accurate-
energyefficient-classification-with-spiking-random-
neural-network/9B4F1A62E8759BF82421831780578887.

Hui:2003:TCM

[HBKK03] K.-P. Hui, N. Bean, M. Kraetzl, and D. Kroese. The tree cut and merge algorithm for estimation of network reliability. Probability in the Engineering and Informational Sciences, 17(1): 23–45, January 2003. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/845E70F473A2FEDF1B9E04E1F1FF852C.

Harchol-Balter:2010:BUL

[HBV10] Mor Harchol-Balter and Rein Vesilo. To balance or unbalance load in size-interval task allocation. Probability in the Engineering and Informational Sciences, 24(2):219–244, April 2010. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
50802CCCB0EAE3E528FF07F7C8F24B9C.

Hristov:2018:TBA

[HBvdMB18] A. Hristov, J. W. Bosman, R. D. van der Mei, and S. Bhulai. Throughput and bottleneck analysis of tandem queues with nested sessions. Probability in the Engineering and Informational Sciences, 32(3):396–408, July 2018. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/throughput-
and-bottleneck-analysis-of-tandem-queues-with-nested-
sessions/B53FC60D12DCEAB6FF1A11821E5C33A9.

Hu:2006:RDL

[HCX06] Taizhong Hu, Jing Chen, and Chaode Xie. Regression dependence in latent variable models. Probability in the Engineering and Informational Sciences, 20(2):363–379, April 2006. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
83FD1C881159383869C5EC2A7AFDD816.

He:1989:OAC

[He89] Shengwu He. Optimization applications of compensators of Poisson random measures. Probability in the Engineering and Informational Sciences, 3(1):149–155, January 1989. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
E97C22653EA023A21007FBD567F020B8.

Heidergott:2001:OPM

[Hei01] Bernd Heidergott. Option pricing via Monte Carlo simulation: a weak derivative approach. Probability in the Engineering and Informational Sciences, 15(3):335–349, July 2001. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
57A242C61555BD51BE8B8CEF6953A248.

Herschkorn:1995:MVF

[Her95] Stephen J. Herschkorn. On the modular value and fractional part of a random variable. Probability in the Engineering and Informational Sciences, 9(4):551–562, October 1995. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
014B195A54A8407658C49C717948C722.

Hesse:1992:HTD

[Hes92] C. H. Hesse. Hitting-time densities of a two-dimensional Markov process. Probability in the Engineering and Informational Sciences, 6(4):561–580, October 1992. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
9D98E84D9C30CBAE4687D485B52DBD4D.

Hassin:1996:OER

[HH96] Refael Hassin and Moshe Haviv. On optimal and equilibrium retrial rates in a queueing system. Probability in the Engineering and Informational Sciences, 10(2):223–227, April 1996. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
1656E080907DED5DC3ABE09F4E62E8E0.

Hu:2000:NCS

[HH00] Taizhong Hu and Fangyi He. A note on comparisons of k-out-of-n systems with respect to the hazard and reversed hazard rate orders. Probability in the Engineering and Informational Sciences, 14(1):27–32, January 2000. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
5AB9F2A34C6A8454B566250C1FC7C7B8.

Hautphenne:2016:BOK

[HH16] Sophie Hautphenne and Moshe Haviv. The bias optimal k in the M/M/1/k queue: an application of the deviation matrix. Probability in the Engineering and Informational Sciences, 30(1):61–78, January 2016. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
3F0B26C33B12AFB0934CCBDED13E33B7.

Heidergott:2007:SEF

[HHvU07] Bernd Heidergott, Arie Hordijk, and Miranda van Uitert. Series expansions for finite-state Markov chains. Probability in the Engineering and Informational Sciences, 21(3):381–400, July 2007. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
E78771D69CB56EC1C9625F22A0963C22.

Hildebrand:1999:CFH

[Hil99] Martin Hildebrand. On a conjecture of Fill and Holst involving the move-to-front rule and cache faults. Probability in the Engineering and Informational Sciences, 13(3):377–385, July 1999. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
E2AD44FFB7FAD91660DB20811F4DE1EE.

Hu:1995:MPD

[HJ95] Taizhong Hu and Harry Joe. Monotonicity of positive dependence with time for stationary reversible Markov chains. Probability in the Engineering and Informational Sciences, 9(2): 227–237, April 1995. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/40E871CEC42CD08ED0FBB84821DFCE23.

Hu:2007:OCD

[HJK07] Taizhong Hu, Wei Jin, and Baha-Eldin Khaledi. Ordering conditional distributions of generalized order statistics. Probability in the Engineering and Informational Sciences, 21(3): 401–417, July 2007. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/680A51B5A865CFAA14151D4D5FD530D0.

Hampshire:2009:TVC

[HJM09] Robert C. Hampshire, Otis B. Jennings, and William A. Massey. A time-varying call center design via Lagrangian mechanics. Probability in the Engineering and Informational Sciences, 23(2):231–259, April 2009. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/D8A32A7C459E67A111EE1706EF02D156.

Hordijk:1990:OGS

[HK90] Arie Hordijk and Ger Koole. On the optimality of the generalized shortest queue policy. Probability in the Engineering and Informational Sciences, 4(4):477–487, October 1990. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
97C7FE557AF1CFAB1F08089B85E56DC0.

Hordijk:1992:ACP

[HK92a] Arie Hordijk and Ger Koole. On the assignment of customers to parallel queues. Probability in the Engineering and Informational Sciences, 6(4):495–511, October 1992. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
293EEEF3E75D600DD487654432DC5B25.

Hordijk:1992:SQP

[HK92b] Arie Hordijk and Ger Koole. On the shortest queue policy for the tandem parallel queue. Probability in the Engineering and Informational Sciences, 6(1):63–79, January 1992. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
940C44730F078DA7AD0EFA56BDFB1F58.

Hordijk:1996:SPM

[HK96] Arie Hordijk and Ger Koole. On suboptimal policies in multiclass tandem models. Probability in the Engineering and Informational Sciences, 10(1):29–39, January 1996. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
6E6E671BEDB240A2CD74687B5A2EE29A.

Huseby:2022:BCI

[HKA22] Arne Bang Huseby, Martyna Kalinowska, and Tobias Abrahamsen. Birnbaum criticality and importance measures for multistate systems with repairable components. Probability in the Engineering and Informational Sciences, 36(1):66–86, January 2022. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
journals/probability-in-the-engineering-and-informational-
sciences/article/birnbaum-criticality-and-importance-
measures-for-multistate-systems-with-repairable-components/
7B17F89B1225A5FFD6669370B8A54859.

Haviv:2010:ESQ

[HKK10] Moshe Haviv, Offer Kella, and Yoav Kerner. Equilibrium strategies in queues based on time or index of arrival. Probability in the Engineering and Informational Sciences, 24(1): 13–25, January 2010. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/7C1D3561B236FB16D37819AE10018D6A.

Hordijk:1994:ACA

[HKL94] A. Hordijk, G. M. Koole, and J. A. Loeve. Analysis of a customer assignment model with no state information. Probability in the Engineering and Informational Sciences, 8(3):419– 429, July 1994. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/9CA985D8EB1818AC57FEBA5D1FDB5D59.

Huang:2007:CIP

[HKS07] Wei Huang, Vidhyadhar Kulkarni, and Jayashankar M. Swaminathan. Coordinated inventory planning for new and old products under warranty. Probability in the Engineering and Informational Sciences, 21(2):261–287, April 2007. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
736C920C52CF5B558924CDFC8BB083D2.

Hwang:1991:RDL

[HL91] F. K. Hwang and Wen-Ching Winnie Li. Reliabilities of double-loop networks. Probability in the Engineering and Informational Sciences, 5(3):255–272, July 1991. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
9D4801BF16AA98CC5A68282BE459AD10.

Hunt:1992:LBA

[HL92] P. J. Hunt and C. N. Laws. Least busy alternative routing in queueing and loss networks. Probability in the Engineering and Informational Sciences, 6(4):439–456, October 1992. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
74F1213808B6A12129D11757CD49E298.

Hwang:2001:ENU

[HL01] F. K. Hwang and Y. C. Liu. The expected numbers of unresolved positive clones for various random pool designs. Probability in the Engineering and Informational Sciences, 15(1): 57–68, January 2001. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/593D1E87D5603F39521052BDAE66BAF8.

Hwang:2004:GAC

[HL04] F. K. Hwang and Y. C. Liu. A general approach to compute the probabilities of unresolved clones in random pooling designs. Probability in the Engineering and Informational Sciences, 18 (2):161–183, April 2004. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/64FC14D3CBC7CC07E6256D86B085202C.

Hashorva:2014:ADT

[HL14] Enkelejd Hashorva and Jinzhu Li. Asymptotics for a discretetime risk model with the emphasis on financial risk. Probability in the Engineering and Informational Sciences, 28(4):573–588, October 2014. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/B2915EF5DB06FF5B7BADB71DFBEA1B4D.

Hofri:2016:CBS

[HLM16] Micha Hofri, Chao Li, and Hosam Mahmoud. On the combinatorics of binary series-parallel graphs. Probability in the Engineering and Informational Sciences, 30(2):244–260, April 2016. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
A0B107F8FF87D8365F50D9BF899CC999.

He:2016:SSS

[HLW16] Beixiang He, Yunan Liu, and Ward Whitt. Staffing a service system with non-Poisson non-stationary arrivals. Probability in the Engineering and Informational Sciences, 30(4):593–621, October 2016. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/0F42FDA80A8B0B197D3D9E0B040A43D2.

Hu:2023:DRM

[HLZ23] Kang Hu, Jingchao Li, and Jieming Zhou. On the dual risk model with Parisian implementation delays under a mixed dividend strategy. Probability in the Engineering and Informational Sciences, 37:442–461, April 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/on-the-
dual-risk-model-with-parisian-implementation-delays-
under-a-mixed-dividend-strategy/193E854165BF29F3419A4A62133F8C3F.

Hagstrom:1987:SRA

[HM87] Jane N. Hagstrom and King-Tim Mak. System reliability analysis in the presence of dependent component failures. Probability in the Engineering and Informational Sciences, 1(4):425– 440, October 1987. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/D6AB65266579FAF6AFA3E9F835983810.

Hurley:1994:AAR

[HM94] Catherine B. Hurley and Hosam M. Mahmoud. Analysis of an algorithm for random sampling. Probability in the Engineering and Informational Sciences, 8(2):153–168, April 1994. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
06213DCC73156D72B6505B026BA07391.

Huang:2015:SIS

[HM15] Kai Huang and Jie Mi. Study of instant system availability. Probability in the Engineering and Informational Sciences, 29(1):117–129, January 2015. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/40A5C09D4086696FC4E632872862627F.

Hussain:2016:RVC

[HM16] Khaled F. Hussain and Ghada S. Moussa. On-road vehicle classification based on random neural network and bag-of-visual words. Probability in the Engineering and Informational Sciences, 30(3):403–412, July 2016. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
E87EECDD5172640C1B2B6E1C01533BBF.

Huillet:2019:RMP

[HM19] Thierry Huillet and Servet Martínez. Regenerative mutation processes related to the self-decomposability of Sibuya distributions. Probability in the Engineering and Informational Sciences, 33(2):291–325, April 2019. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/regenerative-
mutation-processes-related-to-the-selfdecomposability-
of-sibuya-distributions/6476B5741E859014D390FFC266224476.

Huang:2020:ALR

[HM20] Kai Huang and Jie Mi. Applications of likelihood ratio order in Bayesian inferences. Probability in the Engineering and Informational Sciences, 34(1):1–13, January 2020. CODEN ????

ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/applications-
of-likelihood-ratio-order-in-bayesian-inferences/335FFCFEC600F036FB43E9AD6C1FF4A0.

Hazra:2021:RAC

[HM21] Nil Kamal Hazra and Neeraj Misra. On relative aging comparisons of coherent systems with identically distributed components. Probability in the Engineering and Informational Sciences, 35(3):481–495, July 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/on-relative-
aging-comparisons-of-coherent-systems-with-identically-
distributed-components/43CB99BAA516E007F1B32F8F88954A67.

Hampshire:2009:DPC

[HMW09] Robert C. Hampshire, William A. Massey, and Qiong Wang. Dynamic pricing to control loss systems with quality of service targets. Probability in the Engineering and Informational Sciences, 23(2):357–383, April 2009. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
1EFAA5E72664533AB5F2F9F43058FB0E.

Haidari:2019:COR

[HN19] Abedin Haidari and Amir T. Payandeh Najafabadi. Characterization ordering results for largest order statistics from heterogeneous and homogeneous exponentiated generalized gamma variables. Probability in the Engineering and Informational Sciences, 33(3):460–470, July 2019. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/characterization-
ordering-results-for-largest-order-statistics-from-
heterogeneous-and-homogeneous-exponentiated-generalized-
gamma-variables/8CEE45B738ED1B94420013CBCE1F945F.

Hordijk:2001:CQD

[Hor01] Arie Hordijk. Comparison of queues with different discretetime arrival processes. Probability in the Engineering and Informational Sciences, 15(1):1–14, January 2001. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
F07C1FAFC0404FEEBC7DB956C7BB3D04.

Hwang:1989:MRI

[HP89] F. K. Hwang and S. G. Papastavridis. A monotonicity result for inspecting independent items. Probability in the Engineering and Informational Sciences, 3(1):135–140, January 1989. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
FBF49086B67C9DAA9834B7C08D7E8D4C.

Hu:2000:CDS

[HP00] Taizhong Hu and Xiaoming Pan. Comparisons of dependence for stationary Markov processes. Probability in the Engineering and Informational Sciences, 14(3):299–315, July 2000. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
C70F361737DAF12D177BC3A0BC423D15.

Hordijk:2003:LDA

[HP03a] Arie Hordijk and Nikolai Popov. Large deviations analysis of a coupled-processors system. Probability in the Engineering and Informational Sciences, 17(3):397–409, July 2003. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
C6CECAC77CE518139761A8AA7091AFAA.

Hordijk:2003:LDB

[HP03b] Arie Hordijk and Nikolai Popov. Large deviations bounds for face-homogeneous random walks in the quarter-plane. Probability in the Engineering and Informational Sciences, 17(3): 369–395, July 2003. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/98DAA080951D1B76221CDA2A03A77525.

Herschkorn:1996:PMI

[HPR96] Stephen J. Herschkorn, Erol Peköz, and Sheldon M. Ross. Policies without memory for the infinite-armed Bernoulli bandit under the average-reward criterion. Probability in the Engineering and Informational Sciences, 10(1):21–28, January 1996. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
34897756FCD84BA91B458693A6814102.

Hordijk:1999:EPE

[HPS99] A. Hordijk, O. Passchier, and F. M. Spieksma. On the existence of the Puiseux expansion of the discounted rewards: a counterexample. Probability in the Engineering and Informational Sciences, 13(2):229–235, April 1999. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
D0B09939AF26C79C85B1A3937578FEDF.

Hordijk:2008:DSS

[HPS08] Arie Hordijk, Nikolay Popov, and Flora Spieksma. Discrete scattering and simple and nonsimple face-homogeneous random walks. Probability in the Engineering and Informational Sciences, 22(2):163–189, March 2008. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
E29EC0E00483194B347D0C936C2C1D35.

Haas:1988:MPS

[HS88] Peter J. Haas and Gerald S. Shedler. Modeling power of stochastic Petri nets for simulation. Probability in the Engineering and Informational Sciences, 2(4):435–459, October 1988. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
D39DC69996A08EE0E4974F46DEE598BC.

Haas:1991:SPN

[HS91] Peter J. Haas and Gerald S. Shedler. Stochastic Petri nets: Modeling power and limit theorems. Probability in the Engineering and Informational Sciences, 5(4):477–498, October 1991. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
B94E379A01B8AF48A272E9B954190889.

Hofri:1992:ARC

[HS92] Micha Hofri and Hadas Shachnai. The application of restricted counter schemes to three models of linear search. Probability in the Engineering and Informational Sciences, 6(3):371– 389, July 1992. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/AA81A36B5E1F8AAF22D53CCE0A0001E5.

Honhon:2007:ACI

[HS07] Dorothée Honhon and Sridhar Seshadri. Admission control with incomplete information to a finite buffer queue. Probability in the Engineering and Informational Sciences, 21(1): 19–46, January 2007. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/66D0A7E73EE64470C61867269B82C059.

Huang:2013:DAM

[HS13] Jia-Ping Huang and Ushio Sumita. Dynamic analysis of a multivariate reward process defined on the UMCP with application to optimal preventive maintenance policy problems in manufacturing. Probability in the Engineering and Informational Sciences, 27(2):187–208, April 2013. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
9381BAE5C5E950D9296993784662B007.

Haidari:2023:MRL

[HSB23] Abedin Haidari, Mostafa Sattari, and Ghobad Barmalzan. Mean residual life order among largest order statistics arising from resilience-scale models with reduced scale parameters. Probability in the Engineering and Informational Sciences, 37(1):72–85, January 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/mean-
residual-life-order-among-largest-order-statistics-
arising-from-resiliencescale-models-with-reduced-scale-
parameters/0733A084BDE9D3422A1A3056014E9E76.

Hwang:1996:SRB

[HT96] Frank K. Hwang and Dan Trietsch. A simple relation between the pattern probability and the rate of false signals in control charts. Probability in the Engineering and Informational Sciences, 10(3):315–323, July 1996. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
B45A9A2273E9BFB3DA6D93D2A96B02DD.

Huang:2020:GEE

[Hua20] Huilin Huang. The generalized entropy ergodic theorem for nonhomogeneous Markov chains indexed by a homogeneous tree. Probability in the Engineering and Informational Sciences, 34(2):221–234, April 2020. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/generalized-
entropy-ergodic-theorem-for-nonhomogeneous-markov-chains-
indexed-by-a-homogeneous-tree/39C8243C4373A83BDD478476D57BAFC8.

Hunt:1991:ABI

[Hun91] P. J. Hunt. Asymptotic behaviour of an integrated videodata network. Probability in the Engineering and Informational Sciences, 5(4):429–447, October 1991. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
0E3D1130FB204321EC1570A58A33C46F.

Huang:1995:MBD

[HV95] Ying Huang and Arthur F. Veinott. Markov branching decision chains with interest-rate-dependent rewards. Probability in the Engineering and Informational Sciences, 9(1):99–121, January 1995. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
CD3FF7ACB9C579E5AAF5E81C35C41B8C.

Heydenreich:2009:FBB

[HvdHR09] Markus Heydenreich, Remco van der Hofstad, and Georgi Radulov. Functionals of Brownian bridges arising in the current mismatch in D/A converters. Probability in the Engineering and Informational Sciences, 23(1):149–172, January 2009. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
BCE10F089C70E435D2991BA4123C74FD.

Hordijk:2003:NCS

[HvdL03] Arie Hordijk and Dinard van der Laan. Note on the convexity of the stationary waiting time as a function of the density. Probability in the Engineering and Informational Sciences, 17(4):503–508, October 2003. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
60A3DE96BEEF28BB7751422A27814320.

Haviv:1997:ESP

[HvdW97] Moshe Haviv and Jan van der Wal. Equilibrium strategies for processor sharing and random queues with relative priorities. Probability in the Engineering and Informational Sciences, 11(4):403–412, October 1997. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
95D5B74D004E6309280427774A7C3F83.

Haijema:2008:MDA

[HvdW08] René Haijema and Jan van der Wal. An MDP decomposition approach for traffic control at isolated signalized intersections. Probability in the Engineering and Informational Sciences, 22(4):587–602, October 2008. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
A291D2A087A721EE81A99AFFFBF15D9E.

He:1990:TPP

[HW90] Shengwu He and Jiagang Wang. Thinning of point processesmartingale method. Probability in the Engineering and Informational Sciences, 4(1):117–129, January 1990. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
06990709DE0C925E3E4042CBD1093BE0.

Huang:1992:PSS

[HW92] Chueng-Chiu Huang and Gideon Weiss. Preemptive scheduling of stochastic jobs with a two-stage processing time distribution on m + 1 parallel machines. Probability in the Engineering and Informational Sciences, 6(2):171–191, April 1992. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
28AB58B5BE6AC91DB208B7D9DE81C87B.

Hwang:2000:RSP

[Hwa00] F. K. Hwang. Random k-set pool designs with distinct columns. Probability in the Engineering and Informational Sciences, 14(1):49–56, January 2000. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
0E091B775DCD11747CB6AAD956A8B4A5.

Hu:2012:NPT

[HWZ12] Taizhong Hu, Yashi Wang, and Weiwei Zhuang. New properties of the total time on test transform order. Probability in the Engineering and Informational Sciences, 26(1):43–60, January 2012. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
F31D4D8F23E4175F5650323ADDC37571.

Hwang:1990:FRC

[HY90] F. K. Hwang and Y. C. Yao. On the failure rates of consecutive-k-out-of-n systems. Probability in the Engineering and Informational Sciences, 4(1):57–71, January 1990. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
86B8E237514194D1A4FC2E13465BA05B.

Hanukov:2021:ESC

[HY21] Gabi Hanukov and Uri Yechiali. Explicit solutions for continuous-time QBD processes by using relations between matrix geometric analysis and the probability generating functions method. Probability in the Engineering and Informational Sciences, 35(3):565–580, July 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/explicit-
solutions-for-continuoustime-qbd-processes-by-using-
relations-between-matrix-geometric-analysis-and-the-
probability-generating-functions-method/991AFEFDBC6B932AD67FB0B35E2A800A.

Hu:2007:NCB

[HYL07] Taizhong Hu, Junchao Yao, and Qingshu Lu. Nonnegativity of covariances between functions of ordered random variables. Probability in the Engineering and Informational Sciences, 21(4):557–577, October 2007. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
826686E619F1E64D4D7831F943607965.

Hu:2004:LRO

[HZ04] Taizhong Hu and Weiwei Zhuang. Likelihood ratio ordering of the inspection paradox. Probability in the Engineering and Informational Sciences, 18(4):503–510, October 2004.

CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
C6D8F79C0DF76DE5C9C21E62E95B8DF0.

Hu:2005:SPS

[HZ05] Taizhong Hu and Weiwei Zhuang. Stochastic properties of p-spacings of generalized order statistics. Probability in the Engineering and Informational Sciences, 19(2):257–276, April 2005. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
297B4B873D649780928B6F654D624078.

Hu:2006:SOB

[HZ06] Taizhong Hu and Weiwei Zhuang. Stochastic orderings between p-spacings of generalized order statistics from two samples. Probability in the Engineering and Informational Sciences, 20(3):465–479, July 2006. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
98DF9766B5F9AF2CB87FB5B30CC711BF.

Hu:2001:LRM

[HZW01] Taizhong Hu, Zegang Zhu, and Ying Wei. Likelihood ratio and mean residual life orders for order statistics of heterogeneous random variables. Probability in the Engineering and Informational Sciences, 15(2):259–272, April 2001. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
6EFA186A7E93B1F60D8927C867EE1542.

Izadkhah:2023:PPS

[IASB23] Salman Izadkhah, Ebrahim Amini-Seresht, and Narayanaswamy Balakrishnan. Preservation properties of some reliability classes by lifetimes of coherent and mixed systems and their signatures. Probability in the Engineering and Informational Sciences, 37(4):943–960, October 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/preservation-
properties-of-some-reliability-classes-by-lifetimes-
of-coherent-and-mixed-systems-and-their-signatures/
0BE61708F3B576F331E02EC4D1A9E69E.

Iyigun:2008:PDC

[IBI08] Cem Iyigun and Adi Ben-Israel. Probabilistic distance clustering adjusted for cluster size. Probability in the Engineering and Informational Sciences, 22(4):603–621, October 2008. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
31797FC2517E8398CE532535CD166666.

Idriss:2023:NUU

[Idr23] Soumaya Idriss. Nonlinear and unbalanced urn models with two types of strategies: a stochastic approximation point of view. Probability in the Engineering and Informational Sciences, 37(3):711–729, July 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/nonlinear-
and-unbalanced-urn-models-with-two-types-of-strategies-
a-stochastic-approximation-point-of-view/0E14828AF7807BE053A6DB9F7A67A9A5.

Izagirre:2015:HTA

[IVA15] A. Izagirre, I. M. Verloop, and U. Ayesta. Heavy-traffic analysis of a non-preemptive multi-class queue with relative priorities. Probability in the Engineering and Informational Sciences, 29(2):153–180, April 2015. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
DB9E3F8345D764AEE1717679569CC344.

Ishikida:1997:SJS

[IwW97] Takashi Ishikida and Yat wah Wan. Scheduling jobs that are subject to deterministic due dates and have deteriorating expected rewards. Probability in the Engineering and Informational Sciences, 11(1):65–78, January 1997. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
E0E648B940DAB48B85A10D3D710E2E85.

Jayakumar:2006:MOE

[JA06] Archana Jayakumar and Sohrab Asgarpoor. Maintenance optimization of equipment by linear programming. Probability in the Engineering and Informational Sciences, 20(1):183–193, January 2006. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/57C109E28AC1E4337B43C525B19D9118.

Jaffe:1992:EDC

[Jaf92] Steven Jaffe. The equilibrium distribution for a clocked buffered switch. Probability in the Engineering and Informational Sciences, 6(4):425–438, October 1992. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
72607A6A11D09DEFDA57C9B0DE020A4E.

Jennings:2018:BDP

[JC18]  Dale Jennings and Jem N. Corcoran.  A birth and

death process for Bayesian network structure inference. Probability in the Engineering and Informational Sciences, 32(4):615–625, October 2018. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/birth-
and-death-process-for-bayesian-network-structure-inference/
1994473237D58CD9EE9E5888EB48F8C3.

Jianu:2022:MCM

[JCDJ22] Marilena Jianu, Daniel Ciuiu, Leonard Daus, and Mihail Jianu. Markov chain method for computing the reliability of hammock networks. Probability in the Engineering and Informational Sciences, 36(2):276–293, April 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/markov-
chain-method-for-computing-the-reliability-of-hammock-
networks/CD156B1F0E90030F15B77D6F792A424B.

Jouini:2007:MPM

[JD07] Oualid Jouini and Yves Dallery. Monotonicity properties for multiserver queues with reneging and finite waiting lines. Probability in the Engineering and Informational Sciences, 21(3): 335–360, July 2007. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/009DC5699E9BF667E36F868BBCBE5A5E.

Jiang:2019:ACQ

[Jia19] Tao Jiang. Analysis of a clearing queueing system with server maintenance after n negative feedbacks. Probability in the Engineering and Informational Sciences, 33(2):220–240, April 2019. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/analysis-of-a-clearing-queueing-system-with-
server-maintenance-after-n-negative-feedbacks/CD4B0BCEC4B37A87206CAB777F9D4E59.

Jammalamadaka:1998:THN

[JL98] S. Rao Jammalamadaka and Eun-Soo Lee. Testing for harmonic new better than used in expectation. Probability in the Engineering and Informational Sciences, 12(3):409–416, July 1998. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
D281E9B510949CA8E8F21D21EBF9614D.

Jane:2019:DAC

[JL19] Chin-Chia Jane and Yih-Wenn Laih. Decomposition algorithms to compute the quickest time distribution in dynamic networks. Probability in the Engineering and Informational Sciences, 33(1):50–63, January 2019. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/decomposition-
algorithms-to-compute-the-quickest-time-distribution-
in-dynamic-networks/170233C3E6623647E670A287EA6C0B4B.

Javed:2017:RNN

[JLAE17] Abbas Javed, Hadi Larijani, Ali Ahmadinia, and Rohinton Emmanuel. Random neural network learning heuristics. Probability in the Engineering and Informational Sciences, 31(4):436– 456, October 2017. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/2FD3C90B26A129A9551679C47316AD0A. See

corrigendum [JLAE18].

Javed:2018:RNN

[JLAE18] Abbas Javed, Hadi Larijani, Ali Ahmadinia, and Rohinton Emmanuel. Random neural network learning heuristics — corrigendum. Probability in the Engineering and Informational Sciences, 32(3):482, July 2018. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/random-

neural-network-learning-heuristics-corrigendum/84EF9E566838AF51C6083B1BECEA742A.

See [JLAE17].

Jia:2021:SSC

[JLL21] Yisu Jia, Robert Lund, and James Livsey. Superpositioned stationary count time series. Probability in the Engineering and Informational Sciences, 35(3):538–556, July 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/superpositioned-stationary-count-time-series/
40D7C7E214AF1CDF08266009D3057E93.

Joe:2012:BRI

[Joe12] Harry Joe. Book review: Inequalities: Theory of Majorization and Its Applications, 2nd edition by Marshall, A. W., Olkin, I. and Arnold, B. C., Springer. Probability in the Engineering and Informational Sciences, 26(3):449–453, July 2012. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
078EDEB7BF3E9EB673E02BF6A756EC06.

Johannesson:1998:RCS

[Joh98] Pär Johannesson. Rainflow cycles for switching processes with Markov structure. Probability in the Engineering and Informational Sciences, 12(2):143–175, April 1998. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
56D66233C6C5D08EA835746F1B027EC1.

Jang:2023:CMG

[JQZD23] Jiwook Jang, Yan Qu, Hongbiao Zhao, and Angelos Dassios. A Cox model for gradually disappearing events. Probability in the Engineering and Informational Sciences, 37(1):214–231, January 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/cox-
model-for-gradually-disappearing-events/2A8D7DAC5AE9D0CAC7B8E47433D5A8FE.

Jun:1992:SRS

[JR92] Chi Hyuck Jun and Sheldon M. Ross. System reliability by simulation: Random hazards versus importance sampling. Probability in the Engineering and Informational Sciences, 6(1):119– 126, January 1992. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/B611EBD1F69D0E03FC55200854D1090B.

Jagerman:1989:FEA

[JS89] David L. Jagerman and Bhaskar Sengupta. A functional equation arising in a queue with a gating mechanism. Probability in the Engineering and Informational Sciences, 3(3):417– 433, July 1989. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/7F430FBFF016BD07E39739F696F2A302.

Jain:1996:GPK

[JS96] Gautam Jain and Karl Sigman. Generalizing the Pollaczek-Khintchine formula to account for arbitrary work removal. Probability in the Engineering and Informational Sciences, 10(4):519–531, October 1996. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/F544BA2E56581AB47AA24DAB629B0088.

Javidi:2001:EMM

[JST01] Tara Javidi, Nah-Oak Song, and Demosthenis Teneketzis. Expected makespan minimization on identical machines in two interconnected queues. Probability in the Engineering and Informational Sciences, 15(4):409–443, October 2001. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
7F43A32EB90424DAD07FFDCE0051E870.

Jun:1988:HPS

[Jun88] Chi-Hyuck Jun. Heuristic procedures for simultaneous estimation of several normal means. Probability in the Engineering and Informational Sciences, 2(1):95–113, January 1988. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
6D42BBCB0DC2D1D447322CDAA961FC1C.

Jahanshahi:2020:CRE

[JZK20] S. M. A. Jahanshahi, H. Zarei, and A. H. Khammar. On cumulative residual extropy. Probability in the Engineering and Informational Sciences, 34(4):605–625, October 2020. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/

probability-in-the-engineering-and-informational-sciences/
article/on-cumulative-residual-extropy/B24ED7EAB5E9442EC7F9750F8A502BB2.

Jiang:2022:AET

[JZY22] Fan Jiang, Xin Zang, and Jingping Yang. Asymptotic expansion for the transition densities of stochastic differential equations driven by the gamma processes. Probability in the Engineering and Informational Sciences, 36(2):401–432, April 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/asymptotic-expansion-for-the-transition-densities-
of-stochastic-differential-equations-driven-by-the-
gamma-processes/C253249D876ACED22CBD1D24398E896B.

Kayid:2004:MIT

[KA04] M. Kayid and I. A. Ahmad. On the mean inactivity time ordering with reliability applications. Probability in the Engineering and Informational Sciences, 18(3):395–409, July 2004. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
DCCDFA18C9EFCD7B6F1C4CB7F57EBB0D.

Khoolenjani:2016:BID

[KA16] Nayereh Bagheri Khoolenjani and Mohammad Hossein Alamatsaz. A de Bruijn’s identity for dependent random variables based on copula theory. Probability in the Engineering and Informational Sciences, 30(1):125–140, January 2016. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
B8D4336669D61ECB1BC7EC7E3B1728D3.

Konsowa:2012:SRW

[KAA12] Mokhtar Konsowa and Fahimah Al-Awadhi. The speed of random walks on trees and electric networks. Probability in the Engineering and Informational Sciences, 26(1):105–116, January 2012. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
F64D0C8AD8E28861FCB9A8F7969FA332.

Kadioglu:2017:FCE

[Kad17] Yasin Murat Kadioglu. Finite capacity energy packet networks. Probability in the Engineering and Informational Sciences, 31(4):477–504, October 2017. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/F3961D1D13AA45BAC59300D9850365A8.

Konsowa:2007:FRD

[KAJ07] Mokhtar H. Konsowa and Reem A. Al-Jarallah. Fractal and resistance dimensions of random trees. Probability in the Engineering and Informational Sciences, 21(4):623–634, October 2007. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
D02A3EC2B298B0B66FFB258D59D340D4.

Kim:2009:OPP

[KAR09] Jung-Hyun Kim, Hyun-Soo Ahn, and Rhonda Righter. Optimal production policies with multistage stochastic demand lead times. Probability in the Engineering and Informational Sciences, 23(3):515–543, July 2009. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
03398B8B5AF00C1E9EB411FBF869EC95.

Kaynar:2009:OSS

[Kay09] Bahar Kaynar. Optimal stopping in a stochastic game. Probability in the Engineering and Informational Sciences, 23(1): 51–60, January 2009. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/9C3524670E90C703DAAD8948B85274AB.

Kayal:2016:GCE

[Kay16] Suchandan Kayal. On generalized cumulative entropies. Probability in the Engineering and Informational Sciences, 30(4): 640–662, October 2016. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/9606ABD8BCF100EA327EEE711964889F.

Kebir:1994:LTC

[Keb94a] Y. Kebir. Laplace transform characterization of probabilistic orderings. Probability in the Engineering and Informational Sciences, 8(1):69–77, January 1994. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
E3D5716E5F242C5825CFA9747CF61EA1.

Kebir:1994:OPS

[Keb94b] Y. Kebir. Order-preserving shock models. Probability in the Engineering and Informational Sciences, 8(1):125–134, January 1994. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/21D5F45A56E220FA07E9161C47EB6657.

Kennedy:1994:RSF

[Ken94] D. P. Kennedy. On the regularity of skip-free Markov chains. Probability in the Engineering and Informational Sciences, 8 (3):377–384, July 1994. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/FF64F185E3DE5D99472609AB7D64115A.

Key:1990:OCT

[Key90] Peter B. Key. Optimal control and trunk reservation in loss networks. Probability in the Engineering and Informational Sciences, 4(2):203–242, April 1990. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
E0B10EA538E95911F9B13BCCFFE58579.

Kochar:1987:SRW

[KG87] S. C. Kochar and R. P. Gupta. Some results on weighted distributions for positive-valued random variables. Probability in the Engineering and Informational Sciences, 1(4):417– 423, October 1987. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/531DDC05887AAECC9104BA03D46C0114.

Kirmani:1989:RAR

[KG89] S. N. U. A. Kirmani and Ramesh C. Gupta. On repair age residual repair life in the minimal repair process. Probability in the Engineering and Informational Sciences, 3(3):381– 391, July 1989. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/1313B61BE6D35621FEC3801DDDE41CB4.

Kirmani:1992:SMI

[KG92] S. N. U. A. Kirmani and Rmaesh C. Gupta. Some moment inequalities for the minimal repair process. Probability in the Engineering and Informational Sciences, 6(2):245–255, April 1992. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
9F21E3DCCB654469041AE81B40FDEF85.

Korzeniowski:1991:SWI

[KH91] A. Korzeniowski and D. L. Hawkins. On simulating Wiener integrals and their expectations. Probability in the Engineering and Informational Sciences, 5(1):101–112, January 1991. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
B0995149027819A41633C353DD44C586.

Khan:1988:IDT

[Kha88] M. S. Ali Khan. An infinite dam with a time-dependent release rule. Probability in the Engineering and Informational Sciences, 2(4):485–500, October 1988. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
A6F93A72363D3E412DEB0F6C72F95FB6.

Khattree:1999:EPC

[Kha99] Ravindra Khattree. On the estimation of σ and the process capability indices C_{p} and C_{pm}. Probability in the Engineering and Informational Sciences, 13(2):237–250, April 1999. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
C830FE71BD0E4DCEAF187DCE92E7E01F.

Khan:2007:SRB

[Kha07] Rasul A. Khan. Some remarks on Blackwell–Ross martingale inequalities. Probability in the Engineering and Informational Sciences, 21(1):109–115, January 2007. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
3141CC68E6FF02339C252B546F45AAEB.

Khaledi:2013:COP

[Kha13] Baha-Eldin Khaledi. Comments on: “Ordering Properties of Order Statistics from Heterogeneous Populations: a Review with an Emphasis on Some Recent Developments” by N. Balakrishnan and P. Zhao. Probability in the Engineering and Informational Sciences, 27(4):463–464, October 2013. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
EC7A47C0EA6D5C868E7FABB7A46210AB. See [BZ13b].

Kherani:2008:APS

[Khe08] Arzad A. Kherani. Asymptotic properties of sojourn times in multiclass time-shared systems. Probability in the Engineering and Informational Sciences, 22(2):231–259, March 2008. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
1F4CEA2368E6760ACD7A2C337352DE5A.

Kijima:1992:NEU

[Kij92] Masaaki Kijima. A note on external uniformization for finite Markov chains in continuous time. Probability in the Engineering and Informational Sciences, 6(1):127–131, January 1992. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
EF80C622BCF0F6D840ED7F6059ACFFA8.

Kijima:1994:SBD

[Kij94] Masaaki Kijima. On separation for birth–death processes. Probability in the Engineering and Informational Sciences, 8(1):51–68, January 1994. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/311563A7CE208D74AAC7FB20744204B9.

Kim:2022:SAB

[Kim22] Song-Kyoo (Amang) Kim. Strategic alliance for blockchain governance game. Probability in the Engineering and Informational Sciences, 36(1):184–200, January 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/strategic-
alliance-for-blockchain-governance-game/7AC99D9AA58C6C871BE1FBAD12B1C7C0.

Khaledi:2000:DAS

[KK00] Baha-Eldin Khaledi and Subhash Kochar. Dependence among spacings. Probability in the Engineering and Informational Sciences, 14(4):461–472, October 2000. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
D08D56D076DACC49E96E5D2A5A9C4B22.

Khaledi:2001:SPS

[KK01] Baha-Eldin Khaledi and Subhash Kochar. Stochastic properties of spacings in a single-outlier exponential model. Probability in the Engineering and Informational Sciences, 15(3): 401–408, July 2001. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/073ECF7B63E9AAAD819F49181AA6124A.

Khaledi:2005:DDM

[KK05] Baha-Eldin Khaledi and Subhash Kochar. Dependence, dispersiveness, and multivariate hazard rate ordering. Probability in the Engineering and Informational Sciences, 19(4):427–446, October 2005. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/1CA1C4B7ECCB672B6EF8A19E4AB8BD10. See correc-

tion [KK06].

Khaledi:2006:CDD

[KK06] Baha-Eldin Khaledi and Subhash Kochar. Correction to “Dependence, Dispersiveness, and Multivariate Hazard Rate Ordering”. Probability in the Engineering and Informational Sciences, 20(2):381, April 2006. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/25AA57E6AE17100DE4ECE5DE4B0F8F17.

See [KK05].

Kim:2022:PFE

[KKJ22] Kyong-Hui Kim, Jong-Kuk Kim, and Ho-Bom Jo. Pricing formula for exchange option based on stochastic delay differential equation with jumps. Probability in the Engineering and Informational Sciences, 36(2):548–563, April 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/pricing-
formula-for-exchange-option-based-on-stochastic-delay-
differential-equation-with-jumps/900724669CE0264BAD5B6E8316166E47.

Kim:2017:AAS

[KKL17] Hyungsu Kim, Geonwoo Kim, and Sungchul Lee. The analytic approach for the stochastic projection of the public pension fund. Probability in the Engineering and Informational Sciences, 31(2):196–206, April 2017. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
A11C29B541F705E68E4BB4BF40277229.

Koole:1998:SBQ

[KL98] Ger Koole and Zhen Liu. Stochastic bounds for queueing systems with multiple on-off sources. Probability in the Engineering and Informational Sciences, 12(1):25–48, January 1998. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
AC2615DF990D7454C4EC689E00F1BDA8.

Kella:2010:MMG

[KL10] Offer Kella and Andreas Löpker. A Markov-modulated growth collapse model. Probability in the Engineering and Informational Sciences, 24(1):99–107, January 2010. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
411B372FA4BF5A058D386C2860E05692.

Kella:2012:GMP

[KL12] Offer Kella and Andreas Löpker. A generalized memoryless property. Probability in the Engineering and Informational Sciences, 26(3):425–436, July 2012. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
0EF8D37BF3BC01CD2295D6807A2AE87E.

Kerner:2015:GSE

[KL15] Yoav Kerner and Andreas Löpker. On a generalization of the stationary excess operator. Probability in the Engineering and Informational Sciences, 29(2):219–232, April 2015. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
6DB93821F8BDD602FDAF9E097FB5401C.

Klar:2000:BTP

[Kla00]  Bernhard Klar.  Bounds on tail probabilities of dis-

crete distributions. Probability in the Engineering and Informational Sciences, 14(2):161–171, April 2000. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
BEC173F112EC2BF5283B2E651D3663EE.

Kwon:2012:OMP

[KLT12] H. Dharma Kwon, Steven A. Lippman, and Christopher S. Tang. Optimal markdown pricing strategy with demand learning. Probability in the Engineering and Informational Sciences, 26(1):77–104, January 2012. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
41029BADE1FDE508CB82507AB2A63716.

Katehakis:1988:DRA

[KM88] Michael N. Katehakis and Costis Melolidakis. Dynamic repair allocation for a k-out-of-n system maintained by distinguishable repairmen. Probability in the Engineering and Informational Sciences, 2(1):51–62, January 1988. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
ABCAA9F7372BE3F5A00519B5F19701BC.

Koutras:2007:IPP

[KM07] M. V. Koutras and P. E. Maravelakis. On the IFR property preservation for Markov chain imbeddable systems. Probability in the Engineering and Informational Sciences, 21(2): 201–216, April 2007. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/0009DFCB357AC9E1AB6A183D356850AC.

Koot:2012:ASG

[KM12] Matthijs R. Koot and Michel Mandjes. The analysis of singletons in generalized birthday problems. Probability in the Engineering and Informational Sciences, 26(2):245–262, April 2012. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
43C7047F9F23035C08F080B1D11D75D0.

Kalpathy:2016:DPA

[KM16] Ravi Kalpathy and Hosam Mahmoud. Degree profile of m-ary search trees: a vehicle for data structure compression. Probability in the Engineering and Informational Sciences, 30(1): 113–123, January 2016. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/790B4495DD2902C2E6B5A66DAA03F5E7.

Khan:2019:SBS

[KM19] Ruhul Ali Khan and Murari Mitra. Sharp bounds for survival probability when ageing is not monotone. Probability in the Engineering and Informational Sciences, 33(2):205–219, April 2019. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/sharp-bounds-for-survival-probability-when-
ageing-is-not-monotone/2E574E780670775BBEC260A80AD62946.

Kobayashi:2015:SRT

[KMS15] Masahiro Kobayashi, Masakiyo Miyazawa, and Hiroshi Shimizu. Structure-reversibility of a two-dimensional reflecting random walk and its application to queueing network. Probability in the Engineering and Informational Sciences, 29(1): 1–25, January 2015. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/CA527A89856D228B92004F8B2BAA828E.

Kayal:2020:QBS

[KMS20] Suchandan Kayal, Rajesh Moharana, and S. M. Sunoj. Quantile-based study of (dynamic) inaccuracy measures. Probability in the Engineering and Informational Sciences, 34(2): 183–199, April 2020. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/journals/probability-in-the-engineering-and-informational-
sciences/article/quantilebased-study-of-dynamic-inaccuracy-
measures/1722C2FCE3CD80F36CB1D41BC96A8950.

Kuhn:2018:EAS

[KMT18] Julia Kuhn, Michel Mandjes, and Thomas Taimre. Exact asymptotics of sample-mean-related rare-event probabilities. Probability in the Engineering and Informational Sciences, 32(2):207–228, April 2018. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/exact-
asymptotics-of-samplemeanrelated-rareevent-probabilities/
55C1C7539D24AD83B9AC5DCB2D423189.

Kobza:2000:TAP

[KN00] John E. Kobza and Joel A. Nachlas. Transient analysis of permanent customers in a single-server queue with mixed traffic. Probability in the Engineering and Informational Sciences, 14 (3):335–352, July 2000. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/082440B840FD85E6F4F60295F93D7725.

Kiessler:2021:AAB

[KN21] Peter C. Kiessler and Kanoktip Nimitkiatklai. An analytical approach for the breakdown distribution of a thin oxide. Probability in the Engineering and Informational Sciences, 35(3):595–610, July 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/an-analytical-
approach-for-the-breakdown-distribution-of-a-thin-oxide/
E80A1749151DCCB4623A7C82D1F6D5F0.

Koole:2015:OOP

[KNN15] G. M. Koole, B. F. Nielsen, and T. B. Nielsen. Optimization of overflow policies in call centers. Probability in the Engineering and Informational Sciences, 29(3):461–471, July 2015. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
99D1FCC5654FB227E49A9DC8E3F061DF.

Kochar:1996:SRI

[Koc96] Subhash C. Kochar. Some results on interarrival times of nonhomogeneous Poisson processes. Probability in the Engineering and Informational Sciences, 10(1):75–85, January 1996. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
2264D58E6226D6B6A0B91066D5C45E32.

Kochar:2023:DCO

[Koc23] Subhash Kochar. Dependence comparisons of order statistics in the proportional hazards model. Probability in the Engineering and Informational Sciences, 37(3):730–736, July 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/dependence-comparisons-of-order-statistics-
in-the-proportional-hazards-model/30E4B4068CF5CD782E2A9AE7CF364B65.

Kohlas:1988:CBS

[Koh88] Jürg Kohlas. Conditional belief structures. Probability in the Engineering and Informational Sciences, 2(4):415–433, October 1988. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/ABB7F2C65E0E972576D517B102A2B0FA.

Kolonko:1992:GUD

[Kol92] M. Kolonko. Generating uniformly distributed random numbers without floating point operations. Probability in the Engineering and Informational Sciences, 6(1):139–145, January 1992. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
4C1B47A0C509ECB109FC6000F69F7D18.

Kollman:1994:DPC

[Kol94] Craig Kollman. Deep-penetration calculations for scattering neutrons by importance sampling. Probability in the Engineering and Informational Sciences, 8(2):201–211, April 1994. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
AC625F7653A4DC497F0F21E9C0ADB912.

Konsowa:2002:CRT

[Kon02] Mokhtar H. Konsowa. Conductivity of random trees. Probability in the Engineering and Informational Sciences, 16(2): 233–240, April 2002. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/ED8B1D9B9F907E33DC68078E2A395E10.

Konsowa:2003:NCR

[Kon03] Mokhtar H. Konsowa. A note on the conductivity of random trees. Probability in the Engineering and Informational Sciences, 17(4):571–576, October 2003. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
70719BF8FCAE404ED6912DD000231282.

Koole:2004:CTQ

[Koo04] Ger Koole. Convexity in tandem queues. Probability in the Engineering and Informational Sciences, 18(1):13–31, January 2004. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
611F0137618B32828B4EF8E2B1CFCF6D.

Korwar:2003:SOL

[Kor03] Ramesh Korwar. On stochastic orders for the lifetime of a k-out-of-n system. Probability in the Engineering and Informational Sciences, 17(1):137–142, January 2003. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
37A4BBF915018A1C17AE8603AB89907F.

Kuba:2007:WPL

[KP07] M. Kuba and A. Panholzer. On weighted path lengths and distances in increasing trees. Probability in the Engineering and Informational Sciences, 21(3):419–433, July 2007. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
04C448A7D4643C8922E7A276AE10A796.

Kapodistria:2012:SER

[KP12] Stella Kapodistria and Georgios Psarrakos. Some extensions of the residual lifetime and its connection to the cumulative residual entropy. Probability in the Engineering and Informational Sciences, 26(1):129–146, January 2012. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
670C02B9B1DA1319CAF63A1D5446C833.

Kapodistria:2016:LBI

[KPDR16] Stella Kapodistria, Tuan Phung-Duc, and Jacques Resing. Linear birth/immigration-death process with binomial catastrophes. Probability in the Engineering and Informational Sciences, 30(1):79–111, January 2016. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/34EF37F39B715CA5B6E7C43F2A605BFE.

Klar:2010:ZLL

[KPH10] B. Klar, P. R. Parthasarathy, and N. Henze. Zipf and Lerch limit of birth and death processes. Probability in the Engineering and Informational Sciences, 24(1):129–144, January 2010. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
E5719B124DBDEE4FD935A6441849F195.

Kella:2003:SCM

[KPS03] Offer Kella, David Perry, and Wolfgang Stadje. A stochastic clearing model with a Brownian and a compound Poisson component. Probability in the Engineering and Informational Sciences, 17(1):1–22, January 2003. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
86755D1C0FADBC313DEF127E3FBDA483.

Keilson:1987:RTT

[KR87] Julian Keilson and Ravi Ramaswamy. The relaxation time for truncated birth–death processes. Probability in the Engineering and Informational Sciences, 1(4):367–381, October 1987. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
50EB04416EB313B8204F4D0FF1425792.

Kulkarni:1994:FMD

[KR94] Vidyadhar Kulkarni and Tomasz Rolski. Fluid model driven by an Ornstein–Uhlenbeck process. Probability in the Engineering and Informational Sciences, 8(3):403–417, July 1994. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
4ED62752F39F5B6B5D20C63ABE4E152B.

Koole:2001:SBS

[KR01] Ger Koole and Rhonda Righter. A stochastic batching and scheduling problem. Probability in the Engineering and Informational Sciences, 15(4):465–479, October 2001. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
17EA080EB23FC53397224C01057B493A.

Krzesinski:1992:PFS

[KS92] A. E. Krzesinski and R. Schassberger. Product form solutions for multiserver centers with hierarchical concurrency constraints. Probability in the Engineering and Informational Sciences, 6(2):147–156, April 1992. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
2448706E5B83777C40249E98C5820DDE.

Koole:2001:DMB

[KS01] G. M. Koole and F. M. Spieksma. On deviation matrices for birth–death processes. Probability in the Engineering and Informational Sciences, 15(2):239–258, April 2001. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
3A0F2552F1651B2DBA89968EFE67054D.

Kella:2002:ERF

[KS02] Offer Kella and Wolfgang Stadje. Exact results for a fluid model with state-dependent flow rates. Probability in the Engineering and Informational Sciences, 16(4):389–402, October 2002. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
71F26E9103BCAE383D48B779EEE176BD.

Katehakis:2012:SLP

[KS12] Michael N. Katehakis and Laurens C. Smit. A successive lumping procedure for a class of Markov chains. Probability in the Engineering and Informational Sciences, 26(4):483–508, October 2012. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/88684ECDBC16C6F26B52743C65D9D8C3.

Kubat:1988:DPE

[KSM88] Peter Kubat, Ushio Sumita, and Yasushi Masuda. Dynamic performance evaluation of communication/computer systems with highly reliable components. Probability in the Engineering and Informational Sciences, 2(2):185–213, April 1988. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
EAAC6A9CDA4B077CB5454B9F053D8478.

Katehakis:2015:RPT

[KSS15] Michael N. Katehakis, Laurens C. Smit, and Floske M. Spieksma. DES and RES processes and their explicit solutions. Probability in the Engineering and Informational Sciences, 29 (2):191–217, April 2015. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/21C760BA5DF0CC26ACB81E088472E27F.

Karsten:2014:DEE

[KSvH14] Frank Karsten, Marco Slikker, and Geert-Jan van Houtum. Domain extensions of the Erlang loss function: their scalability and its applications to cooperative games. Probability in the Engineering and Informational Sciences, 28(4):473–488, October 2014. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/3BE7649F1FC5E6B40F60D987A3B413A2.

Kolonko:1997:CSA

[KT97] M. Kolonko and M. T. Tran. Convergence of simulated annealing with feedback temperature schedules. Probability in the Engineering and Informational Sciences, 11(3):279–304, July 1997. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
7947320556D0135A7C97326E15B47B15.

Kawanishi:2001:PAS

[KT01] Ken’ichi Kawanishi and Yoshitaka Takahashi. Performance analysis of a server for facsimile communication networks. Probability in the Engineering and Informational Sciences, 15 (3):295–310, July 2001. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/195BA8D18B1BAD853BAFD61C14FAF475.

Katariya:2014:CRL

[KTÇ14] Abhilasha Prakash Katariya, Eylem Tekin, and Sila Çetinkaya. A capacitated replenishment-liquidation model under contractual and spot markets with stochastic demand. Probability in the Engineering and Informational Sciences, 28(2):223–269, April 2014. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/C0BC7A63F8DAA719A1D256A7186469FB.

Kulik:2004:MLN

[Kul04] Rafal Kulik. Monotonicity in lag for nonmonotone Markov chains. Probability in the Engineering and Informational Sciences, 18(4):485–492, October 2004. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/6CEFB7D07C4114BA389D73D97E951F67.

Konstantopoulos:1989:QRA

[KW89] Panagiotis Konstantopoulos and Jean Walrand. A quasireversibility approach to the insensitivity of generalized semi-Markov processes. Probability in the Engineering and Informational Sciences, 3(3):405–415, July 1989. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
CE060E98D11684582028B591C193BC07.

Kesidis:1994:CRF

[KW94] G. Kesidis and J. Walrand. Conservation relations for fully shared ATM buffers. Probability in the Engineering and Informational Sciences, 8(1):147–151, January 1994. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
BA601C3A8510D78C287A898061D07C68.

Kim:2013:EWT

[KW13] Song-Hee Kim and Ward Whitt. Estimating waiting times with the time-varying Little’s Law. Probability in the Engineering and Informational Sciences, 27(4):471–506, October 2013. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
18BF94EF65D06A5E52B72D1F81F3BFE0.

Klutke:1996:AIS

[KWA96] Georgia-Ann Klutke, Martin A. Wortman, and Hayriye Ayhan. The availability of inspected systems subject to random deterioration. Probability in the Engineering and Informational Sciences, 10(1):109–118, January 1996. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
07442767452003F6884919E73AD04AB9.

Kwon:2014:PCF

[Kwo14] H. Dharma Kwon. Prevention of catastrophic failures with weak forewarning signals. Probability in the Engineering and Informational Sciences, 28(1):121–144, January 2014. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
4EAD737C02C17CC8ABEFF28155E4B66A.

Kochar:2007:SCP

[KX07] Subhash Kochar and Maochao Xu. Stochastic comparisons of parallel systems when components have proportional hazard rates. Probability in the Engineering and Informational Sciences, 21(4):597–609, October 2007. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
A46B7D963417044DE35BD4D1048D21C6. See correction [KX08].

Kochar:2008:CSC

[KX08] Subhash Kochar and Maochao Xu. Correction to “Stochastic Comparisons of Parallel Systems When Components have Proportional Hazard Rates”. Probability in the Engineering and Informational Sciences, 22(3):473–474, July 2008. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
26441F04D607A746E50B22603ED988AB. See [KX07].

Kochar:2010:RLS

[KX10] Subhash Kochar and Maochao Xu. On residual lifetimes of k-out-of-n systems with nonidentical components. Probability in the Engineering and Informational Sciences, 24(1):109–127, January 2010. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/CA121E494501A7F8F9454726B296789E.

Kochar:2012:SUR

[KX12] Subhash Kochar and Maochao Xu. Some unified results on comparing linear combinations of independent gamma random variables. Probability in the Engineering and Informational Sciences, 26(3):393–404, July 2012. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
A6626B20C1683A7CFD166DFE4B7E2E47.

Kyriakidis:2003:OCS

[Kyr03] E. G. Kyriakidis. Optimal control of a simple immigration process through the introduction of a predator. Probability in the Engineering and Informational Sciences, 17(1):119–135, January 2003. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/9C696E818F56FAE9643BBCFD481007F5.

Kyriakidis:2004:TSS

[Kyr04] E. G. Kyriakidis. Transient solution for a simple immigration birth–death catastrophe process. Probability in the Engineering and Informational Sciences, 18(2):233–236, April 2004. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
F516641C7D0A058134D732520AB851CA.

Lewis:1999:BOQ

[LAF99] Mark E. Lewis, Hayriye Ayhan, and Robert D. Foley. Bias optimality in a queue with admission control. Probability in the Engineering and Informational Sciences, 13(3):309–327, July 1999. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
6ECE887FDF346B00B9805D8296C55C7F.

Lagnoux:2006:RES

[Lag06] Agnès Lagnoux. Rare event simulation. Probability in the Engineering and Informational Sciences, 20(1):45–66, January 2006. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
A1DFA09B5C6AAC941D46CFCE6978DAC3.

Lalley:1990:TSS

[Lal90] Steven P. Lalley. Travelling salesman with a self-similar itinerary. Probability in the Engineering and Informational Sciences, 4(1):1–18, January 1990. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/84399BBCEA8ED26766CA9CFBD36AFDF0.

Lasserre:1998:LCI

[Las98] Jean B. Lasserre. A Lyapunov criterion for invariant probabilities with geometric tail. Probability in the Engineering and Informational Sciences, 12(3):387–391, July 1998. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
034B795CB15C2E284D8220F34FB7DA50.

Litvak:2001:OPC

[LAWZ01] N. Litvak, I. J. B. F. Adan, J. Wessels, and W. H. M. Zijm. Order picking in carousel systems under the nearest item heuristic. Probability in the Engineering and Informational Sciences, 15(2):135–164, April 2001. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
BC3957EAE342935789E7F20F856D3DE6.

Lee:1997:APK

[LC97] Tian-Shyug Lee and Kwang-Chow Chang. An alternative proof for a known result of noncentral Wishart distribution. Probability in the Engineering and Informational Sciences, 11(4): 523–529, October 1997. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/C86D61722FF0E6FD6422A0DC6B7CDC06.

Liu:2007:RCS

[LC07] Ching-Ti Liu and Yen-Chang Chang. A reliability-constrained software release policy using a non-Gaussian Kalman filter model. Probability in the Engineering and Informational Sciences, 21(2):301–314, April 2007. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/7426A4C0F859A843FE7C1F2F3496CBC2.

Long:2022:REP

[LCYY22] Chunxian Long, Wangxue Chen, Rui Yang, and Dongsen Yao. Ratio estimation of the population mean using auxiliary information under the optimal sampling design. Probability in the Engineering and Informational Sciences, 36(2):449–460, April 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/ratio-estimation-of-the-population-mean-using-
auxiliary-information-under-the-optimal-sampling-design/
FBAE38294EF04335F112D20E9F5BF040.

Liu:2022:PPI

[LDVT22] Xingheng Liu, Yann Dijoux, Jørn Vatn, and Håkon Toftaker. Performance of prognosis indicators for superimposed renewal processes. Probability in the Engineering and Informational Sciences, 36(1):17–40, January 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/performance-
of-prognosis-indicators-for-superimposed-renewal-processes/
D1AEC754A52A3BDCD6B354A762495172.

Lindgren:1994:CZC

[LE94] Georg Lindgren and Weine Eriksson. Clicks in zero-crossing-detecting FM receivers: Theory and experiments. Probability in the Engineering and Informational Sciences, 8(2):265–285, April 1994. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/85228C4606018FD97059A1A5611CDDD6.

Lefebvre:1996:FPD

[Lef96] Mario Lefebvre. First-passage distributions of bidimensional processes. Probability in the Engineering and Informational Sciences, 10(3):325–339, July 1996. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
312233E7961229B51A3265668F35AFB8.

Lefebvre:2001:MTS

[Lef01] Mario Lefebvre. Maximizing the time spent by a diffusion process in an interval. Probability in the Engineering and Informational Sciences, 15(1):69–82, January 2001. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
B0DEB7157AA0F0D665A483133DAF0AF5.

Lefevre:2006:OCF

[Lef06] Claude Lefèvre. On the outcome of a cascading failure model. Probability in the Engineering and Informational Sciences, 20 (3):413–427, July 2006. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/48ECA25265A2F46B28DB70D2C81B2C55.

Li:2023:ODN

[LFNC23] Jingchao Li, Zichen Fang, Ciyu Nie, and Sizhe Chen. Optimal design for network mutual aid. Probability in the Engineering and Informational Sciences, 37:567–596, April 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/optimal-design-for-network-mutual-aid/3ED885C19AF94C6EB0F9BD61F9286605.

Lee:1989:PCI

[LG89] Mel-Ling Ting Lee and Alan J. Gross. Properties of conditionally independent generalized gamma distributions. Probability in the Engineering and Informational Sciences, 3(2): 289–297, April 1989. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/0D6907B1F37BF7777C5028845C28895C.

Li:2008:TDS

[LHL08] Xiaohu Li, Xiaoxiao Hu, and Zhouping Li. Tp 2 dependence of sample spacings with applications. Probability in the Engineering and Informational Sciences, 22(2):287–300, March 2008. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
41D74F9C5B56B3725E213F6E318A25DB.

Lin:2019:OSK

[LHY19] Yi-Shen Lin, Shoou-Ren Hsiau, and Yi-Ching Yao. Optimal selection of the k-th best candidate. Probability in the Engineering and Informational Sciences, 33(3):327–347, July 2019. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/optimal-selection-of-the-kth-best-candidate/
CAA3096A7C7A9323F6DD1E3FF3B7A1C9.

Li:1994:CSA

[Li94] Haijun Li. On a class of stochastic arrangement inequalities arising in optimal allocation of resources. Probability in the Engineering and Informational Sciences, 8(1):113–124, January 1994. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
14143235EFC5B1C3E64010AA06743F20.

Li:2005:PNN

[Li05] Yulin Li. Preservations of NBUC and NBU(2) classes under mixtures. Probability in the Engineering and Informational Sciences, 19(2):277–288, April 2005. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
0ECE1BBFC56DDBFF738FC251345493F5.

Lin:1990:PCA

[Lin90a] Jinn-Tyan Lin. Pocket-calculator approximation to the normal tail. Probability in the Engineering and Informational Sciences, 4(4):531–533, October 1990. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
51C8351F4DFAF4C904FFEC34AF621AF0.

Lin:1990:AKA

[Lin90b] Jonn-Tyan Lin. Alternative to Koehler’s approximation for the percentiles of the t distributions. Probability in the Engineering and Informational Sciences, 4(4):535–537, October 1990. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
ADD5A999C56628BC7A6181400E88535D.

Lin:1992:AST

[Lin92a] Jinn-Tyan Lin. Approximating the Student’s t-tail probability and its inverse. Probability in the Engineering and Informational Sciences, 6(1):133–137, January 1992. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
BF48D36D91865876FCBB701F0CCC87D7.

Lin:1992:PCA

[Lin92b] Jinn-Tyan Lin. A pocket-calculator approximation to the inverse normal tail probability function. Probability in the Engineering and Informational Sciences, 6(3):409–411, July 1992. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
A9CCE82072320358C86ED0E9176032CD.

Lin:1994:DAP

[Lin94a] Jinn-Tyan Lin. A direct approximation for the percentiles of the t distribution. Probability in the Engineering and Informational Sciences, 8(3):443–448, July 1994. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
0B5732F3E4E5A4BA04665B25DB9861D8.

Lin:1994:NAP

[Lin94b] Jinn-Tyan Lin. New approximations for the percentage points of the chi-square distribution. Probability in the Engineering and Informational Sciences, 8(1):135–146, January 1994. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
6772410093E8FDEA04C90EF8487398AF.

Lin:1995:SAB

[Lin95] Jinn-Tyan Lin. A simple approximation for the bivariate normal integral. Probability in the Engineering and Informational Sciences, 9(2):317–321, April 1995. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
9455086EE8DBAB83643B34ADCB9A410D.

Lindvall:1997:SMJ

[Lin97] Torgny Lindvall. Stochastic monotonicities in Jackson queueing networks. Probability in the Engineering and Informational Sciences, 11(1):1–9, January 1997. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
090E63E0479322B63F32BEC73D77E266.

Lindvall:2000:SSO

[Lin00] Torgny Lindvall. On simulation of stochastically ordered life-length variables. Probability in the Engineering and Informational Sciences, 14(1):1–7, January 2000. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
1EC41992ECD9A5CF5C64E5071F200EEC.

Lindvall:2002:CBC

[Lin02]  Torgny Lindvall.  Comparisons between certain queu-

ing systems. Probability in the Engineering and Informational Sciences, 16(1):1–17, January 2002. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
3C8CE2E3002BF39F62DA9DF57DAD06CB.

Lindvall:2004:RP

[Lin04] Torgny Lindvall. On a routing problem. Probability in the Engineering and Informational Sciences, 18(3):329–337, July 2004. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
E63CCF047C822045972FD27BD17A76A7.

Ling:2019:ODS

[Lin19] Man Ho Ling. Optimal design of simple step-stress accelerated life tests for one-shot devices under exponential distributions. Probability in the Engineering and Informational Sciences, 33(1):121–135, January 2019. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/optimal-
design-of-simple-stepstress-accelerated-life-tests-
for-oneshot-devices-under-exponential-distributions/
38F16DE5B3C70AF8316D38A32FCDD99E.

Lee:1996:MMF

[LK96] Tian-Shyug Lee and Georgia-Ann Klutke. Minimizing the makespan and flowtime in two-machine stochastic open shops. Probability in the Engineering and Informational Sciences, 10(4):557–567, October 1996. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
17680F0D90072CEE234EDF5B497EB9C5.

Lember:2007:AVT

[LK07] Jüri Lember and Alexey Koloydenko. Adjusted Viterbi training. Probability in the Engineering and Informational Sciences, 21(3):451–475, July 2007. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/412CFACDB1B3D43B24AC2FF760334CC5.

Liu:2008:BRM

[LK08] Liqiang Liu and Vidyadhar G. Kulkarni. Balking and reneging in M/G/s systems exact analysis and approximations. Probability in the Engineering and Informational Sciences, 22(3): 355–371, July 2008. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/4DA13DABF38ACB70D65EB2C547645CFF.

Liu:2009:DPM

[LK09] Nan Liu and Vidyadhar G. Kulkarni. Decomposition property for Markov-modulated queues with applications to warranty management. Probability in the Engineering and Informational Sciences, 23(3):433–447, July 2009. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
011927BD6FD892422D86BF4BC4BF0509.

Li:2023:RAP

[LK23] Johnny Siu-Hang Li and Joseph H. T. Kim. Rotation in age patterns of mortality decline: statistical evidence and modeling. Probability in the Engineering and Informational Sciences, 37:621–652, April 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/rotation-
in-age-patterns-of-mortality-decline-statistical-evidence-
and-modeling/9DC49B7B2BA73C45B5B0905A4F77BB39.

Liu:2019:BGB

[LKG19] Xin Liu, Vidyadhar G. Kulkarni, and Qi Gong. On bouncing geometric Brownian motions. Probability in the Engineering and Informational Sciences, 33(4):591–617, October 2019. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/on-bouncing-geometric-brownian-motions/6E24497ABF34A98B96F47616AB849B92.

Lee:2014:OSA

[LKH14] Nelson Lee, Vidyadhar G. Kulkarni, and Yasutaka Hirasawa. Optimal static assignment and routing policies for service centers with correlated traffic. Probability in the Engineering and Informational Sciences, 28(3):279–311, July 2014. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
05E7EEADE852D4054AE1AECC9341E269.

Laws:1990:DSF

[LL90] C. N. Laws and G. M. Louth. Dynamic scheduling of a four-station queueing network. Probability in the Engineering and Informational Sciences, 4(1):131–156, January 1990. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
DB545393C835DD6AD92A9D1532A338B6.

Lee:1991:SFS

[LL91] Chung-Yee Lee and Chen-Sin Lin. Stochastic flow-shop scheduling with lateness-related performance measures. Probability in the Engineering and Informational Sciences, 5(2): 245–254, April 1991. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/28AEA5AFB92D4CC8B4A6CCFCF7609F43.

Langseth:1998:UBM

[LL98] Helge Langseth and Bo Henry Lindqvist. Uncertainty bounds for a monotone multistate system. Probability in the Engineering and Informational Sciences, 12(2):239–260, April 1998. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
3E77BBD5F10822AD2141347B4674AE29.

Li:2003:SCR

[LL03] Xiaohu Li and Jingyi Lu. Stochastic comparisons on residual life and inactivity time of series and parallel systems. Probability in the Engineering and Informational Sciences, 17(2): 267–275, April 2003. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/17907C06250F8BFFC5FA3AEB09153C1C.

Lim:2022:AAP

[LLY22] Zhao Ging Lim, Chen-Tuo Liao, and Yi-Ching Yao. Asymptotic analysis of Peres’ algorithm for random number generation. Probability in the Engineering and Informational Sciences, 36(2):341–356, April 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/asymptotic-
analysis-of-peres-algorithm-for-random-number-generation/
3B02FFEA433E0788CFD05B30974E83D3. See Yuval Peres, It-

erating von Neumann’s Procedure for Extracting Random Bits, Annals of Statistics 20(1) 590–597, March 1992, doi:10.1214/aos/1176348543.

Leow:1987:HAR

[LM87] Soo Kar Leow and David F. Mcallister. Heuristic assignments of redundant software versions and processors in faulttolerant computer systems for maximum reliability. Probability in the Engineering and Informational Sciences, 1(4):457–479, October 1987. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/ACE40A3779A1FAC3A6ED4E136FFD8845.

Lippman:1988:WMW

[LM88] Steven A. Lippman and John W. Mamer. When many wrongs make a right: An asymptotic analysis of risk aversion and additive risks. Probability in the Engineering and Informational Sciences, 2(1):115–127, January 1988. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
06A3DE3D343E87A2AE655903E96CDC98.

Lv:2012:PSO

[LMH12] Wenhua Lv, Tiantian Mao, and Taizhong Hu. Properties of second-order regular variation and expansions for risk concentration. Probability in the Engineering and Informational Sciences, 26(4):535–559, October 2012. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
59F6F06714BC337C540AF076B688D471.

Liu:2014:SOR

[LMH14] Qing Liu, Tiantian Mao, and Taizhong Hu. The secondorder regular variation of order statistics. Probability in the Engineering and Informational Sciences, 28(2):209–222, April 2014. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
86F53346808764C61A5441EC6B07F7DA.

Liu:2022:HOC

[LMZ22] Pengfei Liu, Xuejun Ma, and Wang Zhou. High-order conditional distance covariance with conditional mutual independence. Probability in the Engineering and Informational Sciences, 36(1):126–143, January 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/highorder-
conditional-distance-covariance-with-conditional-mutual-
independence/B5B2CF167B24FD2789DC26BBEB4E7AE2.

Li:1992:WTD

[LN92] Jingwen Li and Shun-Chen Niu. The waiting-time distribution for the GI/G/1 queue under the D-policy. Probability in the Engineering and Informational Sciences, 6(3):287–308, July 1992. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
10A977A002C800E27FF04769A9621F79.

Lalropuia:2022:GTM

[LnG22] Kc Lalropuia and Vandana Khaitan (nee Gupta). Game theoretic modeling of economic denial of sustainability (EDoS) attack in cloud computing. Probability in the Engineering and Informational Sciences, 36(4):1241–1265, October 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/game-theoretic-modeling-of-economic-denial-
of-sustainability-edos-attack-in-cloud-computing/0816D54636686667365F72C925BB40E8.

Leite:1987:EML

[LOdBP87] Jose Galvão Leite, Jorge Oishi, and Carlos Alberto de Bragança Pereira. Exact maximum likelihood estimate of a finite population size: Capture/recapture sequential sample data. Probability in the Engineering and Informational Sciences, 1 (2):225–236, April 1987. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/641C6A9061340E6F6945C52D2734185D.

Lomonosov:1994:MCE

[Lom94] M. Lomonosov. On Monte Carlo estimates in network reliability. Probability in the Engineering and Informational Sciences, 8(2):245–264, April 1994. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/433A2D52D38462BDAECAF1BF764D1215.

Lorek:2018:SDM

[Lor18] Pawel Lorek. Siegmund duality for Markov chains on partially ordered state spaces. Probability in the Engineering and Informational Sciences, 32(4):495–521, October 2018. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/siegmund-
duality-for-markov-chains-on-partially-ordered-state-
spaces/9A0361552C9E5BF0B6C7B6D6CEAD7C4A.

Lefevre:2015:RMI

[LP15] Claude Lefèvre and Philippe Picard. Risk models in insurance and epidemics: a bridge through randomized polynomials. Probability in the Engineering and Informational Sciences, 29 (3):399–420, July 2015. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/4930DE9D1BAE903F4A43B51D8E3C827A.

Luo:2023:AAD

[LP23] Liuling Luo and Xingchun Peng. Asset allocation for a DC pension plan with minimum guarantee constraint and hidden Markov regime-switching. Probability in the Engineering and Informational Sciences, 37(4):1035–1062, October 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/asset-allocation-for-a-dc-pension-plan-with-
minimum-guarantee-constraint-and-hidden-markov-regimeswitching/
51EA77D3331A31DA3D8D9817639BB484.

Lim:2015:POD

[LQ15] Thomas Lim and Marie-Claire Quenez. Portfolio optimization in a default model under full/partial information. Probability in the Engineering and Informational Sciences, 29(4):565–587, October 2015. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/E270468DA81108B31A3D48148D60847A.

Lieberman:1991:VHE

[LR91] Gerald J. Lieberman and Sheldon M. Ross. On the variance of the hazard estimator in simulation. Probability in the Engineering and Informational Sciences, 5(3):355–359, July 1991. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
C436C84883385C8486DB6EFEF4161160.

Liu:1997:OSP

[LR97] Zhen Liu and Rhonda Righter. Optimal scheduling on parallel processors with precedence constraints and general costs. Probability in the Engineering and Informational Sciences, 11(1): 79–93, January 1997. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/91A17D1DE629CD2FE6E0523DD7D5C061.

Levikson:1999:PHS

[LRW99] Benny Levikson, Tomasz Rolski, and Gideon Weiss. On a Poisson hyperbolic staircase. Probability in the Engineering and Informational Sciences, 13(1):11–31, January 1999. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
4EC38DB6FDB408191B167EC1753CB260.

Limon-Robles:2009:TDO

[LRW09] Jorge Limón-Robles and Martin A. Wortman. On the timedependent occupancy distribution of the G/G/1 queuing system. Probability in the Engineering and Informational Sciences, 23(2):261–280, April 2009. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/3B650F7B52C5BC74FEF41C3223362741.

Leijdekker:2011:ECF

[LS11] Vincent Leijdekker and Peter Spreij. Explicit computations for a filtering problem with point process observations with applications to credit risk. Probability in the Engineering and Informational Sciences, 25(3):393–418, July 2011.

CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
24154A3ACA573956A5B647EE3F520397.

Lu:2016:SOD

[LSSZ16] Yingdong Lu, Mayank Sharma, Mark S. Squillante, and Bo Zhang. Stochastic optimal dynamic control of GI_{m} /GI_{m} /1_{n} queues with time-varying workloads. Probability in the Engineering and Informational Sciences, 30(3):470–491, July 2016. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
1790DFEF3835BEF325A73CEFEF6975D7.

Liu:2022:IMI

[LST^{+}22] Yue Liu, Zhiyan Shi, Ying Tang, Jingjing Yao, and Xincheng Zhu. An improvement of Markovian integration by parts formula and application to sensitivity computation. Probability in the Engineering and Informational Sciences, 36(2):357–377, April 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/an-improvement-
of-markovian-integration-by-parts-formula-and-application-
to-sensitivity-computation/89C3E0FD6B130260844B7BEDCC20ACFC.

Lindqvist:2021:CPP

[LSW21] Bo H. Lindqvist, Francisco J. Samaniego, and Nana Wang. On the comparison of performance-per-cost for coherent and mixed systems. Probability in the Engineering and Informational Sciences, 35(4):867–884, October 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/on-the-
comparison-of-performancepercost-for-coherent-and-mixed-
systems/FDB12CCCD1873403234470385F675A11.

Lasserre:1996:IPG

[LT96] Jean B. Lasserre and Henk Tijms. Invariant probabilities with geometric tail. Probability in the Engineering and Informational Sciences, 10(2):213–221, April 1996. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
DA71454114268BED45E799A0D07B37BE.

Lott:2000:OIR

[LT00] Christopher Lott and Demosthenis Teneketzis. On the optimality of an index rule in multichannel allocation for single-hop mobile networks with multiple service classes. Probability in the Engineering and Informational Sciences, 14(3):259–297, July 2000. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
5603C6F41B100C826812D4ABA6C9ACC4.

Lu:2009:PEM

[Lu09] Yingdong Lu. Performance estimation of M/D_{K} /1 queue under fair sojourn protocol in heavy traffic. Probability in the Engineering and Informational Sciences, 23(1):61–74, January 2009. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
59789A253AA59E8B3F7DED634955AB83.

Lu:2010:RBA

[Lu10] Yingdong Lu. Ratio between average sojourn times under processor sharing and fair sojourn protocol. Probability in the Engineering and Informational Sciences, 24(4):485–490, October 2010. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
FEDFD0D7218DD907A9D7380DFE434B23.

Lu:2013:FHR

[Lu13] Yingdong Lu. Finite horizon ruin probabilities for random walks with heavy tailed increments. Probability in the Engineering and Informational Sciences, 27(2):237–246, April 2013. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
1D999BA6CD325889017FFE77A540ACB4.

Lu:2021:FLP

[Lu21] Yingdong Lu. A fluid limit for processor-sharing queues weighted by functions of remaining amounts of service. Probability in the Engineering and Informational Sciences, 35(3): 396–403, July 2021. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/journals/probability-in-the-engineering-and-informational-
sciences/article/fluid-limit-for-processorsharing-queues-
weighted-by-functions-of-remaining-amounts-of-service/
A860052583C8794856661EE4465EDBF6.

Lee:2013:CST

[LW13] Wei-Yi Lee and Chia-Li Wang. Conditional sojourn times of processor-sharing queues. Probability in the Engineering and Informational Sciences, 27(1):99–114, January 2013. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
66567B0C1624BC5AC82F96074B9400B3.

Liu:2014:SPN

[LW14] Yunan Liu and Ward Whitt. Stabilizing performance in networks of queues with time-varying arrival rates. Probability in the Engineering and Informational Sciences, 28(4):419–449, October 2014. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/B4C2A71AD264510A61CD0DAF0FD0448D.

Li:2019:NRE

[LW19] Wenxi Li and Zhongzhi Wang. A note on Rényi’s entropy rate for time-inhomogeneous Markov chains. Probability in the Engineering and Informational Sciences, 33(4):579–590, October 2019. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/note-on-renyis-entropy-rate-for-timeinhomogeneous-
markov-chains/E8C54BF5052778A040C211E09C9468FD.

Ling:2019:OHC

[LWL19] Xiaoliang Ling, Yinzhao Wei, and Ping Li. On optimal heterogeneous components grouping in series-parallel and parallelseries systems. Probability in the Engineering and Informational Sciences, 33(4):564–578, October 2019. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/on-optimal-
heterogeneous-components-grouping-in-seriesparallel-
and-parallelseries-systems/86D2BA64D52F69F943EBE8B80F4CD134.

Li:2016:SSB

[LWZ16] Andrew Li, Ward Whitt, and Jingtong Zhao. Staffing to stabilize blocking in loss models with time-varying arrival rates. Probability in the Engineering and Informational Sciences, 30 (2):185–211, April 2016. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/B22F61E8B47DE34664494AF87715932D.

Li:2006:SRA

[LX06] Xiaohu Li and Maochao Xu. Some results about MIT order and IMIT class of life distributions. Probability in the Engineering and Informational Sciences, 20(3):481–496, July 2006. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
AAA5578C2684694CBE3EFB797D4671AE.

Li:2004:PRI

[LXH04] Zhaolin Li, Susan H. Xu, and Jack Hayya. A periodic-review inventory system with supply interruptions. Probability in the Engineering and Informational Sciences, 18(1):33–53, January 2004. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
F75467BA11BD2E0E45E486562581AE39.

Liu:2021:VAR

[LXYL21] Yue Liu, Zhuyun Xie, Jingjing Yao, and Kaodui Li. Volatility analysis of regime-switching models. Probability in the Engineering and Informational Sciences, 35(4):928–941, October 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/volatility-analysis-of-regimeswitching-models/
A5C6FB8CB45E8B79B418AF3F619EF238.

Liu:2004:SSL

[LY04] Wen Liu and Weiguo Yang. Some strong limit theorems for Markov chain fields on trees. Probability in the Engineering and Informational Sciences, 18(3):411–422, July 2004. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
38FEA2CE6C070403F9B8641F4D4C45F8.

Li:2013:COP

[LY13] Xiaohu Li and Yinping You. Comments on “Ordering Properties of Order Statistics from Heterogeneous Populations”. Probability in the Engineering and Informational Sciences, 27(4):455–462, October 2013. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
403CF3F63ECC4CA1478EF848AA98D6AC. See [BZ13b].

Li:2016:WKS

[LYF16] Xiaohu Li, Yinping You, and Rui Fang. On weighted k-out-of-n systems with statistically dependent component lifetimes. Probability in the Engineering and Informational Sciences, 30(4):533–546, October 2016. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
6F6393DCC0F86D81BD0860BBF39158DA.

Lynch:1997:WRP

[Lyn97] James D. Lynch. When is a renewal process convexly parameterized in its mean parameterization? Probability in the Engineering and Informational Sciences, 11(1):43–48, January 1997. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
EF84F6A67A692E09696FBD10AD07EE3C.

Lynch:1999:CMI

[Lyn99] James D. Lynch. On conditions for mixtures of increasing failure rate distributions to have an increasing failure rate. Probability in the Engineering and Informational Sciences, 13(1): 33–36, January 1999. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/6350D3E7774298599E531A33CEAD347D.

Lynch:2003:APC

[Lyn03] James Lynch. Acknowledgment of Priority: “On the conditions for mixtures of increasing failure rate distributions to have an increasing failure rate”. Probability in the Engineering and Informational Sciences, 17(1):153, January 2003. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
86CB9D9C9D37F2965B3655984D5F3E54.

Lin:2000:REL

[LZ00] Daming Lin and Ming J. Zuo. Reliability evaluation of a linear k-within-(r, s)-out-of-(m, n): F lattice system. Probability in the Engineering and Informational Sciences, 14(4):435–443, October 2000. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/8D22BDCC2FD0F3AEE95F13C42117849B.

Li:2010:PDJ

[LZ10] Jun Li and Yiqiang Q. Zhao. On the probability distribution of join queue length in a fork-join model. Probability in the Engineering and Informational Sciences, 24(4):473–483, October 2010. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
5280F7E62E018CB9E0CBE05100066189.

Li:2011:SNR

[LZ11] Xiaohu Li and Shuhong Zhang. Some new results on rényi entropy of residual life and inactivity time. Probability in the Engineering and Informational Sciences, 25(2):237–250, April 2011. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
D5EE57AAE7FA41F2846360FB75820B6C.

Long:2019:NMS

[LZ19] Zhenghua Long and Jiheng Zhang. A note on many-server fluid models with time-varying arrivals. Probability in the Engineering and Informational Sciences, 33(3):417–437, July 2019. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/note-on-manyserver-fluid-models-with-timevarying-
arrivals/851E86A8EBF96D5D9C8D0E1520A86310.

Leng:2016:DOM

[LZH16] Xuan Leng, Jinsen Zhuang, and Taizhong Hu. Dispersive ordering for the multivariate normal distribution. Probability in the Engineering and Informational Sciences, 30(2):141–152, April 2016. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/B82CFEA079833221EB5B51BDC6B2EA84.

Lu:2023:OAC

[LZY23] Bin Lu, Jiandong Zhang, and Rongfang Yan. Optimal allocation of a coherent system with statistical dependent subsystems. Probability in the Engineering and Informational Sciences, 37(1):29–48, January 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/optimal-

allocation-of-a-coherent-system-with-statistical-dependent-
subsystems/6C6C3FB8C1A7BA055B0202AB0D974EB7.

Madigan:1993:NEC

[Mad93] David Madigan. A note on equivalence classes of directed acyclic independence graphs. Probability in the Engineering and Informational Sciences, 7(3):409–412, July 1993. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
DEB661695E7BC6C038FF3C6287A55644.

Mahmoud:1991:LDP

[Mah91] Hosam M. Mahmoud. Limiting distributions for path lengths in recursive trees. Probability in the Engineering and Informational Sciences, 5(1):53–59, January 1991. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
19EF48EFF2BB678ED7FE4091E7B83036.

Mahmoud:2013:DMB

[Mah13a] Hosam M. Mahmoud. Drawing multisets of balls from tenable balanced linear urns. Probability in the Engineering and Informational Sciences, 27(2):147–162, April 2013. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
93BEDDA66B6E92C7FD259C30F72706E1.

Mahmoud:2013:SND

[Mah13b] Hosam M. Mahmoud. Some node degree properties of seriesparallel graphs evolving under a stochastic growth model. Probability in the Engineering and Informational Sciences, 27 (3):297–307, July 2013. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/1C730172DAA5D625EBB6904C2FBBD9C8.

Mahmoud:2014:SPB

[Mah14] Hosam M. Mahmoud. Some properties of binary seriesparallel graphs. Probability in the Engineering and Informational Sciences, 28(4):565–572, October 2014. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
90E62E1769BEBCDFBD50942AF84759C9.

Mahmoud:2019:SSP

[Mah19] Hosam M. Mahmoud. A spectrum of series-parallel graphs with multiple edge evolution. Probability in the Engineering and Informational Sciences, 33(4):487–499, October 2019. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/spectrum-of-seriesparallel-graphs-with-multiple-
edge-evolution/D35F70C5779923D3D93B9AF3BFCDA62D.

Mahmoud:2023:CPU

[Mah23] Hosam Mahmoud. Covariances in Pólya urn schemes. Probability in the Engineering and Informational Sciences, 37(1):60–71, January 2023. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
journals/probability-in-the-engineering-and-informational-
sciences/article/covariances-in-polya-urn-schemes/8E6D3F3355C3368A5CCF980CC405883C.

Malinovsky:2022:DWS

[Mal22] Yaakov Malinovsky. On the distribution of winners’ scores in a round-robin tournament. Probability in the Engineering and Informational Sciences, 36(4):1098–1102, October 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/on-the-distribution-of-winners-scores-in-a-
roundrobin-tournament/6B7D2FC12A87CD0B685F21940C5D3646.

See correction [Mal23].

Malinovsky:2023:CDW

[Mal23] Yaakov Malinovsky. Correction to “On the distribution of winners’ scores in a round-robin tournament”. Probability in the Engineering and Informational Sciences, 37(3):737–739, July 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/correction-to-on-the-distribution-of-winners-
scores-in-a-roundrobin-tournament/5C319A4EBAC62A5F3A60A358F71A0448.

See [Mal22].

Mandjes:1996:BRL

[Man96a] Michel Mandjes. Book review: Large Deviations for Performance Analysis: Queues, Communications, and Computing, Shwartz Adam and Weiss Alan (New York: Chapman and Hall, 1995). Probability in the Engineering and Informational Sciences, 10(4):613–614, October 1996. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
02D9E4A68660157A33542ADFE6FB0C1B.

Mandjes:1996:OAL

[Man96b] Michel Mandjes. Overflow asymptotics for large communications systems with general Markov fluid sources. Probability in the Engineering and Informational Sciences, 10(4):501–518, October 1996. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/9908393047F3A2F75A447E3224023B82.

Marin:2016:PFG

[Mar16] Andrea Marin. Product-form in G-networks. Probability in the Engineering and Informational Sciences, 30(3):345–360, July 2016. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
45284FBBF02B4B6377A037E5BE894AAB.

Mayorga:2006:OCM

[MAS06] Maria E. Mayorga, Hyun-Soo Ahn, and J. George Shanthikumar. Optimal control of a make-to-stock system with adjustable service rate. Probability in the Engineering and Informational Sciences, 20(4):609–634, October 2006. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
F387CF8D58FE9149F084F5656E7443E3.

Matalytski:2017:FNS

[Mat17] Mikhail Matalytski. Finding non-stationary state probabilities of G-network with signals and customers batch removal. Probability in the Engineering and Informational Sciences, 31(4): 396–412, October 2017. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/9BF07D3E9A9415F241F890C2EBC66881.

Matalytski:2019:ANMa

[Mat19a] Mikhail Matalytski. Analysis of the network with multiple classes of positive and negative customers at a transient regime. Probability in the Engineering and Informational Sciences, 33(2):172–185, April 2019. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/analysis-
of-the-network-with-multiple-classes-of-positive-and-
negative-customers-at-a-transient-regime/BF309343691DE1FE71677C3CC1DAD8C3.

Matalytski:2019:FER

[Mat19b] Mikhail Matalytski. Finding expected revenues in Gnetwork with multiple classes of positive and negative customers. Probability in the Engineering and Informational Sciences, 33(1):105–120, January 2019. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/finding-
expected-revenues-in-gnetwork-with-multiple-classes-
of-positive-and-negative-customers/9DC689F960FC87A72B0C98701E3B80E8.

Mamoun:2007:GCM

[MBP07] Mouad Ben Mamoun, Ana Busić, and Nihal Pekergin. Generalized class C Markov chains and computation of closedform bounding distributions. Probability in the Engineering and Informational Sciences, 21(2):235–260, April 2007. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
2E50289CD1442F981FA01711EF677745.

Mukherjee:1991:SPM

[MC91] S. P. Mukherjee and A. Chatterjee. On some properties of the multivariate aging classes. Probability in the Engineering and Informational Sciences, 5(4):523–534, October 1991. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
19238FF501AD6AEBC424FD28615C6AB3.

Miao:2011:VSS

[MC11] Daniel Wei-Chung Miao and Hung Chen. On the variances of system size and sojourn time in a discrete-time DAR(1)/D/1 queue. Probability in the Engineering and Informational Sciences, 25(4):519–535, October 2011. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
23E3F975D956C32BAE6F25D5C63BAF45.

Master:2018:MPN

[MCB18] Neal Master, Carri W. Chan, and Nicholas Bambos. Myopic policies for non-preemptive scheduling of jobs with decaying value. Probability in the Engineering and Informational Sciences, 32(1):1–36, January 2018. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/myopic-
policies-for-nonpreemptive-scheduling-of-jobs-with-
decaying-value/E21D53D5B7FF0C458A613BBC41EC60B5.

Ma:2022:LBA

[MCL22] Jingtang Ma, Zhenyu Cui, and Wenyuan Li. Laplace bounds approximation for American options. Probability in the Engineering and Informational Sciences, 36(2):514–547, April 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/laplace-bounds-approximation-for-american-options/
BC399AE2A11F60BA46F2219B5EE36716.

McNamara:1988:SDG

[McN88] J. M. McNamara. A stochastic differential game with safe and risky choices. Probability in the Engineering and Informational Sciences, 2(1):31–39, January 1988. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
9C962DC96018504BF148A57CC709C6D1.

McNamara:1989:PIN

[McN89] J. M. McNamara. Policy improvement and the Newton– Raphson algorithm for renewal reward processes. Probability in the Engineering and Informational Sciences, 3(3):393– 396, July 1989. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/0F391C79265DB52A771AFB59A7596294.

Maceira:2006:UMS

[MD06] M. E. P. Maceira and J. M. Damázio. Use of the PAR(p) model in the stochastic dual dynamic programming optimization scheme used in the operation planning of the Brazilian hydropower system. Probability in the Engineering and Informational Sciences, 20(1):143–156, January 2006.

CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
A872C9A753EE4BFED702A8A0C7EF729E.

Mandjes:2014:GBP

[ME14] M. Mandjes and Eurandom. Generalized birthday problems in the large-deviations regime. Probability in the Engineering and Informational Sciences, 28(1):83–99, January 2014. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
3D4C0DAA2CDCE48B9E20B8D3CBD2D19A.

Meilijson:1991:SBL

[Mei91] Isaac Meilijson. Sharp bounds on the largest of some linear combinations of random variables with given marginal distributions. Probability in the Engineering and Informational Sciences, 5(1):1–14, January 1991. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/9E0CA31BAD346F2D76173B237FC6BA5A.

Meilijson:2014:ROD

[Mei14] Isaac Meilijson. The re-opening of Dubins and Savage Casino in the era of diversification. Probability in the Engineering and Informational Sciences, 28(1):55–65, January 2014. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
48423B46F21A6354CF0091765F4944C1.

Meng:1994:CCN

[Men94a] Fan Chin Meng. Comparing criticality of nodes via minimal cut (path) sets for coherent systems. Probability in the Engineering and Informational Sciences, 8(1):79–87, January 1994. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
EF805F3093748B0AB0CE97882233B959.

Meng:1994:OAC

[Men94b] Fan Chin Meng. Optimal allocation of components in kout-of-R parallel modules systems. Probability in the Engineering and Informational Sciences, 8(3):435–441, July 1994. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
D920A48FFEFDD10FCD3A1FC867A86EAE.

Meng:2002:NTR

[Men02] Fan C. Meng. A note on two reliability lower bounds for multistate systems. Probability in the Engineering and Informational Sciences, 16(4):485–498, October 2002. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
4FF4AA10F9397922BE8EE14EE1D3052D.

Meng:2004:CBI

[Men04] Fan C. Meng. Comparing Birnbaum importance measure of system components. Probability in the Engineering and Informational Sciences, 18(2):237–245, April 2004. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
774AD90C235DC11EE342A1D90B78FA96.

Misra:2017:SSC

[MFN17] Neeraj Misra, Jisha Francis, and Sameen Naqvi. Some sufficient conditions for relative aging of life distributions. Probability in the Engineering and Informational Sciences, 31(1): 83–99, January 2017. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/05AFA3133D95900AE310F59EAE5636AA.

Ma:2021:EBN

[MG21] Yanbin Ma and Wenhao Gui. Entropy-based and non-entropy-based goodness of fit test for the inverse Rayleigh distribution with progressively type-II censored data. Probability in the Engineering and Informational Sciences, 35(3):631–649, July 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/entropybased-and-nonentropybased-goodness-of-
fit-test-for-the-inverse-rayleigh-distribution-with-
progressively-typeii-censored-data/5C2AF9F06DF4E442EED50817E1E756B1.

Misra:2021:RSP

[MGM21] Amit Kumar Misra, Vaishali Gupta, and Neeraj Misra. Residual stochastic precedence order. Probability in the Engineering and Informational Sciences, 35(3):496–512, July 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/

probability-in-the-engineering-and-informational-sciences/
article/residual-stochastic-precedence-order/AA529D63618FB32806FF21F4C01985B4.

Mao:2010:ECO

[MH10] Tiantian Mao and Taizhong Hu. Equivalent characterizations on orderings of order statistics and sample ranges. Probability in the Engineering and Informational Sciences, 24(2):245– 262, April 2010. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/02F58E68C30C8A3DF2958F5F8118CBF5.

Mao:2010:OCH

[MHZ10] Tiantian Mao, Taizhong Hu, and Peng Zhao. Ordering convolutions of heterogeneous exponential and geometric distributions revisited. Probability in the Engineering and Informational Sciences, 24(3):329–348, July 2010. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
6C2C4EDEBF449374149A25C1449BF6E6.

Mi:1999:MSI

[Mi99] Jie Mi. On measure of system interval availability. Probability in the Engineering and Informational Sciences, 13(3): 359–375, July 1999. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/3244CDF5602DCDB0EC640CEA402DA3A3.

Mi:2000:ANE

[Mi00] Jie Mi. Average number of events and average reward. Probability in the Engineering and Informational Sciences, 14(4): 485–510, October 2000. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/E9B013D8186B1F1EC1A8D0C2292D4912.

Mazalov:2010:EPG

[MI10] Vladimir Mazalov and Anna Ivashko. Equilibrium in n-person game of showcase showdown. Probability in the Engineering and Informational Sciences, 24(3):397–403, July 2010. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
587FD06DD6CC59208DEA2E1DF267379D.

Mi:2021:IIA

[Mi21] Jie Mi. On interval and instant availability of the system. Probability in the Engineering and Informational Sciences, 35(3):581–594, July 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/on-interval-
and-instant-availability-of-the-system/F2A50D8C95E5BC7FC4AEEF5E99753A67.

Manesh:2023:SOA

[MIK23] Sirous Fathi Manesh, Muhyiddin Izadi, and Baha-Eldin Khaledi. On stochastic ordering among extreme shock models. Probability in the Engineering and Informational Sciences, 37(4):961–972, October 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/on-stochastic-
ordering-among-extreme-shock-models/CDD5E91F31B913BE215443135C53EDFD.

Mitrani:2018:MCR

[Mit18] Isi Mitrani. Multi-class resource sharing with preemptive priorities. Probability in the Engineering and Informational Sciences, 32(3):323–339, July 2018. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/multiclass-
resource-sharing-with-preemptive-priorities/D6447FC958978112631C00AE40EE5021.

Miyazawa:1992:LPB

[Miy92] Masakiyo Miyazawa. Loss probability of a burst arrival finite queue with synchronized service. Probability in the Engineering and Informational Sciences, 6(2):201–216, April 1992. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
20E7F757741846754B84D23CC321C4ED.

Miyoshi:2000:AIS

[Miy00] Naoto Miyoshi. Application of IPA to the sensitivity analysis of the leaky-bucket filter with stationary gradual input. Probability in the Engineering and Informational Sciences, 14(2): 219–241, April 2000. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/F117D990B783211CC57E2A300F96E839.

Miyazawa:2002:MRA

[Miy02] Masakiyo Miyazawa. A Markov renewal approach to the asymptotic decay of the tail probabilities in risk and queuing processes. Probability in the Engineering and Informational Sciences, 16(2):139–150, April 2002. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
44971A65CA146F31E37AC350FEB93B80.

Miyoshi:2004:NBM

[Miy04] Naoto Miyoshi. A note on bounds and monotonicity of spatial stationary Cox shot noises. Probability in the Engineering and Informational Sciences, 18(4):561–571, October 2004. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/E029FD384D53DE8AB1ADEBD30DE72232. See correc-

tion [Miy05].

Miyoshi:2005:CNB

[Miy05] Naoto Miyoshi. Correction to “A Note on Bounds and Monotonicity of Spatial Stationary Cox Shot Noises”. Probability in the Engineering and Informational Sciences, 19(3):405–407, July 2005. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/F1E0046EF50022FB04D868833A30DD52. See [Miy04].

MacPhee:1995:OSM

[MJ95] I. M. MacPhee and B. P. Jordan. Optimal search for a moving target. Probability in the Engineering and Informational Sciences, 9(2):159–182, April 1995. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
5AD8698915961F14372C80EBDF77C05C.

Matalytski:2017:FER

[MK17] Mikhail Matalytski and Dmitry Kopats. Finding expected revenues in G-network with signals and customers batch removal. Probability in the Engineering and Informational Sciences, 31(4):561–575, October 2017. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
C9BA539C9176332EBE0218C22DC90640.

Matalytski:2019:ANMb

[MK19] M. Matalytski and D. Kopats. Analysis of the network with multiple classes of positive customers and signals at a nonstationary regime. Probability in the Engineering and Informational Sciences, 33(3):404–416, July 2019. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/analysis-
of-the-network-with-multiple-classes-of-positive-customers-
and-signals-at-a-nonstationary-regime/4146CACABC14500A5BE771C3AA577121.

Matalytski:2021:FNS

[MK21] Mikhail Matalytski and Dmitry Kopats. Finding nonstationary state probabilities of open Markov networks with multiple classes of customers and various features. Probability in the Engineering and Informational Sciences, 35(1):158–179, January 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/finding-
nonstationary-state-probabilities-of-open-markov-networks-
with-multiple-classes-of-customers-and-various-features/
5A6B09569D8253766B7F161AA3DADBFB.

Meziani:2023:EGA

[MK23] S. Meziani and T. Kernane. Extended generator and associated martingales for M/G/1 retrial queue with classical retrial policy and general retrial times. Probability in the Engineering and Informational Sciences, 37(1):206–213, January 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/extended-generator-and-associated-martingales-
for-mg1-retrial-queue-with-classical-retrial-policy-
and-general-retrial-times/E6B74F3B73762A2EC6D4BDDF4832ACD3.

Kil:2022:CFA

[mKK22] Hye mee Kil and Jeong-Hoon Kim. A closed-form approximation formula for pricing European options under a three-factor model. Probability in the Engineering and Informational Sciences, 36(4):1214–1240, October 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-

engineering-and-informational-sciences/article/closedform-
approximation-formula-for-pricing-european-options-
under-a-threefactor-model/57D73BC4188E8AD4E5C6E33C3B1C30FE.

Moutzoukis:2001:TQT

[ML01] Evangelos Moutzoukis and Christos Langaris. Two queues in tandem with retrial customers. Probability in the Engineering and Informational Sciences, 15(3):311–325, July 2001. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
BAC7FCA896E88354CE271F852A26B331.

Misra:2011:ARA

[MMD11] Neeraj Misra, Amit Kumar Misra, and Ishwari Dutt Dhariyal. Active redundancy allocations in series systems. Probability in the Engineering and Informational Sciences, 25(2):219–235, April 2011. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/0D48DEFB66444BCE9F8D8EC4E8E16DA0.

Misra:1993:PSA

[MMS93] Neeraj Misra, M. Manoharan, and Harshinder Singh. Preservation of some ageing properties by order statistics. Probability in the Engineering and Informational Sciences, 7(3): 437–440, July 1993. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/E3BDEBBC9233DA22B4DCE81C84053853.

Miao:2020:ABC

[MMY20] Yu Miao, Huanhuan Ma, and Qinglong Yang. Asymptotic behaviors for correlated Bernoulli model. Probability in the Engineering and Informational Sciences, 34(4):570–582, October 2020. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/asymptotic-behaviors-for-correlated-bernoulli-
model/7126715099019C9A7B16F80DE730AC0E.

Mandjes:2005:STM

[MN05] M. Mandjes and M. Nuyens. Sojourn times in the M/G/1 FB queue with light-tailed service times. Probability in the Engineering and Informational Sciences, 19(3):351–361, July 2005. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
E6850301484C5C601ECA804238217AB2.

Mao:2015:AEG

[MNH15] Tiantian Mao, Kai Wang Ng, and Taizhong Hu. Asymptotic expansions of generalized quantiles and expectiles for extreme risks. Probability in the Engineering and Informational Sciences, 29(3):309–327, July 2015. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
E57E8AD75EC4692A4E58F985DB8D6365.

Mahmoud:1995:TFD

[MO95] Hosam M. Mahmoud and Ratko Orlandić. Toward a formal derivation of the expected behavior of prefix B-trees. Probability in the Engineering and Informational Sciences, 9(2): 183–192, April 1995. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/8732942478D93CE4399F8491D04A0FAD.

Monhor:2007:CIM

[Mon07] Davaadorjin Monhor. A Chebyshev inequality for multivariate normal distribution. Probability in the Engineering and Informational Sciences, 21(2):289–300, April 2007. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
2155D1A7E508E65290DA02710099B4EA.

Monhor:2013:ICB

[Mon13] Davaadorjin Monhor. Inequalities for correlated bivariate normal distribution function. Probability in the Engineering and Informational Sciences, 27(1):115–123, January 2013. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
9B1287EBF7A25359CD4BBD0773EF3AAD.

Miyoshi:2010:SIG

[MOSU10] Naoto Miyoshi, Mariko Ogura, Takeya Shigezumi, and Ryuhei Uehara. Subexponential interval graphs generated by immigration–death processes. Probability in the Engineering and Informational Sciences, 24(2):289–301, April 2010. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
740862BB7B2B9CD38733EBE52E9ED171.

McCormick:1997:APM

[MP97] William P. McCormick and YouSung Park. An analysis of Poisson moving-average processes. Probability in the Engineering and Informational Sciences, 11(4):487–507, October 1997. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
84B03D0726B324ED71B5806AEFB0ABAC.

Mamoun:2002:CFS

[MP02] Mouad Ben Mamoun and Nihal Pekergin. Closed-form stochastic bounds on the stationary distribution of Markov chains. Probability in the Engineering and Informational Sciences, 16(4):403–426, October 2002. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/A23086BA3EA44563CA7AA9CCE0154862.

Mao:2013:OBW

[MPH13] Tiantian Mao, Xiaoqing Pan, and Taizhong Hu. On orderings between weighted sums of random variables. Probability in the Engineering and Informational Sciences, 27(1):85–97, January 2013. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
9DE1738F4273033C05B6BBB0A98BB700.

Muliere:1993:NRE

[MPP93] Pietro Muliere, Giovanni Parmigiani, and Nicholas G. Polson. A note on the residual entropy function. Probability in the Engineering and Informational Sciences, 7(3):413–420, July 1993. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
FA9C6BE0DA7C2F9E2FA8EC26A4982249.

Mallik:1988:SMD

[MR88] Ashim K. Mallik and Jaxk H. Reeves. Sequential methods to detect material losses. Probability in the Engineering and Informational Sciences, 2(4):501–515, October 1988. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
B1DDC69EAD46F732B38EFBFD9A26E477.

Mandjes:1995:FCM

[MR95] Michel Mandjes and Ad Ridder. Finding the conjugate of Markov fluid processes. Probability in the Engineering and Informational Sciences, 9(2):297–315, April 1995. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
77376444D0297660369CB0F4A887DB0A.

Miyazawa:1991:MLP

[MS91] Masakiyo Miyazawa and J. George Shanthikumar. Monotonicity of the loss probabilities of single server finite queues with respect to convex order of arrival or service processes. Probability in the Engineering and Informational Sciences, 5(1): 43–52, January 1991. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/63786FD11B6DEEA1421DE803A85F8C53.

Meester:1993:RSP

[MS93] Ludolf E. Meester and J. George Shanthikumar. Regularity of stochastic processes: A theory based on directional convexity. Probability in the Engineering and Informational Sciences, 7 (3):343–360, July 1993. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/AE90F83AB466EEB14F415C4DCAD4F738.

Meliopoulos:2006:CSE

[MS06] A. P. Sakis Meliopoulos and George K. Stefopoulos. Characterization of state estimation biases. Probability in the Engineering and Informational Sciences, 20(1):157–174, January 2006. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
98D8064A35172B7F54F012D8BB8C10FF.

Miyazawa:1993:CTA

[MT93] Masakiyo Miyazawa and Henk Tijms. Comparison of two approximations for the loss probability in finite-buffer queues. Probability in the Engineering and Informational Sciences, 7(1):19–27, January 1993. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/BD99A58E73F316560A52825EED7484BD.

Mandjes:2016:RML

[MT16] Michel Mandjes and Peter Taylor. The running maximum of a level-dependent quasi-birth-death process. Probability in the Engineering and Informational Sciences, 30(2):212–223, April 2016. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
1E6C9674828E11C0A92D786EEED3F981.

Meszaros:2022:CME

[MT22] András Mészáros and Miklós Telek. Concentrated matrix exponential distributions with real eigenvalues. Probability in the Engineering and Informational Sciences, 36 (4):1171–1187, October 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/concentrated-
matrix-exponential-distributions-with-real-eigenvalues/
2A3A550C0DE41A7AC45801680062ADE1.

Muller:1998:CK

[Mül98] Alfred Müller. On a conjecture of Koole. Probability in the Engineering and Informational Sciences, 12(1):141–142, January 1998. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
56EDF14B8B3F68D9375B482FDEFF5206.

Miyazawa:1997:NBE

[MvD97] Masakiyo Miyazawa and Nico M. van Dijk. A note on bounds and error bounds for nonexponential batch arrival systems. Probability in the Engineering and Informational Sciences, 11 (2):189–201, April 1997. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/FB39DA05C605D6054EF4BF72F3E09F3A.

Massey:1994:SMC

[MW94] William A. Massey and Ward Whitt. A stochastic model to capture space and time dynamics in wireless communication systems. Probability in the Engineering and Informational Sciences, 8(4):541–569, October 1994. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
94257151E6AE8AADC377E7696EC3CC70.

Mao:2018:PLC

[MXH18] Tiantian Mao, Wanwan Xia, and Taizhong Hu. Preservation of log-concavity under convolution. Probability in the Engineering and Informational Sciences, 32(4):567–579, October 2018. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/preservation-of-logconcavity-under-convolution/
8CE16E5EE5885E8BE7950B3AD9487988.

Ma:2011:EAE

[MY11] Hongli Ma and Weiguo Yang. Erratum to “The Asymptotic Equipartition Property for Asymptotic Circular Markov Chains” [Probability in the Engineering and Informational Sciences, 24, 2010, 279–288]. Probability in the Engineering and Informational Sciences, 25(2):265–267, April 2011. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
D8D3C26E5D778A10D87D03711968419E. See [ZYL10].

Mao:2017:HTD

[MZ17] Yong-Hua Mao and Chi Zhang. Hitting time distributions for birth-death processes with bilateral absorbing boundaries. Probability in the Engineering and Informational Sciences, 31 (3):345–356, July 2017. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/1C4A7E153F8648B62D385C2A1A758D7D.

Meng:2016:ODR

[MZS16] Hui Meng, Ming Zhou, and Tak Kuen Siu. Optimal dividendreinsurance with two types of premium principles. Probability in the Engineering and Informational Sciences, 30(2):224–243, April 2016. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/785BFA738B39EC73AEB19DAF1517FF2A.

Nadarajah:2008:NMS

[Nad08] Saralees Nadarajah. A new model for symmetric and skewed data. Probability in the Engineering and Informational Sciences, 22(2):261–271, March 2008. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
A6C7177C0B5946F999985CBC473AF57B.

Nadarajah:2016:WGG

[Nad16] Saralees Nadarajah. The weighted generalized gamma distribution is the generalized gamma distribution. Probability in the Engineering and Informational Sciences, 30(2):298–299, April 2016. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
AC3F1A7F0F9BDEB1F5081C384B489CD1.

Nadarajah:2017:DMP

[Nad17] Saralees Nadarajah. On the distribution of maximal percentages in Pólya’s urn. Probability in the Engineering and Informational Sciences, 31(3):357–365, July 2017. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
7A981B5F56531C9FEEFA7758F98ED46C.

Nakade:2001:EUB

[Nak01] Koichi Nakade. Effective upper bounds for expected cycle times in tandem queues with communication blocking. Probability in the Engineering and Informational Sciences, 15(3): 327–334, July 2001. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/3276DB89A3C767FD51D08525593B9869.

Nayak:1989:MVU

[Nay89] Tapan K. Nayak. Minimum variance unbiased estimation of software reliability. Probability in the Engineering and Informational Sciences, 3(3):335–340, July 1989. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
D82289FDC4717A8FFC5979075D7103E9.

Navarro:1997:NSO

[NBR97] Jorge Navarro, Felix Belzunce, and Jose M. Ruiz. New stochastic orders based on double truncation. Probability in the Engineering and Informational Sciences, 11(3):395–402, July 1997. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
1F7A8F02996F37824291EF26A0AB9648.

Noughabi:2013:RPS

[NBR13] M. Shafaei Noughabi, G. R. Mohtashami Borzadaran, and A. H. Rezaei Roknabadi. On the reliability properties of some weighted models of bathtub shaped hazard rate distributions. Probability in the Engineering and Informational Sciences, 27(1):125–140, January 2013. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
8C02E4E8D63627B2EF8B0650191B114E.

Nadarajah:2018:ACA

[NCJ18] Saralees Nadarajah, Jeffrey Chu, and Xiao Jiang. Aggregation and capital allocation formulas for bivariate distributions. Probability in the Engineering and Informational Sciences, 32(4):556–566, October 2018. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/aggregation-
and-capital-allocation-formulas-for-bivariate-distributions/
5A03548F2CAEE18118713904D608B6B3.

Nanda:2013:DPM

[NDB13] Asok K. Nanda, Suchismita Das, and N. Balakrishnan. On dynamic proportional mean residual life model. Probability in the Engineering and Informational Sciences, 27(4):553–588, October 2013. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/8C3B2607A578263B41779CCDDB12DE0E.

Naqvi:2022:SCP

[NDZ22] Sameen Naqvi, Weiyong Ding, and Peng Zhao. Stochastic comparison of parallel systems with Pareto components. Probability in the Engineering and Informational Sciences, 36(4):950–962, October 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/stochastic-
comparison-of-parallel-systems-with-pareto-components/
E14A9328F273D1408739C7E68F4AB07B.

Navarro:2020:RBI

[NFMFSA20] Jorge Navarro, Pedro Fernández-Martínez, Juan Fernández-Sánchez, and Antonio Arriaza. Relationships between importance measures and redundancy in systems with dependent components. Probability in the Engineering and Informational Sciences, 34(4):583–604, October 2020. CODEN ????

ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/relationships-
between-importance-measures-and-redundancy-in-systems-
with-dependent-components/8A62458E863D79D4BBB5C5CF6DE71C81.

Nguyen:1991:OTR

[Ngu91] Viên Nguyen. On the optimality of trunk reservation in overflow processes. Probability in the Engineering and Informational Sciences, 5(3):369–390, July 1991. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
6EAD0DD5E13BB2D0D33B62294F2D3CB4.

Navarro:2004:HOB

[NH04] Jorge Navarro and Pedro J. Hernandez. How to obtain bathtub-shaped failure rate models from normal mixtures. Probability in the Engineering and Informational Sciences, 18(4):511–531, October 2004. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/BEE453F54AA62EF83BEA4C0BA78DE63A.

Nica:2016:OSG

[Nic16] Mihai Nica. Optimal strategy in “guess who?”: beyond binary search. Probability in the Engineering and Informational Sciences, 30(4):576–592, October 2016. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
8B951CD112BBE7EC63959049C67DD22A.

Nishimura:1995:MGV

[NJ95] Shoichi Nishimura and Yong Jiang. An m/g/l vacation model with two service modes. Probability in the Engineering and Informational Sciences, 9(3):355–374, July 1995. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
AA4674BB8DD47D1652695903B40937CE.

Nadarajah:2005:LCL

[NK05] Saralees Nadarajah and Samuel Kotz. On the linear combination of Laplace random variables. Probability in the Engineering and Informational Sciences, 19(4):463–470, October 2005. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
1C7D0678875E3AE3A30664221FDB93B6.

Nadarajah:2007:BBB

[NK07] Saralees Nadarajah and Samuel Kotz. Block and Basu’s bivariate exponential distribution with application to drought data. Probability in the Engineering and Informational Sciences, 21(1):143–155, January 2007. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/095B6C3462595E08E14B5230157A055D.

Natarajan:2007:MVB

[NL07] Karthik Natarajan and Zhou Linyi. A mean-variance bound for a three-piece linear function. Probability in the Engineering and Informational Sciences, 21(4):611–621, October 2007. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
A199F4E5FB00FF3F1658596A295981F1.

Nahmias:2004:AVP

[NPS04] Steven Nahmias, David Perry, and Wolfgang Stadje. Actuarial valuation of perishable inventory systems. Probability in the Engineering and Informational Sciences, 18(2):219–232, April 2004. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
B1A367A1A069A3B34BC7A3C5C5DB9579.

Nunez-Queija:2001:NGG

[NQ01] Rudesindo Núñez-Queija. Note on the GI/GI/1 queue with LCFS-PR observed at arbitrary times. Probability in the Engineering and Informational Sciences, 15(2):179–187, April 2001. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
3E1A3D00B3D5368F83AD7AE28291779C.

Nakayama:2004:QSM

[NS04] Marvin K. Nakayama and Perwez Shahabuddin. Quick simulation methods for estimating the unreliability of regenerative models of large, highly reliable systems. Probability in the Engineering and Informational Sciences, 18(3):339–368, July 2004. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
E6BC8F8136909803E140F8C4A96173CB.

Naumov:2017:AMR

[NS17] Valeriy Naumov and Konstantin Samouylov. Analysis of multi-resource loss system with state-dependent arrival and service rates. Probability in the Engineering and Informational Sciences, 31(4):413–419, October 2017. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
8632F6C93621F2735D977CCC6428C899.

Naumov:2021:PFM

[NS21] Valeriy Naumov and Konstantin Samouylov. Productform Markovian queueing systems with multiple resources. Probability in the Engineering and Informational Sciences, 35(1):180–188, January 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/productform-
markovian-queueing-systems-with-multiple-resources/
6674EC65E1F4D1E6E4477361BC9B1A10.

Navarro:2022:CRS

[NS22] Jorge Navarro and José María Sarabia. Copula representations for the sum of dependent risks: models and comparisons. Probability in the Engineering and Informational Sciences, 36(2):320–340, April 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/copula-
representations-for-the-sum-of-dependent-risks-models-
and-comparisons/9E364F59A54ED607C5E3A6A3CC4B2A03.

Nadarajah:2013:EMC

[NWB13] S. Nadarajah, C. S. Withers, and S. A. A. Bakar. Expansions for moments of compound Poisson distributions. Probability in the Engineering and Informational Sciences, 27(3): 319–331, July 2013. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/075DA607F193BCA27957A09B3DA2E749.

Ng:2004:RPU

[NYZ04] Kai W. Ng, Hailiang Yang, and Lihong Zhang. Ruin probability under compound Poisson models with random discount factor. Probability in the Engineering and Informational Sciences, 18(1):55–70, January 2004. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
762148BC4F2562A3D7DE2F540D71BCBF.

Nadarajah:2018:SIG

[NZP18] Saralees Nadarajah, Yuanyuan Zhang, and Tibor K. Pogány. On sums of independent generalized Pareto random variables with applications to insurance and cat bonds. Probability in the Engineering and Informational Sciences, 32(2):296–305, April 2018. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/on-sums-of-independent-generalized-pareto-random-
variables-with-applications-to-insurance-and-cat-bonds/
F5BD100B02E7453F8F0F94919DFC1BF9.

Ozekici:2001:GST

[ÖAA01] Süleyman Özekici, I. Kuban Altinel, and Ebru Angün. A general software testing model involving operational profiles. Probability in the Engineering and Informational Sciences, 15(4): 519–533, October 2001. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/440F7A7AF4E9D9606B32E8D5584F49EA.

Ortega:2010:VCB

[OE10] Eva María Ortega and Laureano F. Escudero. Variability for carrier-borne epidemics and Reed–Frost models incorporating uncertainties and dependencies from susceptibles and infectives. Probability in the Engineering and Informational Sciences, 24(2):303–328, April 2010. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
AF091CEF10CE209C4EB47572989D4E09.

Olkin:1995:ENA

[OGP95] Ingram Olkin, Irwin Guttman, and Robert Philips. Estimating the number of aberrant laboratories. Probability in the Engineering and Informational Sciences, 9(1):133–150, January 1995. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
D1F2A1DF2FAFA33AC9590FF4E995BA8D.

Ott:2008:TBP

[OK08] T. J. Ott and J. H. B. Kemperman. Transient behavior of processes in the TCP paradigm. Probability in the Engineering and Informational Sciences, 22(3):431–471, July 2008. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
CC000E05D7DF3D21E08474C1055AA8BD.

Ozkan:2014:OCT

[ÖK14] Erhun Özkan and Jeffrey P. Kharoufeh. Optimal control of a two-server queueing system with failures. Probability in the Engineering and Informational Sciences, 28(4):489–527, October 2014. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
842FE1E1EBC278FB94F49B6BE63C82D8.

Oh:2022:DBM

[OKA22] Rosy Oh, Joseph H. T. Kim, and Jae Youn Ahn. Designing a Bonus–Malus system reflecting the claim size under the dependent frequency-severity model. Probability in the Engineering and Informational Sciences, 36(4):963–987, October 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/designing-a-bonusmalus-system-reflecting-the-
claim-size-under-the-dependent-frequencyseverity-model/
281A56F8B98DE8A8565B4E88EBEC2654.

Ozekici:1995:WTM

[ÒLC95] Süleyman Òzekici, Jingwen Li, and Fee Seng Chou. Waiting time in M/G/1 queues with impolite arrival disciplines. Probability in the Engineering and Informational Sciences, 9(2): 255–267, April 1995. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/B93F03420938ED77C5407421095C9EA4.

Ou:1997:ACD

[OLÖ97] Jihong Ou, Jingwen Li, and Süleyman Özekici. Approximating a cumulative distribution function by generalized hyperexponential distributions. Probability in the Engineering and Informational Sciences, 11(1):11–18, January 1997. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
81007DEF76CF2576E8D9900B2343DCE9.

Olsen:2001:LTP

[Ols01] Tava Lennon Olsen. Limit theorems for polling models with increasing setups. Probability in the Engineering and Informational Sciences, 15(1):35–55, January 2001. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
4D279BD9990869B50CDF321F16B7283E.

Oluyede:1999:ISE

[Olu99] Broderick O. Oluyede. On inequalities and selection of experiments for length biased distributions. Probability in the Engineering and Informational Sciences, 13(2):169–185, April 1999. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
8ED4D11BAE702C47BBF059C2C59FD855.

Oluyede:2006:BAW

[Olu06] Broderick O. Oluyede. On bounds and approximating weighted distributions by exponential distributions. Probability in the Engineering and Informational Sciences, 20(3):517–528, July 2006. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
B62612010C3D4749527864B74BEE915D.

Oluyede:2011:EDU

[Olu11] Broderick O. Oluyede. Exponential dominance and uncertainty for weighted residual life measures. Probability in the Engineering and Informational Sciences, 25(2):251–263, April 2011. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
A9334A1651AB53C2996C187A5A49B05B.

Ott:1990:MMP

[OS90] Teunis J. Ott and J. George Shanthikumar. On maxima and minima of partial sums of strongly interchangeable random variables. Probability in the Engineering and Informational Sciences, 4(3):319–332, July 1990. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
8CAB16DD4D9057DFA84BA4C809321EE4.

Olkin:2006:SCM

[OST06] Ingram Olkin, Thomas Santner, and Y. L. Tong. The scientific contributions of Milton Sobel. Probability in the Engineering and Informational Sciences, 20(3):383–411, July 2006. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
FDB786BDCECF578AAEEC1AC75576CE7C.

Ott:1993:QSB

[Ott93a] Teunis J. Ott. A queuing system with a branching process arrival stream, the case of short-range dependence. Probability in the Engineering and Informational Sciences, 7(4):515–528, October 1993. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/46B2E33673DFD2CCECA935AD47F3EAE5.

Ott:1993:VPA

[Ott93b] Teunis J. Ott. A versatile packet arrival process and its second order properties. Probability in the Engineering and Informational Sciences, 7(4):495–513, October 1993. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
87BC14AB6FD21AE7D34EB625083B1353.

Palacios:1990:RAA

[Pal90] José Luis Palacios. On a result of aleliunas et al. concerning random walks on graphs. Probability in the Engineering and Informational Sciences, 4(4):489–492, October 1990. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
0B388232FE9296F20CF15777FBB1CE15.

Pandelis:2008:OCF

[Pan08] Dimitrios G. Pandelis. Optimal control of flexible servers in two tandem queues with operating costs. Probability in the Engineering and Informational Sciences, 22(1):107–131, January 2008. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
1DE622DB9E172F93B5B5CF714AB6DDF4.

Papadatos:2021:OMI

[Pap21] Nickos Papadatos. Optimal moment inequalities for order statistics from nonnegative random variables. Probability in the Engineering and Informational Sciences, 35(2):316–330, April 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/optimal-moment-inequalities-for-order-statistics-
from-nonnegative-random-variables/51F27D573C9FA284EDB66F2B520605A3.

Pekoz:2006:HTL

[PB06] Erol A. Peköz and Jose Blanchet. Heavy traffic limits via Brownian embeddings. Probability in the Engineering and Informational Sciences, 20(4):595–598, October 2006. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
534F4657BA16B4C270CEDF2EAF7723E2.

Pang:2014:CPI

[PC14] Zhan Pang and Frank Y. Chen. Coordinated pricing and inventory control with batch production and Erlang leadtimes. Probability in the Engineering and Informational Sciences, 28(4): 529–563, October 2014. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/D7E29AC5EEDEF436C0F28CF20712639A.

Peng:2022:DIS

[PC22] Xingchun Peng and Fenge Chen. Deterministic investment strategy in a DC pension plan with inflation risk under meanvariance criterion. Probability in the Engineering and Informational Sciences, 36(1):201–216, January 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/deterministic-
investment-strategy-in-a-dc-pension-plan-with-inflation-
risk-under-meanvariance-criterion/FDF64ACB8510671B19B833A82087E4F2.

Peeters:2022:RLB

[PdB22] Yannik Peeters and Arnoud V. den Boer. A regret lower bound for assortment optimization under the capacitated MNL model with arbitrary revenue parameters. Probability in the Engineering and Informational Sciences, 36 (4):1266–1274, October 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/regret-

lower-bound-for-assortment-optimization-under-the-capacitated-
mnl-model-with-arbitrary-revenue-parameters/AFB64D98533801B1018BA35C0A88519F.

Pekoz:1999:MUF

[Pek99a] Erol A. Peköz. More on using forced idle time to improve performance in polling models. Probability in the Engineering and Informational Sciences, 13(4):489–496, October 1999. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
379A80E8ED27DC3E1E84DF9FF9A4A03F.

Pekoz:1999:NRB

[Pek99b] Erol A. Peköz. On the number of refusals in a busy period. Probability in the Engineering and Informational Sciences, 13(1):71–74, January 1999. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/B0BC03F2EFBB6532DCEEF0BCFF37378C.

Pekergin:2016:FEG

[Pek16] Nihal Pekergin. Festschrift for Erol Gelenbe’s 70th birthday. Probability in the Engineering and Informational Sciences, 30 (3):301–307, July 2016. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/6520D64D168ED4B8AFF7EA6F53462172.

Pekergin:2021:ISI

[Pek21] Nihal Pekergin. Introduction to the special issue on learning, optimization, and theory of g-networks. Probability in the Engineering and Informational Sciences, 35(1):1–5, January 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/introduction-to-the-special-issue-on-learning-
optimization-and-theory-of-gnetworks/CCAA5B0F058EDE0C9942EF25CD4C87A7.

Pendergrass:2010:PGG

[Pen10] Marcus Pendergrass. A path guessing game with wagering. Probability in the Engineering and Informational Sciences, 24 (3):375–396, July 2010. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/6A52FB886D369B7D958A4C4FCDBB5C77.

Pender:2015:NLQ

[Pen15] Jamol Pender. Nonstationary loss queues via cumulant moment approximations. Probability in the Engineering and Informational Sciences, 29(1):27–49, January 2015. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
67950BCC74A6B96711EEED3EB0CD9FF5.

Perry:1997:DBC

[Per97] David Perry. A double band control policy of a Brownian perishable inventory system. Probability in the Engineering and Informational Sciences, 11(3):361–373, July 1997. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
E8A800AC12403DD4EF300B1107F86722.

Pasricha:2022:CFPa

[PG22] Puneet Pasricha and Anubha Goel. A closed-form pricing formula for European exchange options with stochastic volatility. Probability in the Engineering and Informational Sciences, 36(3):606–615, July 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/closedform-
pricing-formula-for-european-exchange-options-with-
stochastic-volatility/929BBD2E9E1F125FF540A843B7D1AFCC.

Pasricha:2022:CFPb

[PGZ22] Puneet Pasricha, Anubha Goel, and Song-Ping Zhu. A closed-form pricing formula for catastrophe equity options. Probability in the Engineering and Informational Sciences, 36(4):1103–1115, October 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/closedform-
pricing-formula-for-catastrophe-equity-options/535EE5AE87EECFFE86C8C9FCA7321758.

Pasricha:2023:SEO

[PH23] Puneet Pasricha and Xin-Jiang He. A simple European option pricing formula with a skew Brownian motion. Probability in the Engineering and Informational Sciences, 37 (4):1029–1034, October 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/simple-
european-option-pricing-formula-with-a-skew-brownian-
motion/09FF2B2A9F43AA64759305CF356B4FE8.

Pinedo:2007:SBS

[Pin07] Michael Pinedo. Stochastic batch scheduling and the “smallest variance first” rule. Probability in the Engineering and Informational Sciences, 21(4):579–595, October 2007. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
9E6737842613BBEFF1A6515872897FD0.

Politis:2006:SNB

[PK06] Konstadinos Politis and Markos V. Koutras. Some new bounds for the renewal function. Probability in the Engineering and Informational Sciences, 20(2):231–250, April 2006. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
8AA2D47DAF7091B536A5FAF88EC6A797.

Pandelis:2013:SVR

[PKK13] Dimitrios G. Pandelis, Constantinos C. Karamatsoukis, and Epaminondas G. Kyriakidis. Single vehicle routing problems with a predefined customer order, unified load and stochastic discrete demands. Probability in the Engineering and Informational Sciences, 27(1):1–23, January 2013. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
B0BDC66C1F9A33021BF0885179F95100.

Park:2020:CMP

[PL20] Jong Jun Park and Kyungsub Lee. Computational method for probability distribution on recursive relationships in financial applications. Probability in the Engineering and Informational Sciences, 34(2):258–278, April 2020. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/computational-
method-for-probability-distribution-on-recursive-relationships-
in-financial-applications/53231C7247F4A309455502EC22F7E4A4.

Platen:1992:HOW

[Pla92] Eckhaard Platen. Higher-order weak approximation of Ito diffusions by Markov chains. Probability in the Engineering and Informational Sciences, 6(3):391–408, July 1992. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
C1B20985468E08865CA8C3640D3C14A4.

Preethi:2023:LSO

[PLS23] K. Preethi, A. Shophia Lawrence, and B. Sivakumar. Lost sales obsolescence inventory systems with positive lead time: a system-point level-crossing approach. Probability in the Engineering and Informational Sciences, 37(3):778–800, July 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/lost-sales-obsolescence-inventory-systems-with-
positive-lead-time-a-systempoint-levelcrossing-approach/
87D757B3A04CD81643F47C2E160A3B1D.

Pender:2017:ASD

[PM17] Jamol Pender and William A. Massey. Approximating and stabilizing dynamic rate Jackson networks with abandonment. Probability in the Engineering and Informational Sciences, 31 (1):1–42, January 2017. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/F6DBC2E39E7D3D8AC801A3863B927543.

Prabhakar:1996:CDT

[PMB96] B. Prabhakar, T. S. Mountford, and N. Bambos. Convergence of departures in tandem networks of · /GI/∞ queues. Probability in the Engineering and Informational Sciences, 10(4): 487–500, October 1996. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/9ED2ADB0721E050A1AEC542560E4F0EB.

Priyadarshani:2015:TPW

[PO15] Hewa A. Priyadarshani and Broderick O. Oluyede. Theoretical properties of the weighted generalized gamma and related distributions. Probability in the Engineering and Informational Sciences, 29(3):421–432, July 2015. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
72808B2CAC01D21719FB16A76FD39DA2.

Pascual-Ortigosa:2021:ARM

[POdCW21] Patricia Pascual-Ortigosa, Eduardo Sáenz de Cabezón, and Henry P. Wynn. Algebraic reliability of multi-state k-out-of-n systems. Probability in the Engineering and Informational Sciences, 35(4):903–927, October 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/algebraic-
reliability-of-multistate-koutofn-systems/D90BD22B886AD38E0536B29B3A47219E.

See corrigendum [Ano21a].

Perry:1989:CPT

[PP89] D. Perry and M. J. M. Posner. Control policies for two classes of inventory systems via a duality-equivalence relationship. Probability in the Engineering and Informational Sciences, 3(4): 561–579, October 1989. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/F581F1B5DA8FA0F0C47164E79035DAD5.

Perry:2002:PIM

[PP02] David Perry and M. J. M. Posner. Production-inventory models with an unreliable facility operating in a two-state random environment. Probability in the Engineering and Informational Sciences, 16(3):325–338, July 2002. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
8357E4046B1559C8D3EE94A32644B2E4.

Pinedo:1988:NSS

[PR88] Michael Pinedo and Elias Rammouz. A note on stochastic scheduling on a single machine subject to breakdown and repair. Probability in the Engineering and Informational Sciences, 2(1):41–49, January 1988. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
7F8973D00CA6F00DBDEBBAD7FF3962D5.

Pekoz:1994:IPA

[PR94] Erol A. Peköz and Sheldon M. Ross. Improving Poisson approximations. Probability in the Engineering and Informational Sciences, 8(4):449–462, October 1994. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
637263CF7EA362DAFAA9CF57923CB9AA.

Pekoz:1997:EMC

[PR97] Erol Peköz and Sheldon M. Ross. Estimating the mean cover time of a semi-Markov process via simulation. Probability in the Engineering and Informational Sciences, 11(2):267–271, April 1997. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
197B81E69F570E20A4F224365AC229B7.

Pekoz:2004:CRV

[PR04] Erol Peköz and Sheldon M. Ross. Compound random variables. Probability in the Engineering and Informational Sciences, 18(4):473–484, October 2004. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/2585C3B5DEC3027F67E0394433DB970D.

Pacheco:2008:MDB

[PR08] António Pacheco and H. Ribeiro. Moments of the duration of busy periods of M^{x} /g/1/n systems. Probability in the Engineering and Informational Sciences, 22(3):347–354, July 2008. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
21EC97581F7541C57FAE1C1F00FCD9EB.

Preater:1994:MSP

[Pre94] J. Preater. A multiple stopping problem. Probability in the Engineering and Informational Sciences, 8(2):169–177, April 1994. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
B6E91F5E50F310AB217360E7BB344D7B.

Preater:2002:EFM

[Pre02] J. Preater. An emergency facility with minimal admission. Probability in the Engineering and Informational Sciences, 16(4):499–511, October 2002. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/0D4FA09F02834A251F70F3FB8075C846.

Pekoz:2022:DBP

[PRZ22] Erol Peköz, Sheldon M. Ross, and Zhengyu Zhang. Dueling bandit problems. Probability in the Engineering and Informational Sciences, 36(2):264–275, April 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/dueling-
bandit-problems/73C2EE7B3A324E8C9F0F097148B8D654.

Pellerey:1993:SCS

[PS93] Franco Pellerey and Moshe Shaked. Stochastic comparison of some wear processes. Probability in the Engineering and Informational Sciences, 7(3):421–435, July 1993. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
A2AA5CAD4308C899BE52811CB39DE6F3.

Perry:2002:EDJ

[PS02] David Perry and Wolfgang Stadje. Exact distributions in a jump-diffusion storage model. Probability in the Engineering and Informational Sciences, 16(1):19–27, January 2002. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
B87A3C51709ADDB374E504C9A5CC726D.

Pozdnyakov:2010:BBD

[PS10] Vladimir Pozdnyakov and J. Michael Steele. Bugs on a budget: distributed sensing with cost for reporting and nonreporting. Probability in the Engineering and Informational Sciences, 24(4):525–534, October 2010. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
F51D061CE6DF18D1764BE8A7E77FCE07.

Pinedo:1992:OPS

[PSC92] Michael Pinedo, Dequan Shaw, and Xiuli Chao. On optimal permutation scheduling in stochastic proportionate flowshops. Probability in the Engineering and Informational Sciences, 6(4):513–523, October 1992. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
485900BFC67E0B765FCDC9A7B6F96BC2.

Pellerey:2000:NPP

[PSZ00] Franco Pellerey, Moshe Shaked, and Joel Zinn. Nonhomogeneous Poisson processes and logconcavity. Probability in the Engineering and Informational Sciences, 14(3):353–373, July 2000. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
69CC6A8807A519F241EFCECD04ED3BB4.

Pandelis:1993:SSP

[PT93a] Dimitrios G. Pandelis and Demosthenis Teneketzis. Stochastic scheduling in priority queues with strict deadlines. Probability in the Engineering and Informational Sciences, 7(2): 273–289, April 1993. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/AF4D7305A4C389E1E2CEF50ED8C5705A.

Pollett:1993:PEE

[PT93b] P. K. Pollett and P. G. Taylor. On the problem of establishing the existence of stationary distributions for continuoustime Markov chains. Probability in the Engineering and Informational Sciences, 7(4):529–543, October 1993. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
2021963A5946F2B6C727E3437A5526CB.

Pandelis:1994:OSD

[PT94] Dimitrios G. Pandelis and Demosthenis Teneketzis. Optimal stochastic dynamic scheduling in multi-class queues with tardiness and/or earliness penalties. Probability in the Engineering and Informational Sciences, 8(4):491–509, October 1994. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
20F4011BEB26384FB1968293CFC66F98.

Pandelis:1996:OSM

[PT96] Dimitrios G. Pandelis and Demosthenis Teneketzis. Optimal sequencing in multiserver systems. Probability in the Engineering and Informational Sciences, 10(3):377–396, July 1996. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
15E124EB24D20130FF8B08E073286C55.

Psarrakos:2011:RMT

[PT11] Georgios Psarrakos and Michael Tsatsomeros. Ratio monotonicity for tail probabilities in the renewal risk model. Probability in the Engineering and Informational Sciences, 25(2): 171–185, April 2011. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/481C64887A58B3DF41A20219B0F5983C.

Psarrakos:2020:EEI

[PT20] Georgios Psarrakos and Abdolsaeed Toomaj. On the elasticity of expected interepoch intervals in a non-homogeneous Poisson process under small variations of hazard rate. Probability in the Engineering and Informational Sciences, 34(4):550–569, October 2020. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/on-the-
elasticity-of-expected-interepoch-intervals-in-a-nonhomogeneous-
poisson-process-under-small-variations-of-hazard-rate/
BFAE1D1BCCC76BDC181668F7359F7717.

Perez:2011:MGT

[PV11] Juan F. Pérez and Benny Van Houdt. The M/G/1-type Markov chain with restricted transitions and its application to queues with batch arrivals. Probability in the Engineering and Informational Sciences, 25(4):487–517, October 2011. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
F4A45829C200EF53BD97AAC8FA558B8D.

Puhalskii:1998:FLD

[PW98] Anatolii A. Puhalskii and Ward Whitt. Functional large deviation principles for waiting and departure processes. Probability in the Engineering and Informational Sciences, 12(4):479–507, October 1998. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/B34B7A1E3E9E85945DD81B758A5F111B.

Pang:2012:ISQa

[PW12a] Guodong Pang and Ward Whitt. Infinite-server queues with batch arrivals and dependent service times. Probability in the Engineering and Informational Sciences, 26(2):197–220, April 2012. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
1C6673F764AA58C0D09B9DEDD9C777F5.

Pang:2012:ISQb

[PW12b] Guodong Pang and Ward Whitt. Infinite-server queues with batch arrivals and dependent service times-corrigendum. Probability in the Engineering and Informational Sciences, 26(3): 455, July 2012. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/7373EA98857036E9B043E522582D3B32.

Perel:2014:IQI

[PY14] Nir Perel and Uri Yechiali. The Israeli queue with infinite number of groups. Probability in the Engineering and Informational Sciences, 28(1):1–19, January 2014. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
1D1317AB03A9A65492EF16E295100EA9.

Perel:2016:FTL

[PY16] Efrat Perel and Uri Yechiali. Finite two layered queueing systems. Probability in the Engineering and Informational Sciences, 30(3):492–513, July 2016. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
87382A18C3629335331913FD61A605A1.

Pinotsi:2004:SCG

[PZ04] Dimitra Pinotsi and Michael A. Zazanis. Stability conditions for gated M/G/∞ queues. Probability in the Engineering and Informational Sciences, 18(1):103–110, January 2004. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
058111F30C3B6A712FB85EE7B609D667.

Pirjol:2019:SMA

[PZ19] Dan Pirjol and Lingjiong Zhu. Short maturity Asian options for the CEV model. Probability in the Engineering and Informational Sciences, 33(2):258–290, April 2019. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/short-
maturity-asian-options-for-the-cev-model/3BFF756DC98C002EB3F4A18E3C5056A7.

Pirjol:2021:ATD

[PZ21] Dan Pirjol and Lingjiong Zhu. Asymptotics of the timediscretized log-normal SABR model: the implied volatility surface. Probability in the Engineering and Informational Sciences, 35(4):942–974, October 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/asymptotics-
of-the-timediscretized-lognormal-sabr-model-the-implied-
volatility-surface/BA4705506BFAB99D27E5A4B1AE09BBA2.

Qiu:2019:FRQ

[Qiu19] Guoxin Qiu. Further results on quantile entropy in the past lifetime. Probability in the Engineering and Informational Sciences, 33(1):146–159, January 2019. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/further-
results-on-quantile-entropy-in-the-past-lifetime/13A4BE3B11CC79D933D4262F2311DA37.

Quan:2023:HTB

[QWGV23] Zhiyu Quan, Zhiguo Wang, Guojun Gan, and Emiliano A. Valdez. On hybrid tree-based methods for short-term insurance claims. Probability in the Engineering and Informational Sciences, 37:597–620, April 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/on-hybrid-
treebased-methods-for-shortterm-insurance-claims/6578C721FEB7D7470C0C159234650C23.

Qiu:2019:EPM

[QWW19] Guoxin Qiu, Lechen Wang, and Xingyu Wang. On extropy properties of mixed systems. Probability in the Engineering and Informational Sciences, 33(3):471–486, July 2019. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/on-extropy-properties-of-mixed-systems/0E3870010D9EAF9FA95225C0625DDE20.

Rao:2015:OPP

[Rao15] B. L. S. Prakasa Rao. Option pricing for processes driven by mixed fractional Brownian motion with superimposed jumps. Probability in the Engineering and Informational Sciences, 29(4):589–596, October 2015. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
BBDB6D4EC814CAE8DE1B558663273293.

Ratsaby:2019:DFS

[Rat19] Joel Ratsaby. On deterministic finite state machines in random environments. Probability in the Engineering and Informational Sciences, 33(4):528–563, October 2019. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/on-deterministic-
finite-state-machines-in-random-environments/BDA3B15306DB4287FB314C5770BFE16A.

Ravid:2021:NLS

[Rav21] Rachel Ravid. A new look on the shortest queue system with jockeying. Probability in the Engineering and Informational Sciences, 35(3):557–564, July 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/new-
look-on-the-shortest-queue-system-with-jockeying/1481E80AF83789F88CFEE86B1B47C8C2.

Rajadhyaksha:2005:CTR

[RB05a] Gauresh Rajadhyaksha and Vivek S. Borkar. Correction to “Transmission Rate Control over Randomly Varying Channels”. Probability in the Engineering and Informational Sciences, 19(4):471, October 2005. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/D79AA8D75B793A1F619E1021C133700A.

See [RB05b].

Rajadhyaksha:2005:TRC

[RB05b] Gauresh Rajadhyaksha and Vivek S. Borkar. Transmission rate control over randomly varying channels. Probability in the Engineering and Informational Sciences, 19(1):73–82, January 2005. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/A9F6DE357ECE5B303FBB01F1F1666D4B. See correc-

tion [RB05a].

Rukhin:2006:SMN

[RB06] Andrew L. Rukhin and Ionut Bebu. Stochastic model for the number of atoms in a magneto-optical trap. Probability in the Engineering and Informational Sciences, 20(2):351–361, April 2006. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
2086D85C41D482536A4F8F0D45DCBA06.

Roubos:2010:ADP

[RB10] Dennis Roubos and Sandjai Bhulai. Approximate dynamic programming techniques for the control of timevarying queuing systems applied to call centers with abandonments and retrials. Probability in the Engineering and Informational Sciences, 24(1):27–45, January 2010. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
100930330D140491A3434DDFED3C9A54.

Roubos:2012:ADP

[RB12] D. Roubos and S. Bhulai. Approximate dynamic programming techniques for skill-based routing in call centers. Probability in the Engineering and Informational Sciences, 26(4):581–591, October 2012. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/5C27CBCFF9C2A0865B16F8A1451F7973.

Robini:2002:CMT

[RBM02] Marc C. Robini, Yoram Bresler, and Isabelle E. Magnin. On the convergence of Metropolis-type relaxation and annealing with constraints. Probability in the Engineering and Informational Sciences, 16(4):427–452, October 2002. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
DFF15C155E37AA7B9209315126F9F179.

Rego:1990:BBT

[Reg90] Vernon Rego. A band and bound technique for simple random algorithms. Probability in the Engineering and Informational Sciences, 4(3):333–344, July 1990. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
00CD6746C410D1BCD785AACC2E21DE35.

Reiman:1990:APR

[Rei90] Martin I. Reiman. Asymptotic properties of a recursive procedure for simultaneous estimation. Probability in the Engineering and Informational Sciences, 4(4):461–475, October 1990. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
17D6EB21DF474C4F9249FE1E6533FBBA.

Reuveni:2014:CT

[Reu14] Shlomi Reuveni. Catalan’s trapezoids. Probability in the Engineering and Informational Sciences, 28(3):353–361, July 2014. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
963790DCBC9E604B19CCA7831AAFE5F9.

Righter:1987:SSA

[Rig87] Rhonda Righter. The stochastic sequential assignment problem with random deadlines. Probability in the Engineering and Informational Sciences, 1(2):189–202, April 1987. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
1ACFFA55CE8865CE6255AB6A7B8E8C7A.

Righter:1992:LSP

[Rig92] Rhonda Righter. Loading and sequencing on parallel machines. Probability in the Engineering and Informational Sciences, 6 (2):193–199, April 1992. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/0912999F818B07DABBFBB12214E8E3FE.

Righter:1996:OSM

[Rig96] Rhonda Righter. Optimal scheduling of multiclass stochastic systems. Probability in the Engineering and Informational Sciences, 10(2):229–241, April 1996. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/5B8C922E4EE69D80D9C5C3503756460F.

Righter:1998:OCD

[Rig98] Rhonda Righter. Optimal computer disk access. Probability in the Engineering and Informational Sciences, 12(2):211–220, April 1998. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/4606E9A7036D57362986272F60FBF2F1.

Righter:2002:OMO

[Rig02] Rhonda Righter. Optimal maintenance and operation of a system with backup components. Probability in the Engineering and Informational Sciences, 16(3):339–349, July 2002. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
2ED0855903012CDA42BE016C50DD2341.

Righter:2011:SSA

[Rig11] Rhonda Righter. Stochastic sequential assignment problem with arrivals. Probability in the Engineering and Informational Sciences, 25(4):477–485, October 2011. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
4A1A45392DA81800BD678238EC22595C.

Rykov:2021:RFK

[RKFI21] Vladimir Rykov, Dmitry Kozyrev, Andrey Filimonov, and Nika Ivanova. On reliability function of a k-out-of-n system with general repair time distribution. Probability in the Engineering and Informational Sciences, 35(4):885–902, October 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/on-reliability-function-of-a-koutofn-system-
with-general-repair-time-distribution/647CDAFFE7EB990CF4A2288B78A6091D.

Rychlik:1993:CTF

[RL93] Igor Rychlik and Georg Lindgren. CROSSREG — a technique for first passage and wave density analysis. Probability in the Engineering and Informational Sciences, 7(1):125–148, January 1993. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/4AECC71666AB6BE4D0D496E30058625C.

Ross:1999:PCS

[RL99] Sheldon M. Ross and Kyle Y. Lin. A pseudo counterexample to a simulation metatheorem. Probability in the Engineering and Informational Sciences, 13(3):329–332, July 1999. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
BDF930168F63D08994D3B4421968974C.

Ross:2001:AVR

[RL01] Sheldon M. Ross and Kyle Y. Lin. Applying variance reduction ideas in queuing simulations. Probability in the Engineering and Informational Sciences, 15(4):481–494, October 2001. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
A098543B8FA34B78E2FB5B9D668791E3.

Ryan:2003:OEP

[RL03] Reade Ryan and Steven A. Lippman. Optimal exit from a project with noisy returns. Probability in the Engineering and Informational Sciences, 17(4):435–458, October 2003. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
0BE66864D8D7B67B99D734E101F697BE.

Ryan:2005:OED

[RL05] Reade Ryan and Steven A. Lippman. Optimal exit from a deteriorating project with noisy returns. Probability in the Engineering and Informational Sciences, 19(3):327–343, July 2005. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
716726C9268DE8A8781D580A3AFFB151.

Reeves:1987:PP

[RMM87] Jaxk H. Reeves, Ashim Mallik, and William P. McCormick. The pricing problem. Probability in the Engineering and Informational Sciences, 1(3):349–366, July 1987. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
5D0137143A64AC0213B0A25E3C2A74F0.

Rajgopal:1994:SPP

[RMS94] Jayant Rajgopal, Mainak Mazumdar, and Thomas H. Savits. Some properties of the Poisson distribution with an application to reliability testing. Probability in the Engineering and Informational Sciences, 8(3):345–354, July 1994. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
0A551D6B6864F4F2E149C723098C8F9D.

Rolski:2005:NID

[Rol05] Tomasz Rolski. A note on the increasing directionally concave monotonicity in queues. Probability in the Engineering and Informational Sciences, 19(1):33–43, January 2005. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
0CAE1C0F7E791D67FBA15CA79E32E051.

Rollin:2022:KBN

[Röl22] Adrian Röllin. Kolmogorov bounds for the normal approximation of the number of triangles in the Erdős–Rényi random graph. Probability in the Engineering and Informational Sciences, 36(3):747–773, July 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/kolmogorov-
bounds-for-the-normal-approximation-of-the-number-of-
triangles-in-the-erdosrenyi-random-graph/115EB2811D3F0AF2947FF3B654B4D91A.

Ross:1987:ATP

[Ros87a] Sheldon M. Ross. Approximating transition probabilities and mean occupation times in continuous-time Markov chains. Probability in the Engineering and Informational Sciences, 1 (3):251–264, July 1987. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/B11AC7D07575CBEA92CB73346C3A0012.

Ross:1987:ART

[Ros87b] Sheldon M. Ross. Approximations in renewal theory. Probability in the Engineering and Informational Sciences, 1(2): 163–173, April 1987. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/77710DA428C8A55E4D26E1DF3A1E7655.

Ross:1987:MER

[Ros87c] Sheldon M. Ross. Are mass extinctions really periodic? Probability in the Engineering and Informational Sciences, 1(1): 61–64, January 1987. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/248F2270F479FF624999D46EB3805E8A.

Ross:1987:E

[Ros87d] Sheldon M. Ross. Editorial. Probability in the Engineering and Informational Sciences, 1(1):1, January 1987. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
62AF8454918B7CFC281229BD88907DBC.

Ross:1987:PRD

[Ros87e] Sheldon M. Ross. Peaks from random data. Probability in the Engineering and Informational Sciences, 1(1):65–68, January 1987. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
2715F0860A1C6DD636339C44E2F3F613.

Ross:1988:NAM

[Ros88a] Sheldon M. Ross. A note on approximating mean occupation times of continuous-time Markov chains. Probability in the Engineering and Informational Sciences, 2(2):267–268, April 1988. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
1E0D9F5BA6E491E26ECCCC02512900D0.

Ross:1988:SPI

[Ros88b] Sheldon M. Ross. A simple proof of instability of a randomaccess communication channel. Probability in the Engineering and Informational Sciences, 2(3):383–385, July 1988. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
518F04DDE0B1E7188F72CECA5526D182.

Ross:1988:SAD

[Ros88c] Sheldon M. Ross. Simulating average delay-variance reduction by conditioning. Probability in the Engineering and Informational Sciences, 2(3):309–312, July 1988. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
252F08046EC2BD8C604B79CBCFE72ABE.

Ross:1989:EMN

[Ros89] Sheldon M. Ross. Estimating the mean number of renewals by simulation. Probability in the Engineering and Informational Sciences, 3(3):319–321, July 1989. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
B4E36645CF6B209C294961FE2E7098F6.

Ross:1990:VRS

[Ros90a] Sheldon M. Ross. Variance reduction in simulation via random hazards. Probability in the Engineering and Informational Sciences, 4(3):229–309, July 1990. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/151DC09B3BE8A41477B16DF9F753D602.

Ross:1990:PRR

[Ros90b] Shledon M. Ross. Processor reordering rules. Probability in the Engineering and Informational Sciences, 4(2):181–186, April 1990. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
CA0536B801B5037A7FD5B47D8B4ED7C5.

Ross:1995:EBS

[Ros95] Sheldon M. Ross. Extending Blackwell’s strengthening of Azuma’s martingale inequality. Probability in the Engineering and Informational Sciences, 9(3):493–496, July 1995. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
747771D51741197E4709A13A8FF0E9B3.

Ross:1998:UIS

[Ros98] Sheldon M. Ross. Using the importance sampling identity to bound tail probabilities. Probability in the Engineering and Informational Sciences, 12(4):445–452, October 1998. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
7EA396A3BF49C923D2830D5E34F8D914.

Ross:1999:ARL

[Ros99a] Sheldon M. Ross. Average run lengths for moving average control charts. Probability in the Engineering and Informational Sciences, 13(2):209–220, April 1999. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
602718175AAC7532C86FC1CD8DFBA93F.

Ross:1999:MWT

[Ros99b] Sheldon M. Ross. The mean waiting time for a pattern. Probability in the Engineering and Informational Sciences, 13(1): 1–9, January 1999. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/EF6753111A35D98824BC69DCA6A561D9.

Ross:2003:IP

[Ros03a] Sheldon M. Ross. The inspection paradox. Probability in the Engineering and Informational Sciences, 17(1):47–51, January 2003. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
542D0FC4CC95D0B66AA8321EE9212B15.

Ross:2003:NIR

[Ros03b] Sheldon M. Ross. A note on the insurance risk problem. Probability in the Engineering and Informational Sciences, 17(2): 199–203, April 2003. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/643B3C6DCFF445859053B55F575E29EB.

Ross:2006:BSD

[Ros06a] Sheldon M. Ross. Bounding the stationary distribution of the M/G/1 queue size. Probability in the Engineering and Informational Sciences, 20(4):571–574, October 2006. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
978836CDBE200DCAC641D645B036CD26.

Ross:2006:MSM

[Ros06b] Sheldon M. Ross. Making simulations more efficient when analyzing Poisson arrival systems and means of monotone functions. Probability in the Engineering and Informational Sciences, 20(2):251–256, April 2006. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
CF5179274AD294C04DA1921BF01D0016.

Ross:2011:MPA

[Ros11] Sheldon M. Ross. The multiple-player ante one game. Probability in the Engineering and Informational Sciences, 25(3): 343–353, July 2011. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/B5F11ED6546694199D110843A6739031.

Ross:2013:MCC

[Ros13] Sheldon M. Ross. A Markov chain choice problem. Probability in the Engineering and Informational Sciences, 27(1):53–55, January 2013. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/4A9668C19570F16A700DF208335DEA25.

Ross:2015:ISO

[Ros15] Sheldon M. Ross. An individual and socially optimal policy minimizing expected flow times. Probability in the Engineering and Informational Sciences, 29(2):147–151, April 2015. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
E5A0B616029935D15953BBFC5C47343D.

Ross:2022:TSW

[Ros22] Sheldon M. Ross. Team’s seasonal win probabilities. Probability in the Engineering and Informational Sciences, 36(4):988–998, October 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/teams-
seasonal-win-probabilities/C04CBFFFD277BE65A85BE6AE48C7E798.

Ross:2023:UCS

[Ros23] Sheldon M. Ross. Using a Chen–Stein identity to obtain low variance simulation estimators. Probability in the Engineering and Informational Sciences, 37(1):232–244, January 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/using-a-chenstein-identity-to-obtain-low-variance-
simulation-estimators/BA57100880C4D632945B3C0C01471117.

Rotondi:1993:TRB

[Rot93] Renata Rotondi. Tests of randomness based on the K-NN distances for data from a bounded region. Probability in the Engineering and Informational Sciences, 7(4):557–569, October 1993. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
98EA6C346A9FB36A6665A084BE506208.

Roy:1990:CLS

[Roy90] Dilip Roy. A characterization of Lindley and Singpurwalla’s bivariate Lomax distribution. Probability in the Engineering and Informational Sciences, 4(4):431–435, October 1990. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
A75689C815D07C09773F0EB5192967A5.

Roy:2002:CML

[Roy02] Dilip Roy. Classification of multivariate life distributions based on partial ordering. Probability in the Engineering and Informational Sciences, 16(1):129–137, January 2002. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
D89A7C887799174F1DF0F5AA99B508E3.

Righter:1989:SMS

[RS89] Rhonda Righter and J. George Shanthikumar. Scheduling multiclass single server queueing systems to stochastically maximize the number of successful departures. Probability in the Engineering and Informational Sciences, 3(3):323–333, July 1989. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
87D87477242A039E0ACBEFDEF2F80B2B.

Rom:1990:MDD

[RS90] Raphael Rom and Moshe Sidi. Message delay distribution in generalized time division multiple access (TDMA). Probability in the Engineering and Informational Sciences, 4(2):187– 202, April 1990. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/BE60E3397C8D6ADB58A678862109C209.

Richter:1994:BST

[RS94a] Rhonda Richter and J. George Shanthikumar. Bounds for stopping times with application to the approximation of distribution functions. Probability in the Engineering and Informational Sciences, 8(1):21–31, January 1994. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
F48107703A6E5D8BBF74A49AB510AF65.

Romeijn:1994:SAA

[RS94b] H. Edwin Romeijn and Robert L. Smith. Simulated annealing and adaptive search in global optimization. Probability in the Engineering and Informational Sciences, 8(4):571–590, October 1994. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/42C76DB626E83B60C9C97733CB70FBCE.

Ravishankar:1995:CLT

[RS95] Krishnamurthi Ravishankar and Suresh Singh. Central limit theorem for time to broadcast in radio networks. Probability in the Engineering and Informational Sciences, 9(2):201– 209, April 1995. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/90E6B5B77D2AA110DA3ED1EC857A7B76.

Righter:1998:IEM

[RS98a] Rhonda Righter and J. George Shanthikumar. Independently expiring multiarmed bandits. Probability in the Engineering and Informational Sciences, 12(4):453–468, October 1998. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
E2F69971CAB971AF64A8C1F8665DD94D.

Rowell:1998:RAD

[RS98b] Ginger Rowell and Kyle Siegrist. Relative aging of distributions. Probability in the Engineering and Informational Sciences, 12(4):469–478, October 1998. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
CCB093EB4763B199C05A0890F90FB71B.

Ross:2000:PEO

[RS00] Sheldon M. Ross and J. George Shanthikumar. Pricing exotic options. Probability in the Engineering and Informational Sciences, 14(3):317–326, July 2000. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
11CFA673D92BD388D5B6BD67BC01DB40.

Ross:2002:HTE

[RS02] Sheldon M. Ross and Sridhar Seshadri. Hitting time in an Erlang loss system. Probability in the Engineering and Informational Sciences, 16(2):167–184, April 2002. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
EFB4781247213B5C1625FE65A8A4FB6C.

Righter:2009:IAC

[RSS09] Rhonda Righter, Moshe Shaked, and J. George Shanthikumar. Intrinsic aging and classes of nonparametric distributions. Probability in the Engineering and Informational Sciences, 23(4):563–582, October 2009. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
B2087ABAD760285DB247713AB2D9C45D.

Ross:1993:SPB

[RSZ93] Sheldon M. Ross, J. George Shanthikumar, and Xiang Zhang. Some pitfalls of black Box queue inference: The case of state-dependent server queues. Probability in the Engineering and Informational Sciences, 7(2):149–157, April 1993. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
C9A09BC59903099FC69C2DD40E9A51C2.

Ross:2009:MLG

[RTW09] Sheldon M. Ross, Henk Tijms, and Shinyi Wu. A model for locking in gains with an application to clinical trials. Probability in the Engineering and Informational Sciences, 23(4): 637–647, October 2009. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/46127E7E354FADEC407C8F5A0E9253CF.

Ridder:1992:SLD

[RW92a] Ad Ridder and Jean Walrand. Some large deviations results in Markov fluid models. Probability in the Engineering and Informational Sciences, 6(4):543–560, October 1992. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
AB35CF2B09DCAA258C538675238A887A.

Ross:1992:MCS

[RW92b] Keith W. Ross and Jie Wang. Monte Carlo summation applied to product-form loss networks. Probability in the Engineering and Informational Sciences, 6(3):323–348, July 1992. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
5652398E149E18E2C7F1A71821E4A36D.

Rieder:1995:CSI

[RW95] Ulrich Rieder and Jürgen Weishaupt. Customer scheduling with incomplete information. Probability in the Engineering and Informational Sciences, 9(2):269–284, April 1995.

CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
FB89CB5B542452815D6004B1E7E421FD.

Shah:1988:SAA

[SA88] Arvind K. Shah and Robert D. Allen. Simpler approximations for the areas under the t distribution. Probability in the Engineering and Informational Sciences, 2(1):139–141, January 1988. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
B1EC572FEB05A516DC9FDDB6CDFC9CAD.

Singham:2018:BCP

[SA18] Dashi I. Singham and Michael P. Atkinson. Boundary crossing probabilities for the cumulative sample mean. Probability in the Engineering and Informational Sciences, 32(2):275–295, April 2018. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/boundary-crossing-probabilities-for-the-cumulative-
sample-mean/F4EC9D6DAFA13D19DDFD4AAE6E4E1CC8.

Sadeh:1996:UDC

[Sad96] Ilan Sadeh. Universal data compression algorithm based on approximate string matching. Probability in the Engineering and Informational Sciences, 10(4):465–486, October 1996. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
F03252060CD00649B02534E0CF74CD68.

Sadeh:1998:BDC

[Sad98] Ilan Sadeh. Bounds on data compression ratio with a given tolerable error probability. Probability in the Engineering and Informational Sciences, 12(2):189–210, April 1998. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
A2EDAF726A22B80AE89F80B7B15BC62B.

Smits:2004:WTC

[SAdK04] Sanne R. Smits, Ivo Adan, and Ton G. de Kok. Waiting time characteristics in cyclic queues. Probability in the Engineering and Informational Sciences, 18(3):299–313, July 2004. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
E690EEB8E1C55CAF8A0D28106B69B7A7.

Ahn:2002:OCT

[sADL02] Hyun soo Ahn, Izak Duenyas, and Mark E. Lewis. Optimal control of a two-stage tandem queuing system with flexible servers. Probability in the Engineering and Informational Sciences, 16(4):453–469, October 2002. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
8241FDC9CA815272FFAA09847615F8EF.

Shojaee:2022:SPG

[SAF22] Omid Shojaee, Majid Asadi, and Maxim Finkelstein. Stochastic properties of generalized finite α-mixtures. Probability in the Engineering and Informational Sciences, 36 (4):1055–1079, October 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/stochastic-
properties-of-generalized-finite-mixtures/D81AA9C7EA4636774C4F553F1C9D4827.

Saito:1998:OSP

[Sai98] Tsuyoshi Saito. Optimal stopping problem with controlled recall. Probability in the Engineering and Informational Sciences, 12(1):91–108, January 1998. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
B2E905E98818BA6FC37E5BE55CF259A8.

Siegrist:1989:CLT

[SAS89] Kyle T. Siegrist, Ashok T. Amin, and Peter J. Slater. The central limit theorem and the law of large numbers for pairconnectivity in Bernoulli trees. Probability in the Engineering and Informational Sciences, 3(4):477–491, October 1989. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
545928BC61FE1BCF298321DCD40851F3.

Sato:1997:OAU

[Sat97] Masahiro Sato. On optimal ammunition usage when hunting fleeing targets. Probability in the Engineering and Informational Sciences, 11(1):49–64, January 1997. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
8BDEC0646E7E96915B8B0176B2FD9457.

Sarhangian:2013:WTA

[SB13] Vahid Sarhangian and Baris Balcioğlu. Waiting time analysis of multi-class queues with impatient customers. Probability in the Engineering and Informational Sciences, 27(3): 333–352, July 2013. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/E4E124E683875BA7031D87CDEB6944B1.

Song:2016:OPC

[SBZ16] Jingjing Song, Xiuchun Bi, and Shuguang Zhang. Optimal portfolio and consumption models under loss aversion in infinite time horizon. Probability in the Engineering and Informational Sciences, 30(4):553–575, October 2016. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
91DD38E7B4CD438563D55FBA7FC64AA7.

Shin:2003:QPN

[SC03] Yang Woo Shin and Bong Dae Choi. A queue with positive and negative arrivals governed by a Markov chain. Probability in the Engineering and Informational Sciences, 17(4):487–501, October 2003. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/63D039CCEBBB91426857AD4AB2E19A8B.

Sun:2023:EBT

[SC23] Hongfang Sun and Yu Chen. Extreme behaviors of the Tail Gini-Type Variability Measures. Probability in the Engineering and Informational Sciences, 37(4):928–942, October 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/extreme-behaviors-of-the-tail-ginitype-variability-
measures/67B059F8EAECAB50B8A2AEA26BE4C29E.

Schassberger:1995:SPN

[Sch95] R. Schaßberger. Stability of polling networks with statedependent server routing. Probability in the Engineering and Informational Sciences, 9(4):539–550, October 1995. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
1C69BD83EDEFCE7558FF4E874E5F37B1.

Schweitzer:1996:BRS

[Sch96] Paul Schweitzer. Book review: Stochastic Models, an Algorithmic Approach, by Tijms Henk C. (Chichester: Wiley, 1994), 375 pages, paperback. Probability in the Engineering and Informational Sciences, 10(3):463–464, July 1996. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
846B3EC92BEA0F7FF189B2F34F00AB95.

Schalekamp:1998:SMP

[Sch98] Frans Schalekamp. Simulation of the matching problem of Montmort. Probability in the Engineering and Informational Sciences, 12(3):325–328, July 1998. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
E8606B9D29DC47F056D6CC991DDECB9B.

Smith:1997:SMR

[SD97] Marcel A. J. Smith and Rommert Dekker. On the (S − 1, S) stock model for renewal demand processes: Poisson’s poison. Probability in the Engineering and Informational Sciences, 11 (3):375–386, July 1997. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/DC72A6A5524D27192806FFFE5EAA49A6.

Saito:2022:NES

[SD22a] Yasuhiro Saito and Tadashi Dohi. Nash equilibrium strategies revisited in software release games. Probability in the Engineering and Informational Sciences, 36(1):105–125, January 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/nash-equilibrium-strategies-revisited-in-software-
release-games/918E313DA0F9709A9BCDA2BC120D6DF2.

Shi:2022:CSD

[SD22b] Zhiyan Shi and Chengjun Ding. A class of small deviation theorems for functionals of random fields on a tree with uniformly bounded degree in random environment. Probability in the Engineering and Informational Sciences, 36(1):169–183, January 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/class-
of-small-deviation-theorems-for-functionals-of-random-
fields-on-a-tree-with-uniformly-bounded-degree-in-random-
environment/1368C3BF069033FFCC6C2C200882EC3E.

Sennott:1989:ACS

[Sen89] Linn I. Sennott. Average cost semi-Markov decision processes and the control of queueing systems. Probability in the Engineering and Informational Sciences, 3(2):247–272, April 1989. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
590B2E28A07A0FBE6FFFB413F439C29B.

Sennott:1991:CDM

[Sen91] Linn I. Sennott. Constrained discounted Markov decision chains. Probability in the Engineering and Informational Sciences, 5(4):463–475, October 1991. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
0E92DED9321D666F4BC944AE481381D7.

Sennott:1993:ACO

[Sen93a] Linn I. Sennott. The average cost optimality equation and critical number policies. Probability in the Engineering and Informational Sciences, 7(1):47–67, January 1993. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
667D8C1ADCA798BF5497E273C80AE61F.

Sennott:1993:CAC

[Sen93b] Linn I. Sennott. Constrained average cost Markov decision chains. Probability in the Engineering and Informational Sciences, 7(1):69–83, January 1993. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
541480633345391975A7D320B59347E9.

Sennott:2001:CAO

[Sen01] Linn I. Sennott. Computing average optimal constrained policies in stochastic dynamic programming. Probability in the Engineering and Informational Sciences, 15(1):103–133, January 2001. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
EC7D23E3094186407B417EDC8EB4F5C1.

Shen:2020:ATP

[SGF20] Xinmei Shen, Mingyue Ge, and Ke-Ang Fu. Approximation of the tail probabilities for bidimensional randomly weighted sums with dependent components. Probability in the Engineering and Informational Sciences, 34(1):112–130, January 2020. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/approximation-of-the-tail-probabilities-for-
bidimensional-randomly-weighted-sums-with-dependent-
components/519A4DA8EC7C34F0BAD049633F712376.

Sakaguchi:2000:GMP

[SH00] Minoru Sakaguchi and Toshio Hamada. A gold-mining problem: optimal backup strategy in computer programs. Probability in the Engineering and Informational Sciences, 14(2): 151–160, April 2000. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/AE380D85BEB80452AA5A665186561A3F.

Sun:2019:SVC

[SH19] Fei Sun and Yijun Hu. Set-valued cash sub-additive risk measures. Probability in the Engineering and Informational Sciences, 33(2):241–257, April 2019. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/setvalued-
cash-subadditive-risk-measures/87F4966E2E7EC02C02D23C04C41CDCCE.

Sahoo:2023:OAP

[SH23] Tanmay Sahoo and Nil Kamal Hazra. Ordering and aging properties of systems with dependent components governed by the Archimedean copula. Probability in the Engineering and Informational Sciences, 37(1):1–28, January 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/ordering-
and-aging-properties-of-systems-with-dependent-components-
governed-by-the-archimedean-copula/93658D6C16CDCE097D7D8EABC494B323.

Shalmon:1988:AGG

[Sha88] Michael Shalmon. Analysis of the GI/GI/1 queue and its variations via the LCFS preemptive resume discipline and its random walk interpretation. Probability in the Engineering and Informational Sciences, 2(2):215–230, April 1988. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
D659F4C3561A1FB09004985CA8AAA957.

Shahabuddin:1998:BRR

[Sha98] Perwez Shahabuddin. Book review: Reliability Engineering, 1996, by Elsayed Elsayed A. (Reading, Mass.: Addison Wesley Longman, Inc.), 737 pp. ISBN 0-201-63481-3. Probability in the Engineering and Informational Sciences, 12(4):533–534, October 1998. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/1440C59AAE958B24C0E471004BD16A6E.

Shaked:2013:CSB

[Sha13] Moshe Shaked. Comments on the survey by Balakrishnan and Zhao. Probability in the Engineering and Informational Sciences, 27(4):445–449, October 2013. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
7401D8FD16020B32DD4E7450B89C28B8. See [BZ13b].

Steckley:2009:FES

[SHM09] Samuel G. Steckley, Shane G. Henderson, and Vijay Mehrotra. Forecast errors in service systems. Probability in the Engineering and Informational Sciences, 23(2):305–332, April 2009. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
55A9DFC32EEDA01B1D3C90089FF1CD8D.

Shioda:2004:MRS

[SI04] Shigeo Shioda and Daisuke Ishii. Monotonicity results for single-server finite-capacity queues with respect to directionally convex order. Probability in the Engineering and Informational Sciences, 18(3):369–393, July 2004. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
0E7D71C7A68B4AC5D19EA33CEA50833C.

Siegrist:1993:OOC

[Sie93] Kyle Siegrist. Optimal occupation in the complete graph. Probability in the Engineering and Informational Sciences, 7(3): 369–385, July 1993. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/CBC3B093BED8BE87BB8728032BC06D5D.

Siegrist:2007:RFS

[Sie07] Kyle Siegrist. Random finite subsets with exponential distributions. Probability in the Engineering and Informational Sciences, 21(1):117–132, January 2007. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
0729E46A17C88C56FA28C891A20B7FC0.

Sherman:2011:OBR

[SK11] Nathan P. Sherman and Jeffrey P. Kharoufeh. Optimal Bernoulli routing in an unreliable M/G/1 retrial queue. Probability in the Engineering and Informational Sciences, 25(1): 1–20, January 2011. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/A176FCF325798962BF7776083B7F1DCF.

Shi:2020:VSB

[SK20]  Daoyuan Shi and Lynn Kuo.  Variable selection for

Bayesian survival models using Bregman divergence measure. Probability in the Engineering and Informational Sciences, 34(3):364–380, July 2020. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/variable-
selection-for-bayesian-survival-models-using-bregman-
divergence-measure/32EE0A310AED9E563F9A8C289658D03E.

Sharma:2023:VDT

[SK23] Akash Sharma and Chanchal Kundu. Varentropy of doubly truncated random variable. Probability in the Engineering and Informational Sciences, 37(3):852–871, July 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/varentropy-of-doubly-truncated-random-variable/
B9327BAAC01ABC23DD77473E603EE01C.

Sherman:2009:MGR

[SKA09] Nathan P. Sherman, Jeffrey P. Kharoufeh, and Mark A. Abramson. An M/G/1 retrial queue with unreliable server for streaming multimedia applications. Probability in the Engineering and Informational Sciences, 23(2):281–304, April 2009. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
93D5E85CE75BE62F797FD3352ECD2318.

Steuber:2012:TRM

[SKL12] Tara L. Steuber, Peter C. Kiessler, and Robert Lund. Testing for reversibility in Markov chain data. Probability in the Engineering and Informational Sciences, 26(4):593–611, October 2012. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
D6B8663EC6F4CAB44A0432AA91562884.

Simchi-Levi:2004:VIS

[SLZ04] David Simchi-Levi and Yao Zhao. The value of information sharing in a two-stage supply chain with production capacity constraints: the infinite horizon case. Probability in the Engineering and Informational Sciences, 18(2):247–274, April 2004. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
99FF76E9EBE8A42C7FBB6949B4346F25.

Smith:1997:AIA

[Smi97a] Marcel A. J. Smith. An approximation of the interval availability distribution. Probability in the Engineering and Informational Sciences, 11(4):451–467, October 1997. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
7F75F6F9FB0B609A914A55FE6C25DE96.

Smith:1997:SLB

[Smi97b] Peter J. Smith. Sequences of linear birth, death, and immigration processes with applications to cascades of optical amplifiers. Probability in the Engineering and Informational Sciences, 11(3):387–394, July 1997. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
DCDD87729EE4CE543243173526DBC878.

Sudyko:2019:AWT

[SNS19] E. Sudyko, A. A. Nazarov, and J. Sztrik. Asymptotic waiting time analysis of a finite-source M/M/1 retrial queueing system. Probability in the Engineering and Informational Sciences, 33(3):387–403, July 2019. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/asymptotic-
waiting-time-analysis-of-a-finitesource-mm1-retrial-
queueing-system/7125FD1EDD6480292216846C9B7D0E9C.

Strauss:1994:POI

[SO94] Todd Strauss and Shmuel Oren. Priority ordering in interruptible electric power service with irreversible early notification. Probability in the Engineering and Informational Sciences, 8 (2):287–290, April 1994. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/743CDE93AFF9BDEA72FC6013E8C4E3EF.

Sobel:1995:LSS

[Sob95] Matthew J. Sobel. Lot sizes in serial manufacturing with random yields. Probability in the Engineering and Informational Sciences, 9(1):151–157, January 1995. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
B0F606BB312D3A786283F48759D0E4CB.

Seshadri:1998:BMD

[SP98] Sridhar Seshadri and Michael Pinedo. Bounds on the mean delay in multiclass queueing networks under shortfallbased priority rules. Probability in the Engineering and Informational Sciences, 12(3):329–350, July 1998. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
36A9D9D1BC093BB638345F4D16DB73A5.

Spieksma:1991:GEA

[Spi91] F. M. Spieksma. Geometric ergodicity of the ALOHA-system and a coupled processors model. Probability in the Engineering and Informational Sciences, 5(1):15–42, January 1991. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
387D0C782ACC4AEF8CB6D0C064249FAD.

Spieksma:2015:CSM

[Spi15] F. M. Spieksma. Countable state Markov processes: nonexplosiveness and moment function. Probability in the Engineering and Informational Sciences, 29(4):623–637, October 2015. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
665C6E373DB000F6E7A46F2EF7D8B7F8.

Sumita:1990:NAC

[SR90] Ushlo Sumita and Maria Rieders. A new algorithm for computing the ergodic probability vector for large Markov chains. Probability in the Engineering and Informational Sciences, 4(1):89–116, January 1990. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
524767C2A74C390EFCFD73ADC6A8CE9D.

Shaked:1987:CSF

[SS87] Moshe Shaked and J. George Shanthikumar. Characterization of some first passage times using log-concavity and log-convexity as aging notions. Probability in the Engineering and Informational Sciences, 1(3):279–291, July 1987. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
94BDC822102419C22FCD03FEA2EDAABA.

Shaked:1989:SRP

[SS89] Moshe Shaked and J. George Shanthikumar. Some replacement policies in a random environment. Probability in the Engineering and Informational Sciences, 3(1):117–134, January 1989. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
103F631366F446FA1BD94E33F8BA3F16.

Sadowsky:1992:MQL

[SS92] John S. Sadowsky and Wojciech Szpankowski. Maximum queue length and waiting time revisited: Multserver G/G/c queue. Probability in the Engineering and Informational Sciences, 6(2):157–170, April 1992. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
1C4A0AB2C785CF3525AC7AC6047958C4.

Shaked:1998:TVO

[SS98] Moshe Shaked and J. George Shanthikumar. Two variability orders. Probability in the Engineering and Informational Sciences, 12(1):1–23, January 1998. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
027C6C671C66E0428C1AC1562AB6852E.

Spreij:2020:DLM

[SS20] Peter Spreij and Jaap Storm. Diffusion limits for a Markov modulated binomial counting process. Probability in the Engineering and Informational Sciences, 34(2):235–257, April 2020. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/diffusion-limits-for-a-markov-modulated-binomial-
counting-process/A4B53513A1C483B314C771EE4CE90CCF.

Shaked:2002:NBP

[SSS02] Moshe Shaked, Fabio Spizzichino, and Florentina Suter. Nonhomogeneous birth processes and l_{∞}-spherical densities, with applications in reliability theory. Probability in the Engineering and Informational Sciences, 16(3):271–288, July 2002. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
7DAE9BA69F3F796193005FE013FE4F7C.

Shaked:2004:UOS

[SSS04] Moshe Shaked, Fabio Spizzichino, and Florentina Suter. Uniform order statistics property and ℓ_{∞}-spherical densities. Probability in the Engineering and Informational Sciences, 18(3): 275–297, July 2004. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/5F01A911DD2C2F015F2CCC92496017B8.

Shaked:2008:UOS

[SSS08] Moshe Shaked, Fabio Spizzichino, and Florentina Suter. Uniform order statistics property and l_{∞}-spherical densities. Probability in the Engineering and Informational Sciences, 22(1): 161–162, January 2008. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/87414B7D8D81DB8CC11D3617AC4F89BE.

Sparaggis:1994:ORS

[ST94] Panayotis D. Sparaggis and Don Towsley. Optimal routing and scheduling of customers with deadlines. Probability in the Engineering and Informational Sciences, 8(1):33–49, January 1994. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
209F1F27527922FAE7C0C61748581929.

Stadje:2007:SSS

[Sta07] Wolfgang Stadje. A storage system with sporadic and continuous clearings. Probability in the Engineering and Informational Sciences, 21(4):539–549, October 2007. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
AEAC1F43A06EC76EED38B25342354F65.

Stark:2011:OPM

[Sta11] Dudley Stark. Oil production models with normal rate curves. Probability in the Engineering and Informational Sciences, 25 (2):205–217, April 2011. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/14817FBA2DA314CA99D8258C286FC12F.

Steele:1987:GMC

[Ste87] Michael Steele. Gibbs’ measures on combinatorial objects and the central limit theorem for an exponential family of random trees. Probability in the Engineering and Informational Sciences, 1(1):47–59, January 1987. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
CA6955E0F7064FEB1B20E6CCFFBCDA54.

Stirzaker:2007:PRR

[Sti07] David Stirzaker. Processes with random regulation. Probability in the Engineering and Informational Sciences, 21(1):1–17, January 2007. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/788AFC98B359BB5DE8F28B040E7AFE7D.

Saunders:2018:LTI

[STI18] David Saunders, Lung Kwan Tsui, and Satish Iyengar. Lower tail independence of hitting times of two-dimensional diffusions. Probability in the Engineering and Informational Sciences, 32(4):522–535, October 2018. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/lower-
tail-independence-of-hitting-times-of-twodimensional-
diffusions/071FA4F7FF567115F76999D61A740F95.

Stolyar:2005:ORO

[Sto05] Alexander L. Stolyar. Optimal routing in output-queued flexible server systems. Probability in the Engineering and Informational Sciences, 19(2):141–189, April 2005. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
E2733AC40E2FF226D8D19284DEB21AE8.

Sharma:2002:PAR

[SVL02] V. Sharma, J. Virtamo, and P. Lassila. Performance analysis of the random early detection algorithm. Probability in the Engineering and Informational Sciences, 16(3):367–388, July 2002. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
4B51666C2001AF0CF25D097C645CF360.

Speijker:2000:OMD

[SvNKC00] Lennaert J. P. Speijker, Jan M. van Noortwijk, Matthijs Kok, and Roger M. Cooke. Optimal maintenance decisions for dikes. Probability in the Engineering and Informational Sciences, 14(1):101–121, January 2000. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
EBEE7E88ECC0FFC5E6113395998B5B24.

Shaked:1995:PSO

[SW95] Moshe Shaked and Tityik Wong. Preservation of stochastic orderings under random mapping by point processes. Probability in the Engineering and Informational Sciences, 9(4):563–580, October 1995. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/CBD7D4A7B2C15733E0F056057B7EC6CC.

Song:2022:TCM

[SX22] Ruiyang Song and Kuang Xu. Temporal concatenation for Markov decision processes. Probability in the Engineering and Informational Sciences, 36(4):999–1026, October 2022.

CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/temporal-concatenation-for-markov-decision-
processes/9F799719DCA333A132DCA3D957B2FD80.

Su:2022:LRM

[SXWW22] Xiaonan Su, Yu Xing, Wei Wang, and Wensheng Wang. Locally risk-minimizing hedging for European contingent claims written on non-tradable assets with common jump risk. Probability in the Engineering and Informational Sciences, 36(3):581–605, July 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/locally-
riskminimizing-hedging-for-european-contingent-claims-
written-on-nontradable-assets-with-common-jump-risk/
1F2D9F6A0CC99C83F48BD9FDB9893AB8.

Shanthikumar:1992:SCS

[SY92a] J. George Shanthikumar and David D. Yao. Spatiotemporal convexity of stochastic processes and applications. Probability in the Engineering and Informational Sciences, 6(1):1–16, January 1992. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/11CBEC850E47B3F71BEF7A54179228AD.

Sigman:1992:FMB

[SY92b] Karl Sigman and Genji Yamazaki. Fluid models with burst arrivals: A sample path analysis. Probability in the Engineering and Informational Sciences, 6(1):17–27, January 1992. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
FAC5DCBC0F7D1C11D8645C2E8DA774D3.

Shen:2015:BPR

[SY15] Yan Shen and Zhisheng Ye. On the behaviors of the percentile residual life for components and for k-out-of-n systems. Probability in the Engineering and Informational Sciences, 29(2): 291–308, April 2015. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/64C290DFFA903D075CFD5BFB4B224869.

Santha:1987:CST

[SZ87] Miklos Santha and Yanjun Znang. Consecutive-2 systems on trees. Probability in the Engineering and Informational Sciences, 1(4):441–456, October 1987. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
A154C9EAC3F9B7653FECAE5C4C4CF628.

Shanthikumar:1999:IBE

[SZ99] J. George Shanthikumar and Michael A. Zazanis. Inequalities between event and time averages. Probability in the Engineering and Informational Sciences, 13(3):293–308, July 1999. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
DF48E5A453F9C526A76CAC64378DE355.

Scheinhardt:2002:TFQ

[SZ02] Werner R. W. Scheinhardt and Bert Zwart. A tandem fluid queue with gradual input. Probability in the Engineering and Informational Sciences, 16(1):29–45, January 2002. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
494F02651FFC29F5CCA09EC138FAD682.

Stadje:2003:UFE

[SZ03] W. Stadje and S. Zacks. Upper first-exit times of compound Poisson processes revisited. Probability in the Engineering and Informational Sciences, 17(4):459–465, October 2003. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
CB38F83918BF903E8A41C38545D7C185.

Stadje:2005:UFE

[SZ05] W. Stadje and S. Zacks. On the upper first-exit times of compound G/M processes. Probability in the Engineering and Informational Sciences, 19(3):397–403, July 2005. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
CBBB7AFE3F3F135842A503D09821E187.

Shi:2018:SMT

[SZF18] Zhiyan Shi, Pingping Zhong, and Yan Fan. The Shannon– McMillan theorem for Markov chains indexed by a Cayley tree in random environment. Probability in the Engineering and Informational Sciences, 32(4):626–639, October 2018. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/shannonmcmillan-theorem-for-markov-chains-indexed-
by-a-cayley-tree-in-random-environment/1ACEE66C4EE262F8827AB468A691CC20.

Tavangar:2011:SAP

[TA11] Mahdi Tavangar and Majid Asadi. On stochastic and aging properties of generalized order statistics. Probability in the Engineering and Informational Sciences, 25(2):187–204, April 2011. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
5C68825A9D7D3FF6E1A16414EDF0A28A.

Tagliani:2000:ITM

[Tag00] Aldo Tagliani. Inverse z transform and moment problem. Probability in the Engineering and Informational Sciences, 14(3): 393–404, July 2000. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/3E501B4479311578EA02AD53C9BAA4A9.

Takahashi:1990:CPS

[Tak90] Haruhisa Takahashi. Coupled processor: A second-order continuous-state-space model. Probability in the Engineering and Informational Sciences, 4(2):277–298, April 1990. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
7E5609BEB6E192BA78E89EB64B9239EE.

Takahara:1996:FPA

[Tak96] Glen K. Takahara. Fixed point approximations for retrial networks. Probability in the Engineering and Informational Sciences, 10(2):243–259, April 1996. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
8B7A897D7B7C2BCF23E99D0872841412.

Tamaki:1998:NDT

[Tam98] Mitsushi Tamaki. A note on the distribution of the time of the first k-record index. Probability in the Engineering and Informational Sciences, 12(3):321–323, July 1998.

CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
B98D0479363FD382347216C5524BE47C.

Tahir:2021:HHD

[TAM^{+}21] Ahsen Tahir, Jawad Ahmad, Gordon Morison, Hadi Larijani, Ryan M. Gibson, and Dawn A. Skelton. Hrnn4F: hybrid deep random neural network for multi-channel fall activity detection. Probability in the Engineering and Informational Sciences, 35(1):37–50, January 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/hrnn4f-
hybrid-deep-random-neural-network-for-multichannel-
fall-activity-detection/11B025BCE0080B8BF6D8E4D5AD4467D0.

Tang:2004:UET

[Tan04] Qihe Tang. Uniform estimates for the tail probability of maxima over finite horizons with subexponential tails. Probability in the Engineering and Informational Sciences, 18(1):71–86, January 2004. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/8F7246CB5292B70FCAC17416FEE1B54E.

Tang:2006:ARP

[Tan06] Qihe Tang. Asymptotic ruin probabilities in finite horizon with subexponential losses and associated discount factors. Probability in the Engineering and Informational Sciences, 20(1): 103–113, January 2006. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/E9F72367B4D2C3151166763963E173C8.

Tardelli:2011:UMP

[Tar11] Paola Tardelli. Utility maximization in a pure jump model with partial observation. Probability in the Engineering and Informational Sciences, 25(1):29–54, January 2011. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
CF48F7C62C27A54081F208AE7A622DC8.

Tardelli:2017:RBS

[Tar17] Paola Tardelli. Recursive backward scheme for the solution of a BSDE with a non Lipschitz generator. Probability in the Engineering and Informational Sciences, 31(2):207–225, April 2017. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
E0AF1D341A18AC7ACAD414CCF839D23D.

Taylor:1987:MFP

[Tay87] Howard M. Taylor. A model for the failure process of semicrystalline polymer materials under static fatigue. Probability in the Engineering and Informational Sciences, 1(2):133–162, April 1987. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/8EE31B9C5CB8E259E83E0DFE7455666D.

Tzavelas:2015:EVD

[TE15] George Tzavelas and Polychronis Economou. Extreme value distributions for biased samples. Probability in the Engineering and Informational Sciences, 29(2):277–290, April 2015. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
D43F20DF32A4168E995F3CAC1E5C5AC0.

Temme:1994:SAI

[Tem94] N. M. Temme. A set of algorithms for the incomplete gamma functions. Probability in the Engineering and Informational Sciences, 8(2):291–307, April 1994. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
F5677268895A0805A0BDF31E4B20A106.

Tandra:2004:JMC

[THM04] Rahul Tandra, N. Hemachandra, and D. Manjunath. Join minimum cost queue for multiclass customers: stability and performance bounds. Probability in the Engineering and Informational Sciences, 18(4):445–472, October 2004. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
38681D2C69FCDA5AAEE771DAF2A9E60E.

Tijms:1992:HFB

[Tij92] Henk Tijms. Heuristics for finite-buffer queues. Probability in the Engineering and Informational Sciences, 6(3):277–285, July 1992. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/61DC6CF8708033D5F04E6D5E13B4AC42.

Tijms:2008:NAM

[Tij08] Henk Tijms. Note on approximations for the multiserver queue with finite buffer and deterministic services. Probability in the Engineering and Informational Sciences, 22(4):653–658, October 2008. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/73AD875F9466DAC141B1C4FFAB9E8A2C.

Timotheou:2016:FNN

[Tim16] Stelios Timotheou. Fast non-negative least-squares learning in the random neural network. Probability in the Engineering and Informational Sciences, 30(3):379–402, July 2016. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
A3F14D593C7A80CDF554E8F44DB01B55.

Triantafyllou:2008:SCS

[TK08] Ioannis S. Triantafyllou and Markos V. Koutras. On the signature of coherent systems and applications. Probability in the Engineering and Informational Sciences, 22(1):19–35, January 2008. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
2B7AB68D18D3EB2563C0DC436537DADD.

Tao:2021:SAB

[TP21] Shuang Tao and Jamol Pender. A stochastic analysis of bikesharing systems. Probability in the Engineering and Informational Sciences, 35(4):781–838, October 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/stochastic-
analysis-of-bikesharing-systems/6D9B41167B5921F35B095EA507C9374E.

Tamaki:1996:FIB

[TS96] Mitsushi Tamaki and J. George Shanthikumar. A fullinformation best-choice problem with allowance. Probability in the Engineering and Informational Sciences, 10(1):41–56, January 1996. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/E8902915E3997C047200C20886D01736.

Tijms:2007:NPW

[TS07] Henk Tijms and Koen Staats. Negative probabilities at work in the M/D/1 queue. Probability in the Engineering and Informational Sciences, 21(1):67–76, January 2007. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
0CD1AB17D5FC1CCDE84914F4E4484A9A.

Toyoizumi:1997:TEA

[TSW97] Hiroshi Toyoizumi, J. George Shanthikumar, and Ronald W. Wolff. Two extremal autocorrelated arrival processes. Probability in the Engineering and Informational Sciences, 11(4): 441–450, October 1997. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/A218D9885A07ED2A93A7DF30D888B4AE.

Turner:1998:EIR

[Tur98] Stephen R. E. Turner. The effect of increasing routing choice on resource pooling. Probability in the Engineering and Informational Sciences, 12(1):109–124, January 1998. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
7150BB39E2DFBC93EA93C645D00F1402.

Tijms:1989:AAB

[TV89] H. C. Tijms and J. C. W. Van Ommeren. Asymptotic analysis for buffer behavior in communication systems. Probability in the Engineering and Informational Sciences, 3(1):1–12, January 1989. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/37F8CEE0897FE188F206A5F33778168E.

Tijms:1991:SNA

[TV91] Henk C. Tijms and Michel C. T. Van De Coevering. A simple numerical approach for infinite-state Markov chains. Probability in the Engineering and Informational Sciences, 5(3): 285–295, July 1991. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/B982BB0DF1150CA3567BAF25E5D95C48.

Tijms:2006:RWS

[TvdW06] Henk Tijms and Jan van der Wal. A real-world stochastic two-person game. Probability in the Engineering and Informational Sciences, 20(4):599–608, October 2006. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
42135CADE3C1BA79B3FEA7C6955C964C.

Tang:2019:SLT

[TYZ19] Ying Tang, Weiguo Yang, and Yue Zhang. The strong limit theorem for relative entropy density rates between two asymptotically circular Markov chains. Probability in the Engineering and Informational Sciences, 33(2):161–171, April 2019. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/strong-limit-theorem-for-relative-entropy-density-
rates-between-two-asymptotically-circular-markov-chains/
CD94A1D55CB759CD57A3C1F9286EAF08.

Tian:2022:MQS

[TZS22] Ruiling Tian, Zhe George Zhang, and Siping Su. On Markovian queues with single working vacation and Bernoulli interruptions. Probability in the Engineering and Informational Sciences, 36(3):616–643, July 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/on-markovian-
queues-with-single-working-vacation-and-bernoulli-interruptions/
69EA21DC2E34E0711A412557FF49884C.

Ulukus:2012:ORP

[UKM12] M. Yasin Ulukus, Jeffrey P. Kharoufeh, and Lisa M. Maillart. Optimal replacement policies under environmentdriven degradation. Probability in the Engineering and Informational Sciences, 26(3):405–424, July 2012. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
C43ED775D3C4DF98542F71AF876431EE.

Vlasiou:2005:ASP

[VA05] M. Vlasiou and I. J. B. F. Adan. An alternating service problem. Probability in the Engineering and Informational Sciences, 19(4):409–426, October 2005. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
82C4EDB8297002336F13899C62A42644.

VanDerHeijden:1988:TMA

[Van88] M. C. Van Der Heijden. On the three-moment approximation of a general distribution by a Coxian distribution. Probability in the Engineering and Informational Sciences, 2(2): 257–261, April 1988. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/090CBFBFA9994D46F2422DB278E83F89.

VanDijk:1992:EBT

[Van92a] Nico M. Van Dijk. An error-bound theorem for approximate Markov chains. Probability in the Engineering and Informational Sciences, 6(3):413–424, July 1992. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
4A29D651AEDBFB35CCF5045D387DECF0.

Vanneste:1992:GAR

[Van92b] Stephan G. Vanneste. A generalized age-replacement model. Probability in the Engineering and Informational Sciences, 6(4):525–541, October 1992. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/5E8188C8C16658A4288141B3ED399E72.

VanDijk:1996:NTM

[Van96a] Nico M. Van Dijk. A note on time monotonicity for performability models. Probability in the Engineering and Informational Sciences, 10(2):299–310, April 1996. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
0182308196333C7E55B0FC4F520F6946.

VanderWiel:1996:ODA

[Van96b] Scott A. Vander Wiel. Optimal discrete adjustments for short production runs. Probability in the Engineering and Informational Sciences, 10(1):119–151, January 1996. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
BFC6C32096437BFEDC50881191272E3B.

VanOyen:1997:MOP

[Van97] Mark P. Van Oyen. Monotonicity of optimal performance measures for polling systems. Probability in the Engineering and Informational Sciences, 11(2):219–228, April 1997.

CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
6D86A8B31E43389532D9AC7C177753BF.

VanMieghem:2001:PSR

[Van01] Piet Van Mieghem. Paths in the simple random graph and the Waxman graph. Probability in the Engineering and Informational Sciences, 15(4):535–555, October 2001. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
48F8F004D949614F52FA1A8E41B5B1EE.

VanLaarhoven:1989:BAS

[VBAK89] P. J. M. Van Laarhoven, C. G. E. Boender, E. H. L. Aarts, and A. H. G. Rinnooy Kan. A Bayesian approach to simulated annealing. Probability in the Engineering and Informational Sciences, 3(4):453–475, October 1989. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
EDD870AF02A72E297D6DFCF604002299.

vanDijk:1988:NEB

[vD88] Nico M. van Dijk. A note on error bounds for approximating transition probabilities in continuous-time Markov chains. Probability in the Engineering and Informational Sciences, 2(4):471–474, October 1988. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/B29C8D0A2F21F0F4D26FBBA584AB2277.

vanDijk:1995:SPE

[vD95] Nico M. van Dijk. A simple performability estimate for Jackson networks with an unreliable output channel. Probability in the Engineering and Informational Sciences, 9(4):603–614, October 1995. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/43B20AE0EA3F2C48FB2C5460BA0EEC76.

vanderHeijden:1987:IAD

[vdH87] M. C. van der Heijden. Interval availability distribution for a 1-out-of-2 reliability system with repair. Probability in the Engineering and Informational Sciences, 1(2):211–223, April 1987. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
07CF8AEE2AB55760A0EF51895C6F3D0B.

vanderHofstad:2001:FPP

[vdHHV01] Remco van der Hofstad, Gerard Hooghiemstra, and Piet Van Mieghem. First-passage percolation on the random graph. Probability in the Engineering and Informational Sciences, 15 (2):225–237, April 2001. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/A9CA85176F5205348B2BCD50C69389F0.

vanderHeijden:2001:WTP

[vdHvHE01] Matthieu van der Heijden, Aart van Harten, and Mark Ebben. Waiting times at periodically switched one-way traffic lanes: a periodic, two-queue polling system with random setup times. Probability in the Engineering and Informational Sciences, 15(4):495–517, October 2001. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
F6D1E70931CB5829913071DF7304DFD6.

VanOmmeren:1987:ARB

[VdK87] J. C. W. Van Ommeren and A. G. de Kok. Asymptotic results for buffer systems under heavy load. Probability in the Engineering and Informational Sciences, 1(3):327–348, July 1987. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
33B3B5027B862D5FD0B28BCBD378389B.

vanderLaan:2011:OMM

[vdL11] Dinard van der Laan. Optimal mixing of Markov decision rules for MDP control. Probability in the Engineering and Informational Sciences, 25(3):307–342, July 2011. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
3E7FE3FF3948292C3012216927760A8B.

vanderMei:2008:PST

[vdMR08] R. D. van der Mei and J. A. C. Resing. Polling systems with two-phase gated service: heavy traffic results for the waiting time distribution. Probability in the Engineering and Informational Sciences, 22(4):623–651, October 2008. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
DC146E055CCD8D047342F60AA6038790.

vanDijk:1988:SBM

[vDTW88] Nico M. van Dijk, Pantelis Tsoucas, and Jean Walrand. Simple bounds and monotonicity: the call congestion of finite multiserver delay systems. Probability in the Engineering and Informational Sciences, 2(1):129–138, January 1988. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
857CADBF0091834F4901F7CFB8C61F86.

vanderWeide:2008:RTE

[vdWSvN08] J. A. M. van der Weide, Suyono, and J. M. van Noortwijk. Renewal theory with exponential and hyperbolic discounting. Probability in the Engineering and Informational Sciences, 22(1):53–74, January 2008. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/6DEA97B9F4D007CD9A7EFEB084310138.

vanderWal:2003:DVO

[vdWY03] Jan van der Wal and Uri Yechiali. Dynamic visit-order rules for batch-service polling. Probability in the Engineering and Informational Sciences, 17(3):351–367, July 2003. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
C5549254144096E30D0CF504F9316B1E.

Vesilo:2022:SPQ

[VHBSW22] Rein Vesilo, Mor Harchol-Balter, and Alan Scheller-Wolf. Scaling properties of queues with time-varying load processes: extensions and applications. Probability in the Engineering and Informational Sciences, 36(3):690–731, July 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/scaling-properties-of-queues-with-timevarying-
load-processes-extensions-and-applications/84573EC2464DD83FA935071B89C13BAC.

Villar:2016:IOI

[Vil16] Sofía S. Villar. Indexability and optimal index policies for a class of reinitialising restless bandits. Probability in the Engineering and Informational Sciences, 30(1):1–23, January 2016. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
895527D87CB65AFFD3A65F266BF4A9A9.

Villar:2018:BSE

[Vil18] Sofía S. Villar. Bandit strategies evaluated in the context of clinical trials in rare life-threatening diseases. Probability in the Engineering and Informational Sciences, 32(2):229–245, April 2018. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/bandit-strategies-evaluated-in-the-context-
of-clinical-trials-in-rare-lifethreatening-diseases/
D43001FE3A525853A6725D6145091E0E.

Vanichpun:2004:WSS

[VM04] Sarut Vanichpun and Armand M. Makowski. When are on-off sources SIS?: conditions and applications. Probability in the Engineering and Informational Sciences, 18(4):423–443, October 2004. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/492318184677112126B4B9E21EF56A17.

Vanderperre:2012:RAR

[VM12] Edmond J. Vanderperre and Stanislav S. Makhanov. Risk analysis of a robot-safety device system subjected to a priority rule. Probability in the Engineering and Informational Sciences, 26 (2):295–306, April 2012. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/EBDCFF48470043C0367FF1DA1DADC843.

Verwijmeren:2000:AEB

[VMB00] Stefan Verwijmeren, Michel Mandjes, and Richard J. Boucherie. Asymptotic evaluation of blocking probabilities in a hierarchical cellular mobile network. Probability in the Engineering and Informational Sciences, 14(1):81–99, January 2000. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
3B09A78307C60ADE3F42A2E0E7F1D8E8.

Voice:2009:SSF

[Voi09] Thomas Voice. Stochastically scalable flow control. Probability in the Engineering and Informational Sciences, 23(4):675–698, October 2009. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/AA578B0D0266ED89A0FA3A2AAE502ED4.

VanDerDuynSchouten:1989:CAT

[VR89] Frank A. Van Der Duyn Schouten and Tjerk Ronner. Calculation of the availability of a two-unit parallel system with cold standby: An illustration of the embedding technique. Probability in the Engineering and Informational Sciences, 3(3): 341–353, July 1989. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/0AC83383281B6E8F2FF3883F6C541D9B.

VanDerWal:1987:IBE

[VS87] Jan Van Der Wal and Paul J. Schweitzer. Iterative bounds on the equilibrium distribution of a finite Markov chain. Probability in the Engineering and Informational Sciences, 1(1):117– 131, January 1987. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/B42A37C0337F53F9D2191E6CC2C4400D.

vanSambeeck:2020:BGP

[vSvDdK^{+}20] Joost H. J. van Sambeeck, Nico M. van Dijk, Wim L. A. M. de Kort, Henk Schonewille, and Mart P. Janssen. Blood group probabilities by next of kin. Probability in the Engineering and Informational Sciences, 34(3):343–363, July 2020. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/blood-group-probabilities-by-next-of-kin/9C72A8B0485EDC9C3544E5C495B6E43B.

VanOyen:1992:OSF

[VT92] Mark P. Van Oyen and Demosthenis Teneketzis. Optimal scheduling of a finite capacity Shuttle under delayed information. Probability in the Engineering and Informational Sciences, 6(1):29–61, January 1992. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
6F3BF1D87343EC942D6E1B7F80C0BD4E.

VanMieghem:2008:WSP

[VT08] P. Van Mieghem and S. Tang. Weight of the shortest path to the first encountered peer in a peer group of size m. Probability in the Engineering and Informational Sciences, 22(1):37–52, January 2008. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/C63256F76EB312395B479149C19A17B9.

VanDijk:1988:PFS

[VV88] N. M. Van Dijk and J. P. Veltkamp. Product forms for stochastic interference systems. Probability in the Engineering and Informational Sciences, 2(3):355–376, July 1988. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
30BBDA9C6B6B84F97F8F5F27C45E2E0C.

Vlasiou:2008:MGP

[VY08] M. Vlasiou and U. Yechiali. M/G/∞ polling systems with random visit times. Probability in the Engineering and Informational Sciences, 22(1):81–106, January 2008. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
CDFC5CA9E058DFC83A1CFD24CF13E542.

vanZuylen:2004:AHG

[vZS04] Anke van Zuylen and Frans Schalekamp. The Achilles’ heel of the GSR shuffle: a note on New Age Solitaire. Probability in the Engineering and Informational Sciences, 18(3):315–328, July 2004. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
08D061078271423B3ECEB5C9AABCBA44.

Winch:1995:OTC

[WAI95] Janice Kim Winch and Benjamin Avi-Itzhak. Ordering of tandem constant-service stations to minimize inprocess stock cost. Probability in the Engineering and Informational Sciences, 9(3):457–473, July 1995. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
FDC4DD55ED285B012368F9F99D49D047.

Wakuta:1998:DCM

[Wak98] Kazuyoshi Wakuta. Discounted cost Markov decision processes with a constraint. Probability in the Engineering and Informational Sciences, 12(2):177–187, April 1998. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
1B6C8C66BA9413C6E63C346DD00225A0.

Wakuta:2000:VVM

[Wak00] Kazuyoshi Wakuta. Vector-valued Markov decision processes with average reward criterion: the multichain case. Probability in the Engineering and Informational Sciences, 14(4):533–548, October 2000. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/EF6F0FD21D60067AF15CDA76E5C5BE73.

Wang:2002:IGG

[Wan02] Chia-Li Wang. An identity of the GI/G/1 transient delay and its applications. Probability in the Engineering and Informational Sciences, 16(1):47–66, January 2002. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
8B2C111D196A8E0B7B90383552C03363.

Wang:2017:RNN

[Wan17a] Lan Wang. The random neural network for cognitive traffic routing and task allocation in networks and the cloud. Probability in the Engineering and Informational Sciences, 31(4): 540–560, October 2017. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/CDE23C4C102B031FFCFF4E66F103095B.

Wang:2017:AVV

[Wan17b] Xingchun Wang. Analytical valuation of vulnerable options in a discrete-time framework. Probability in the Engineering and Informational Sciences, 31(1):100–120, January 2017. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
17C4D4DE15552217EB7040DFA9CFAD10.

Wang:2018:LRO

[Wan18a] Jiantian Wang. Likelihood ratio ordering of parallel systems with heterogeneous scaled components. Probability in the Engineering and Informational Sciences, 32(3):460–468, July 2018. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/likelihood-ratio-ordering-of-parallel-systems-
with-heterogeneous-scaled-components/862F392759323EEDCF3CA4DBD7C5972B.

Wang:2018:PVE

[Wan18b] Xingchun Wang. Pricing vulnerable European options with stochastic correlation. Probability in the Engineering and Informational Sciences, 32(1):67–95, January 2018. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/pricing-vulnerable-european-options-with-stochastic-
correlation/6727C5E9A91826ACA4C57632841D67AD.

Wang:2021:NOS

[Wan21] Yi Wang. G-networks and the optimization of supply chains. Probability in the Engineering and Informational Sciences, 35(1):62–74, January 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/gnetworks-
and-the-optimization-of-supply-chains/96576C5EE1D470C7DFB73EFAA5B82BAA.

Winands:2009:SDP

[WAvHD09] E. M. M. Winands, I. J. B. F. Adan, G. J. van Houtum, and D. G. Down. A state-dependent polling model with k-limited service. Probability in the Engineering and Informational Sciences, 23(2):385–408, April 2009. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
D0F2100FACF3AC2A90440B603733B684.

Wasserman:1996:OSA

[WB96] K. M. Wasserman and Nicholas Bambos. Optimal server allocation to parallel queues with finite-capacity buffers. Probability in the Engineering and Informational Sciences, 10(2): 279–285, April 1996. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/98C00A6DF415CEA4557DD9891A1CF09E.

Willmot:2000:CLD

[WC00] Gordon E. Willmot and Jun Cai. On classes of lifetime distributions with unknown age. Probability in the Engineering and Informational Sciences, 14(4):473–484, October 2000. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
B47C8280A94A2AF1D03E8A118353DA08.

Wei:2022:ARS

[WÇC22] Bo Wei, Sila Çetinkaya, and Daren B. H. Cline. Analytical results on the service performance of stochastic clearing systems. Probability in the Engineering and Informational Sciences, 36(2):217–236, April 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/analytical-
results-on-the-service-performance-of-stochastic-clearing-
systems/D3DF7FCAB755613B76A993FDFF20A1EA.

Wee:1990:OMS

[Wee90] Nam-Sook Wee. Optimal maintenance schedules of computer software. Probability in the Engineering and Informational Sciences, 4(2):243–255, April 1990. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
C94D5F24FAEEBBE17FE8783056D0D223.

Weiss:1988:BBP

[Wei88] Gideon Weiss. Branching bandit processes. Probability in the Engineering and Informational Sciences, 2(3):269–278, July 1988. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
6B4CFFF14103A8B0344300C31A4315E6.

Walraevens:2009:USA

[WFM09] Joris Walraevens, Dieter Fiems, and Marc Moeneclaey. Using singularity analysis to approximate transient characteristics in queueing systems. Probability in the Engineering and Informational Sciences, 23(2):333–355, April 2009. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
85CBC9572391ACFED717EBCB72455271.

Wong:1999:TSE

[WGI99] Eugene W. Wong, Peter W. Glynn, and Donald L. Iglehart. Transient simulation via empirically based coupling. Probability in the Engineering and Informational Sciences, 13(2): 147–167, April 1999. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/6D8E88D31B6F70F64CA376E5EE99800C.

Wei:2020:CAM

[WH20] Linxiao Wei and Yijun Hu. Capital allocation with multivariate risk measures: an axiomatic approach. Probability in the Engineering and Informational Sciences, 34(2):297–315, April 2020. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/capital-allocation-with-multivariate-risk-measures-
an-axiomatic-approach/3533298E7AD050A3D8039CAE0CF67EE6.

Whittle:1988:ALS

[Whi88] P. Whittle. Approximation in large-scale circuit-switched networks. Probability in the Engineering and Informational Sciences, 2(3):279–291, July 1988. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/6204ABE3D679BD7AA3AE507D7C266810.

Whittle:1989:CDP

[Whi89] P. Whittle. A class of decision processes showing policyimprovement/Newton–Raphson equivalence. Probability in the Engineering and Informational Sciences, 3(3):397–403, July 1989. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
DB67934B6AFB23EBEAA4F86B34532921.

Whitt:2000:LCI

[Whi00] Ward Whitt. Limits for cumulative input processes to queues. Probability in the Engineering and Informational Sciences, 14 (2):123–150, April 2000. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/21EACF1FDDB4EA5FF172A49C2987E147.

Winch:1997:BBA

[Win97] Janice Kim Winch. Branch and bound algorithm for ordering of tandem constant-service stations. Probability in the Engineering and Informational Sciences, 11(4):413–439, October 1997. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
6F449249957F13419C21FCE8EE383E2B.

Wu:2018:SLB

[WJ18] Qunying Wu and Yuanying Jiang. Some limiting behavior for asymptotically negative associated random variables. Probability in the Engineering and Informational Sciences, 32(1):58–66, January 2018. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-

engineering-and-informational-sciences/article/some-
limiting-behavior-for-asymptotically-negative-associated-
random-variables/87AF048D0DA84705621B2D2678168C27.

Wu:2022:PVS

[WJYL22] Huojun Wu, Zhaoli Jia, Shuquan Yang, and Ce Liu. Pricing variance swaps under double Heston stochastic volatility model with stochastic interest rate. Probability in the Engineering and Informational Sciences, 36(2):564–580, April 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/pricing-variance-swaps-under-double-heston-
stochastic-volatility-model-with-stochastic-interest-
rate/505E19B0BB1132DC28DFCECE00A7864C.

Whittle:1988:PIN

[WK88] P. Whittle and N. Komarova. Policy improvement and the Newton–Raphson algorithm. Probability in the Engineering and Informational Sciences, 2(2):249–255, April 1988. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
AD1E75CFD423A4912F7DBFE2C299E4D9.

Withers:2009:ESR

[WN09] Christopher S. Withers and Saralees Nadarajah. Expansions for sums of Rayleighs. Probability in the Engineering and Informational Sciences, 23(3):481–487, July 2009. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
971C42DE368A6FFBCFACF3E4BE5EBA92.

Withers:2010:MSR

[WN10] Christopher S. Withers and Saralees Nadarajah. Methods for symmetrizing random variables. Probability in the Engineering and Informational Sciences, 24(4):549–559, October 2010. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
0463CDCCE7ADC963BED9E442B288FEEC.

Withers:2015:GCC

[WN15] Christopher S. Withers and Saralees Nadarajah. A general class of cusum statistics. Probability in the Engineering and Informational Sciences, 29(3):361–384, July 2015.

CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
841AA5BBCC73C70BD537BD238961D2CC.

Wang:2023:VIE

[WNT23] Jiesen Wang, Yoni Nazarathy, and Thomas Taimre. The value of information and efficient switching in channel selection. Probability in the Engineering and Informational Sciences, 37(4):907–927, October 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/value-
of-information-and-efficient-switching-in-channel-selection/
8F62500CFEF282110566193C5531FAF1.

Weichbold:2006:OCG

[WS06] Josef Weichbold and Klaus Schiefermayr. The optimal control of a general tandem queue. Probability in the Engineering and Informational Sciences, 20(2):307–327, April 2006. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
53DD9802AF952EA8C22EA6A655606629.

Wang:2018:VES

[WSX18] Xingchun Wang, Zhiwei Su, and Guangli Xu. The valuation of executive stock options under GARCH models. Probability in the Engineering and Informational Sciences, 32(3):409–433, July 2018. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/valuation-
of-executive-stock-options-under-garch-models/9AE4FED658D32B45BC62481C08DE9F84.

Wang:2005:NEE

[WW05] Chia-Li Wang and Ronald W. Wolff. New estimators for efficient GI/G/1 simulation. Probability in the Engineering and Informational Sciences, 19(2):219–239, April 2005. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
168E35F8C1875D4430DD49459EC81530.

Wu:2018:MDP

[WWH18] Yi Wu, Xuejun Wang, and Shuhe Hu. Moderate deviation principle of sample quantiles and order statistics.

Probability in the Engineering and Informational Sciences, 32(4):603–614, October 2018. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/moderate-
deviation-principle-of-sample-quantiles-and-order-statistics/
95B1FCE6E3FA065321B07089A0EA0A7B.

Wang:2017:PVA

[WWL17] Guanying Wang, Xingchun Wang, and Zhongyi Liu. Pricing vulnerable American put options under jump-diffusion processes. Probability in the Engineering and Informational Sciences, 31(2):121–138, April 2017. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/DB7F2F4E4A8F3164B5F6A67D98742AE6.

Wu:2018:CMC

[WWS18] Yi Wu, Xuejun Wang, and Soo Hak Sung. Complete moment convergence for arrays of rowwise negatively associated random variables and its application in non-parametric regression model. Probability in the Engineering and Informational Sciences, 32(1):37–57, January 2018. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/complete-
moment-convergence-for-arrays-of-rowwise-negatively-
associated-random-variables-and-its-application-in-
nonparametric-regression-model/49277821B2FC0A2893BA3B51FD94E798.

Wan:2004:PCO

[wWW04] Yat wah Wan and Ronald W. Wolff. Picking clumpy orders on a carousel. Probability in the Engineering and Informational Sciences, 18(1):1–11, January 2004. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
6E8F675BE48E940E4DF27B656AA9C40C.

Weicai:2015:SLL

[WWZ15] Peng Weicai, Yang Weiguo, and Shi Zhiyan. Strong law of large numbers for Markov chains indexed by spherically symmetric trees. Probability in the Engineering and Informational Sciences, 29(3):473–481, July 2015. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
F04510A716D3B30FB62C1EC651E6F1E3.

Wang:2022:NAB

[WWZ22] Feng Wang, Xian-Yuan Wu, and Rui Zhu. A note on the asymptotic behavior of the height for a birth-and-death process. Probability in the Engineering and Informational Sciences, 36(1):41–48, January 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/note-
on-the-asymptotic-behavior-of-the-height-for-a-birthanddeath-
process/54D55DCB871F9692332CCC1B7864EF43.

Wang:2023:OAP

[WWZZ23] Chuchu Wang, Wei Wang, Yiying Zhang, and Peng Zhao. Optimal allocation of policy limits in layer reinsurance treaties. Probability in the Engineering and Informational Sciences, 37:546–566, April 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/optimal-
allocation-of-policy-limits-in-layer-reinsurance-treaties/
906BD8F078DA17A472D689E3D3E79EB6.

Wang:2020:CFG

[WXL20] Xingchun Wang, Guangli Xu, and Dan Li. A closed-form GARCH valuation model for power exchange options with counterparty risk. Probability in the Engineering and Informational Sciences, 34(2):279–296, April 2020. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/closedform-
garch-valuation-model-for-power-exchange-options-with-
counterparty-risk/34593DAD751E2D85B52A28D4225D2C0D.

Wang:2018:CEE

[WXWH18] Xuejun Wang, Mengmei Xi, Hongxia Wang, and Shuhe Hu. On consistency of LS estimators in the errors-in-variable regression model. Probability in the Engineering and Informational Sciences, 32(1):144–162, January 2018. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-

engineering-and-informational-sciences/article/on-consistency-
of-ls-estimators-in-the-errorsinvariable-regression-
model/E46F8B2C35D59BE2DF0B45AD8795BDDC.

Xu:2009:DPI

[XC09] Yifan Xu and Xiuli Chao. Dynamic pricing and inventory control for a production system with average profit criterion. Probability in the Engineering and Informational Sciences, 23(3): 489–513, July 2009. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/507BF0767877F5F13D2096327F049C87.

Xia:2010:AOL

[XC10] Cathy H. Xia and Alix L. H. Chow. Asymptotically optimal load distribution for multipath streaming under FEC. Probability in the Engineering and Informational Sciences, 24(4): 509–523, October 2010. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/41464A27243353289F6120D443C10622.

Xu:2015:EDI

[XDZ15] Yifan Xu, Shyamal K. De, and Shelemyahu Zacks. Exact distribution of intermittently changing positive and negative compound Poisson process driven by an alternating renewal process and related functions. Probability in the Engineering and Informational Sciences, 29(3):385–397, July 2015. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
020CC5EA1117DEF26F5A3E2791AD6B60.

Xie:2008:COG

[XH08] Hongmei Xie and Taizhong Hu. Conditional ordering of generalized order statistics revisited. Probability in the Engineering and Informational Sciences, 22(3):333–346, July 2008. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
52B0B8804F9C9389DDFEA79F10B7DB4C.

Xie:2009:OSG

[XH09] Hongmei Xie and Taizhong Hu. Ordering p-spacings of generalized order statistics revisited. Probability in the Engineering and Informational Sciences, 23(1):1–16, January 2009.

CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
A6401E6A8F06370A5C2003279080C462.

Xu:2011:OSH

[XH11] Maochao Xu and Taizhong Hu. Order statistics from heterogeneous negative binomial random variables. Probability in the Engineering and Informational Sciences, 25(4):435–448, October 2011. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/91E6CFC900554074219B0682D5931176.

Xia:2004:LBL

[XLSZ04] Cathy H. Xia, Zhen Liu, Mark S. Squillante, and Li Zhang. Lower bounds for LRD/GI/1 queues with subexponential service times. Probability in the Engineering and Informational Sciences, 18(1):87–101, January 2004. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
1CE9714FD53936AD7C9C7C364EAB71BF.

Xia:2002:OCP

[XMBG02] Cathy H. Xia, George Michailidis, Nicholas Bambos, and Peter W. Glynn. Optimal control of parallel queues with batch service. Probability in the Engineering and Informational Sciences, 16(3):289–307, July 2002. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
ACA1E2A8CEDB44599867417FC97DF8BA.

Xia:2021:PLC

[XMH21] Wanwan Xia, Tiantian Mao, and Taizhong Hu. Preservation of log-concavity and log-convexity under operators. Probability in the Engineering and Informational Sciences, 35(3):451–464, July 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/preservation-
of-logconcavity-and-logconvexity-under-operators/350041454E2DB86802A0D20AAD3BEB38.

Xu:1991:SMT

[Xu91] Susan H. Xu. Stochastically minimizing total delay of jobs subject to random deadlines. Probability in the Engineering and Informational Sciences, 5(3):333–348, July 1991.

CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
47B8A561384D19D80013FA55275B224F.

Xu:2013:COP

[Xu13] Maochao Xu. Comments on “Ordering Properties of Order Statistics from Heterogeneous Populations: a Review with an Emphasis on Some Recent Developments”. Probability in the Engineering and Informational Sciences, 27(4):451–454, October 2013. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/43847B4DCA6BA97B996345ECE9382514. See [BZ13b].

Xu:2023:OSD

[Xu23] Ran Xu. Optimal singular dividend control with capital injection and affine penalty payment at ruin. Probability in the Engineering and Informational Sciences, 37:462–490, April 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/optimal-singular-dividend-control-with-capital-
injection-and-affine-penalty-payment-at-ruin/749A8F777109E45905CDD34178556FF2.

Xia:2018:CWE

[XWTW18] Hao Xia, Yi Wu, Xinran Tao, and Xuejun Wang. The consistency for the weighted estimator of non-parametric regression model based on widely orthant-dependent errors. Probability in the Engineering and Informational Sciences, 32(3):469–481, July 2018. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/consistency-
for-the-weighted-estimator-of-nonparametric-regression-
model-based-on-widely-orthantdependent-errors/5434F65A63C32061FC411886E67A7FB0.

Xing:2021:EVC

[XXN21] Yu Xing, Yuhua Xu, and Huawei Niu. Equilibrium valuation of currency options under a discontinuous model with co-jumps. Probability in the Engineering and Informational Sciences, 35(3):432–450, July 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/equilibrium-

valuation-of-currency-options-under-a-discontinuous-
model-with-cojumps/556ACAA576F34FDAF79C75D8AC06E53C.

Xie:2023:GSA

[XYZC23] Jiayi Xie, Wenguang Yu, Zhimin Zhang, and Zhenyu Cui. Gerber–Shiu analysis in the compound Poisson model with constant inter-observation times. Probability in the Engineering and Informational Sciences, 37:324–356, April 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/gerbershiu-analysis-in-the-compound-poisson-
model-with-constant-interobservation-times/44597BF1D4FA4892D0E7897D6DD0E693.

Xie:2011:SNR

[XZ11] Hongmei Xie and Weiwei Zhuang. Some new results on ordering of simple spacings of generalized order statistics. Probability in the Engineering and Informational Sciences, 25(1): 71–81, January 2011. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/B1532F684950DD3AC85D5476574B614D.

Xu:2018:BAO

[XZP18] Yuqian Xu, Jiawei Zhang, and Michael Pinedo. Budget allocations in operational risk management. Probability in the Engineering and Informational Sciences, 32(3):434–459, July 2018. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/budget-allocations-in-operational-risk-management/
7899FE0AF39CCC14E1D38417D14E4C56.

Yamazaki:1999:QPF

[Yam99] Genji Yamazaki. Quasi-product form to a multinode queueing system with a common exponential setup server. Probability in the Engineering and Informational Sciences, 13(1):103–119, January 1999. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/AAA3296A2CA6FFB05C78C2469E5E9117.

Yamazaki:2000:QPF

[Yam00] Genji Yamazaki. Quasi-product form to a multinode queueing system sharing a state-dependent setup server. Probability in the Engineering and Informational Sciences, 14(4):425–433, October 2000. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/ED0D9A2E4655B6906536E6FA5C8CE399.

Yang:1998:AEP

[Yan98] Weiguo Yang. The asymptotic equipartition property for a nonhomogeneous Markov information source. Probability in the Engineering and Informational Sciences, 12(4):509–518, October 1998. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/A669ADE5705144C3651DF5B9D14AC6BD.

Yang:2020:NSP

[Yan20] Zhaoqiang Yang. A new stopping problem and the critical exercise price for American fractional lookback option in a special mixed jump-diffusion model. Probability in the Engineering and Informational Sciences, 34(1):27–52, January 2020. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/new-stopping-problem-and-the-critical-exercise-
price-for-american-fractional-lookback-option-in-a-
special-mixed-jumpdiffusion-model/9084956C4F00D63B52F72E07635438ED.

Yao:1988:IIL

[Yao88] Y. C. Yao. An improvement over the information lower bound in binomial group testing. Probability in the Engineering and Informational Sciences, 2(3):313–320, July 1988. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
D106EA35B2B0A41D43D6CADB08342AC4.

Yao:1989:OCS

[Yao89] Y. C. Yao. On optimal consecutive k-out-of-n: F systems subject to a fixed product of failure probabilities. Probability in the Engineering and Informational Sciences, 3(2):165–173, April 1989. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
355EACAFAF8B507458B7F51DFF0A6855.

Yi:2022:EJS

[YBL22a] He Yi, Narayanaswamy Balakrishnan, and Xiang Li. Equivalency in joint signatures for binary/multi-state systems of different sizes. Probability in the Engineering and Informational Sciences, 36(4):1027–1054, October 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/equivalency-
in-joint-signatures-for-binarymultistate-systems-of-
different-sizes/13F877901B4E766193A296158439D072.

Yi:2022:SMS

[YBL22b] He Yi, Narayanaswamy Balakrishnan, and Xiang Li. Signatures of multi-state systems based on a series/parallel/ recurrent structure of modules. Probability in the Engineering and Informational Sciences, 36(3):824–850, July 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/signatures-of-multistate-systems-based-on-a-
seriesparallelrecurrent-structure-of-modules/6CA8EB6B60070C34C452BBC2EC734AA2.

Yi:2023:EMS

[YBL23a] He Yi, Narayanaswamy Balakrishnan, and Xiang Li. Equivalency of multi-state survival signatures of multi-state systems of different sizes and its use in the comparison of systems. Probability in the Engineering and Informational Sciences, 37(3):801–817, July 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/equivalency-
of-multistate-survival-signatures-of-multistate-systems-
of-different-sizes-and-its-use-in-the-comparison-of-
systems/CD7D624E305CA651267C31C7B78CF909.

Yi:2023:OMS

[YBL23b] He Yi, Narayanaswamy Balakrishnan, and Xiang Li. Ordered multi-state system signature and its dynamic version in evaluating used multi-state systems. Probability in the Engineering and Informational Sciences, 37(4):873–887, October 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/ordered-multistate-system-signature-and-its-
dynamic-version-in-evaluating-used-multistate-systems/
213EA9FCF1706136D568F8C6B021645F.

Yu:2023:NBA

[YDJ23] Letian Yu, Fraser Daly, and Oliver Johnson. A negative binomial approximation in group testing. Probability in the Engineering and Informational Sciences, 37(4):973–996, October 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/negative-binomial-approximation-in-group-testing/
5473D37A26D956C93812370034A86966.

Yechiali:1988:SSP

[Yec88] Uri Yechiali. Sequencing an N-stage process with feedback. Probability in the Engineering and Informational Sciences, 2 (2):263–265, April 1988. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/9A54D2333D9271D06DD17F3958B2BE94.

Yao:1988:ITI

[YH88] Y. C. Yao and F. K. Hwang. Individual testing of independent items in optimal group testing. Probability in the Engineering and Informational Sciences, 2(1):23–29, January 1988. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
56D8180F0C5B92ADF2D8323462284BB6.

Yu:2006:MCS

[YHZ06] Hai-Bo Yu, Qi-Ming He, and Hanqin Zhang. Monotonicity and convexity of some functions associated with denumerable Markov chains and their applications to queuing systems. Probability in the Engineering and Informational Sciences, 20(1): 67–86, January 2006. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/8DEBEE47659A6E91E2B27C56A87362C0.

Yin:2017:OEE

[Yin17] Yonghua Yin. Optimum energy for energy packet networks. Probability in the Engineering and Informational Sciences, 31(4):516–539, October 2017. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://www.

cambridge.org/core/product/D02F5AABC7A8C30CC573CAE604ADAC03.

Yin:2021:RNN

[Yin21] Yonghua Yin. Random neural network methods and deep learning. Probability in the Engineering and Informational Sciences, 35(1):6–36, January 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/random-
neural-network-methods-and-deep-learning/4D2FDD954B932B2431F4E4A028AA44E0.

Yang:2000:IIS

[YK00] Yoonjung Yang and Georgia-Ann Klutke. Improved inspection schemes for deteriorating equipment. Probability in the Engineering and Informational Sciences, 14(4):445–460, October 2000. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
05C69376D67961B669FAA461A0925B4E.

Yoshida:1996:BEN

[YLB96] Olga S. Yoshida, José G. Leite, and Heleno Bolfarine. Bayes’ estimation of the number of component processes of a superimposed process. Probability in the Engineering and Informational Sciences, 10(3):443–461, July 1996. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
B06D8588302988344156083985CB95A4.

Yue:1992:ODP

[YM92] Wuyi Yue and Yutaka Matsumoto. Output and delay processes in a slotted ALOHA multichannel packet radio network with capture. Probability in the Engineering and Informational Sciences, 6(4):471–493, October 1992. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
CEAFCB0B57F59A5178DE2BACF0724A99.

Yoon:1989:BAT

[YS89] Bok Sik Yoon and J. George Shanthikumar. Bounds and approximations for the transient behavior of continuoustime Markov chains. Probability in the Engineering and Informational Sciences, 3(2):175–198, April 1989. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
845B5BB5E86A9A1ABB5A173A07A7A86D.

Yamazaki:2008:IFE

[YS08] Genji Yamazaki and Tamotsu Sogo. Insensitivity of a front-end Web system. Probability in the Engineering and Informational Sciences, 22(2):273–286, March 2008. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
EB01C4AAEA7412150A67EA9C53D2EB40.

Yu:2006:AMC

[YSM06] Wang Yu, Gerald B. Sheblé, and Manuel António Matos. Application of Markov chain models for short-term generation assets valuation. Probability in the Engineering and Informational Sciences, 20(1):127–141, January 2006. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
4A98FE0890BCB2F1A689737011392A5D.

Yu:2023:ASH

[YT23] Miaomiao Yu and Yinghui Tang. Analyzing a single hyperexponential working vacation queue from its governing difference equation. Probability in the Engineering and Informational Sciences, 37(4):997–1019, October 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/analyzing-
a-single-hyperexponential-working-vacation-queue-from-
its-governing-difference-equation/A8507E2325835C6DBC7B5762DE17E326.

Yu:2021:SCO

[Yu21] Yaming Yu. On stochastic comparisons of order statistics from heterogeneous exponential samples. Probability in the Engineering and Informational Sciences, 35(3):532–537, July 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/on-stochastic-comparisons-of-order-statistics-
from-heterogeneous-exponential-samples/377B2760E28C0CFD2FA5A5927526D9DB.

Ye:2023:PVD

[YWC23] Wuyi Ye, Bin Wu, and Pengzhan Chen. Pricing VIX derivatives using a stochastic volatility model with a flexible jump structure. Probability in the Engineering and Informational Sciences, 37(1):245–274, January 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/pricing-
vix-derivatives-using-a-stochastic-volatility-model-
with-a-flexible-jump-structure/6C4C46FE4B9703FF790CB351D2E909C2.

Yan:2022:OCL

[YWL22] Rongfang Yan, Junrui Wang, and Bin Lu. Orderings of component level versus system level at active redundancies for coherent systems with dependent components. Probability in the Engineering and Informational Sciences, 36(4):1275–1297, October 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/orderings-
of-component-level-versus-system-level-at-active-redundancies-
for-coherent-systems-with-dependent-components/FF67D9DD7BF5819B54D793B2702E3EAE.

Yang:2021:MME

[YWZH21] Xiangyu Yang, Yanfeng Wu, Zeyu Zheng, and Jian-Qiang Hu. Method of moments estimation for Lévy-driven Ornstein–Uhlenbeck stochastic volatility models. Probability in the Engineering and Informational Sciences, 35 (4):975–1004, October 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/method-
of-moments-estimation-for-levydriven-ornsteinuhlenbeck-
stochastic-volatility-models/CCE3B90AD579769CF4A8080FA3E941EA.

Yang:2019:BEV

[YXH19] Jianping Yang, Wanwan Xia, and Taizhong Hu. Bounds on extropy with variational distance constraint. Probability in the Engineering and Informational Sciences, 33(2):186–204, April 2019. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/bounds-on-extropy-with-variational-distance-
constraint/1A5CCB39DCCDB1E64B6FF7E261F20639.

Yang:2023:NES

[YXZ23] Yang Yang, Jiayi Xie, and Zhimin Zhang. Nonparametric estimation of some dividend problems in the perturbed compound Poisson model. Probability in the Engineering and Informational Sciences, 37:418–441, April 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/nonparametric-
estimation-of-some-dividend-problems-in-the-perturbed-
compound-poisson-model/B29DB2B2B862822586CE0867F5758C20.

Yin:1989:WPC

[YZ89] G. Yin and Y. M. Zhu. On W.P.1 convergence of a parallel stochastic approximation algorithm. Probability in the Engineering and Informational Sciences, 3(1):55–75, January 1989. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
10C9B3885A88F4F9A37FE55B498A2B9D.

Yao:1998:MDP

[YZ98] David D. Yao and Shaohui Zheng. Markov decision programming for process control in batch production. Probability in the Engineering and Informational Sciences, 12(3):351–371, July 1998. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
C1A10E7806264B34BED0E07EF22EDC5C.

Yang:2003:MMR

[YZ03] Hailiang Yang and Lihong Zhang. Martingale method for ruin probability in an autoregressive model with constant interest rate. Probability in the Engineering and Informational Sciences, 17(2):183–198, April 2003. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
B83CF586D87E3A4C2A13841B68F57AE9.

Yin:2000:EBO

[YZL00] G. Yin, Q. Zhang, and Q. G. Liu. Error bounds for occupation measure of singularly perturbed Markov chains including transient states. Probability in the Engineering and Informational Sciences, 14(4):511–531, October 2000. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
EC51695EE7FCC25EAD6B4DEC1249FB62.

Yang:2023:AFO

[YZZ23] Jianping Yang, Tian Zhou, and Weiwei Zhuang. Almost first-order stochastic dominance by distorted expectations. Probability in the Engineering and Informational Sciences, 37(4):888–906, October 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/almost-
firstorder-stochastic-dominance-by-distorted-expectations/
66B543670D240684F1A3CE85480B577D.

Zarezadeh:2019:SBI

[ZAE19] S. Zarezadeh, M. Asadi, and S. Eftekhar. Signature-based information measures of multi-state networks. Probability in the Engineering and Informational Sciences, 33(3):438–459, July 2019. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/signaturebased-information-measures-of-multistate-
networks/0959BD6C6F10FFB2AA124218D15C0029.

Zare:2021:TEP

[ZAMB21] Ata G. Zare, Hossein Abouee-Mehrizi, and Oded Berman. Two-echelon production inventory systems with strategic customers. Probability in the Engineering and Informational Sciences, 35(2):258–275, April 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/twoechelon-
production-inventory-systems-with-strategic-customers/
66117ABFD73AA5B9A38A2699CE25CE9A.

Zardasht:2023:SOE

[Zar23] V. Zardasht. On the second-order excess wealth order and its properties. Probability in the Engineering and Informational Sciences, 37(1):135–153, January 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/on-the-
secondorder-excess-wealth-order-and-its-properties/
EF7E445C8259C83130074221B5A1B5B8.

Zazanis:2004:TET

[Zaz04] Michael A. Zazanis. The time-to-empty for tandem Jackson networks. Probability in the Engineering and Informational Sciences, 18(4):493–502, October 2004. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
389C3486A764EEF3246173A5EAFF4BC4.

Zazanis:2007:AVP

[Zaz07] Michael A. Zazanis. Asymptotic variance of passage time estimators in Markov chains. Probability in the Engineering and Informational Sciences, 21(2):217–234, April 2007. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
8012D9225458A62F1FC6827000CD292E.

Zhao:2012:SCL

[ZB12] Peng Zhao and N. Balakrishnan. Stochastic comparisons of largest order statistics from multiple-outlier exponential models. Probability in the Engineering and Informational Sciences, 26(2):159–182, April 2012. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
E78B9295F311496A569B4C83F6324344.

Zhang:2005:SSF

[ZBA05] Jianfeng Zhang, James A. Bucklew, and Fernando L. Alvarado. Search strategies for failure cascade paths in power system graphs. Probability in the Engineering and Informational Sciences, 19(4):507–517, October 2005. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
38E927269FC7A5FF19408797E8B6A29B.

Zayas-Caban:2018:DCS

[ZCA18] Gabriel Zayas-Cabán and Hyun-Soo Ahn. Dynamic control of a single-server system when jobs change status. Probability in the Engineering and Informational Sciences, 32(3):353–395, July 2018. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/dynamic-control-of-a-singleserver-system-when-
jobs-change-status/CAEDF1DA6BC87BF6D37B0D0495BFD416.

Zayas-Caban:2023:SST

[ZCC23] Gabriel Zayas-Cabán and Amy L. Cochran. Scheduling servers in a two-stage queue with abandonments and costs. Probability in the Engineering and Informational Sciences, 37(3):833–851, July 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/scheduling-
servers-in-a-twostage-queue-with-abandonments-and-costs/
07872168CE73AAAD817DD1D255268844.

Zheng:2004:TAL

[ZCJ04] Yuxi Zheng, Xiuli Chao[dagger], and Xiaomei Ji. Transient analysis of linear birth–death processes with immigration and emigration. Probability in the Engineering and Informational Sciences, 18(2):141–159, April 2004. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
3F23520400C7DE3DE7C2BDC6BD22351B.

Zhang:2023:CNZ

[ZCJL23] Jiannan Zhang, Ping Chen, Zhuo Jin, and Shuanming Li. A class of non-zero-sum stochastic differential games between two mean-variance insurers under stochastic volatility. Probability in the Engineering and Informational Sciences, 37:491–517, April 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/class-
of-nonzerosum-stochastic-differential-games-between-
two-meanvariance-insurers-under-stochastic-volatility/
69DDFFF87729E60D251086576D85CE31.

Zhang:2019:DPG

[ZD19] Panpan Zhang and Dipak K. Dey. The degree profile and Gini index of random caterpillar trees. Probability in the Engineering and Informational Sciences, 33(4):511–527, October 2019. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/degree-profile-and-gini-index-of-random-caterpillar-
trees/B9298F6D9FEFA9ED5026F0C0627BCD64.

Zou:2020:EVA

[ZDS20] Jingjing Zou, Richard A. Davis, and Gennady Samorodnitsky. Extreme value analysis without the largest values: what can be done? Probability in the Engineering and Informational Sciences, 34(2):200–220, April 2020. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/extreme-
value-analysis-without-the-largest-values-what-can-
be-done/7CE1039B912AE0232A92362EFB8CDD2D.

Zhang:2020:VSP

[ZDZ20] Yiying Zhang, Weiyong Ding, and Peng Zhao. On variability of series and parallel systems with heterogeneous components. Probability in the Engineering and Informational Sciences, 34(4):626–645, October 2020. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/on-variability-
of-series-and-parallel-systems-with-heterogeneous-components/
3A60653F5BB8050991CDBFE386B041D8.

Zhuang:2007:MSC

[ZH07] Weiwei Zhuang and Taizhong Hu. Multivariate stochastic comparisons of sequential order statistics. Probability in the Engineering and Informational Sciences, 21(1):47–66, January 2007. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
19C71B1A74DEA918D7DA59F22F9D67A2.

Zhuang:2009:SOC

[ZH09] Weiwei Zhuang and Taizhong Hu. Stochastic ordering of a class of symmetric distributions. Probability in the Engineering and Informational Sciences, 23(4):583–595, October 2009. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
60A540A4B81796F7DE6F2E25CCAF224A.

Zhao:1999:CCB

[Zha99] Yiqiang Q. Zhao. Counterexamples to a conjecture on bounds for the loss probability in finite-buffer queues. Probability in the Engineering and Informational Sciences, 13(2):221–227, April 1999. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
6E53F5EC2D3543B9A541BB972359D4F9.

Zhao:2011:PSH

[Zha11]  Peng Zhao.  On parallel systems with heterogeneous

gamma components. Probability in the Engineering and Informational Sciences, 25(3):369–391, July 2011. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
C6A660973622137C4C33DE4AB2F22443.

Zhang:2021:SPC

[Zha21a] Panpan Zhang. On several properties of a class of preferential attachment trees-plane-oriented recursive trees. Probability in the Engineering and Informational Sciences, 35(4):839–857, October 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/on-several-
properties-of-a-class-of-preferential-attachment-treesplaneoriented-
recursive-trees/2DB6C8CDB56B3E35AC0176E3E15E39EC.

Zhang:2021:OED

[Zha21b] Yunxiao Zhang. Optimal energy distribution with energy packet networks. Probability in the Engineering and Informational Sciences, 35(1):75–91, January 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/optimal-
energy-distribution-with-energy-packet-networks/B734DD448FA77BE50AD690944544AE9B.

Zhao:2015:SNR

[ZHZ15] Peng Zhao, Yanni Hu, and Yiying Zhang. Some new results on the largest order statistics from multiple-outlier gamma models. Probability in the Engineering and Informational Sciences, 29(4):597–621, October 2015. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
B668997C7E01CCE1EDAF1A0EED284DEC.

Ziedins:1996:OAC

[Zie96] Ilze Ziedins. Optimal admission controls for Erlang’s loss system with phase-type arrivals. Probability in the Engineering and Informational Sciences, 10(3):397–414, July 1996. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
D5825F042B2DF499D11EB683258AFD9A.

Zhao:2006:SOO

[ZK06] Yao Zhao and Michael N. Katehakis. On the structure of optimal ordering policies for stochastic inventory systems with minimum order quantity. Probability in the Engineering and Informational Sciences, 20(2):257–270, April 2006. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
6F2869AF67BF1C8C7A681DDA0D493460.

Zhang:2002:SCE

[ZL02] Hanqin Zhang and Wei Li. On the speed of convergence to an exponential distribution for the time of the first occurrence of a rare event in a regenerating process. Probability in the Engineering and Informational Sciences, 16(1):121–128, January 2002. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
8410EFCEDB0E790435891AE10995A1B1.

Zhao:2009:SOS

[ZL09] Peng Zhao and Xiaohu Li. Stochastic order of sample range from heterogeneous exponential random variables. Probability in the Engineering and Informational Sciences, 23(1):17–29, January 2009. CODEN ???? ISSN 0269-9648 (print), 1469-

8951 (electronic). URL https://www.cambridge.org/core/
product/834C34CA2DAC0A2C3CC471BBE103736B.

Zografos:2023:RED

[Zog23] Konstantinos Zografos. On reconsidering entropies and divergences and their cumulative counterparts: Csiszár’s, DPD’s and Fisher’s type cumulative and survival measures. Probability in the Engineering and Informational Sciences, 37(1):294–321, January 2023. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/on-reconsidering-
entropies-and-divergences-and-their-cumulative-counterparts-
csiszars-dpds-and-fishers-type-cumulative-and-survival-
measures/1B8FC76B950450B4AF91F368068103D3.

Zaman:2016:AAO

[ZP16] Nabil Zaman and Nicholas Pippenger. Asymptotic analysis of optimal nested group-testing procedures. Probability in the Engineering and Informational Sciences, 30(4):547–552, October 2016. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
66B4CBCADF7490E3BE0190453571C14A.

Zhou:2006:PWR

[ZPD06] Yujia Zhou, Anil Pahwa, and Sanjoy Das. Prediction of weather-related failures of overhead distribution feeders. Probability in the Engineering and Informational Sciences, 20(1): 117–125, January 2006. CODEN ???? ISSN 0269-9648 (print),

1469-8951 (electronic). URL https://www.cambridge.org/
core/product/7FD7DE46EC90F68A3D0AFC9DD0955C3C.

Zuo:2022:TCE

[ZY22] Baishuai Zuo and Chuancun Yin. Tail conditional expectations for generalized skew-elliptical distributions. Probability in the Engineering and Informational Sciences, 36(2):500–513, April 2022. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/
probability-in-the-engineering-and-informational-sciences/
article/tail-conditional-expectations-for-generalized-
skewelliptical-distributions/4726B8F0D23067AD952B394529593FAC.

Zhong:2010:AEP

[ZYL10] Pingping Zhong, Weiguo Yang, and Peipei Liang. The asymptotic equipartition property for asymptotic circular Markov chains. Probability in the Engineering and Informational Sciences, 24(2):279–288, April 2010. CO-DEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
84EE959D0585000EB70A93DBC4177785. See erratum [MY11].

Zhong:2021:SLT

[ZYSZ21] Pingping Zhong, Weiguo Yang, Zhiyan Shi, and Yan Zhang. Some limit theorems of delayed averages for countable nonhomogeneous Markov chains. Probability in the Engineering and Informational Sciences, 35(2):297–315, April 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/journals/

probability-in-the-engineering-and-informational-sciences/
article/some-limit-theorems-of-delayed-averages-for-
countable-nonhomogeneous-markov-chains/C6A25539F0969569C740731A50AF5002.

Zhou:2006:CPD

[ZYWN06] M. Zhou, D. Yang, Y. Wang, and S. Nadarajah. Conditional Poisson distributions. Probability in the Engineering and Informational Sciences, 20(1):95–102, January 2006. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
B596EC6C7E2D6B0BDECA645AD9A9890D.

Zhang:2013:QIP

[ZZ13] Lihong Zhang and Lin Zhao. Quantifying the impact of partial information on Sharpe ratio optimization. Probability in the Engineering and Informational Sciences, 27(3):375–402, July 2013. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (elec-

tronic). URL https://www.cambridge.org/core/product/
023FC7292D42E193B07CFA9FE0EA7DE3.

Zhang:2021:CLO

[ZZZ21] Ting Zhang, Yiying Zhang, and Peng Zhao. Comparisons on largest order statistics from heterogeneous gamma samples. Probability in the Engineering and Informational Sciences, 35(3):611–630, July 2021. CODEN ???? ISSN 0269-9648 (print), 1469-8951 (electronic). URL https://

www.cambridge.org/core/journals/probability-in-the-
engineering-and-informational-sciences/article/comparisons-
on-largest-order-statistics-from-heterogeneous-gamma-
samples/01C05F0DEEE924242AE7609A079DD9A8.